A Bibliography of Publications in Theoretical Computer Science: 2000–2004

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

24 April 2020 Version 1.30

Title word cross-reference

(0, ±1) [122]. (1 + 1) [869]. (1 + ϵ) [1136]. (2 + p) [649]. (2 − 2/(k + 1))^{n} [1116]. (3^{n} − 1) [1618]. (A, B) [1278]. (h, k) [1755]. + [1780, 1272]. 1 [626, 1701]. 1/ϵ [1889]. 2 [1609, 1443, 2146, 2105, 1804, 1687, 452, 1587, 1376, 1399, 1795, 650]. 2^{≤ω} [2046]. 2^{n} [1513, 1618]. 2k [2207]. 2 × 2 × 2 × J [2145]. 3 [652, 1443, 1979, 1128, 615, 1866, 2004, 674]. 4 [1699]. 5 [1380]. 50 [2055]. 8 [37, 1863]. 8 × 8 [15]. = k [1460]. ^{∗} [2089]. ^{,} [2178]. ^{0} [1374]. _{lgol} [1733]. α, β [385]. A ⊗ x = B ⊗ y [1272]. AX − XB^{T} = C [530]. b [78]. β [2037]. C [1005, 606]. d [1357]. δ [1938]. E [254]. ℓ_{1} [1972]. ϵ [1820]. ∃ [403]. ∀∃^{5} [857]. H [2044, 945, 1946, 1755]. K

[240, 1666, 6, 2108, 516, 1116, 128, 945, 323, ; 213, 1392, 1354, 210, 2109, 1755, 453, 1444, ; 776, 823, 604]. l [680]. L(A) = L(B) [374]. ; L^{1}/L^{∞} [1274]. λ [340, 1905, 808, 1769, 690, ; 810, 1712, 727, 166, 803]. ≤ k [1460]. M ; [1380, 564, 1853]. M^{4} − 1 [1380]. μ ; [496, 1169, 1177, 267]. N ; [446, 213, 159, 2218]. ≠ [191]. O(n^{4.03}) ; [1356]. O(n log n) [1568]. Ω ; [426, 2209, 2045, 58, 2091, 1295, 1829]. ⊕ ; [2195]. P [6, 3, 95, 2092, 77]. P (n)/n [818]. π ; [2117, 902]. p × p [1214]. Q [448, 1971]. ; qAC^{0}[2] [565]. R [447]. R^{N} [1141]. R_{1} [1150]. ; R_{1} ⊆ R ⊆ R_{2} [1150]. R_{2} [1150]. σ [2009, 8]. t ; [1759, 1460]. τ [175]. Z [1277]. Z^{2} ; [579, 2064]. Z^{k} [1229]. Z^{n} [978, 1972, 1618]. ; Z_{m} [66].

* [154].

-adic [77]. -algorithm [1568]. -Approximation [1136]. -Arbiters [1755]. -ary [516]. -Automata [8, 2009]. -Balanced [1853]. -body [446]. -bounded [2044]. -calculi [166]. -calculus [340, 2117, 902, 1769, 690, 810, 727, 496, 1169, 1177, 267]. -center [240]. -colorability [1979]. -colorings [945]. -connected [2146, 128, 1618]. -cycles [1820]. -dense [453]. -deterministic [1513]. -diameter [323]. -dim [1376]. -dimension [1804]. -dimensional [1357]. -disconnected [1618]. -dominating [604]. -domination [604]. -enumeration [1971]. -equational [857]. -exclusion [680]. -factorizations [1399]. -function [818]. -functions [1005]. -gram [2218]. -grammars [231]. -hierarchy [171]. -in-a-row [776]. -inertness [175]. -interval [1795]. -invariance [1278]. -languages [2045, 58, 2091, 1829]. -level [452]. -manifolds [1609]. -matching [78]. -matrices [122]. -mode [1460]. -models [1905, 808]. -NFAs [2195]. -oriented [606]. -out-of- [1755]. -point-based [1759]. -power-free [823]. -pruning [385]. -Recursion [1523]. -regular [1666, 1699, 2209]. -Role [1587]. -SAT [652, 649, 1116, 650, 2207]. -satisfiability [2004]. -search [1141]. -selective [95, 3]. -semirings [2089]. -server [2108, 1128, 2105, 2109]. -shifts [2037]. -stage [1866]. -state [37, 159, 1863]. -subdomination [1354]. -subsets [1277]. -sufficient [6]. -system [1273]. -term-DNF [6]. -theory [1712]. -tough [626]. -trees [128, 945, 210, 1687, 1295, 803]. -unification [254]. -Universal [403]. -valued [674]. -wise [1392].

/subsequence [1999]. /substring [1999].

0L [74, 428].

14 [1873]. 16 [1873].

2002 [1848]. 251-260 [549]. 26-surfaces [979]. 272 [2229]. 2D [986, 980].

3 [2220]. 3-SAT [2220]. 3D [89].

3D-manifolds [89].

=L [955].

A. [1549]. abc} [1890]. Abel [617]. abelian [1977]. absolutely [782]. abstract [856, 859, 892, 1180, 1190, 902, 1305, 396, 1296, 1477, 84, 2049, 1298, 659, 413]. abstraction [664, 2118, 2173]. Abstractions [2071, 1249]. accelerated [776]. accelerating [1925]. acceptance [26, 1821]. accepting [1323, 1150, 660]. acceptors [700]. access [2000, 1645, 296, 1871, 2072, 1974, 1092, 202]. Accidental [1694]. accuracy [1497]. Accurate [330, 334]. ACI [164]. ACI-unification [164]. action [1038, 198, 1191, 1779]. actions [728, 492, 183]. activation [287]. active [1777]. Actor [1041]. Acyclic [2188, 214, 9, 605, 2191, 1297]. ad [2158]. Adapting [1712]. adaptive [777, 1163]. Adaptively [1090]. add [1215]. addition [1137, 450, 1803]. additional [520]. Additive [1213, 1575, 858, 1976, 2143, 1506]. Additives [1318, 1312]. adiabatic [1924]. adic [77]. adjacency [978]. adjoint [1908]. adjunctions [1768]. admissibility [1021]. admissions [2057]. admit [633]. admitting [2031]. ADMs [868]. Advanced [1408]. advancing [1175]. adversarial [380]. affine [2203, 459, 1013, 1953, 1945]. affines [511]. against [14, 833, 2078, 1497, 970]. agenda [292]. agent [1553]. agents [228]. aggregation [1371, 2133, 1893]. agreement [1207, 2156]. ahead [560, 2211]. Alarm [259]. aléatoires [436]. aleatory [436]. algebra [1280, 732, 485, 1293, 2166, 1281, 1282, 949, 1279, 837, 688, 1896, 678, 1479, 407, 2148, 124]. Algebraic [1054, 1292, 182, 406, 1373, 1554, 585, 1437, 1962, 1243, 1184, 423, 948, 1642, 1422, 1899, 764, 911, 107, 342, 1290, 720, 1221]. algebras [1772, 761, 1168, 831, 989, 1056, 1057, 18, 832, 1723, 909, 173, 2093, 1275]. algorithm [1976, 1379, 1254, 2186, 1639, 1983, 1203, 2108, 974, 491, 359, 1856, 2220, 1580, 455, 2110, 2130, 1382, 1116, 578, 252, 1569, 1773, 869, 1658, 68, 76, 534, 1511, 1861, 1253, 1263, 78, 368, 917, 311, 1576, 2219, 973, 2021, 1444, 714, 1536, 608, 1991, 1793, 2192, 731, 1468, 151, 1515, 235, 1166, 1568, 1691, 1455, 1845, 446]. Algorithmic [1921, 1848, 2006, 1071, 951]. Algorithms [296, 2005, 1464, 1150, 2003, 210, 228, 966, 1746, 110, 1146, 87, 1081, 697, 1083, 422, 1443, 1085, 1810, 1066, 523, 1752, 2017, 52, 1921, 1919, 48, 624, 1026, 2177, 2226, 9, 1125, 1442, 2136, 2207, 1544, 663, 1487, 1572, 1353, 209, 513, 1071, 712, 1115, 1064, 619, 1892, 1000, 1072, 229, 744, 561, 600, 261, 174, 530, 1051, 1216, 964, 86, 2060, 1491, 1397, 1414, 590, 386, 1297, 501, 2094, 1031, 1401]. alignment [502, 2130, 1359]. all-optical [687, 1085]. all-pairs [1787]. all-to-all [1398]. allocation [228, 1209]. allusion [1255]. Almost [1643, 1392, 1296, 1579, 1028, 39]. Almost-certain [1296]. almost-periodicity [39]. almost-polynomial [1028]. along [1938, 2001]. alphabet [1600]. Alphabetic [438]. alphabets [1642]. Alternating [1446, 85, 1482, 1429, 296, 129, 30, 522]. Alternation [522, 742, 1340]. alternative [2042, 2113]. always [1682]. amalgamation [662]. Amazons [1844]. ambient [1609]. ambients [2069, 201, 1711]. Ambiguity [1501]. ambiguous [2182]. amidst [755]. Ammann [1968]. among [272, 1289, 1267]. amplitude [1552]. analog [2008]. Analogies [572]. analogs [279]. analogue [686]. analyse [1066]. analyses [334]. Analysis [848, 1389, 2027, 1194, 1521, 87, 1254, 1141, 1443, 1939, 170, 1008, 2004, 869, 1658, 274, 1115, 1064, 330, 1745, 1906, 1843, 1581, 1895, 2219, 1776, 1161, 895, 993, 1401, 266]. analytic [935, 806, 268]. Ancestor [2015]. Anderson [2219]. animation [807]. anniversary [2055]. annotation [333]. annotation-directed [333]. annotations [1801, 335]. anonymous [1507]. anthologies [1255]. anti [1399]. anti-Oberwolfach [1399]. anticipating [381]. ants [2058]. any [36]. appeal [1837]. Appendix [1848, 1847]. applicable [1532]. Application [1281, 1528, 382, 340, 613, 1567, 627, 1530, 702, 358, 1582, 341, 407]. Applications [816, 511, 283, 1958, 309, 362, 1006, 1681, 1645, 1683, 102, 1057, 719, 1496, 81, 806, 771, 1058, 668, 2192, 301, 88, 844]. applicative [173]. applied [382, 2054]. Applying [380]. approach [2166, 923, 2167, 999, 423, 1971, 1089, 13, 629, 1598, 236, 1251, 1286, 648, 1787, 1043, 745]. approaches [1072, 888]. Approximability [1656, 1797, 1014, 1359, 509, 1344]. Approximate [2133, 1072, 347, 1530, 571, 1449, 855, 596, 1650, 1655, 1151, 1556, 156, 2032, 1899, 1600, 1626]. Approximating [1743, 1155, 1889, 1028, 306, 764, 111]. Approximation [422, 1085, 1442, 1136, 515, 600, 1493, 60, 1609, 2098, 1886, 1203, 998, 1025, 2207, 1544, 2135, 1662, 1576, 2222, 1528, 136, 5, 692, 1192]. Approximations [577, 573, 1352, 524, 762, 1112]. APX [137]. APX-completeness [137]. Arbiters [1755]. arbitrary [1892, 278, 93, 1536, 1746]. Arborescence [1511]. arborescent [1683]. arc [1801]. Archimedean [34, 2062]. Architecture [1496]. arcs [2077, 848, 585, 862, 1047]. area [1686, 1731, 1266]. argument [1928, 153]. arising [617]. arithmetic [1936, 615, 1326, 984, 475, 915, 919, 1841, 1216, 740, 2052, 1891]. arithmetic-periodicity [1841]. Arithmetical [1661, 2047, 478, 480]. Arithmetics [471, 472]. arrangements [1355]. array [1868]. Arrays [969, 145, 1383, 619, 1529, 1537, 106, 2027, 2023, 852]. artificial [657]. ary [516]. Ascending [1581]. asking [556]. Aspects [721, 1292, 1919, 1804, 2090, 361]. Assembling [1840]. assembly [947, 1236]. assertional [467]. assertions [1485]. assignment [618, 1084, 277, 1442, 1467, 808, 600, 813, 1761]. assignments [1587]. assistance [961]. assistants [221]. associated [1265, 1605]. association [1857]. Associative [757, 70]. asymmetric [1426, 1866, 1771]. Asymptotic [1512, 759, 663, 1458, 1662, 1503, 656]. Asymptotically [140, 1376]. asymptotics [617]. Asynchronous [318, 31, 404, 408, 1209, 862, 842]. asynchronously [836]. ATM [752, 1205]. atomic [404, 1208]. ATOMIX [1840]. atoms [503]. attendants [778]. auctions [2097, 1745]. Augmenting [1121]. Authenticated [2156]. Authentication [991, 2068, 989, 1219, 726]. Author [419, 461, 470, 483, 500, 536, 548, 568, 602, 635, 658, 693, 705, 725, 741, 843, 865, 889, 897, 913, 920, 1023, 785, 801, 814, 834, 929, 957, 976, 994, 1044, 1060, 1095, 1113, 1164, 1222, 1270, 1299, 1321, 1345, 1377, 1403, 1433, 1472, 1488, 1517, 1546, 1578, 1606, 1630, 1663, 1702, 1717, 1736, 1764, 1782, 1808, 1849, 1872, 1901, 1917, 1933, 1955, 1980, 2012, 2040, 2073, 2086, 2111, 2138, 2162, 2184, 2206, 2224]. autobiography [1694]. Automata [31, 1545, 1483, 400, 8, 1486, 1949, 1076, 1323, 2176, 1482, 11, 2121, 854, 1228, 2008, 572, 241, 1429, 2033, 2122, 1644, 239, 695, 62, 1381, 66, 280, 2123, 696, 1, 2124, 1813, 1352, 1450, 1522, 1814, 1605, 944, 2054, 392, 291, 1533, 318, 512, 701, 307, 1456, 2063, 882, 1453, 1591, 1506, 1333, 59, 1768, 463, 347, 1049, 1446, 2181, 2009, 293, 771, 1532, 1531, 1335, 1819, 1492, 1336, 507, 1013, 1878, 2189, 704, 1734, 1586, 1744, 106, 2183, 504, 623, 2064, 1714, 702, 146, 1268, 2125, 863, 1973, 1807, 2036, 1703, 43, 85, 30, 569]. automata [2050, 401, 1368, 1343, 2126, 2193, 1498, 1794, 1295, 27, 460, 1601, 922, 522, 1160, 1858, 28, 1821, 26]. Automata-driven [400]. Automated [1939, 1200]. automates [944]. Automatic [1327, 369, 1423, 962, 468, 1503, 285, 1144]. automaticity [518]. automaton [23, 1119, 1994, 1367, 1121, 37, 2182, 2127, 110]. aux [946]. auxiliary [1446]. avalanche [1267]. Average [2032, 1254, 2004, 556, 229, 2149]. average-case [2004, 229]. avoidance [1148]. Avoiding [1002, 1977, 1584]. axiom [184]. axiomatic [733]. axiomatizability [547]. Axiomatization [1185]. Axiomatizing [828]. Axioms [840, 50, 1324, 1137]. axis [1938].

B [1848, 955]. backtrack [2004, 750]. backtranslation [2060]. backups [1027]. Bahncard [716]. Balance [827]. Balanced [1853, 1127, 971, 1748]. Balances [1668]. balancing [751]. Banach [554]. Bancomat [1209]. banded [1127]. banded-system [1127]. bandwidth [100, 1697, 1516, 1701]. bandwidths [351]. base [2151, 1237]. based [1903, 1482, 1428, 491, 18, 1116, 303, 589, 1759, 1575, 1071, 2201, 199, 174, 583, 1622, 1477, 397, 173, 2035, 1414, 42, 212, 1528, 855, 220, 1252, 1406]. bases [866, 608]. Basic [935, 678, 668, 980, 2212]. basis [1280, 1976, 101, 763]. batch [1996]. batching [1650]. Bayes [1250]. Bayesian [1072]. BB’IW [1952]. BB’IW-logic [1952]. BC [223]. BDDs [393]. be [884, 2205, 960, 1842]. beaver [2141]. Beenker [1968]. behaved [2042, 1779]. behavior [1553, 2201, 759, 1400].

Behavioral [2173]. behaviors [31]. behaviour [663]. Behavioural [1905, 1703, 2072]. behaviours [232]. Bendix [738]. Berarducci [1419]. Bernoulli [824]. Bernstein [1887]. Berstel [1224, 1229]. Besicovitch [1506]. best [383, 2110, 1253, 558, 1108]. best-first [383]. beta [1969, 1575]. beta-integers [1575]. beta-lattices [1969]. Better [1722, 1649]. Better-quasi-orderings [1722]. between [572, 1050, 1685, 790, 800, 1124, 766, 417, 871, 40, 325, 1909, 55, 1316, 1879, 797]. beyond [928]. Bezoutian [1887]. BFS [81]. bi [818, 1238, 1640, 1883]. bi-decoding [1640]. bi-infinite [818, 1238]. bibliography [1033]. bicolored [1978]. bicomponent [2179]. bicriteria [1650, 1743]. Bideterministic [2193]. bidimensional [1439, 1239]. bidirected [122]. bidirectional [703]. bidirectionality [1340]. bijection [1685, 1689]. Bijections [1690]. bijective [1683, 1686]. Bilateral [1469]. bin [708, 1158, 1662]. bin-stretching [708]. Binary [874, 1129, 1856, 365, 793, 1749, 1977, 1522, 779, 2208, 534, 2159, 1970, 576, 1142, 980, 1214, 829, 1029, 1069, 119, 349, 886, 656, 1400, 1215, 2195]. binding [330, 2171]. binding-time [330]. binoid [1592, 1796]. binoids [1796]. bio [1633, 1986]. bio-operations [1633, 1986]. BioAmbients [2118]. biochemical [2115, 2117]. biographie [932]. Biography [1549, 932]. biological [1281, 627, 2118]. biology [2116, 1799]. biology-inspired [1799]. biomolecular [1063, 2114]. bipartite [753, 609, 1567, 1260, 633, 301, 1273]. Birkhoff [686]. bisection [1993, 1699]. bisimilarity [170, 766, 1585]. bisimilarity-checking [1585]. Bisimulation [844, 2115, 1729, 1773, 493, 2074, 398, 163, 841]. bisimulation-like [493]. Bisimulations [677]. bit [1810, 1214, 616]. bit-vector [1810]. bits [2024]. bivariate [1326]. black [1919, 590]. BLL [1941]. BLL-like [1941]. block [577, 336, 2021]. blocking [1264]. Board [1210, 1052, 1079, 1105, 1153, 1196, 1244, 1283, 1310, 1346, 1360, 1404, 1417, 1473, 1490, 1518, 1547, 1563, 1607, 1631, 1664, 1678, 1718, 2199, 2225, 1884, 1918, 1934, 1947, 1981, 1992, 2025, 2053, 2087, 2095, 2119, 2163, 2174, 1737, 1766, 1783, 1791, 1822, 1862]. body [446]. Böhm [1712]. boolean [1415, 903, 186, 676, 566, 1493, 1482, 101, 1860, 1288, 758, 278, 1392, 2131, 571, 872, 762, 142, 2024, 599, 1800]. boosting [1252]. Borel [476, 1447]. boring [350]. bottleneck [153]. bound [1152, 707, 1759, 113, 1109, 529, 919, 1260, 425, 148, 171, 301]. boundary [143, 563, 614]. Bounded [1815, 1171, 2042, 430, 1349, 2044, 875, 1269, 1577, 2034, 554, 141, 1338, 82, 149, 358, 222, 91, 970, 1568, 133, 522, 38]. Bounded-connect [1815]. bounded-width [149, 222]. boundedness [2208, 1771]. Bounding [1997, 351]. Bounds [2000, 1699, 652, 527, 996, 1324, 72, 1993, 653, 1457, 1451, 651, 59, 2139, 159, 1354, 1103, 219, 885, 964, 2035, 1788, 1561, 886, 2022, 567, 1740, 247]. boxes [304]. Boyer [1166]. BPA [766]. BPP [642, 766, 286]. Branch [425, 276]. Branching [1729, 2189, 839, 642, 1393, 1163, 171]. branching-time [1393]. breadth [363]. breadth-first [363]. Breaking [1454, 1839]. brève [932]. bridge [1316]. brief [932]. Brier [553]. Broadcast [726, 1535, 1467, 298, 1115, 1208]. broadcasting [1086]. broadcasts [2137]. Brownian [1394]. Bruijn [619]. BSS [41]. Büchi [1381]. buffer [2099, 583]. build [1877]. Building [1874]. bundles [991]. buses [561]. busy [2141]. bytecode [1432]. Byzantine [1445].

C [333, 885]. c.e [788]. cache [442]. caching [2101, 2153, 713, 2103, 2100]. cacti [1683].

CafeOBJ [1037]. calculating [630]. calculi [1038, 809, 166, 1558, 1912, 5, 692, 685]. Calculs [746]. calculus [255, 340, 2117, 1481, 1419, 423, 1719, 496, 486, 53, 904, 677, 412, 1188, 1704, 529, 666, 902, 1770, 1769, 690, 810, 727, 1169, 1315, 1177, 267, 1703, 342, 256, 109, 1727, 413, 488, 1478]. call [1085, 94, 421, 808, 1177]. call-by-name [94]. call-by-value [808, 1177]. Can [117, 207, 960]. cancellation [1893]. cancellative [142]. cannot [2205]. Canonical [1119, 1968, 780]. Cantor [474]. capabilities [360, 1927]. capital [1520]. card [990]. cardinals [2049]. cards [533]. carriers [770]. carte [309]. cartesian [1695, 1170, 899]. cas [367]. cascade [1810]. cascades [1816]. case [1254, 1130, 2004, 1213, 1864, 229, 1497, 2035, 80]. cases [1520, 1553, 1756, 1214]. CASL [1054, 1059]. Catalan [2143]. Catalytic [1803]. catch [809]. catch/throw [809]. Categorical [1714, 1857, 673, 1305, 321, 2079, 1562]. categories [747, 1303, 1304, 1705, 542, 909, 911]. Category [2170, 1183, 1170]. caterpillars [1701]. causal [265, 851, 1114, 177]. Cayley [281, 1761]. CCS [1475, 851]. CCS-like [1475]. Cell [1629, 1799]. cells [89]. cellulaires [944]. Cellular [2063, 1453, 1368, 11, 2121, 2122, 1644, 66, 280, 2123, 2124, 1605, 944, 2054, 318, 512, 307, 1456, 1591, 1506, 1367, 37, 771, 1336, 1586, 504, 623, 1268, 2125, 1973, 2126, 1498, 1794, 2127]. center [240]. centroskew [1886]. centrosymmetric [1886]. certain [845, 578, 1296]. certainly [350]. certificates [769]. chain [581, 784]. chain-free [581]. chains [1085, 1480, 1146]. Chaitin [426, 1009]. challenges [1837]. challenging [820]. change [1001, 745]. Channel [616, 600, 1761]. channels [1645]. chaos [280]. Characterisations [1748]. characteristic [1334, 571, 1256]. characteristics [1414, 185].

Characterization

[62, 2144, 788, 357, 23, 1522, 1117, 948, 1642, 129, 565, 1320, 1625, 2212, 1916, 2197]. Characterizations [1807, 406, 1248, 355]. characterized [960]. Characterizing [609, 282, 1922, 1684, 1790]. chart [704]. charts [1735]. Chebyshev [2009, 8]. Chebyshev-polynomials [8]. checker [528]. checking [1172, 1140, 1711, 578, 1553, 496, 467, 1585, 465, 167, 267]. Checkpointing [1198]. checkpoints [1198]. checks [672]. chemical [523]. Chemins [436]. chessboard [1762]. Chew [487]. chi [1188]. Chilean [2154]. chip [75, 773]. choice [1771, 1710, 1626]. chord [972]. chordal [639, 1752, 1684]. chromatic [1457, 1835]. Chu [1316]. Church [1318]. ciliate [1633, 1986]. ciliates [1236]. circuit [873]. circuits [691, 1073, 207, 142]. circulants [272]. circular [1355, 106]. circular-ones [1355]. clairvoyant [2107]. class [1436, 735, 392, 316, 2058, 435, 331, 2094, 2022, 1401]. class-constrained [2022]. classes [1903, 1502, 357, 1749, 1633, 1603, 1850, 513, 540, 3, 95, 1422, 2031, 355, 226, 120, 1112, 737, 1827]. classical [1076, 1026, 1313, 55, 1708, 2075, 56, 1255, 1922]. classifiability [11]. classification [1303, 512, 1207, 1687, 1003]. Classifying [1985]. clausal [661, 268]. clause [1135]. clause-variable [1135]. climbing [1991]. clique [2142, 1457, 1464, 604]. clique-width [1464]. cliques [867]. Clobber [1847, 1832]. clock [1710]. Clos [1866, 567]. closed [1637, 1010, 850, 2031, 41, 1170]. Closedness [724]. closing [1735]. Closure [1633, 2045, 1627, 34, 448, 1374, 294, 1449, 742, 1461, 1020]. Clustering [1706, 1051]. CNN [2109]. co [1459]. co-idempotent [1459]. Coalgebra [926, 1903, 1774, 2168, 65, 925, 547, 339]. Coalgebraic [1725, 1474, 2166, 923, 539, 928, 686, 924, 2172]. Coalgebras [1220, 1301, 1774, 2167, 540, 1705, 183, 542, 2169, 543, 546, 737, 922]. coalitions [833, 967]. cocycle [579]. code [1200, 784, 1018, 2061]. Codes [2213, 443, 1645, 1975, 316, 833, 825, 1673, 1437, 587, 1509, 794, 427, 826, 119, 631, 1145, 1640, 824]. codicity [1571]. Coding [937, 1680, 99, 617, 1756, 1002]. coefficients [917]. cographs [613, 1317]. Coherence [1314, 1937]. coinduction [1722]. coinductive [861, 1769, 1703]. Collage [1416]. collapsing [2115, 1364, 1792]. collection [1430, 1431]. collective [67]. Colonies [287]. Color [480, 120]. colorability [1979]. colored [1959, 277, 1683]. colored-edges [1959]. coloring [1129, 753, 1384, 655, 421, 1530, 1338]. colorings [9, 945, 628]. coloured [831]. colouring [2159, 1385]. colourings [610]. column [1685]. column-convex [1685]. Combination [662]. combinations [1174]. combinator [1188, 1192]. Combinatorial [624, 1605, 1848, 795, 176, 458, 205, 2207, 817, 1759, 345, 1837, 629, 756, 1493, 824]. Combinatorics [1567, 1390, 935]. combinatorist [1694]. combinators [326, 842]. combinatory [1772]. Combined [1999]. Combining [2101, 1304, 927, 1460]. comma [1509]. comma-free [1509]. commerce [1218]. committee [997]. Common [1054, 60, 1463, 1552, 1655, 353, 1491]. communicating [444, 875, 413, 428, 836]. Communication [640, 1426, 67, 444, 2063, 68, 1485, 726, 1078]. communications [1203, 1115]. commutation [820, 2094]. commutative [1312, 1137, 1320]. comonad [927]. Compact [781, 2200, 1612, 1813, 367, 1909, 1915]. companion [2100]. comparable [1541]. comparative [163]. Comparator [365, 297]. Comparing [197, 1038]. Comparison [790, 1658, 1219, 1721, 723, 1178, 2021, 987, 888]. comparisons [1281]. compartments [2118]. compatible [1348, 940, 1745, 487]. competition [1836, 2141]. Competitive [87, 1757, 2104, 1745, 2103, 2108, 2136, 1763]. competitiveness [244]. compilation [332, 963]. compiler [333]. compilers [1221]. Compiling [1293]. complematarity [361]. complementary [622]. complements [168]. Complete [91, 44, 1643, 241, 2146, 2, 96, 1603, 291, 778, 450, 494, 541, 478, 1051, 1701, 826, 349, 1452, 374, 1148, 1399, 1914, 674, 7]. completely [1474]. Completeness [412, 808, 402, 813, 469, 137, 861, 503, 1978, 630, 1308, 1708, 952, 1725, 670]. Completing [1509]. Completion [1640, 905, 906]. complex [1900]. complexes [114, 1629]. complexité [367]. complexities [2215, 345, 1825, 798]. Complexity [1436, 1098, 344, 657, 1837, 2147, 1064, 1558, 1497, 150, 1667, 1225, 1857, 2115, 238, 1140, 1939, 761, 2157, 1005, 1326, 502, 170, 101, 1752, 1921, 696, 2044, 2160, 104, 2221, 254, 512, 790, 763, 791, 1330, 2063, 1117, 2142, 1101, 421, 237, 2034, 1655, 1421, 1363, 1261, 1661, 63, 799, 2188, 1598, 2047, 1465, 279, 1374, 1334, 1988, 1001, 464, 2149, 278, 800, 1531, 1162, 598, 82, 1942, 1585, 2131, 1393, 1584, 980, 268, 721, 1651, 1753, 794, 2032, 873, 1967, 795, 1946, 300, 2002, 310, 775, 796, 1017, 2126, 1541, 2194, 407, 792]. complexity [1795, 133, 2051, 1078, 88]. Complexity-theoretic [344, 279]. component [581, 514, 1466]. components [1441, 1470, 287]. composed [1452]. Composing [262]. Composition [1240, 679, 2217, 1779]. Compositional [928, 1183, 2167]. comprehensive [2065]. compressed [1416]. Compression [96, 2, 1532, 1819, 1582, 212, 1528]. Comput [418, 2229, 2226, 1545, 96, 1781, 95, 2227, 692, 460, 1471, 2228]. Computability [1611, 83, 2148, 1930, 1799, 637, 1015, 789, 593, 1019]. computable [782, 2043, 1010, 791, 1011, 1013, 1878, 917, 1020, 266]. computably [1022]. computation [190, 1936, 1636, 340, 497, 1063, 1860, 1927, 1987, 2150, 1075, 37, 2131, 557, 1926, 1068, 1070, 811, 410, 1134, 1018]. computation-universal [37]. Computational [1609, 271, 761, 1313, 1804, 873, 1140, 1005, 104, 1330, 63, 2047, 1988, 1942, 980, 138, 545, 1946, 975, 1795, 133]. computationally [2183]. Computations [533, 1226, 2046, 448, 490, 1190, 1049, 690, 312, 1395, 1016, 1788, 207]. Computer [476, 915, 1831, 1058, 2049]. Computers [1211]. Computing [1801, 1351, 1507, 769, 1075, 871, 2024, 1062, 1082, 2157, 1984, 64, 2071, 1773, 1088, 1206, 1261, 1792, 1197, 608, 42, 1065, 1264, 1146, 1922, 446]. concatenation [525, 1291, 1243, 70]. concept [722, 1930]. concerning [2208]. Concrete [489]. Concurrency [1734, 939, 2070, 1911]. Concurrent [517, 4, 418, 938, 949, 908, 2074, 411, 842]. condensation [1958]. condensing [699]. condition [884, 1895]. Conditional [794, 793, 1036, 586, 1868, 792, 798, 130]. Conditioning [13]. Conditions [1273, 186, 1494, 26, 1728]. Conference [473]. confidence [1071]. configurations [2121, 1139, 984, 1963, 1961]. confined [1682]. conflict [1645, 1048]. conflict-free [1048]. Confluence [1769, 166, 175]. Confluency [1177]. confluent [598]. conformance [2078]. conformon [1799]. conformon-P [1799]. congested [880]. congruences [761, 1168, 578, 1492, 2093, 729, 1646, 841, 1601, 1479]. Congruential [168]. conjecture [204, 456, 1890, 1871, 1439, 120, 480]. conjectures [2123, 2208]. Conjugacy [1126, 1519, 1231, 1749]. conjunctions [2034]. Conjunctive [189, 1461, 1536, 2007]. connect [1815]. connected [2146, 128, 353, 1618, 2161, 1854, 2222, 604].

Connection [54, 2056, 323, 423, 1165]. Connections [1267, 1056]. connectives [1480, 1288]. Connectivity [1966, 1602, 983, 654, 148]. connectors [939]. coNP [642]. conquer [1265, 247]. consecutive [81]. consequence [1725]. conservative [1189, 2125]. conserving [1456, 1591, 1268]. consistency [1107]. Constant [690, 886, 2209, 1468, 1990, 1259]. constant-round [1259]. constant-space [1468]. constrained [1203, 1156, 1589, 850, 2152, 2022]. constraint [491, 1100, 2113, 433, 513, 908, 278, 2210, 1263, 1069, 1298, 662, 495, 656, 409, 2061, 1215]. constraint-based [491]. constraints [2203, 753, 1288, 600, 668, 331, 721]. Constructible [114, 771, 1356]. Constructing [656, 2146, 1125, 427]. construction [365, 696, 1480, 54, 574, 727, 5, 692]. Constructions [1645, 2000, 1119, 538, 1605, 445, 53, 1365, 1462, 256]. Constructive [892, 1613, 178, 1006, 1008, 917, 151, 1628]. constructor [1428]. constructor-based [1428]. consumption [260]. containment [189]. content [1515]. Context [1869, 697, 1639, 19, 1671, 444, 1470, 2178, 499, 783, 601, 734, 1501, 1447, 666, 457, 330, 245, 449, 1286, 138, 315, 152, 331, 789, 308, 857, 139, 1504]. context-free [697, 1639, 19, 1869, 1671, 444, 1470, 2178, 499, 601, 734, 457, 245, 449, 138, 308]. context-free-language [331]. Context-freeness [1869]. contiguous [774]. contingency [314, 2145]. Continuation [199]. Continuation-based [199]. continues [367]. continuity [356, 887]. Continuous [1927, 1951, 1301, 1176, 974, 1874, 1906, 1907, 633, 1483, 1652, 83]. Continuous-time [1927, 1652]. continuum [925]. contract [993]. contractile [2026]. contracting [1788]. contraction [1181]. Contractivity [918]. Contrast [213].

Contrast-optimal [213]. Control [749, 1274, 196, 2072, 971, 2158, 202, 276]. control-driven [196]. controllers [1550, 839]. convention [1973]. convergence [996, 1647, 2208, 1616, 1896, 1003, 1746, 1031, 1798]. Conversation [2187]. conversion [687, 18, 1237]. Converting [1333]. convex [2017, 1580, 1125, 1351, 1233, 1685, 1971, 986, 117, 118, 121]. Convexity [274, 620]. convolution [236]. Conway [1143]. Cook [103]. cooperating [581, 519, 1460]. cooperative [228]. coordinated [237]. Coordinating [198]. coordination [196, 199]. copies [2024]. coping [743]. Coq [2082, 2083, 811]. corecursion [544, 1307]. cores [641]. correctly [1890]. Correctness [990, 1166, 675]. Correlation [1800, 1267]. Correspondence [849, 874, 38, 1485, 1879]. Correspondences [55]. corridor [453]. Corrigendum [418, 2229, 692, 1471]. coset [119]. cost [707, 72, 1233, 1588, 2161, 446, 896]. cost-effective [896]. Costs [521, 2189, 1092]. countable [781, 1102, 1626]. countably [1622]. Counter [1120, 1139, 1994, 1585, 1160]. counterexample [1672]. counters [1130, 772, 1802]. Counting [90, 1811, 1291, 945, 477, 1974, 230, 643, 1846, 238, 2221, 314, 157, 2139, 1103, 1584, 153]. covariances [1394]. Covarieties [541]. cover [362, 695]. cover-automata [695]. Covering [10, 654, 1662, 632, 119, 631]. covers [588, 606]. CPS [94, 1709]. CREW [446]. Crick [535, 1543, 954, 1499, 1134, 1638]. criteria [1533, 723, 2035]. critical [242]. Cross [2161, 1494]. Cross-monotonic [2161]. crossing [301]. crossings [273]. Cryptographic [719, 1186, 534, 1641]. Cryptographically [872]. cryptography [360, 1865, 213]. cryptosystems [439]. CSP [922]. Cube [1396, 293]. cube-product [293]. cubes [2051]. cubic [137, 1699]. Cumulative [1230]. curve [1743]. curves [2019, 1899, 107, 108]. custom [915]. customizable [1202]. Cut [49, 385, 2076, 1909, 952, 255]. Cut-elimination [49, 952]. Cut-free [255]. cuts [384]. cutsets [1514]. cutting [2202]. cycle [117, 415]. cycle-sum [415]. cycles [2056, 612, 1820, 1356]. cyclic [947, 378, 2064, 1701, 2061]. cyclicity [570]. Cylindric [739].

d [1051, 1682]. D-finite [1682]. D0L

[1661, 275, 535, 1543, 954, 1499, 1134, 1638]. DAC [1046]. Dagstuhl [1848]. d’arbres [942]. dark [105]. dart [2026]. dart-free [2026]. Data [96, 1572, 1249, 849, 856, 99, 804, 2229, 1130, 1430, 223, 1785, 2, 1372, 1101, 274, 679, 489, 1049, 1121, 442, 1908, 1582, 910, 1161, 162, 226, 2132, 498, 225, 1413, 2228]. data-dependent [1049]. data-independent [1049]. database [1777, 498, 414]. databases [1371, 739, 409]. datalog [1707, 1555, 191]. dataspaces [1430]. dates [1996]. David [532]. DAWGs [1495]. dcpos [1616]. deadlines [2137, 1337]. deadlocking [678]. decentralized [261]. Decidability [474, 1487, 254, 1543, 465, 877, 420, 482, 1777, 1952, 1633, 1456, 1953, 40, 1268, 1725, 374, 955]. Decidable [854, 357, 1167, 1352, 429, 766, 1756, 1189, 669, 2080]. decide [356, 887]. Deciding [459, 1817, 493, 2182, 1227, 29]. Decimation [518, 1118]. Decimation-invariant [518]. deciphering [427]. Decision [1482, 758, 1409, 762, 1098, 2044, 526, 1249, 885, 43, 558, 1108, 1252]. deck [533]. declarative [1476, 1781, 736]. declustered [2132]. declustering [1376]. decodable [2218]. decoding [311, 1640]. decomposability [232]. Decomposable [1387, 826, 1161]. Decomposition [92, 134, 947, 563, 1466, 257].

Decompositions [1385, 1810]. decorated [406]. decrements [772]. Dedication [900]. Dédié [1229]. deduction [1058, 812, 860]. Deductive [391]. default [1140]. defaults [2081]. defeasible [1179]. defect [1230, 2090, 821, 1238]. definability [903, 2047, 478]. definable [238]. defined [2121, 1502, 1986, 1492]. defining [2167]. définis [70]. definite [100]. definition [1342, 1929]. definitions [49]. Deformation [1898, 633]. degree [1338]. degrees [434, 529, 1368]. Delaunay [1357]. Delay [290, 825, 1305, 427]. Delay-insensitivity [290]. delays [1798, 1640]. Deleting [2180]. Deletion [2001, 1870, 1389, 294]. deletions [2129]. demonic [675]. dénombrable [746]. denotational [1937, 902, 396, 1286, 413, 484]. denoting [1796]. Dense [349, 408, 453]. Density [1212, 1093]. Dependable [1197, 1202]. dependence [1776]. dependencies [164, 739, 417]. dependency [127]. Dependent [1710, 667, 1107, 735, 1049, 2153, 671, 1634]. dependent-type [671]. depth [381, 295, 742]. deque [1871]. Derivation [738]. derivational [598]. derivations [1559, 167, 1562, 674]. Derivatives [1812, 1005, 1119]. describing [368]. Description [1234]. descriptions [757, 2195]. Descriptive [791, 1598]. design [752, 1063, 892, 2201, 24, 151]. designs [1397]. destinies [2044]. Detecting [321, 1077, 881]. detection [1881, 1751, 774]. determinacy [1754]. determined [1617, 1688]. Determining [508, 63, 1150]. determinism [799]. Deterministic [750, 1124, 2220, 1116, 1642, 950, 2078, 2181, 1513, 1263, 78, 809, 1560, 1495, 2214, 1498, 482, 110, 836]. deterministically [395]. Determinization [1123, 2186]. developable [117]. development [655]. devices [767]. DFA [2192]. DFAs [159]. diagnosis [1199]. diagonal [943, 118]. diagram [985]. diagrams [627, 321, 121, 1182]. dial [711]. dial-a-ride [711]. diameter [323, 1088, 116]. Dichotomies [1850]. dictionaries [1813]. difference [382, 1135, 381]. differences [572]. different [1175, 800, 1092]. differentiable [1525]. differential [652, 1719, 1703]. difficult [1067]. difficulties [1647]. Diffusion [1445]. digit [1118, 1237]. digital [981, 983, 236, 17, 980, 1627, 1629]. digraph [1091]. digraphs [1602, 298, 363, 1335]. dilation [82]. dim [1376]. dimension [1379, 1107, 1738, 554, 1804, 1151, 1599]. dimensional [1855, 145, 422, 1141, 1964, 827, 1965, 2122, 432, 2124, 2062, 1088, 1158, 2058, 2144, 1838, 744, 37, 561, 1529, 1537, 505, 1586, 1357, 1584, 1394, 17, 2125, 300, 775, 121, 1498, 1794, 88, 1440, 628]. dimensions [755, 1051, 304]. Diophantine [763, 1802, 2198]. Direct [701, 1365, 1000]. Directed [1205, 935, 1685, 333, 1492, 2191, 1914]. direction [1156, 296, 1206, 1507]. directions [1974]. discipline [218]. disconnected [1618]. Discovering [1256, 1255]. discovery [1246, 1257, 150]. discret [511]. Discrete [1175, 2054, 1964, 305, 1966, 503, 1352, 1450, 1614, 2019, 951, 1279, 1121, 986, 962, 839, 1435, 1753]. discretization [987]. discriminating [1815]. discriminating-reverse [1815]. discrimination [1419]. disjoint [640, 1341, 349, 1174]. disjunction [12]. disjunctions [1101, 1000, 668]. Disjunctive [1275, 497]. disk [1383, 2152]. dispenser [1636]. display [237]. Disproving [456]. disruptions [1415]. Distance [1972, 2140, 1100, 613, 2152, 59, 1556, 1744, 325, 798, 604]. distance-hereditary [604]. Distances [1050, 2129]. Distributed [1535, 681, 1184, 2000, 1382, 581, 519, 1460, 1206, 1090, 263, 1390, 404, 2072, 228, 1581, 964, 498, 1202]. Distribution [1970, 577, 1107, 2000, 1142, 2027, 1503, 108]. distribution-dependent [1107]. distributional [91]. distributions [2143, 1437, 962, 1435, 1069, 119]. Distributivity [1705]. disturbed [1141]. Divergence [665]. Divide [1265, 247]. Divide-and-conquer [1265, 247]. dividing [1380]. division [114, 1214, 2092]. divisor [1552, 1491]. DNA [1062, 1082, 1551, 1064, 38, 361, 42, 1600, 1264]. DNF [6, 179, 220]. do [1260, 1896]. Does [883, 1604, 633]. Domain [2077, 324, 1006, 1911, 1183, 1904, 1569, 907, 563, 1435, 2035, 1414, 912, 1240, 1309]. domain-models [1183]. domain-partitioning [2035]. domain-theoretic [1309]. domains [1961, 1620, 1625]. Dominance [1288]. dominating [2142, 2222, 604]. domination [1502, 1648, 1354, 2010, 1508, 604]. Domineering [15]. Domino [1961, 1978, 441, 1565, 2066, 1568]. domino-shuffling [1565]. Dot [594, 742, 295]. Dot-depth [742, 295]. Double [366, 1976, 622, 1916]. double-loop [1976, 622]. doubling [28]. doubly [1889]. down [1593, 1252]. DPDA [420]. drawability [748]. drawing [1266, 1440]. drawings [273]. drift [722, 1133]. driven [196, 400]. dropping [336]. dual [701, 119, 631]. duality [1428, 1093]. duals [845]. due [442]. duotaire [1838]. durational [728, 492]. dyadic [1293]. DyC [333]. Dynamic [939, 560, 513, 312, 2064, 1375, 2133, 880, 1382, 363, 333, 1374, 971, 574, 558, 2156, 836]. Dynamical [1401, 1325, 459, 1573, 2205, 1574]. Dynamically [453]. Dynamics [2058, 280, 1605, 1963, 1591]. DynTC [1374].

E. [2054]. easier [1140]. easy [1603, 1104]. EATCS [933]. eco [1634, 1852]. eco-grammar [1634]. Edalat [1620]. Edge [1695, 753, 211, 1697, 273, 1434, 2159, 514, 641, 1389, 1338, 134, 349, 1385]. edge-bandwidth [1697]. edge-coloring [1338]. edge-colouring [2159]. edge-disjoint [349]. Edge-isoperimetric [1695]. edges [1156, 1959, 1689]. edit [1658]. Editorial [22, 33, 46, 98, 112, 181, 203, 217, 1052, 1079, 1105, 1153, 1196, 1210, 1244, 1283, 1310, 1346, 1360, 1404, 1417, 1473, 1490, 1518, 1547, 1563, 1607, 1631, 1664, 1678, 1718, 2199, 2225, 1884, 1918, 1934, 1947, 1981, 1992, 2025, 2053, 2087, 2095, 2119, 2163, 2174, 645, 988, 2185, 1935, 1284, 835, 890, 1737, 1766, 1783, 1791, 1822, 1862]. Edmonds [1511]. EDT0L [877]. EF [465]. effect [821]. Effective [1012, 1014, 232, 1845, 896]. Effectively [1010]. Effects [1414, 545]. Efficiency [1864, 862, 1922]. Efficient [67, 523, 2122, 51, 9, 597, 619, 561, 1508, 30, 2132, 1081, 1254, 1227, 752, 2017, 1026, 2130, 2207, 1773, 534, 140, 1263, 2158, 311, 257, 2021, 1444, 29, 400]. efficiently [393, 556]. eigenproblem [1886]. eigenvalue [643, 1273]. eight [1577]. eight-vertex [1577]. ELAN [1034]. election [1973]. Electronic [2020, 1218, 1864, 2187]. element [1552]. elementary [1441, 816, 280, 1953, 1408, 1407]. elements [1538, 1503]. Eliminating [1462]. elimination [2076, 1684, 49, 952]. eliminative [1288]. Elliot [608]. Elliot-MacMahon [608]. elliptic [2019]. embedded [1331]. Embedding [1156, 349, 1018]. embeddings [2218, 272, 1868]. Emerging [1400]. EMPA [418, 4]. empirical [387]. empty [1533, 1372, 1446]. EMSO [2050]. EMSO-logic [2050]. enclosure [919]. Encoding [691, 1315, 1880]. end [475]. Endgame [1148]. energy [2158]. enforcing [1262]. engine [1064]. engineering [1043]. engines [2103]. enhanced [2117]. Enhancements [1819]. Enhancing [2202].

Ensuring [469]. Entailment [1915, 718]. entangled [1260]. entanglement [1984, 1077]. entities [1816]. entropy [1551, 1670, 789]. enumerable [426, 1470, 1870, 1022, 551, 41, 355]. Enumerating [353, 1811, 1958]. enumeration [6, 1971, 1477]. environment [1445]. Episturmian [879, 1519, 445]. epsilon [1559]. equation [1888, 1272, 530, 248, 1068, 1660, 2198]. Equational [1427, 540, 547, 1042, 1280, 126, 1302, 923, 1526, 1510, 1307, 1342, 857]. Equations [1391, 652, 1811, 1137, 336, 1786, 1674, 1802, 443, 912, 1703, 685]. équilibrées [233]. equilibrium [1864]. Equilogical [1875]. Equivalence [1796, 1189, 1729, 497, 1431, 861, 578, 1773, 1386, 719, 1721, 284, 293, 321, 161, 877, 44, 1562, 420]. equivalences [406, 1653, 493, 464]. equivalent [159]. erasing [224]. Ergodicity [66]. Eric [2055]. Errata [1545]. Erratum [2226, 96, 1781, 95, 2227, 460, 2228]. Error [881, 1843, 1759, 1003, 1202]. Error-detecting [881]. errors [574, 743]. espace [367]. Espaces [367]. essence [905]. Esterel [733]. estimates [1003]. estimation [643, 1497, 27, 460, 896]. estimations [110]. estimator [2157]. ET0L [1657, 308]. Euclidean [1505, 1088, 1401]. Eulerian [1335]. European [1833]. evaluating [407]. evaluation [527, 2215, 940, 837, 2158, 894]. evaluations [571]. evaluator [332]. Even [1131]. event [2077, 13, 1279, 1048, 839, 2050, 1779]. eventualities [1296]. every [1552, 1569, 179]. everywhere [466]. evidence [205]. Evolution [2029, 1141, 1647, 767, 830]. Evolutionary [1068, 205, 1066, 869, 218]. evolutive [1751]. ex [724]. ex-identification [724]. Exact [1323, 221, 1016, 2005, 1087, 1658, 917, 1216, 565, 2032]. exactly [1569]. example [782, 1209]. examples [807, 1768, 551]. exceptions [896]. exchange [435]. exclusion [680, 1755]. exclusive [907, 220]. exclusive-or [220]. excursion [1394]. exercise [941]. Exhausting [1308]. exhaustive [2122, 1844]. exist [1260]. existence [282, 1626, 1273]. existential [61]. expansion [1767, 256, 641]. expansions [220]. expected [964]. Experience [1923]. expert [2110]. expiration [713]. Expired [1430]. Explaining [684]. Explicit [1193, 729, 166, 671, 335]. exploration [2155]. explore [1251]. exponent [242]. Exponential [1660, 1652, 2031, 267, 1798]. exponentially [1762, 1977, 1151]. exponentiation [1097]. expression [682]. expressions [1812, 1526, 1592, 1796, 1465, 71, 1256, 407]. Expressive [1370, 2069, 333, 676]. Expressiveness [1132, 1471, 1046, 1424, 409]. expressivity [1393]. Extended [50, 360, 1021, 379, 697, 1365, 2049]. Extending [332, 1747]. extension [271, 1532]. extensional [166]. Extensions [1144, 1590, 475]. external [519]. extract [1816]. extracted [2075]. Extraction [1820, 402]. extremal [1214, 119, 1275]. Extremality [115]. extremity [1552].

F [1874, 1649]. FA [1802]. face [117]. face-cycle [117]. facility [1757, 2161]. factor [347, 2217]. factorials [2160]. Factoring [93]. factorization [819, 1760, 1528, 1676, 1900]. factorizations [821, 1399]. Factorizing [1201]. factors [1595, 506, 597, 1028, 781, 822]. faculty [2057]. failure [1190]. failures [1724, 681]. fairness [1305, 261]. false [1445]. Families [246, 1986, 1292, 1048, 873, 84]. family [1513, 1996]. fan [116]. fanin [207]. Fanout [433]. farthest [1351]. Fast [408, 530, 1336, 1894, 304, 144, 1891, 1983, 212, 1515]. Faster [1087, 1893]. Fault [1199, 240, 455, 2071, 1881, 580, 259, 1201, 1202].

Fault-diagnosis [1199]. fault-free [580]. fault-tolerant [455, 2071]. fault/error [1202]. faults [72, 514, 641, 580, 207]. faulty [1389, 2024, 1868]. feared [1274]. feasible [176, 277, 1553, 2212]. feedback [1533, 306]. festoon [1971]. few [2055]. fibers [1888]. Fibonacci [145, 2151]. fibrations [1768]. field [34, 647, 446]. fields [2009, 283, 93, 41]. fifty [743]. file [713]. files [68]. filled [359]. filter [172]. filters [1459]. FIN [555]. FIN-learning [555]. final [1301, 65, 925]. find [1253]. Finding [1750, 1785, 510, 1556, 614, 1358, 1254, 1463, 52, 1919, 1655, 1162, 917, 1753, 1568, 110, 1406]. fine [476, 754, 1239]. Finitary [669]. Finite [239, 1, 1545, 1813, 1738, 1533, 1816, 587, 829, 847, 993, 1295, 184, 628, 1280, 1076, 1123, 241, 1429, 1682, 695, 1760, 1119, 820, 1786, 477, 828, 431, 2045, 1453, 1331, 1817, 1190, 825, 654, 1305, 1049, 2181, 515, 2009, 493, 1532, 1531, 1335, 1819, 1820, 283, 1013, 1878, 766, 93, 2183, 24, 1074, 427, 826, 863, 703, 2036, 398, 1452, 43, 85, 30, 1240, 569, 1720, 110, 27, 460, 1640, 1276, 1821]. finite-connectivity [654]. Finite-dimensional [628]. finite-memory [43]. Finite-state [1738, 1816, 993, 1817, 493, 766, 24, 703]. finitely [271, 1502, 18, 2182, 1479]. finiteness [1494]. firewalls [992]. firing [1997, 75, 773, 505, 1863, 281, 870]. first [854, 383, 1293, 940, 1709, 1775, 363, 478, 742, 1708, 322, 952, 2080, 676, 1847]. first-order [1293, 940, 1709, 478, 742, 1708, 322, 676]. fitness [1746]. five [1470, 872]. fixed [1668, 1379, 1851, 2142, 1135, 1464, 1102, 507, 177, 410, 1515, 765, 812]. fixed-order [410]. Fixpoint [901, 169, 912, 1713, 327]. flat [627, 407]. flexible [585, 774]. flips [118]. floating [1212, 1214, 1215]. floating-point [1212]. florilegium [2054]. floundering [674]. Flow [2016, 330, 1882, 992, 1788, 136, 2050, 1031, 1779]. Folding [2059]. fonctions [367]. Forbidden [1854, 822]. Forbidding [1262, 1657, 139]. force [56, 446]. Forced [120]. Forcing [1570]. forests [1760, 629]. Foreword [1823, 187, 1548, 1258, 1024, 898, 2041, 694, 914, 329, 1693, 1154, 787, 1564, 2175, 2096, 1294, 995, 636, 1271, 559, 289, 2088, 537, 921, 958, 338, 1004, 1608, 550, 965, 1361, 195, 1096, 1223, 1322, 1405, 1369, 1956, 717, 376]. form [697, 50, 1683, 905, 2204, 487]. Formal [2116, 1236, 1246, 575, 1099, 2029, 1131, 1285, 1408, 963, 740, 1407, 1624, 44, 374, 2196, 1043]. Formalism [125, 2187]. formalisms [1287]. Formalizing [2082, 2083, 80]. format [1721]. formation [1246]. forms [1056, 444, 348, 670]. formula [1683, 1960, 193]. formulas [6, 1234, 1659, 1135, 1554, 220]. Foundation [1011, 442]. Foundations [1873, 340, 1037, 1714]. founded [497]. four [2005, 480]. fourth [1977]. Fractal [1599]. fractals [589]. fraction [1214]. Fractional [437, 918, 1530]. fragment [1952, 1510, 2074, 1320]. fragments [496]. frame [2196]. Frameproof [833]. framework [673, 53, 1476, 1781, 1190, 1416, 736, 1763, 1715, 633]. Frankenstein [969]. free [697, 1669, 1639, 255, 19, 241, 1869, 1671, 444, 1470, 1613, 2178, 581, 499, 601, 734, 1501, 1447, 1391, 1796, 1207, 570, 457, 580, 245, 449, 1048, 1448, 1509, 1189, 248, 138, 331, 308, 2093, 29, 71, 823, 2026, 85, 1276]. freedom [434]. freeness [1869, 1286]. French [309, 932, 942, 944, 946, 436]. Frequency [2179, 1226]. frequent [1185, 1785]. Fringe [1443]. front [244]. Frontier [40]. Frozen [655, 2210]. fruitful [121]. FST [25]. full [1176, 1635, 1308]. fully [1382, 902, 1305, 396, 1662, 413]. Function [1634, 1083, 2108, 474, 790, 818, 449, 704, 1491, 1746]. Function-dependent [1634]. Functional [1551, 1254, 1780, 489, 1944, 417, 1561, 1946, 670, 1989]. functionality [1227]. functionals [2143, 554, 1016, 1910, 2212]. functions [1186, 2056, 979, 1482, 1325, 340, 1006, 1005, 101, 485, 903, 1010, 1233, 1067, 758, 1850, 1818, 723, 771, 224, 1411, 1392, 1825, 115, 1013, 1878, 571, 219, 565, 872, 762, 1753, 397, 1015, 177, 452, 764, 356, 887, 676, 1240, 310, 775, 1493, 672, 5, 692, 1112, 1990, 2024, 599, 1309, 12, 83, 1800]. functor [1019]. functorial [1057]. functors [1301, 135, 806]. Fusion [545]. fuzzy [1292, 1922].

G [859]. G-Log [859]. GA [1069]. Gaining [434]. Gale [532]. Game [1848, 386, 1152, 1757, 1831, 382, 385, 1864, 1834, 1835, 345, 377, 1596, 1597, 773, 1648, 810, 1712, 776, 1844, 150, 553, 1851]. game-tree [385]. Games [1550, 191, 1790, 1754, 1833, 16, 378, 969, 1837, 387, 13, 1839, 531, 1841, 2161, 630, 1308, 276, 743, 73, 2049, 1148, 1846]. gap [1560]. gaps [1735]. garbage [1431]. gas [1644, 2127]. Gauß [1891]. Gemmating [1364]. Gene [947, 1415, 2006, 1236]. General [1481, 800, 577, 2000, 521, 1350, 1864, 293, 1440, 674, 1891]. generalised [1898, 1788]. Generalization [285, 1680, 327, 118, 1853, 142]. Generalizations [1529, 1537, 1241, 2084, 111, 2052]. Generalized [362, 1538, 1352, 440, 1617, 612, 1394, 172, 1565, 1821, 2106, 1645, 959, 1265, 2082, 1614, 2019, 874, 236, 210, 2141, 1215, 2223]. generate [2029, 1470, 1569, 36, 534, 1515]. generated [271, 208, 1069, 1660, 1479]. Generating [1684, 1357, 162, 2121, 831, 767, 1818, 1150, 1691]. generation [1681, 303, 589, 2145, 1215]. generations [1923]. Generative [315, 1175, 1285, 1638]. generative-reactive [1175]. generic [1114, 1770, 1309]. genericity [850]. generously [778]. Genetic [2060, 1746, 1415, 736, 1991, 501, 2061]. genome [767, 524]. Genomic [2129, 2060]. Gentzen [255, 1909, 256, 1562]. Gentzen-style [256, 1562]. Gentzen-type [255]. Genuine [404]. genus [108]. geometric [1436, 1954, 2104, 999, 612, 80]. geometrically [1390, 1581]. geometries [607]. Geometry [956, 1006, 1629]. Getting [2075]. GHI [383]. giant [514]. given [1150, 632, 797]. glasses [647]. Global [1647, 1798, 2134, 2071, 570, 1973, 620, 1746, 2202]. Glueing [1306]. Glushkov [62]. go [379, 1632, 380]. goal [1298]. goal-independent [1298]. goals [167]. Gödel [1009]. Goles [2054, 2055]. Golomb [629]. good [2155, 551, 1683]. Gossiping [639]. Graceful [982]. gram [2218]. grammar [1180, 444, 875, 1470, 519, 1460, 138, 1287, 1528, 1634, 1455]. grammar-based [1528]. grammars [697, 1639, 1291, 2178, 581, 1852, 882, 1363, 366, 1586, 231, 315, 1657, 234, 146, 308, 1536, 2007, 2080, 1413, 2228, 1504]. grammatical [1131]. Graph [2028, 610, 2056, 2134, 1979, 1429, 867, 64, 1684, 884, 655, 613, 2155, 832, 790, 947, 1632, 1972, 81, 1516, 2003, 1530, 632, 1320, 586, 120, 1148, 1108, 1440, 273]. graph-theoretic [1320]. graphical [1958]. Graphs [1688, 137, 1441, 673, 626, 1695, 1993, 1750, 359, 2146, 277, 1502, 2104, 1752, 1463, 1644, 753, 1167, 1362, 1010, 9, 456, 1589, 132, 273, 1434, 613, 499, 1699, 113, 609, 625, 2152, 1089, 1567, 1596, 1597, 983, 1464, 514, 641, 1389, 75, 1839, 654, 510, 580, 2010, 1492, 353, 1514, 1266, 1508, 2191, 621, 967, 1787, 1854, 281, 120, 2026, 1297, 1587, 1399, 122, 616, 628, 604, 351, 1761, 1402, 1385]. Gray [1018]. greatest [1491, 812]. greedy [1254, 2222]. Greibach [366]. Grid [495, 1199, 2147]. grids [2009, 1584, 742]. Gröbner [866]. Grothendieck [1778]. ground [364, 168, 1601, 1524]. groundness [895]. group [238, 528, 1645, 726, 1690]. group-definable [238]. Grouped [308]. Groups [936, 1669, 241, 1969, 1994, 1786, 1391, 2156, 1276]. grows [1977]. Growth [1671, 449, 627, 2031]. Growth-sensitivity [1671]. Grzegorczyk [266]. Guaranteeing [417]. guard [614]. Guarded [1102]. guide [1065].

H [1365]. Hadamard [155]. hair [1701]. half [1726]. half-monadic [1726]. Halfliar [1834]. halting [482]. Hamiltonian [691]. Hamming [1100, 1516, 1556]. hand [152]. HandiAS [702]. handling [1149, 1797]. Handsome [1317]. Hanoi [151]. Hard [878, 1762, 1025, 204, 1138, 1028, 1420, 1087, 1818, 1151, 1840, 515, 1392, 2183, 243, 91, 1646]. Hardness [998, 1588, 2146, 277, 2139, 1030]. hardware [915]. Harmonic [2099]. Hartmanis [204]. Hashiguchi [1744]. Hausdorff [1615, 987]. having [1441, 1167, 1907]. HDT0L [284]. head [1049]. heap [365, 377, 759]. heaps [1265]. height [1760, 1685, 779, 2208, 1142, 886]. help [883]. Herbrand [1193]. hereditarily [1016]. hereditary [613, 1162, 604]. heterogeneous [849]. Heuristic [2004]. heuristics [1067, 1763]. hexadecimal [1841]. hexagonal [1692]. hexagons [623]. hidden [1750, 292]. hiding [2173]. Hierarchical [1574, 1203, 1948]. hierarchically [1266]. Hierarchies [555, 1867, 1910, 1364, 582, 1792, 436, 436]. hierarchy [2165, 892, 581, 734, 1332, 1598, 771, 1557, 171, 266]. High [1895, 1141, 1088, 1735, 304]. high-dimensional [1141]. high-level [1735]. Higher [730, 1939, 910, 952, 80, 659, 413, 685]. Higher-order [730, 1939, 910, 80, 659, 413, 685]. highness [566]. Highway [2154]. Hilbert [763, 608]. hill [1991]. hill-climbing [1991]. hints [1831]. Hoare [836]. hoc [2158]. holds [286]. holes [1961, 1540, 510, 1568]. home [2121]. homogeneous [696, 2050]. homological [1494]. Homomorphic [355]. homomorphism [1850, 1677]. homomorphisms [2221]. homotopical [1494]. Homotopy [17]. Hood [1604]. hop [1467, 1398]. Hopcroft [368]. horizons [1566]. Horrocks [634]. Hour [778]. houses [510]. Huffman [574, 311]. hull [2017, 121]. hulls [1125]. hundred [2039]. Hybrid [519, 1460, 1191, 1182, 578, 183, 1039]. hybridization [519]. Hydra [1484]. Hyper [582]. Hyper-polynomial [582]. hyperbolic [1367, 504]. hypercoherences [319]. Hypercomputation [1932, 1924, 1929, 1919]. hypercomputing [1928]. hypercubes [455, 744, 1584, 272, 1868]. hyperdecidability [447]. hypergraph [134]. hypergraphs [673, 903, 306, 1338]. hypermaps [1680]. hyperplanes [983]. hyperrobust [1002]. hypertext [156]. hypotheses [1406]. Hypothesis [1406, 103, 749, 1246].

ICALP [933]. Ideal [1621, 271, 905].

idealized [1733]. ideals [866]. Idempotent [1906, 1459]. identical [1172]. Identification [1415, 12, 724, 206, 723, 1413, 2228]. Identifying [1975]. idioms [703]. if [1151]. II [1666, 255, 102, 2070, 1303, 1591, 1597, 1768, 1494, 1944, 1746, 569, 662]. Image [1669, 212, 1532, 1819]. images [17, 980]. imitation [2060]. IMLL [1308, 1562]. immune [1540]. immunity [1800, 1267]. impact [1086]. impacts [1075]. impartial [378, 73]. imperative [330, 1942, 910, 173]. imperfect [387]. implementable [392]. implementation [1254, 1235, 2078, 30]. Implementing [586]. implicational [52, 540]. implicit [1193]. implies [735]. Improved [1831, 1324, 455, 2136, 219, 156, 1297, 2220]. improvement [634]. Improving [1788, 384]. in-place [1527, 140]. Inapproximability [1786, 1138]. incentive [1745]. include [1618]. inclusion [19, 486, 417]. incompleteness [1009, 1720]. incompressibility [102]. inconsistent [1371]. incorporating [193]. increasing [1936, 2038]. incremental [1111]. incrementally [1878]. increments [772]. independence [1484, 1674, 1392]. independencies [670]. Independent [1502, 632, 797, 1750, 1496, 1159, 1049, 760, 1298]. Index [20, 21, 45, 57, 79, 97, 123, 131, 160, 180, 194, 215, 216, 227, 250, 270, 288, 302, 317, 328, 337, 352, 370, 371, 375, 388, 419, 461, 470, 483, 500, 536, 548, 568, 602, 635, 658, 693, 705, 725, 741, 549, 843, 865, 889, 897, 913, 920, 1023, 785, 801, 814, 834, 929, 957, 976, 994, 1044, 1060, 1095, 1113, 1164, 786, 930, 1222, 1270, 1299, 1321, 1345, 1377, 1403, 1433, 1472, 1488, 1517, 1546, 1578, 1606, 1630, 1663, 1702, 1717, 1736, 1489, 1764, 1782, 1808, 1849]. index [1872, 1901, 1917, 1933, 1955, 1980, 2012, 2040, 2073, 2086, 2111, 2138, 2162, 2184, 2206, 2224, 1765, 2013, 1235, 1835]. indexed [144]. indexes [844]. indices [630]. indifference [1385]. indirect [1000]. induced [1601]. inducing [1857]. Induction [1478, 2042, 683, 2043, 1931, 1249, 1944, 49, 1074, 2035, 2084]. Inductive [804, 2229, 2089, 225, 644, 1482, 1921, 157, 723]. Inductive-data-type [804, 2229]. inductively [607]. inequalities [2177, 2226, 893, 1397]. inequality [2146]. inertness [175]. inferable [226]. Inference [1829, 644, 255, 2203, 491, 805, 1131, 723, 1767, 225, 689, 1728]. Inferring [205]. Infinitary [1419]. Infinite [1474, 378, 469, 1123, 2046, 1167, 1963, 1330, 584, 828, 1447, 1642, 1190, 818, 1238, 520, 1448, 829, 1523, 356, 887, 2049, 765]. infinite-state [469]. infinitely [1324, 1974]. inflated [2010]. influence [749]. Information [188, 798, 1086, 387, 520, 1882, 1257, 2173, 1252]. Inheritance [745, 736]. inhibitor [2077, 848]. injective [439, 1728]. innermost [185]. Innocent [810, 1712]. input [1820, 2024]. input- [1820]. insecurity [1452]. insensitivity [290]. insertion [1443, 583, 1495]. insertions [2129]. inspired [1799]. instance [1414]. instance-based [1414]. instances [2140, 2004, 2219, 243, 1255]. instantiation [53]. institution [1059]. institutions [1730, 924, 1778]. insure [1896]. Integer [1698, 1327, 108]. integers [1732, 474, 2151, 1575, 1970, 792]. integral [107, 1146]. Integrating [1257]. integration [1927, 1620, 188]. intelligence [1920]. Interaction [320, 2114, 1147, 1386, 261, 417]. interactive [959, 961, 1259]. interconnection [432, 349]. interconnections [1089]. Interconvertibility [331]. interior [1689]. intermediate [1368]. internal [1441, 519, 1723, 236]. internalized [1931]. International [1847]. Interpolation [1778, 1019, 176, 2215]. interpretation [1954, 458, 892, 1180, 1190, 1945, 847, 795, 1298]. Interpreted [2079]. interpreting [825]. interprocedural [1776]. interreduction [494]. intersecting [1531]. Intersection [1904, 981, 1479, 803, 154, 528, 808, 1767, 120, 813]. Interval [1478, 1750, 1742, 880, 1117, 2144, 299, 81, 919, 82, 435, 300, 1740, 1795]. intervalle [367]. intervals [937, 1552, 1907]. INTEX [25]. Intractability [43]. Intrinsic [1944]. Introducing [2190]. Introduction [1809, 47, 484]. intuitionistic [55, 1708, 56, 812, 1170]. invariance [1278, 1989]. invariant [518, 1118, 1653, 1103]. invariants [1688, 1566]. Inverse [1886, 241, 2124, 1905]. inversion [951, 1894, 1897, 2127]. inversions [2130]. invertible [1212]. Investigating [280]. investigation [387]. investment [1520]. involution [1391]. involving [761, 1850]. irreducible [1453]. irregular [162]. Ising [1858]. Isomorphism [41, 578]. isomorphisms [1313]. isoperimetric [1695]. isotopic [1609]. issues [848, 1932]. items [1785, 1158]. itemsets [1185]. Iterated [876, 1964, 845, 1602, 1870]. iterating [1732]. Iterative [1250, 1897, 1474, 1247]. iterators [732].

JAGs [148]. Japanese [1255]. Java [990, 1220]. Jay [416]. Jean [1224, 1229]. job [2136]. jobs [710, 1996]. join [677]. join-calculus [677]. joker [1308]. jugs [968]. jump [582]. jury [2020]. JVM [1187].

Kautz [619]. keep [971]. Kernels [1008]. key [2000, 1090, 439, 2156]. keys [1452]. Khalimsky [1618]. Kirby [1484]. Kleene [2166, 18]. KMP [1468]. knapsack [439]. knowledge [341]. Knuth [738]. Kolmogorov [1269, 512, 799, 789, 795, 796, 1017, 792]. kos [379]. Kripke [2079]. Krohn [92]. Krohn-Rhodes [92]. Kurtz [2030].

L [1649, 955]. label [1353, 1740]. labeling [1353, 2152]. labelled [1950]. labels [1551]. lack [1540]. Lagrange [1683]. Lambda [336, 1419, 1719, 904, 342, 1720, 109, 5, 692, 488]. lambda-calculus [1719]. Lambda-dropping [336]. landscape [618]. langage [746, 23]. Langages [70, 944]. Language [1054, 264, 361, 1680, 196, 393, 552, 757, 859, 1633, 767, 1131, 36, 1733, 1150, 1448, 1424, 331, 1410, 910, 355, 703, 29, 285, 740, 953]. Language-theoretic [361]. Languages [1545, 238, 19, 1292, 458, 1539, 698, 1869, 197, 1876, 695, 357, 246, 23, 1475, 2029, 1670, 1671, 2209, 1050, 1, 1470, 165, 1986, 942, 944, 2062, 1828, 1870, 192, 1420, 589, 556, 783, 1131, 431, 601, 734, 1501, 1447, 2045, 1642, 47, 313, 1469, 1818, 699, 1109, 1805, 1150, 1592, 1796, 1941, 457, 330, 245, 449, 294, 199, 224, 551, 354, 467, 1422, 141, 784, 208, 881, 1942, 1466, 149, 58, 1059, 74, 1560, 2091, 355, 1461, 91, 84, 61, 85, 30, 2002, 1741, 225, 1289, 2214, 2193, 184, 1221, 409, 1880, 424, 1829]. languages [484, 2195, 139]. large [1380, 1372, 1251, 207, 2049, 107, 792]. large-scale [1251]. largest [60]. latency [1114]. lattice [935, 1580, 793, 1975, 2058, 1974, 454, 341, 2066, 1145, 2127]. Lattices [1965, 1081, 2054, 1969, 2065, 1623, 310, 775, 1146]. law [1269, 286]. laws [701, 1704, 852]. lax [1309]. layout [432, 2003]. layouts [1205, 2132]. laziness [492]. lazy [1780]. LD [1954]. LD-resolution [1954]. Leader [1973]. leads [2043, 262]. leads-to [262]. learnabilities [1248]. Learnability [1659, 551, 718, 220]. learners [552, 999]. Learning [1828, 556, 224, 1826, 1424, 1407, 720, 1112, 555, 6, 434, 221, 1107, 382, 2097, 1110, 1923, 223, 521, 722, 749, 657, 1109, 1000, 1001, 1411, 520, 1111, 1247, 219, 557, 1435, 1412, 721, 1410, 222, 1070, 1414, 276, 1002, 2214, 1252, 1827, 1989]. Least [777, 2153, 812, 1552, 2031, 1342]. Least-recently-used [2153]. Left [1908, 69, 313, 249, 152, 1973]. left-hand [152]. left-linear [249]. Leftmove [141]. Leftmove-bounded [141]. legacy [334]. Legendre [1825]. Lemma [1195, 783, 605, 139, 451, 2223]. Lempel [1528]. length [1544, 1881, 236, 116, 1356, 1073, 1701, 1660, 119, 797]. lengths [1839, 245, 738, 1146]. Letichevsky [1533]. letter [1639]. level [1876, 2003, 1735, 1059, 452, 395, 1845, 2015]. levels [5, 692]. leveraging [999]. Levy [1790]. Lex [81]. Lex-BFS [81]. Liapunov [1652]. liar [14, 970]. liars [743]. library [24]. Lie [832]. lies [1851]. Lifting [909, 1888, 1302]. light [1937, 2203, 1319, 1945]. lighting [979]. like [197, 1475, 406, 1480, 1941, 493, 417, 973, 1148, 853, 1880]. limit [2056, 1905]. limitations [433, 719]. limited [687, 560, 833, 287, 179, 428]. limitedness [59, 1744]. Limiting [1772, 2143, 1142]. limits [1923, 1523, 410, 1020]. Linda [173, 197]. Linda-based [173]. Linda-like [197]. Lindström [852]. line [2106, 708, 1084, 639, 710, 1350, 1602, 1149, 1520, 1451, 711, 1237, 1158, 1000, 1337, 751, 1745, 1514, 219, 1763, 1444, 1498]. Linear [2098, 2150, 529, 1868, 2198, 1937, 458, 1732, 1860, 66, 2123, 51, 1467, 942, 490, 1086, 763, 1131, 997, 54, 2144, 699, 773, 554, 249, 1598, 619, 918, 1306, 800, 1319, 1494, 1124, 668, 691, 1943, 1189, 2027, 320, 1320, 1461, 2007, 1316, 1468, 161, 1877, 1741, 235, 1318, 1726, 409, 2148]. linear-constraint [409]. Linear-time [2150, 1124, 1468]. Linearisability [1707]. linéarité [942]. Linearity [2204]. linearly [970]. lines [982, 1889, 1899, 2039]. link [1602, 68, 620]. link-connectivity [1602]. linkage [1051]. lion [532]. list [707, 1083, 478]. listen [640]. listen-in [640]. Listing [867]. listings [2122]. lists [758, 526, 1409]. literary [1256]. liveness [1450, 1771]. LL [698]. LL-languages [698]. LLL [1379]. Load [1127, 971, 751, 2027]. Load-balanced [1127]. Local [1856, 192, 570, 967, 1743, 2220, 984, 1594, 1116, 1480, 584, 2150, 1157, 2074, 1394, 1753, 1725, 1967, 718, 110, 27, 460, 2194]. locale [1914]. locales [1915]. Localic [1623]. Locally [431, 826, 246, 2045, 313, 1469, 1915]. location [1757, 560, 2161, 681]. Löf [1170]. Log [1356, 129, 859]. Log-space [1356]. logarithm [2019]. logarithmic [1889, 1599]. LOGCFL [769]. logic [1312, 1218, 1780, 1937, 2203, 854, 1428, 1034, 126, 1952, 2070, 1431, 51, 1179, 1035, 1036, 164, 132, 273, 1434, 399, 1313, 851, 490, 946, 901, 899, 1038, 54, 1555, 1617, 1306, 1001, 1319, 691, 543, 1943, 1953, 1715, 1173, 320, 1033, 660, 465, 49, 964, 1296, 1307, 1945, 1320, 1708, 1039, 2172, 1725, 1316, 80, 847, 402, 718, 1912, 546, 415, 403, 1298, 1741, 2050, 672, 1041, 1318, 1042, 853, 674, 836]. Logical [1055, 1037, 638, 1161, 895, 796, 50, 178, 990, 274, 1422, 58, 188]. Logicality [130]. logics [1140, 52, 2167, 1480, 1038, 1102, 1246, 1393, 1449, 1103, 992, 322, 952, 1877, 840, 414]. logique [946]. Long [525]. Longest [2038, 580, 682]. Longley [1946]. Longo [1790]. look [560, 2211]. look-ahead [560, 2211]. lookahead [1149]. lookup [990]. Loop [167, 672, 1976, 1207, 622, 1438, 2202]. loopless [1691]. Lorentz [1644]. Loss [1825, 1599]. lot [778]. low [880, 104, 1970, 744]. low-congested [880]. low-dimensional [744]. low-rank [104]. low-weight [1970]. Lower [652, 996, 1451, 2139, 1354, 1103, 885, 567, 1152, 527, 707, 1993, 1457, 651, 529, 919, 148, 1561, 171, 301, 83]. lowness [566]. lozenges [1569]. LR [699, 231]. Luca [445]. Łukasiewicz [13]. Lyndon [831, 1675, 1676].

M [1481, 152, 1477]. M-based [1477].

M-net [1481]. machine

[1650, 1923, 296, 1451, 1246, 2078, 664, 1996]. Machines [1120, 1139, 1009, 1583, 1362, 1963, 515, 1124, 479, 2141, 1074, 873, 714, 703, 1925, 482, 1922, 768]. MacMahon [608]. made [1140]. mail [1864]. maintaining [453]. majorities [967]. majority [630, 498]. makespan [1651, 1996]. making [359]. Mal’cev [451]. man [532]. management [51, 2099, 1165]. Manhattan [1030]. manifesto [975]. MANIFOLD [196]. manifolds [1609, 89]. Many [2090, 1172, 1324, 1329, 1974, 2031]. map [309, 1293, 1303, 1030]. Maple [23]. mappings [1860, 1279, 2132]. maps [90, 747, 132, 2082, 2155, 926, 69]. Marked [429]. Markov [959, 1950, 781]. Markovian [1168]. Marriage [1195, 1654, 1656, 2135, 878]. Martin [1152, 1170]. Martin-Löf [1170]. mask [1215]. massive [1463]. Master [549, 786, 930, 1489, 1765, 2013]. matchgates [1132, 1471]. matching [1855, 422, 1441, 2076, 1995, 76, 1817, 347, 78, 174, 1416, 575, 1449, 156, 2032, 715, 144, 235, 1166, 1795, 1455, 730]. matchings [1567, 1958, 1317, 1402]. mathematical [1115, 2084, 1873]. mathematics [1316, 80]. Matiyasevich [482]. matrices [1886, 1887, 1328, 206, 104, 1892, 283, 1529, 1537, 1894, 158, 943, 1897, 1747, 122]. Matrix [1896, 577, 1131, 1091, 1893, 2131, 2217, 1859]. matroid [2121, 1387]. matroids [1629]. Maude [1035, 1040]. Maurice [932, 2055, 933]. max [1280, 1272, 1699, 1279, 563, 2182, 1459, 1275, 1273, 1281]. max-plus [1280, 1279, 2182, 1281]. max-separable [1275]. MAX3SAT [1151]. maxima [1594]. maximal [867, 2140, 825, 1673, 353, 587, 1468, 1402]. maximin [247]. maximizing [1534, 2188]. Maximum [128, 422, 1753, 760, 1788]. Mazurkiewicz [1619]. MCD [2157]. MCD-estimator [2157]. mean [1754, 1251, 647]. means [458, 2177, 2226, 764]. measure [430, 1168, 1269, 397, 91]. measures [1098, 1071, 276]. Measuring [968, 1789, 1998]. mechanics [654, 646]. mechanisms [2202]. Median [7]. MELL [405, 1727]. Mellin [236]. membership [126, 719, 29, 1541]. Membrane [770, 1988, 1065, 2092]. membranes [1792]. memory [2099, 43]. Memoryless [1754]. merge [1265]. merging [1350, 1527, 140, 2216]. mesh [561]. meshes [1155, 305, 1357]. message [1735, 1171]. message-passing [1171]. MetaML [335]. method [527, 125, 102, 990, 1744, 975]. methodology [953]. methods [469, 54, 2161, 646, 1989]. Metric [411, 502, 1611, 1610, 16, 1012, 906, 484]. Metrics [1950, 2016, 1617, 1879]. métriques [367]. metrisability [1613]. metrizability [1625]. MFPS [1873]. microprocessors [915]. migration [709]. Min [563, 1699, 1459, 1273]. Min-max-boundary [563]. Mind [1001]. minds [1928]. Minimal [695, 1742, 1881, 26, 359, 2146, 607, 1538, 1135, 755, 1334, 32, 1863, 402, 684, 2193, 2198]. Minimality [842, 239]. Minimax [1274, 1843]. Minimization [1940, 86, 1651, 1701, 2192, 28]. minimize [868, 1421, 1996]. Minimizing [1232, 574, 2183]. minimum [100, 573, 592, 1467, 298, 306, 116, 748, 2216, 2222, 797]. Mining [1372]. minor [64]. minus [1354, 1508]. mismatch [1188]. mistake [1109]. mix [1026]. Mixed [1338]. mixing [1573]. ML [2074, 1880, 896]. ML-like [1880]. Mobile [201, 2069, 1711, 1200, 842]. mobility [200, 2072]. Mod [615]. modal [1928, 1480, 543, 2172, 1725, 546]. mode [640, 1460]. Model [1172, 1711, 1415, 707, 1176, 639, 1140, 1184, 1288, 1114, 1577, 496, 2113, 1573, 1799, 381, 1109, 1861, 396, 425, 467, 595, 963, 759, 1214, 465, 1962, 276, 1040, 267, 5, 692, 1202, 1440, 674]. model-eliminative [1288]. model-theoretic [1962]. Modeling [2114]. Modelization [950]. modelled [1761]. Modelling [2117, 200, 1775, 1011, 80]. models [1183, 2115, 1874, 340, 722, 1475, 1480, 1961, 1905, 475, 344, 1331, 773, 2065, 1305, 1941, 1306, 808, 810, 1712, 727, 1048, 524, 1882, 1173, 1926, 172, 1621, 1622, 2079, 1913, 342, 1877, 1746, 1720, 647, 858, 1728, 2179, 484]. modes [519, 1460]. Modified [1029]. modular [1648, 77, 924, 321]. modulo [477]. Molecular [1147, 2116, 1799]. molecules [1840]. Moments [1749, 886]. monad [1302, 927]. Monadic [853, 1167, 132, 273, 940, 1434, 1102, 1557, 742, 847, 1726]. monads [1220, 846]. mono [1708]. mono-sorted [1708]. monoid [2181, 183, 248]. monoids [1951, 1494, 1731, 1144, 1677]. monopolies [967]. Monotone [526, 2056, 176, 178, 2125, 142, 220]. monotonic [2176, 1713, 1802, 2161]. monotonicity [721]. Moore [1166]. morphism [1241]. morphismes [942]. morphisms [1964, 1328, 942, 879, 208, 926, 39, 823, 1519, 765]. Morse [1241]. mortality [459]. most [1685, 1358]. Motif [1094]. motion [1614]. motivated [578]. move [244]. move-to-front [244]. movement [2069]. moving [755]. MSP [583]. Muller [26]. Multi [385, 1049, 214, 1141, 827, 1965, 1467, 1057, 1553, 2113, 1387, 2144, 1861, 2003, 1438, 1797, 1529, 1537, 1394, 1398, 395, 300, 151, 335, 1740]. multi-agent [1553]. multi-algebras [1057]. Multi-cut [385]. multi-dimensional [827, 1965, 2144, 1529, 1537, 1394, 300]. Multi-head [1049]. multi-hop [1467, 1398]. multi-label [1740]. multi-level [2003, 395]. multi-loop [1438]. multi-parameter [1387]. multi-peg [151]. multi-recombination [1141]. multi-scale [2113]. multi-stage [335]. multi-tree [214, 1861]. multi-vehicle [1797]. multiagent [2079]. multicast [1588, 1861, 404, 562, 517]. multidimensional [1542, 247]. multiequations [856]. multifunctions [1940]. multigraphs [1338]. multilevel [1896]. multimedia [237]. Multiple [821, 637, 1831, 831, 1645, 502, 1881, 1359, 2032, 308, 1521, 1689, 379, 2024]. multiple-fault [1881]. multiplication [1615, 1893]. multiplicative [101, 1510, 1575, 691, 1320]. multiplicatives [1312]. Multiplicities [588, 701]. multipoint [2215]. multiprocessor [1534, 136, 591]. Multirate [562]. multiset [463, 1691]. multivariate [1900]. musical [1739]. must [2205]. mutable [1198]. MUTEX [862]. MUTEX-problem [862].

Mutilated [1762]. mutual [1755].

N [1448]. N-free [1448]. name [94]. named [1816, 148]. names [2074, 1030]. nasty [1110]. Natural [1926, 1280, 1983, 474, 1552, 703, 812, 860]. Naundorf [327]. NC [528, 1938]. near [2219]. near-optimal [2219]. nearest [1444]. Nearly [1520, 1297, 558, 1108]. nearness [1628]. necessary [884]. necklaces [1515]. need [1513]. negation [2204, 415]. negative [1213]. neighbor [1410]. Neighborhood [625]. neighborhoods [1463, 2124, 1794]. neighbors [1351, 1444]. nested [1801, 407]. Nesting [516]. net [499, 1481, 2201, 464]. nets [265, 2077, 848, 165, 1386, 54, 405, 1314, 463, 1771, 1706, 585, 1169, 320, 1317, 251]. network [998, 624, 425, 257, 1788, 567]. networks [1976, 1415, 687, 67, 1205, 365, 432, 2114, 1866, 1535, 1467, 1086, 1987, 323, 1117, 657, 1507, 2144, 1861, 2188, 619, 622, 1438, 562, 260, 468, 2158, 517, 2216, 297, 1040, 349, 616, 1798, 1761, 836]. Neural [1987, 1073, 1989, 998, 657, 1072, 557, 1070, 1798]. neural-Bayesian [1072]. neuronal [1660]. neurons [1074]. News [1339, 2227]. NFA [2226, 2177]. NFAs [159, 1513, 2195, 1653]. NFL [1067]. Nim [1596, 1597]. Nine [36]. Nivat [934, 932, 152, 2055, 933]. NL [204]. NL-hard [204]. NMR [1442]. No [176, 1280, 1618, 417]. Node [673, 1358, 211, 148]. node-named [148]. noise [1141, 1110, 552]. noise-tolerant [552]. noisy [223]. Nominal [2085]. Non [732, 1641, 439, 1359, 142, 1312, 1936, 661, 735, 178, 1138, 2029, 1713, 805, 942, 1213, 1533, 70, 269, 799, 1615, 1616, 2078, 809, 166, 1286, 69, 1926, 2171, 1477, 1320, 1112, 1174, 674]. Non-approximability [1359]. non-associative [70]. Non-cancellative [142]. non-clausal [661]. non-commutative [1312, 1320]. non-constructive [178]. non-context-freeness [1286]. Non-cryptographic [1641]. non-dependent [735]. non-determinism [799]. non-deterministic [2078, 809]. non-disjoint [1174]. non-empty [1533]. non-extensional [166]. non-floundering [674]. non-Hausdorff [1615]. Non-injective [439]. non-linear [942]. non-linéarité [942]. non-monotone [178]. non-monotonic [1713]. non-negative [1213]. non-NP-hard [1138]. non-recursive [1112]. Non-regular [732]. non-separable [69]. non-size-increasing [1936]. non-standard [2029]. non-standard-type [805]. non-terminating [269]. non-termination [1477]. non-topological [1616]. non-Turing [1926]. non-wellfounded [2171]. nonadaptive [528]. nonblocking [1866, 567]. noncanonical [1815]. nondeterminism [4, 418]. Nondeterministic [1526, 239, 1333]. nonexpansive [1279]. Noninterference [938]. nonlinear [1279]. nonlinearity [1267]. nonparametric [996]. Nonprimitive [464]. nonregular [1543]. Nonterminal [1363, 315]. nonterminals [1150, 2007]. nonuniform [1242]. norm [1972]. Normal [697, 1056, 348, 670, 1728, 50, 782, 1179, 444, 2204, 487]. normalizability [809, 1177]. Normalization [1192, 735, 185, 1318]. normalize [416]. normalizing [269]. normed [766]. Notations [1097]. Note [884, 1835, 885, 621, 887, 886, 883, 888, 1525, 845, 554, 1285, 1998]. Notes [605, 1149]. notion [1025, 1302]. notions [593]. NP [642, 1138, 503, 1940, 1603, 1028, 1087, 450, 1978, 1151, 515, 504, 630, 358, 1701, 1341, 1646, 1452, 1859, 7]. NP-complete [1603, 450, 1701, 1452, 7]. NP-completeness [503, 1978, 630]. NP-hard [1028, 1087, 515, 1646]. null [29]. null-free [29]. Number [1456, 1591, 1152, 780, 2134, 782, 1539, 1007, 816, 607, 573, 2209, 1977, 1851, 1960, 1457, 981, 1150, 159, 1895, 648, 1268, 1016, 2007, 1452, 179, 301, 481, 1568, 1018, 1692, 1845, 1846]. Number-conserving [1456, 1268]. numberings [593]. Numbers [914, 1211, 1280, 1380, 426, 1009, 1552, 10, 1212, 1354, 612, 1570, 611, 1017, 1020]. Numeration [740, 969, 285]. Numerical [1900]. nymphs [966].

O [1549]. OBDD [452]. OBDDs

[1995, 222]. Oberwolfach [1399]. Object [183, 1852, 736, 1040, 1912, 1043]. Object-oriented [183, 736, 1040, 1043]. Objects [488, 1006, 538, 237]. observable [1184]. observation [1550, 2029]. Observational [1428, 864, 923, 396]. observations [398]. obstacles [755]. obstruction [64]. obtain [527]. occurrences [1539, 179, 2179]. octagon [1960]. Off [1084, 915, 1165]. Off-line [1084]. offsets [1742]. OLAP [1375]. omega [525, 601, 734, 1501, 1447]. omega-power [525]. On-line [2106, 708, 710, 1350, 711, 1237, 1337, 751, 1763, 1149, 1520, 1451, 1158, 1000, 1745, 219, 1498]. Onbeta [592]. Onbeta-skeleton [592]. once [642, 1569, 171]. One [1261, 1160, 1639, 1327, 2122, 1329, 2124, 1269, 1709, 1605, 254, 556, 2058, 269, 1838, 282, 1336, 598, 1585, 1586, 295, 2125, 1499, 852, 2039, 286]. one-counter [1585, 1160]. one-dimensional [2124, 2058, 1838, 1586, 2125]. one-hundred-and-one [2039]. one-letter [1639]. one-pass [1709]. one-rule [269, 598]. one-sided [1605]. one-variable [556]. One-way [1160, 282, 1336]. ones [81, 1355, 764]. Online [2097, 1027, 712, 1159, 713, 2100, 715, 2104, 2136, 2102, 78, 1339, 2227, 714, 2022]. only [1973]. Open [1848, 850, 928, 926, 1194]. operation [1215]. Operational [859, 1386, 1286, 2170, 58, 484]. Operations [767, 1633, 1986, 113, 1908, 740, 796, 1627]. operator [366]. Operators [737, 1904, 594, 912, 1746, 1459, 1020]. opponent [381]. opponent-model [381]. optical [687, 1085, 67, 1861, 1398, 1761]. optima [1746]. Optimal [1949, 432, 14, 1383, 2155, 1051, 1495, 498, 27, 460, 1440, 1379, 2186, 1265, 777, 1520, 1457, 213, 78, 505, 2219, 714, 1896, 1104, 2094, 1376, 446]. optimality [895]. Optimistic [1208]. Optimization [1067, 1025, 1138, 1876, 624, 1387, 1511, 515, 278, 646, 358, 760, 620, 1275]. Optimizing [1527, 1373]. orbits [1512]. Order [1720, 854, 1939, 1590, 1293, 132, 273, 940, 1434, 1709, 1733, 478, 738, 742, 910, 1708, 322, 952, 2080, 676, 80, 847, 410, 659, 413, 2127, 110, 27, 460, 1047, 853, 685, 730]. Order-incompleteness [1720]. ordered [642, 1122, 2018, 1685, 1804]. ordering [100, 167]. orderings [2213, 1684, 1383, 1722, 1748]. orders [2178, 738, 2091]. ordinal [397]. orientation [81]. orientations [214, 605]. oriented [183, 736, 606, 1040, 1629, 1043]. original [2140]. orthogonal [1729, 682, 441, 1375, 1440]. orthogonality [1306]. other [709, 100, 277, 2221, 1190, 1121, 2109, 1059, 987]. outperforms [1991]. overexpressions [1415]. overview [1197].

P [96, 1364, 2, 1500, 1799, 1540, 1803, 354, 1454, 1366, 1998, 1859]. P-complete [96, 2]. P-immune [1540]. PA [575, 838]. PA-matching [575]. PA-processes [838]. PAC [6, 1110, 1435]. package [23]. packet [260]. Packing [2023, 128, 1158, 1396, 1397, 2022]. page [709, 2102, 712]. paging [87, 713]. pair [891]. pair-sharing [891]. Paired [2010]. Paired-domination [2010]. pairing [474, 2048, 2019]. pairs [127, 1654, 2159, 1970, 1787, 1341]. pairwise [2147]. Palindrome [1225, 1542]. palindromes [1525]. pancyclic [1399]. Papers [1873]. Paracategories [1723, 1768]. paradigm [613]. Parallel [875, 319, 590, 428, 1650, 1443, 710, 1125, 1589, 296, 1127, 444, 556, 76, 1496, 679, 489, 750, 690, 1263, 1586, 312, 149, 1763, 1776, 308, 2094, 509, 446, 1996, 1779, 944]. parallèle [944]. parallelism [1713, 2189, 311]. parallelizing [719]. parameter [1327, 2142, 1387]. parameterised [1481]. Parameterized [1162, 468]. Parametric [544, 1421, 248, 2197]. Parametrisation [248]. Parametrization [1899]. Parametrized [1715]. Pareto [1743]. Parikh [457, 1747]. Paris [1484]. parity [1754, 1331, 529, 2132, 12]. parity-declustered [2132]. parking [778]. parsable [1586]. parsers [1815, 231]. parsimonious [1979]. Parsing [1180, 405, 703, 698, 699, 1536, 1290, 2196]. part [531, 102, 1494, 1318]. Partial [754, 675, 1047, 1772, 707, 1550, 1126, 2213, 1856, 1006, 1119, 2136, 747, 1303, 1057, 128, 945, 828, 332, 1723, 210, 134, 1559, 308, 1625, 1879, 12, 922]. partial-edge [134]. partially [2018, 1804]. particle [2125]. partition [868, 81, 531, 1146]. partitionable [162, 1209]. Partitioning [626, 974, 1464, 1819, 648, 2035, 1003, 28]. partitions [1698, 610]. Pascal [2121]. pass [1709]. passing [170, 1171]. passport [971]. Past [1814]. patch [1612]. path [1129, 1205, 1889, 1384, 1282, 1117, 421, 1458, 2147, 1373, 236, 1797, 82, 1358, 1793, 1297, 150]. paths [1949, 935, 1156, 626, 640, 1382, 1686, 755, 580, 1974, 505, 436, 1787]. pathways [2117]. Pattern [2076, 174, 1855, 996, 422, 556, 1817, 1109, 1416, 1073, 1449, 156, 144, 235, 225, 1166, 1795, 1455]. Pattern-matching [174, 1817]. patterns [1436, 1109, 1253, 1409]. pavages [511, 946]. PAXOS [252]. payoff [1754]. PC [1470, 1355, 138]. PCF [669, 676]. PCP [429]. peak [1442]. Peano [475]. pebbling [72]. peer [2156]. Peg [2036, 1838, 151]. Peg-solitaire [2036]. Peirce [860]. perceptron [1436]. Perfect [122, 1441, 277, 1684, 1567, 1317, 120]. perfectly [2026]. perform [207]. Performance [1141, 1168, 2158, 837, 1069]. perhaps [350]. period [1600]. Periodic [983, 297, 1139, 1742, 1963, 1579]. Periodicity [1328, 1838, 1841, 1967, 584, 39]. periods [1856, 584, 2150, 596, 1891]. permitting [783]. permutation [255, 528, 1384, 744, 1641, 2031, 943, 616]. permutation-path [1384]. permutations [1637, 282, 1675, 1691]. perpetual [682]. perspective [654, 1142, 2079]. PERT [1233]. pessimal [1383]. Petri [265, 2077, 848, 499, 463, 2201, 464, 1706, 251]. Petri-net-based [2201]. PEZ [1636]. Pham [1898]. Phase [1319, 1953, 1538, 344, 654, 646, 656]. phenomena [1081, 651]. philosophical [1932]. phylogeny [277]. physicist [648]. pi [486, 1704, 729, 1770]. pi-calculus [486, 1770]. pi-congruences [729]. picture [303, 141, 784, 61, 1568, 1504]. pictures [742]. pieces [759]. piecewise [459, 627]. Pigeonhole [1324, 1561]. piles [595]. pipeline [1651]. pirates [833]. place [1527, 1114, 140]. placement [259, 498]. placing [1030]. plan [511]. Planar [258, 1680, 2017, 456, 132, 944, 1687, 1266, 69, 633, 1753, 1108]. planarity [2003]. plane [1979, 1571, 1128, 1567, 1367, 1687, 504, 1973, 107, 1568]. planes [1964, 982, 1966, 984]. planning [624, 380]. plans [944]. play [1831]. played [1596, 1597]. player [73]. Playing [970, 382]. PLC [392]. PLC-automata [392]. plumber [1842]. plus [1280, 1279, 183, 2182, 424, 1281]. poems [1255]. poetic [1255]. point [60, 1034, 1613, 1759, 1353, 1088, 1575, 1102, 1212, 229, 507, 1214, 177, 606, 1342, 1628, 1215]. point-free [1613]. point-set [1628]. points [1668, 1349, 573, 1689, 107, 1624, 765, 108, 812, 2039]. Polarized [1169]. policy [2099]. polyatomic [503, 63]. polygon [884, 143, 614, 118, 1689, 2066]. polygons [1155, 1157, 606]. polyhedron [117]. Polymorphic [904]. polymorphism [1176]. Polynat [1913]. Polynomial [1888, 866, 1353, 314, 1135, 718, 2214, 1413, 2228, 1936, 1379, 527, 357, 1544, 578, 1480, 128, 1028, 582, 893, 1248, 1465, 1334, 1662, 986, 766, 1943, 1895, 1688, 136, 916]. polynomial-query [1248]. Polynomial-time [314, 1135, 718, 1413, 2228, 1936, 1544, 578]. polynomially [2043]. polynomials [1349, 1632, 2009, 77, 93, 571, 917, 1900, 8]. polyominoes [1685, 1971, 1692]. polytope [1351]. pomsets [318, 949, 1448]. population [1990]. populations [1746]. Poset [1877]. Poset-valued [1877]. Posets [1907, 114, 1914]. position [696, 1440]. positional [16]. positive [2151, 1805, 1151, 1558, 2172, 226, 225, 1413, 2228]. Positivstellensatz [529]. Possible [1883]. possibly [2046]. Post [38, 424, 35, 874]. Post-plus [424]. potential [867]. Power [1724, 260, 1486, 1380, 1326, 525, 1869, 2069, 275, 1111, 1370, 138, 1424, 630, 660, 315, 2092, 676, 823, 740, 1703, 1499, 569, 1638, 1343, 1922, 2081, 1276]. powerdomain [2200, 1789]. powerlocale [1916]. powers [1695, 1977, 1285, 437]. powerset [407]. PRA [481]. PRA-provable [481]. practical [359, 1253, 1893, 731]. practice [2115]. pragmatic [1208]. PRAM [446]. pre [2186]. pre-determinization [2186]. precedence [1203, 1263]. Precise [1776, 1298]. precisely [1618]. precision [1349]. preconditioners [1896]. precongruence [264, 396]. predicative [357]. Predicting [558, 1108]. Prediction [1071, 702]. Predictive [722, 800].

Preemptive [591, 1530]. Preface [1300, 1245, 931, 1679, 389, 977, 1902, 462, 603, 802, 1784, 1885, 1053, 1830, 2014, 1311, 2164, 1425, 343, 2120, 1211, 1204, 706, 1032, 815, 1665, 1080, 2112, 1061, 1982, 1045, 2067, 1217, 1378, 1347, 2128, 1106, 1418]. prefix [574, 1145]. prefixed [1342]. prefixes [1829]. Premonoidal [911]. preorder [264, 184]. preorders [828, 960]. preprocessing [1166]. Presburger [1811, 1450]. presence [1139, 1431, 1836, 207]. presentation [539]. Presentations [1022, 1774, 508]. presented [856]. présentés [367]. Presenting [1914]. preservation [981, 2201]. preserve [2180]. preserving [1593, 249, 1839, 5, 692]. Presheaf [1475, 1305]. Presorting [229]. primal [2080]. Primality [1380, 2051]. prime [516, 1552, 481]. primes [1324]. Primitive [659, 1668, 1512, 663, 1487, 1641, 1944, 1491, 2052]. primitive-recursion [1944]. Primitives [989]. principal [326]. Principality [1767]. principle [683, 1561, 1324]. principles [2042, 1775, 24]. printing [2188]. prior [1251]. priori [760]. priorities [4, 418, 576]. Priority [576, 1496]. Private [2068, 1636]. Probabilistic [644, 1793, 768, 2200, 555, 2165, 1228, 2008, 1175, 68, 13, 1064, 1768, 960, 1581, 1789, 675, 964, 721, 760, 1160, 858]. probabilities [1323, 4, 418, 1296]. Probability [553, 515, 962, 1435, 1895, 322]. problem [1762, 707, 1082, 618, 1743, 19, 2106, 1777, 1203, 2108, 1571, 1128, 1997, 1025, 1429, 383, 1889, 143, 820, 2105, 1467, 1384, 1544, 817, 2005, 578, 972, 1654, 1994, 716, 1978, 1158, 2019, 2135, 284, 971, 1143, 1797, 1494, 1931, 1731, 1278, 505, 2109, 454, 1756, 748, 1744, 1576, 1635, 1755, 1701, 760, 1863, 631, 1600, 281, 1646, 1859, 151, 136, 532, 1677, 1209, 482, 1344, 111, 862, 2061, 1996, 1761, 2015, 874, 38, 2154]. Problems [96, 1120, 1650, 2098, 362, 1325, 998, 1156, 761, 1695, 1647, 241, 135, 100, 277, 523, 1138, 560, 1538, 1100, 2, 1603, 1848, 2054, 1806, 1087, 344, 1850, 1387, 711, 1304, 450, 513, 1511, 1464, 1656, 515, 278, 2210, 240, 507, 1558, 885, 980, 504, 646, 1395, 230, 358, 1999, 1069, 91, 211, 43, 44, 1148, 656, 1638, 150, 1795, 2198, 28, 1275, 650, 745]. procedure [1227, 397, 29, 674]. procedures [1482, 1888, 1327, 1174]. Process [165, 837, 732, 1168, 989, 949, 688, 173, 276, 678, 184, 685]. Processes [251, 1172, 1186, 934, 4, 418, 170, 1175, 959, 492, 1950, 907, 1389, 679, 493, 960, 1924, 1585, 468, 838, 688, 395, 681, 858, 842, 836]. processing [485, 953]. processor [1763]. processors [1263]. product [293, 1243]. productions [152]. products [1304]. profile [1892]. Program [1431, 852, 1876, 412, 442, 1189, 963, 964, 334, 674, 730]. programmed [1363]. Programming [1873, 1179, 1035, 941, 2113, 1476, 1781, 901, 908, 489, 1942, 736, 415, 509, 335, 558, 218, 484, 124]. Programs [295, 642, 1707, 1780, 938, 164, 336, 1617, 1001, 704, 1944, 394, 1715, 660, 675, 1776, 545, 402, 2075, 226, 718, 403, 171, 1298, 672, 797, 896]. projections [441]. Prolog [226]. Proof [52, 47, 961, 671, 811, 1754, 126, 861, 1484, 51, 634, 503, 1195, 92, 53, 54, 1554, 405, 1314, 1169, 487, 268, 1559, 1708, 952, 1317, 1104, 955, 2212, 1259, 1845]. proof-nets [1314, 1169, 1317]. proof-number [1845]. Proof-search [47]. Proof-term [671]. proof-theoretic [1484, 2212]. Proofs [490, 1312, 50, 1647, 864, 1683, 1099, 529, 807, 2075, 670]. propagating [1445]. propagation [259]. properties [724, 1184, 827, 1007, 1595, 2123, 1179, 104, 1633, 192, 1131, 601, 2045, 1219, 1575, 1190, 1540, 2139, 1157, 1494, 1162, 575, 1277, 881, 262, 358, 1962, 1461, 61, 987, 1858, 1020]. property [1226, 1698, 249, 149, 487, 1560, 1318, 1276, 1504]. Propositional [1140, 255, 1104]. protein [2060]. proteins [1551]. Protocol [1452, 498]. protocols [1218, 2187, 1485, 2020, 1194, 993]. Prouhet [1676]. provability [55]. provable [481]. provably [1200]. proved [2205]. prover [961]. Proving [704, 661]. pruning [385, 558, 1108]. Pseudo [32]. Pseudo-minimal [32]. pseudoknotted [1985]. pseudorandom [1641, 829]. pseudorandomness [430, 888]. PSPACE [241, 778, 1148, 1259]. PSPACE-complete [241, 778, 1148]. PTAS [1589]. Public [1130, 1090, 439]. public-key [1090, 439]. Publications [1224]. pullback [1914]. pumping [783]. pure [683, 735, 2069]. purpose [521]. purposes [534]. Pushdown [1522, 463, 1814, 522]. PVS [1216].

qTL [1296]. quadrant [1682]. quadratic [618, 1802, 509]. quanta [1075]. Quantales [398]. quantifiable [1625]. quantifications [1434]. quantified [1659]. Quantifier [1331, 1557, 742]. quantifiers [852, 813]. Quantitative [964, 1857, 905, 1393, 1296]. Quantum [1984, 830, 146, 1078, 1076, 1323, 572, 1583, 1261, 1075, 1924, 643, 873, 1807, 1077, 1259, 1160]. Quasi [1963, 1610, 511, 2178, 1722, 906, 479, 2091, 167]. quasi-affines [511]. quasi-metric [1610, 906]. quasi-orders [2178]. Quasi-periodic [1963]. quasi-unilateral [479]. quasigroups [2121]. quasiperiodic [348]. quaternion [1615]. Queries [190, 1824, 1107, 2209, 556, 719, 191, 1373, 1103, 1375, 1407, 1376, 2214]. Query [997, 222, 863, 1541, 189, 859, 192, 1248, 2103, 1444, 409]. querying [2114]. questions [345]. queues [1496, 576]. quickest [1793]. QuickHeapsort [1026]. quickselect [1521]. quorums [258]. quote [1710].

R [1150]. radii [119]. radio [624, 1535, 1086, 260]. radius [631]. Radon [951]. ramified [1888, 129]. Ramsey [612, 611, 621, 1148]. Random [1081, 641, 2145, 650, 652, 649, 1681, 788, 2018, 2004, 2207, 1699, 783, 997, 719, 1089, 1604, 514, 1389, 1390, 654, 1581, 656, 1402, 139, 1504]. random-self-reductions [719]. Randomized [1129, 2135, 87, 1067, 561, 2027]. randomly [1714, 1069]. Randomness [1865, 2030, 1071, 133]. range [1291, 1467, 919, 1124, 1375, 1376]. Rank [1260, 1397, 2133, 104, 1447, 1331, 1892]. rankings [210]. Rasiowa [1058]. Rasiowa-Sikorski [1058]. rate [996, 1896, 1003]. ratio [2108]. rational [1229, 1812, 944, 983, 950, 1465, 764, 356, 887, 662]. rationnellement [367]. rationnels [944, 70]. Rauzy [445]. rays [63, 564, 503]. Re [368, 36]. Re-describing [368]. Reachability [1325, 2197, 1486, 1522, 1462, 980, 331, 2198]. reactive [1175, 182, 2187, 1948]. read [642, 2077, 171, 862, 1047]. read-once [642, 171]. readiness [665]. Real [914, 2062, 1013, 1878, 1211, 1990, 1018, 391, 1007, 1010, 1742, 448, 392, 1237, 723, 411, 1124, 961, 962, 963, 917, 394, 41, 1016, 1039, 1787, 1859, 840, 1017, 2197, 83, 1020]. Real-time [2062, 391, 1742, 392, 1124, 961, 962, 963, 394, 1039, 840, 2197]. real-valued [723]. real-weighted [1787]. realistic [1067]. Realizability [1941, 1735]. realization [499]. realizations [2043]. reals [426, 788, 1022, 450, 1012, 1014, 230, 1910, 71]. Realtime [147]. rearrangement [524]. reason [1190]. Reasoning [2081, 176, 668, 1944, 855]. rebound [768]. recently [2153]. reciprocal [1890]. recoding [1820]. recognition [996, 1436, 1542, 613, 944, 2062, 1135, 81, 1073, 1536]. recognizability [249, 740, 1242]. recognizable [1732, 1277, 942]. Recognizing [2210, 763]. recombination [1141]. recomputing [72]. reconfiguration [619]. Reconnaissance [944]. Reconstructing [503, 943, 1580].

Reconstruction [986, 441]. recovering [1181]. recruitment [2057]. rectangle [775]. rectangles [531, 1957, 1493]. rectilinear [143]. Recurrence [584, 129, 1660]. recurrences [458, 1265, 247]. recurrent [1798]. Recursion [2046, 1307, 1523, 1481, 1944, 1945, 161, 659]. Recursive [1610, 272, 685, 1327, 1921, 1919, 861, 336, 663, 1487, 114, 723, 464, 688, 1491, 545, 1112, 2052]. Recursively [426, 1470, 1870, 551, 41, 355]. recursivity [1758]. red [590]. red-black [590]. redex [1048]. reduce [442]. reduced [1385]. Reducibility [3, 95, 1540, 1341]. reducible [854]. Reducing [1235, 1653, 2194, 1669]. reduction [1081, 1186, 2177, 2226, 682, 1708, 1720, 109, 942]. Reductions [661, 135, 1979, 719, 682]. redundancies [452]. redundant [891]. réelles [367]. refined [5, 692]. Refinement [390, 664, 1481, 1724, 81, 394, 910, 1406, 1779]. refinements [2201]. reflecting [2189]. Reflection [1036]. reflexivity [1050]. Refutable [1410]. refutably [1411]. Refuting [1412]. region [491, 423]. register [277]. Regular [1637, 1592, 1466, 414, 1666, 1226, 458, 732, 1539, 1695, 23, 1670, 2177, 2226, 1137, 581, 1526, 1828, 1699, 1642, 1818, 1733, 514, 641, 1389, 1373, 1150, 1796, 535, 1422, 1409, 838, 1622, 234, 2091, 29, 285, 740, 30, 2002, 235, 2193, 1455, 2195, 2209]. regular-language [1733]. regularity [66, 1758, 625, 2204, 2180, 1909]. regularizing [1639]. regulated [1792]. regulation [2113]. rejection [591]. relabeling [876]. related [1812, 1265, 104, 254, 1961, 758, 651, 773, 1558, 2050, 745]. Relating [1179, 1048, 1059]. relation [793, 423, 1724, 1391]. relation-algebraic [423]. relational [1775, 739, 412, 407]. relations [1254, 831, 516, 1050, 990, 950, 800, 978, 358, 253, 1181, 1728]. relationship [1909, 593]. Relative [903, 1016]. relaxations [100]. Relaxed [1173, 709]. release [1797, 714, 1996]. relevance [384, 557]. reliable [207]. Renyi [1851]. repairman [1339, 2227]. repeats [1751, 774]. Repetitions [346, 145, 1556]. Repetitive [350]. Repetitiveness [208]. replicas [498]. replication [486, 2102, 712]. Report [1847]. representability [842]. representation [1612, 1813, 1091, 283, 2125, 452, 341, 1746, 1181]. Representations [593, 324, 1006, 1007, 2151, 1970, 1012, 1013, 1193, 894, 2193]. representative [2214]. Representing [1870, 393]. reproduction [2027]. request [2101, 715]. requests [2153]. requirements [1266]. requires [267]. research [651]. Residue [916, 1406]. resiliency [1800]. Resolution [204, 1561, 1762, 1954, 1645, 2004, 2034, 855, 2202]. Resource [907, 133, 430, 51, 1269, 949, 91, 1209]. resource-bounded [430, 1269, 91]. resources [875, 907, 1883]. respect [1633, 113, 478, 674]. restatements [480]. Restricted [364, 1927, 1654, 1318, 1524]. Restriction [747, 1303, 2069]. restrictions [710, 1974]. result [1797]. Results [651, 649, 137, 1323, 998, 382, 1697, 1487, 254, 1786, 1588, 1686, 2102, 1672, 345, 1970, 1656, 1212, 1465, 1543, 771, 278, 784, 668, 2103, 1165, 909, 42, 2075, 1163, 823, 374, 1844, 647, 111, 842, 824, 650]. retracts [154]. retrieval [263, 1257, 212]. return [330]. reversal [1458, 75]. reversals [2130, 1136, 524, 2021, 2002]. reverse [1815]. reversed [30]. reversibility [2009]. Reversible [307, 2122, 2124, 37, 2126, 1160]. review [967]. revised [1332]. revisited [1824, 754, 189, 751, 1744, 1412, 1388]. Revisiting [252]. rewrite [249, 174, 841, 168, 1726, 130]. Rewriting [1033, 127, 673, 1034, 864, 1035, 1036, 851, 1500, 1304, 1792, 1038, 269, 494, 2180, 598, 1173, 638, 487, 586, 1039, 2036, 451, 1041, 1524, 1042]. rewritings [1373]. RFSAs [1828]. Rhodes [92]. rhombi [1960]. Rice [279]. rich [467]. ride [711]. Riemann [615, 1620]. Right [313, 1726, 1908, 1973]. Right-linear [1726]. rigid [254]. rigidity [158]. Rigorous [649, 647]. ring [1666, 1997, 1868]. rings [271, 639, 1085, 868, 712, 2027, 623, 1398, 1859]. Riordan [1690]. RNA [1985, 1658, 871]. road [1990]. roadmap [1033]. Robin [1604]. robot [1763]. Robson [2208]. Robust [1125, 520, 2196]. robustness [1089]. Role [1587]. root [1231, 2131]. root-conjugacy [1231]. rooted [90, 1680, 1683, 69]. roots [1890, 917, 155]. Rosser [1318]. rotatable [1158]. Rotation [2159, 435]. rotations [1855, 937, 10]. round [1259]. rounded [1890]. Rounding [985, 1806, 1214]. roundings [2134]. roundoff [574]. Routing [1861, 214, 1085, 752, 2104, 1866, 880, 1117, 2144, 299, 2147, 744, 561, 82, 300, 1740]. routings [1398]. row [776]. rowed [314]. royal [1990]. RSTP [583]. RSTP-MSP [583]. RTILE [1576]. ruggedness [618]. rule [269, 598, 244, 1909, 1003, 860]. Rules [1780, 1857, 193, 948, 256, 841, 1042]. rumors [1445]. run [1390]. runs [1581]. runtime [710]. Rush [778].

s [22, 33, 46, 98, 112, 181, 203, 217]. safe [1200, 1945]. safety [1352, 1522, 1814]. salesman [1025, 1149]. salesmen [1836]. same [1307]. sample [2214]. sampling [1759, 314]. sand [595]. Sandpile [2065]. sandpiles [866]. SAT [652, 649, 2220, 2207, 1116, 1064, 650]. satisfaction [1100, 513, 278, 2210, 1069, 656, 2061]. satisfiability [2004, 1538, 2005, 653, 651, 1174]. satisfies [1667]. Satisfying [893, 1181]. saturated [1723]. SC-hyperdecidability [447]. Scalar [1371]. scale [2113, 554, 1251]. scale-sensitive [554]. scaling [1746]. scans [440]. scattered [294, 315, 152]. SCCS [1305]. scenarios [1067]. schedulability [1742]. schedulers [1423]. schedules [1133, 1421]. Scheduling [105, 1818, 2137, 1650, 2098, 1203, 1085, 2107, 710, 2136, 1233, 1544, 1451, 421, 2101, 1530, 1263, 1797, 1337, 312, 394, 1763, 714, 591, 1031, 1996]. Schéma [309]. schemas [1189]. scheme [214, 2043, 1662, 136, 558, 2127]. schemes [2098, 752, 2000, 1865, 880, 412, 2144, 213, 300, 161, 1376, 852, 987]. Schröder [436, 1686, 436]. Sci [418, 2229, 2226, 1545, 96, 1781, 95, 2227, 692, 460, 1471, 2228]. science [476, 1058, 2049]. scientific [932]. scientifique [932]. scopes [850]. score [502]. scoring [193]. Search [48, 1743, 1141, 385, 383, 2220, 51, 777, 1116, 779, 2208, 1067, 344, 53, 363, 387, 47, 381, 750, 384, 1142, 2103, 564, 56, 886, 1844, 1400, 1845]. Searching [2018, 743, 1983, 597, 1092, 1763, 1029, 1444, 970, 211]. second [132, 273, 940, 1434, 1733, 279, 1708, 2127, 853]. second-order [132, 273, 940, 1434, 1733, 1708, 2127, 853]. secondary [1658]. Secrecy [1426]. Secret [316, 1865, 1329, 213]. Secret-sharing [316]. secrets [1329]. sections [1494]. Secure [263, 2000, 1090]. security [1194]. seen [2221]. segmenting [583]. selected [206]. selection [1496]. selective [3, 95]. Self [680, 1716, 1282, 106, 719, 1540, 1584, 119, 631, 109]. self-avoiding [1584]. self-dual [119, 631]. self-reducibility [1540]. self-reduction [109]. Self-stabilization [1282, 106]. Self-stabilizing [680, 1716]. Semantic [538, 1882, 325, 1721, 952, 1316]. Semantics [1912, 1312, 265, 1937, 2077, 196, 665, 94, 497, 197, 728, 1057, 859, 892, 1180, 851, 901, 949, 1733, 902, 169, 1305, 1220, 1319, 411, 1953, 2170, 1715, 1178, 1883, 413, 733, 1309, 1047, 674, 1192, 484, 1873]. Semi [2107, 100, 363, 959, 793, 494, 83]. Semi-clairvoyant [2107]. semi-complete [494]. semi-continuous [83]. Semi-definite [100]. Semi-dynamic [363]. semi-lattice [793]. semi-Markov [959]. semigroups [369, 1533, 477, 1391, 570, 295, 2093]. semilattices [1906]. semilinear [1803]. Seminar [1848]. seminorms [1008]. Semiperiodic [1231]. Semiretracts [1672]. semiring [578]. semirings [1427, 2089, 1951, 451]. semivaluations [1879]. Sense [1206, 1866, 1507, 567]. sensitive [1168, 554]. sensitivity [1671, 1506, 330]. separable [1183, 69, 1275]. separate [332, 1826]. Separating [1938, 26]. Separation [1827, 997, 1530, 1397, 842]. separators [134]. sequence [502, 2130, 1281, 284, 2149, 1735, 1359, 2021]. sequences [1801, 1666, 518, 1118, 1595, 791, 534, 2159, 1356, 1581, 1579, 1503, 2217, 285, 1240]. sequent [1315, 256]. sequentiability [1817]. Sequential [1860, 882, 1325, 1871, 232, 704, 86, 395]. sequentiality [1227, 2182]. serial [319]. Series [149, 1122, 1326, 1589, 1867, 275, 1251, 1278, 740, 1703]. Series-parallel [149]. server [2108, 1128, 2105, 711, 2109, 715]. servers [1649]. service [971]. services [2187]. ses [309]. sessions [1452]. Set [891, 1349, 1434, 2142, 53, 948, 1088, 2048, 1237, 331, 760, 2066, 689, 1628, 28]. set-pairing [2048]. Set-sharing [891]. sets [1152, 60, 1637, 1643, 1118, 2121, 1750, 1732, 1580, 204, 2018, 64, 1010, 820, 476, 1372, 1447, 1575, 306, 1804, 1159, 3, 95, 1540, 1803, 1143, 1618, 986, 26, 41, 826, 2222, 1877, 1541, 824]. setting [1350]. Seventeen [2039]. shadows [1696]. Shape [1593]. shaped [1357]. shared [1430, 2099]. Sharing [1329, 891, 1865, 316, 1588, 907, 1314, 213, 442, 2161]. sharpened [2146]. Shasha [1658]. shelf [915]. shellability [599]. shelling [605]. shepherds [966]. shifting [974]. shifts [2037, 588, 781, 1453, 508]. shogi [382]. shop [136, 1031]. short [1149, 1133, 284]. shortest [1889, 1382, 1117, 2147, 82, 454, 614, 1358, 1787, 1297]. shortest-path [1117]. should [778]. shrinking [139].

Shuffle [819, 1805, 294, 354, 831]. shuffling [1565]. side [2221]. sided [1605, 765]. sides [152]. Signed [524, 1508]. significant [872]. signing [993]. signs [1570]. Sikorski [1058]. Sim [1148]. Sim-like [1148]. similarity [1801, 871, 855, 1289]. similarity-based [855]. Simone [184]. Simple [257, 1863, 1708, 774, 1254, 176, 1647, 1754, 884, 793, 613, 1131, 614, 688, 2214, 1413, 2228]. simpler [1333]. Simplification [186]. simplified [2015, 955]. simply [1618]. Simulation [1586, 407, 290, 1583, 307, 1724, 169, 623, 239]. Simulations [2168, 960, 446]. simultaneous [254, 1014]. single [1451, 711, 219, 1996]. single-machine [1451]. single-server [711]. Size [452, 1936, 2124, 1700, 2011, 2181, 231, 756, 1990, 1343]. skeleton [592]. sketch [309]. skew [1889]. skewsymmetric [1892]. SLD [167, 855]. SLD-derivations [167]. slender [457, 74]. sliding [2038]. Slissenko [1549]. SLS [2202]. SLS-resolution [2202]. Small [2141, 1402, 1700, 65, 1073, 355]. Smaller [870]. smallest [1150]. smartest [2110]. smooth [1573, 219]. smooth-functions [219]. Smoothness [1251]. snapshot [731]. snooping [883]. social [1987]. sofic [588, 508]. Soft [1943]. software [334, 1043]. Sokoban [384]. solid [1011, 587]. solitaire [2036, 1832]. Solovay [1009]. Solution [1082, 2123, 532, 383, 284, 1278, 1863, 386, 1399, 2197]. solutions [1811, 1526, 1087, 505, 248, 1626, 870, 247, 2198]. solvable [1511, 1494]. solve [38, 507, 1898, 151]. solved [1994]. solvers [1127]. Solving [15, 1215, 1888, 1031]. Some [137, 458, 1595, 104, 1686, 1465, 744, 1277, 480, 2216, 322, 226, 185, 1163, 823, 1624, 1858, 824, 19, 761, 1647, 1683, 2123, 1633, 1975, 1087, 445, 1672, 515, 617, 1756, 595, 1558, 885, 980, 1241, 2052]. SONET [868]. sorted [1708]. Sorting [1122, 1115, 28, 1026, 455, 1136]. SOS [928].

Sound [2084, 674]. Soundness [910, 1725]. Source [75, 121]. SP [502]. SP-score [502]. Space [2017, 1332, 2130, 1235, 578, 1480, 1961, 307, 1088, 893, 1614, 1356, 906, 311, 504, 148, 1468, 522]. space-bounded [522]. Space-efficient [2017, 2130, 311]. space-time [307]. spaces [1183, 2200, 1141, 1937, 1875, 324, 1611, 1610, 845, 749, 1612, 1372, 554, 1616, 1012, 1908, 846, 1621, 1622, 1316]. spanners [1752]. spanning [2146]. Sparse [687, 1700, 450, 204, 1434, 1895, 1397, 894]. sparsification [1594]. spatial [1006, 2070, 621, 409]. special [1130, 506, 1520, 1974]. specialization [334]. Specification [1054, 126, 1039, 1221, 182, 538, 1035, 2187, 1059, 1040, 484]. specifications [1055, 1293, 1422, 924, 325, 321]. specified [854]. Specifying [543, 923]. spectra [1987, 872]. spectral [1993, 1442]. spectrum [1700]. speed [261]. Speeding [1751]. speeds [1175]. Sperner [1698]. spernerity [634]. spherical [1647]. Spiking [1074]. spin [647]. splay [1871]. Splicing [84, 2113, 35, 637, 42]. Split [1892]. Splittable [868]. sporadic [1818]. SPP [2011]. Spring [209]. SQL [1370]. squad [505, 1863, 281, 870]. square [1890, 1975, 2009, 155]. Squaring [1227]. Śrī [956]. st [148]. st-connectivity [148]. stabbing [304]. stability [459, 1025, 2205]. stabilization [1282, 106]. stabilizing [680, 1716]. stable [1527, 1654, 76, 2205, 1656, 2135, 727, 1048, 1514, 878, 674]. stably [2200, 1612, 1915]. stack [1446, 756, 1459]. stack-size [756]. stacks [1122]. stage [1866, 136, 335]. stamping [1171]. standard [816, 2029, 805, 896]. Star [71, 580, 1731, 85]. star-free [85]. state [469, 1738, 578, 781, 1816, 1817, 13, 37, 159, 493, 1531, 1820, 766, 664, 1477, 24, 1074, 1863, 703, 2002, 993]. state-space [578]. Statecharts [1178]. statements [2048]. states [1076, 1150, 1260, 159, 1513, 678]. Static [880, 497, 1029]. Statistical [654, 646, 89]. statistics [6, 1390, 1094]. Steiner [2106, 573, 1589, 143, 1383, 1635, 1344, 111]. step [279, 1178, 391]. steps [1974]. still [1379]. stochastic [572, 1031, 2179]. Stone [1183, 2169]. stop [1570]. Storage [2211, 263, 1462]. Storage-to-tree [2211]. Strand [1063]. strategic [1415]. Strategies [1851, 381, 1780, 1141, 1647, 1866, 14, 1027, 1520, 1308, 970, 2049]. strategy [1092, 564]. Stratified [1937, 415]. stream [485, 1350, 2078]. Streams [1099, 1785, 1703]. street [1030]. stretching [708]. Strictly [177]. Strictness [805]. String [1995, 269, 2180, 347, 598, 638, 660, 2032, 2036, 1746, 7]. String-matching [1995]. string-rewriting [598]. string-tensor [1746]. strings [1099, 1551, 793, 1505, 346, 348, 596, 797, 7]. Strip [234, 973]. strip-like [973]. Strong [579, 809, 979, 735, 2043, 205, 1530, 1285, 261, 1177, 185, 398, 120]. stronger [1269]. strongly [1125, 1453]. Structural [486, 358, 276, 1289, 1226, 104, 877, 1273]. Structure [2201, 595, 395, 2134, 2077, 336, 1658, 476, 298, 773, 991, 542, 1251, 911, 2066, 1104]. Structured [1730, 1055, 539, 466, 1897]. structures [99, 2000, 746, 1130, 1199, 749, 696, 2044, 406, 1265, 503, 923, 1985, 1572, 739, 63, 489, 1121, 478, 1574, 871, 1161, 162, 495, 2050, 720, 1727, 853, 1779]. Structuring [1441]. Student [2057]. study [89, 1458, 756, 80, 163]. Sturmian [816, 1595, 2037, 346, 437, 1574, 242]. style [256, 1562]. styles [2060]. subclass [1423, 509]. subclasses [19]. subdomination [1354]. subgraph [592]. subgraphs [2146, 1975, 1162, 353, 1584, 1854]. subgroups [1669, 241]. sublinear [76]. sublogarithmic [522]. submodular [362, 1091, 115]. submonoids [1144]. subregion [212]. Subresultants [1388]. subsequence [1253, 1999]. subsequences [2038]. subsequential [1232]. Subset [696, 2051]. subsets [1611, 1229, 1277, 1753, 1193, 1104]. subshifts [579, 147]. Substitution [2171, 940, 1559]. Substitutions [435, 1668, 1512, 1968, 166, 671, 729]. substring [1999, 1166]. substring-preprocessing [1166]. substrings [1655, 699]. substructures [516]. subsume [926]. subsumption [828]. subsymbolic [434]. subsystems [475]. subterm [400]. Subtractive [399]. subtrees [60]. subtype [893]. Subtyping [667, 904]. subword [828]. successful [334]. succession [948]. successor [474]. succinct [2195]. succinctness [1821]. sufficient [6, 1150]. suffix [1529, 1537, 1582, 774]. suffixes [1468]. suggested [767]. Suites [233]. sum [415, 2051]. sums [617, 574, 943, 684]. sup [1623]. sup-lattices [1623]. Super [1602, 1925, 1919, 1999, 1922]. super- [1999]. super-/subsequence [1999]. super-/substring [1999]. super-recursive [1919]. Super-tasks [1925]. super-Turing [1922]. superimposed [1645]. supernondeterministic [1821]. support [832]. supporting [1117, 2144]. supports [332]. surface [90]. surfaces [979, 615, 981, 627]. survey [644, 1098, 1232, 901, 299, 2065, 622, 1438, 40]. SVP [1028]. swans [1919]. Swinging [253]. switches [2099]. switching [562]. Sylvester [530]. symbol [1539, 2179]. Symbolic [467, 267, 469, 1186, 1089]. symbolical [1976]. symmetric [2056, 1866, 1661, 565, 1800]. symmetrically [1400]. Symmetries [1118, 1213]. Symmetry [1969, 209, 1839, 1341]. Synchronization [1739, 934, 1997, 444, 505, 257, 1863, 281]. synchronized [1443, 438, 395].

Synchronizing [2176, 1335]. synchronous [1673]. syntactic [1551, 2181, 702, 670]. syntactical [1195]. syntax [1057, 2171, 659]. syntenic [2140]. synthesis [1550, 1423, 671]. Synthesizing [552]. system [2165, 1874, 178, 196, 1127, 892, 763, 1573, 2205, 1799, 2201, 1770, 971, 1416, 1187, 702, 1410, 334, 811, 1104, 1209, 1273]. systems [469, 1172, 780, 1777, 1325, 709, 683, 735, 1184, 1364, 391, 804, 2229, 459, 1055, 938, 182, 1742, 1383, 539, 928, 444, 875, 1470, 578, 1526, 1553, 291, 1852, 1262, 1236, 433, 519, 1460, 35, 1500, 466, 440, 969, 1090, 1365, 269, 494, 1554, 404, 1279, 249, 198, 2072, 2180, 535, 1543, 1988, 1803, 808, 183, 408, 1285, 261, 174, 682, 637, 598, 961, 1058, 1454, 1423, 766, 962, 259, 1408, 1574, 839, 1948, 1895, 138, 487, 770, 1366, 1194, 1315, 1397, 1171, 671, 308, 1896, 1039, 253, 1898, 2092, 2079, 2036, 1623]. systems [285, 740, 1191, 339, 1407, 954, 1499, 662, 44, 374, 1652, 1134, 1638, 413, 916, 184, 168, 1726, 1915, 862, 1202, 2197, 428, 1634, 1259, 2198, 130, 813, 1192]. Systolic [58].

T [1957, 628]. T-colorings [628]. T-tetrominoes [1957]. T0 [846]. T0-spaces [846]. tableaux [268]. tables [831, 314, 2145, 1029]. tabling [2202]. Tabular [1290]. tackling [745]. taking [1974]. Taming [2115]. tandem [1751]. tape [1542, 1462]. tardy [1421]. task [709, 228]. tasks [1084, 1025, 1742, 1421, 1534, 1818, 1207, 751, 1925]. Tau [1704]. TAUT [1104, 846]. tautologies [179]. teaching [221]. team [555]. teams [2006, 1634]. Techniques [2207, 1200, 1893, 1819, 1731, 1898, 380, 28]. telecommunication [1040]. Temporal [382, 597, 1555, 668, 1393, 465, 1296, 415, 1741]. temporary [1084, 751]. tensor [1746]. tenure [1833]. Term [1524, 6, 127, 494, 526, 174, 487, 671, 586, 168, 1479, 1726]. terminal [546]. terminals [143].

Terminating [152, 269]. Termination [127, 598, 1777, 1559, 1477, 397, 403]. terms [1524]. Ternary [2191, 290, 316, 69, 631]. tessalation [1498]. test [1380, 36, 1881, 415, 1215]. testability [110, 27, 460, 2194]. testable [246, 2209, 313, 1469]. Testing [2078, 960, 401, 1486, 523, 1645, 665, 1349, 492, 81, 807, 2003, 77, 1437, 2194, 184, 2051]. tests [2166, 777, 991]. tetrominoes [1957]. Text [1594, 720]. texts [1816, 144]. th [2055]. their [145, 934, 1886, 518, 458, 1428, 866, 1391, 771, 2125, 1582, 272, 1794, 488, 844]. them [1811]. theorem [661, 754, 2083, 1067, 1871, 1332, 686, 327, 1238, 65, 961, 487, 41, 1239, 1320, 1708, 481, 1181, 1484, 92, 279]. theorems [50, 1812, 2090, 621, 842]. Theoret [418, 2229, 2226, 1545, 96, 1781, 95, 2227, 692, 460, 1471, 2228]. theoretic [1484, 344, 47, 279, 361, 1320, 1962, 2212, 1309]. theoretical [387, 442, 620]. theories [1427, 1474, 1730, 1287, 495, 1041, 1174]. theorist [1261]. Theory [501, 1062, 2115, 271, 99, 4, 418, 525, 959, 816, 48, 1167, 474, 206, 1713, 1510, 1071, 1102, 1279, 2072, 1768, 1075, 617, 2074, 1712, 1920, 1448, 2170, 789, 1911, 341, 2080, 395, 339, 689, 1170, 1628, 857, 553, 1848, 1746]. thermography [379]. Thin [305, 277]. those [1618]. thrashing [442]. thread [938, 2062]. threads [1187]. Three [1919, 2048, 531, 73, 121, 88, 197, 432, 503, 1685, 1353, 1500, 1219, 755, 1674, 1143, 315, 1440]. Three-dimensional [121, 88, 432, 1440]. three-label [1353]. three-nonterminal [315]. Three-part [531]. Three-player [73]. Three-variable [2048]. three-word [1143]. Threshold [99, 772, 1081, 1325, 523, 653, 651, 514, 2194, 12]. throughput [1534]. throw [809]. Thue [1241]. ties [1656]. Tight [159, 2022, 247, 1337, 2219, 72]. Tile [1566, 1038]. tiling [1569, 946, 1573, 2059, 234, 973, 441, 1568].

Tilings

[1959, 1758, 941, 1957, 936, 1572, 1963, 1961, 1960, 1567, 1978, 1968, 1958, 1962, 2066]. Time [755, 1936, 2098, 265, 854, 1729, 4, 418, 391, 1175, 390, 1889, 1927, 393, 1742, 1544, 578, 2062, 128, 392, 1086, 1353, 307, 2150, 314, 1135, 699, 330, 1662, 408, 690, 411, 1494, 1251, 1124, 505, 961, 986, 766, 962, 1943, 1393, 129, 963, 394, 557, 839, 964, 1171, 1863, 1039, 267, 1483, 1788, 718, 1468, 840, 136, 1652, 2214, 2194, 2197, 1413, 2228]. Time-minimal [755]. time-stamping [1171]. timed [1172, 1486, 1949, 854, 1228, 2033, 1352, 1450, 1522, 1814, 1121, 1423, 704, 1734, 1714, 401, 1295, 251, 1478]. times [1149, 1797, 713, 1394, 714]. timestamps [1716]. tip [778]. Tissue [1366]. Toeplitz [609, 1892, 1894, 1896]. token [702]. tolerance [1201]. tolerant [552, 455, 2071, 240]. tomographical [1978]. tool [1893]. toolbox [25]. tools [1243]. Top [1252, 1593]. Top-down [1252, 1593]. Topics [1062]. topologic [2158]. Topological [1007, 601, 89, 978, 987, 324, 845, 280, 2123, 1987, 1614, 1616]. topologies [1700, 1506, 1618]. Topology [7, 1609, 2046, 1535, 1612, 1613, 981, 1619, 980, 1624, 1627]. total [323, 191, 1073, 1016, 1910]. totality [805]. tough [626, 1842]. tour [206]. Tournament [1847]. tournaments [611]. towers [151]. trace [406, 1731, 1756, 1677]. traces [1724, 907, 1619]. Tractability [255]. Tractable [668, 504]. Trade [1165]. Trade-off [1165]. tradeoff [1650]. trading [2083]. traffic [868]. trajectories [2001]. transcendence [2037]. transducer [1816, 32, 24]. transducers [2186, 1123, 1227, 1232, 364, 876, 1593, 1817, 2211, 1820, 86, 700, 438]. transform [94, 951]. transformation [1179, 1709, 1157, 2028, 1825, 730]. transformations [1732, 1867, 918, 199, 442, 435, 1290, 1601]. transforming [1780, 336]. transient [1660]. transients [1652]. transition [196, 1167, 1362, 539, 892, 466, 654, 1048, 1315, 253, 1499, 656]. transitional [954]. transitions [1538, 344, 646]. transitive [81, 1374, 1449]. transitivity [66]. translate [797]. translation [849]. transportation [1651]. transpositions [1136, 524]. Trapezoidal [817]. Traveling [1836, 1025, 1339, 2227]. travelling [1149]. traversal [2028, 1356]. treatment [1288]. Tree [1752, 303, 589, 1557, 599, 1774, 214, 687, 974, 385, 1098, 406, 143, 581, 942, 1480, 1658, 364, 876, 1593, 1867, 2159, 1109, 1861, 2211, 1712, 1142, 1249, 417, 1635, 762, 58, 1560, 386, 438, 161, 235, 774, 558, 1252, 1344, 1601, 111, 853]. Tree-based [303, 589]. tree-languages [942]. tree-like [406, 1480, 417, 853]. Tree-shellability [599]. Tree-width [1557]. Trees [124, 1474, 1129, 205, 1683, 573, 1589, 1234, 1384, 128, 1685, 1419, 945, 779, 2208, 1871, 1835, 2143, 1102, 1675, 1355, 236, 574, 210, 1687, 2027, 1843, 69, 443, 1582, 863, 1790, 590, 211, 386, 349, 886, 1400, 1295, 803, 1443]. triangle [2146]. triangles [1959, 632]. triangular [1155, 305, 37, 1894]. triangulated [615, 1587]. triangulation [592]. triangulations [1348, 748, 1689]. trichromatic [1959]. tricks [1002]. tridiagonal [440]. tries [756, 1163]. triggers [535, 1543]. triple [1383, 1397]. triviality [579]. tropical [1427]. tropological [1623]. true [442, 411]. Truly [908, 949]. Truncated [1582]. TSP [1743]. tubes [36]. tuple [1813]. Turing [1139, 34, 1583, 1362, 1963, 1124, 1926, 479, 2141, 873, 1925, 1922, 768]. turn [2127]. tutorial [4, 418]. Tutte [1234]. Two [1076, 1855, 1964, 1683, 1876, 1149, 1158, 806, 261, 784, 1731, 2094, 1498, 1794, 1801, 145, 422, 1122, 640, 1429, 2062, 314, 1333, 1260, 1802, 37, 561, 1263, 505, 353, 1584, 759, 17, 714, 970, 441, 775, 136, 765, 993, 593, 888, 1845, 569]. Two-dimensional [1855, 1964, 1158, 1498, 1794, 145, 422, 2062, 37, 561, 505, 1584, 17, 775]. Two-level [1876, 1845]. two-rowed [314]. two-sided [765]. two-stage [136]. Two-way [1076, 640, 1429, 1333, 1802, 569]. Type [2203, 164, 689, 255, 683, 735, 804, 2229, 48, 805, 1453, 47, 1968, 1770, 808, 1767, 1187, 671, 1030, 1170, 813]. type-theoretic [47]. typed [666]. Types [202, 1426, 856, 1904, 667, 2165, 326, 861, 800, 1767, 507, 1908, 738, 1880, 803]. types=functions [253]. Typical [1806]. Typing [1485].

ukasiewicz [1680]. Ulam [1851, 13].

ultimate [2160, 564]. unambiguity [2182]. Unambiguous [1381]. Unary [1545, 2093, 1, 1850, 1333, 660]. unbounded [261, 225, 1996]. uncaught [896]. uncompletable [427]. Uncomputability [1931, 1925]. Uncomputably [107]. unconditionally [2000]. uncountably [2031]. Undecidability [1756, 61, 464, 784, 40]. Undecidable [1494, 1395, 1571, 1869, 1963, 495, 1638, 857]. Underspecification [688]. Understanding [1188]. Uni [954]. Uni-transitional [954]. Unification [1532, 164, 254, 400, 2085, 857]. unified [163, 953]. Uniform [1005, 1248, 617, 1398, 1242, 1155, 1681, 679, 55, 873, 952, 1827]. uniformément [367]. Uniformly [1420, 1434, 428]. unifying [1416, 1715]. unilateral [479]. Union [1696, 808, 813]. Unions [1174, 625, 1109, 225]. unipolar [1944]. Unique [487, 1626, 1979, 819, 454, 177]. Uniquely [2218, 1617, 1586]. Unit [1308]. unity [155]. Universal [403, 339, 1134, 1365, 1367, 37, 637, 1356, 479, 1916]. Universality [1268, 35, 42]. universally [1603]. universe [1193]. unknown [1535]. unpredictability [1599]. unreliable [777, 1395]. unsatisfiable [1135].

Unsupervised [1070]. untyped [810, 727]. Updatable [2033]. update [707, 1083, 414]. updated [1400]. updates [1445, 684]. updating [68]. Upper [793, 653, 113, 59, 83]. upward [1406]. urgency [492]. used [2153]. Using [38, 1216, 700, 214, 127, 1636, 1542, 391, 2044, 164, 597, 1828, 381, 949, 347, 1819, 1767, 961, 1013, 1454, 704, 1559, 992, 30, 774, 188].

Vacillatory [223]. valences [882]. Valiant [103]. validate [1216]. Validating [992]. valuation [1229]. value [1274, 170, 808, 1177]. value-passing [170]. valued [723, 872, 1877, 674]. values [831, 2136]. Vandermonde [158]. variability [261, 847]. variable [1544, 254, 556, 53, 1135, 2048, 219, 2171, 179, 1798]. variables [1390, 1674, 2210, 1767, 1581, 718]. variance [779, 236]. variant [157]. Variants [1247, 2005, 1500, 2109, 878, 1521]. varieties [1805]. variety [18, 686]. various [1866, 952, 5, 692]. VAUCANSON [2190]. vector [1810, 1803, 454]. Vectorial [1741]. vectors [1881]. Vegas [569]. vehicle [2098, 1797]. Verification [1120, 394, 469, 854, 391, 1327, 1352, 1450, 1522, 1814, 2187, 186, 962, 468, 1948]. Verified [1432]. verifiers [1432]. versions [1468]. versus [642, 103, 2110, 915, 519, 1555, 1049, 1531, 1193, 593]. vertex [100, 640, 1577, 306, 1464, 654, 580, 210]. vertex-disjoint [640]. vertex-ordering [100]. vertex-partitioning [1464]. vertex-rankings [210]. vertices [626, 632]. Verus [393]. very [556, 355, 1413, 2228]. via [652, 221, 2203, 340, 903, 1594, 851, 990, 1748, 574, 1048, 671, 1093, 608, 2214, 1914]. view [1474, 2006, 1034, 928, 1261, 1614, 1616, 229, 58, 2172, 163]. viewpoint [838, 1208]. views [188]. virtual [1205]. visibility [456, 884, 1157, 88]. visual [360, 1865, 213]. visualization [1257]. vital [1358]. volume [419, 461, 470, 483, 500, 536, 548, 568, 602, 635, 658, 693, 705, 725, 741, 843, 865, 889, 897, 913, 920, 1023, 1006]. volumes [549]. Voronoi [1757, 985, 627]. voting [2020, 498]. vs [1329].

Wadge [734]. wait [1207]. wait-free [1207]. Walks [1682, 997, 1584]. Walsh [872, 1015]. Watermelon [1681]. Watson [535, 1543, 954, 1499, 1134, 1638]. wavelength [687]. wavelet [212]. wavelet-based [212]. way [1076, 640, 1429, 2029, 1333, 282, 1802, 1336, 2189, 569, 1160]. wci [1181]. WDM [868]. WDM/SONET [868]. Weak [979, 735, 1590, 766, 2192, 471, 175, 448, 999, 1285, 1561, 472, 88, 1324]. Weak-order [1590]. weaker [1324]. weakly [1573]. web [190, 2101, 597]. webbed [1874, 340]. weight [1589, 592, 1970, 1437, 748, 119]. Weighted [604, 2186, 1949, 2105, 448, 1649, 1532, 1819, 1359, 517, 630, 24, 1787]. weights [193, 278]. Weil [2019]. Well [2042, 2178, 466, 2091, 1779, 497, 1357, 167, 558, 1108]. Well-behaved [2042, 1779]. well-founded [497]. well-quasi-ordering [167]. Well-quasi-orders [2091]. well-shaped [1357]. Well-structured [466]. wellfounded [2171]. which [1447, 2205, 332, 1513]. whose [1667, 1552, 1994, 2205, 1618, 1568]. Wide [1866, 116, 567]. Wide-diameter [116]. Wide-sense [1866, 567]. widest [453]. width [1993, 1464, 149, 1557, 222]. Wilf [754, 1239]. win [1604]. windows [2038]. Winning [2049]. Wire [583, 1073]. wireless [1467, 2158]. wise [1392]. within [1028, 981, 1289]. without [1984, 1811, 1137, 2155, 1445]. wMSO [1287]. Woll [2219]. word [135, 1994, 1304, 1143, 1494, 2191, 1646, 2094]. Words [1667, 506, 822, 1669, 2203, 1123, 827, 1965, 1539, 831, 1126, 754, 2213, 1856, 525, 816, 1231, 1739, 2037, 1749, 1977, 817, 1963, 1533, 445, 1330, 584, 1439, 828, 1661, 818, 1675, 245, 1748, 437, 879, 821, 1238, 830, 829, 1239, 427, 2055, 356, 887, 1853, 1676, 242]. work [1083, 2108]. workflows [745]. works [1256]. worlds [1883]. worst [2035]. worst-case [2035]. Wright [1068]. WWW [1257].

x [793, 503, 63, 2078]. X-machine [2078]. X-rays [503, 63]. XII [132]. XIII [273]. XIV [1434]. XOR [2223].

Yanov [1510]. Yantra [956]. years [743, 372, 373]. yield [2207]. yielding [1989]. Yoneda [906].

Zero [77, 1349, 1269, 108, 852, 286]. zero-one [1269, 852, 286]. zeta [831]. Zhang [1658]. Ziv [1528]. ZPC [696]. ZPC-structures [696].

References

Chrobak:1986:FAU

[1] M. Chrobak. Finite automata and unary languages. Theoretical Computer Science, 47(2):149–158, ???? 1986. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See errata [1545].

DeAgostino:1994:PCP

[2] Sergio De Agostino. P-complete problems in data compression. Theoretical Computer Science, 127(1):181–186, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1994&volume=127&issue=1&aid=1568.

See erratum [96].

Hemaspaandra:1996:RCS

[3] Lane A. Hemaspaandra, Albrecht Hoene, and Mitsunori Ogihara. Reducibility classes of P -selective sets. Theoretical Computer Science, 155(2):447–457, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=2035.

See erratum [95].

Bernardo:1998:TET

[4] Marco Bernardo and Roberto Gorrieri. A tutorial on EMPA: A theory of concurrent processes with nondeterminism, priorities, probabilities and time. Theoretical Computer Science, 202(1– 2):1–54, July 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/202/1-2/2663.pdf.  See

corrigendum [418].

Spreen:1999:FPL

[5] Dieter Spreen. On functions preserving levels of approximation: A refined model construction for various lambda calculi. Theoretical Computer Science, 212(1–2):261–303, February 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
212/1-2/2964.pdf. See corrigendum

[692].

Apolloni:2000:SSP

[6] B. Apolloni and C. Gentile. P -sufficient statistics for PAC learning k-term-DNF formulas through enumeration. Theoretical Computer Science, 230(1– 2):1–37, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/21/article.pdf.

delaHiguera:2000:TSM

[7] C. de la Higuera and F. Casacuberta. Topology of strings: Median string is NP-complete. Theoretical Computer Science, 230(1–2): 39–48, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/22/article.pdf.

Sutner:2000:ACP

[8] Klaus Sutner. σ-automata and Chebyshev-polynomials. Theoretical Computer Science, 230(1–2):49– 73, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/23/article.pdf.

Chen:2000:EAA

[9] Zhi-Zhong Chen. Efficient algorithms for acyclic colorings of graphs. Theoretical Computer Science, 230(1–2): 75–95, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

                                        164/21/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/164/21/24/article.pdf.

Chekhova:2000:CNR

[10] Nataliya Chekhova. Covering numbers of rotations. Theoretical Computer Science, 230(1–2):97–116, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        164/21/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/164/21/25/article.pdf.

Baldwin:2000:CCA

[11] John T. Baldwin and Saharon Shelah. On the classifiability of cellular automata. Theoretical Computer Science, 230(1–2):117–129, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        164/21/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/164/21/26/article.pdf.

Uehara:2000:IPD

[12] Ryuhei Uehara, Kensei Tsuchida, and Ingo Wegener. Identification of partial disjunction, parity, and threshold functions. Theoretical Computer Science, 230(1–2):131–147, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        164/21/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/164/21/27/article.pdf.

Gerla:2000:CSL

[13] Brunella Gerla. Conditioning a state by a Łukasiewicz event: a probabilistic approach to Ulam games. Theoretical Computer Science, 230(1–2): 149–166, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/28/article.pdf.

Cicalese:2000:OSA

[14] Ferdinando Cicalese and Ugo Vaccaro. Optimal strategies against a liar. Theoretical Computer Science, 230(1–2): 167–193, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/29/article.pdf.

Breuker:2000:SD

[15] D. M. Breuker, J. W. H. M. Uiterwijk, and H. J. van den Herik. Solving 8 × 8 domineering. Theoretical Computer Science, 230(1–2):195– 206, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/30/article.pdf.

Ettinger:2000:MPG

[16] J. Mark Ettinger. A metric for positional games. Theoretical Computer Science, 230(1–2):207–219, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/31/article.pdf.

Malgouyres:2000:HTD

[17] Rémy Malgouyres. Homotopy in twodimensional digital images. Theoretical Computer Science, 230(1–2): 221–233, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/32/article.pdf.

Crvenkovic:2000:VKA

[18] S. Crvenković, I. Dolinka, and Z. Ésik. The variety of Kleene algebras with conversion is not finitely based. Theoretical Computer Science, 230(1–2): 235–245, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/33/article.pdf.

Asveld:2000:IPS

[19] Peter R. J. Asveld and Anton Nijholt. The inclusion problem for some subclasses of context-free languages. Theoretical Computer Science, 230(1–2): 247–256, January 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/34/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/164/21/34/article.pdf.

Anonymous:2000:Ia

[20] Anonymous.  Index.  Theoreti-

cal Computer Science, 230(1–2):291, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/35/article.pdf.

Anonymous:2000:Ib

[21] Anonymous.  Index.  Theoretical

Computer Science, 230(1–2):293–299, January 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
164/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/164/21/36/article.pdf.

Anonymous:2000:Ea

[22] Anonymous. Editorial(s). Theoretical Computer Science, 231(1):1– 3, January 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/21/21/article.pdf.

Caron:2000:LMP

[23] Pascal Caron. LANGAGE: A Maple package for automaton characterization of regular languages. Theoretical Computer Science, 231(1):5–15, January 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/21/22/article.pdf.

Mohri:2000:DPW

[24] Mehryar Mohri, Fernando Pereira, and Michael Riley. The design principles of a weighted finite-state transducer library. Theoretical Computer Science, 231(1):17–32, January 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/21/23/article.pdf.

Silberztein:2000:IFT

[25] Max Silberztein. INTEX: an FST toolbox. Theoretical Computer Science, 231(1):33–46, January 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/21/24/article.pdf.

Lescow:2000:MSS

[26] Helmut Lescow and Jens Vöge. Minimal separating sets for acceptance conditions in Muller Automata. Theoretical Computer Science, 231(1): 47–57, January 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/25/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/163/21/25/article.pdf.

Trahtman:2000:OEO

[27] A. N. Trahtman. Optimal estimation on the order of local testability of finite automata. Theoretical Computer Science, 231(1):59–74, January 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/21/26/article.pdf. See er-

ratum [460].

Ziadi:2000:SDT

[28] D. Ziadi. Sorting and doubling techniques for set partitioning and automata minimization problems. Theoretical Computer Science, 231(1): 75–87, January 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/21/27/article.pdf.

Ponty:2000:ENF

[29] J.-L. Ponty. An efficient nullfree procedure for deciding regular language membership. Theoretical Computer Science, 231(1):89– 101, January 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/21/28/article.pdf.

Salomaa:2000:EIR

[30] K. Salomaa, X. Wu, and S. Yu. Efficient implementation of regular languages using reversed alternating finite automata. Theoretical Computer Science, 231(1):103–111, January 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        163/21/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/163/21/29/article.pdf.

Brzozowski:2000:AAB

[31] J. A. Brzozowski and R. Negulescu. Automata of asynchronous behaviors. Theoretical Computer Science, 231(1): 113–128, January 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        163/21/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/163/21/30/article.pdf.

Maurel:2000:PMT

[32] Denis Maurel. Pseudo-minimal transducer. Theoretical Computer Science, 231(1):129–139, January 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/163/21/31/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/163/21/31/article.
                                        pdf.

Anonymous:2000:Eb

[33] Anonymous. Editorial(s). Theoretical Computer Science, 231(2):141– 142, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/21/article.pdf.

Boldi:2000:TCA

[34] Paolo Boldi and Sebastiano Vigna. The Turing closure of an Archimedean field. Theoretical Computer Science, 231(2): 143–156, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/22/article.pdf.

Ferretti:2000:UPS

[35] C. Ferretti, G. Mauri, S. Kobayashi, and T. Yokomori. On the universality of Post and splicing systems. Theoretical Computer Science, 231(2): 157–170, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/23/article.pdf.

Ferretti:2000:NTT

[36] C. Ferretti, G. Mauri, and C. Zandron. Nine test tubes generate any RE language. Theoretical Computer Science, 231(2):171–180, January 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/24/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/163/22/24/article.pdf.

Imai:2000:CUT

[37] Katsunobu Imai and Kenichi Morita. A computation-universal two-dimensional 8-state triangular reversible cellular automaton. Theoretical Computer Science, 231(2):181–191, January 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        163/22/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/163/22/25/article.pdf.

Kari:2000:UDS

[38] Lila Kari, Greg Gloor, and Sheng Yu. Using DNA to solve the Bounded Post Correspondence Problem. Theoretical Computer Science, 231(2):193– 203, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        163/22/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/163/22/26/article.pdf.

Maes:2000:MMA

[39] Arnaud Maes. More on morphisms and almost-periodicity. Theoretical Computer Science, 231(2):205– 215, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        163/22/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/163/22/27/article.pdf.

Margenstern:2000:FBD

[40] Maurice Margenstern. Frontier between decidability and undecidability: a survey. Theoretical Computer Science, 231(2):217–251, January 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/28/article.pdf.

Michaux:2000:ITB

[41] C. Michaux and C. Troestler. Isomorphism theorem for BSS recursively enumerable sets over real closed fields. Theoretical Computer Science, 231(2): 253–273, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/29/article.pdf.

Paun:2000:DCB

[42] Gheorghe Păun. DNA computing based on splicing: universality results. Theoretical Computer Science, 231(2): 275–296, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/30/article.pdf.

Sakamoto:2000:IDP

[43] Hiroshi Sakamoto and Daisuke Ikeda. Intractability of decision problems for finite-memory automata. Theoretical Computer Science, 231(2):297– 308, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/31/article.pdf.

Senizergues:2000:CFS

[44] Géraud Sénizergues. Complete formal systems for equivalence problems. Theoretical Computer Science, 231(2): 309–334, January 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/32/article.pdf.

Anonymous:2000:Ic

[45] Anonymous.  Index.  Theoretical

Computer Science, 231(2):335–336, January 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
163/22/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/163/22/33/article.pdf.

Anonymous:2000:Ec

[46] Anonymous. Editorial(s). Theoretical Computer Science, 232(1–2): 1–4, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/21/article.pdf.

Galmiche:2000:PST

[47] Didier Galmiche and David J. Pym. Proof-search in type-theoretic languages: an introduction. Theoretical Computer Science, 232(1–2):5– 53, February 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/22/article.pdf.

Caldwell:2000:SAT

[48] James L. Caldwell, Ian P. Gent, and Judith Underwood. Search algorithms in type theory. Theoretical Computer Science, 232(1–2):55– 90, February 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/23/article.pdf.

McDowell:2000:CEL

[49] Raymond McDowell and Dale Miller. Cut-elimination for a logic with definitions and induction. Theoretical Computer Science, 232(1–2):91– 119, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/24/article.pdf.

Arai:2000:ENF

[50] Toshiyasu Arai and Grigori Mints. Extended normal form theorems for logical proofs from axioms. Theoretical Computer Science, 232(1–2): 121–132, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/25/article.pdf.

Cervesato:2000:ERM

[51] Iliano Cervesato, Joshua S. Hodas, and Frank Pfenning. Efficient resource management for linear logic proof search. Theoretical Computer Science, 232(1–2):133– 163, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/26/article.pdf.

Bunder:2000:PFA

[52] M. W. Bunder. Proof finding algorithms for implicational logics. Theoretical Computer Science, 232(1–2): 165–186, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/27/article.pdf.

Felty:2000:CCF

[53] Amy Felty. The calculus of constructions as a framework for proof search with set variable instantiation. Theoretical Computer Science, 232(1–2): 187–229, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/28/article.pdf.

Galmiche:2000:CML

[54] D. Galmiche. Connection methods in linear logic and proof nets construction. Theoretical Computer Science, 232(1– 2):231–272, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/29/article.pdf.

Nadathur:2000:CBC

[55] Gopalan Nadathur. Correspondences between classical, intuitionistic and uniform provability. Theoretical Computer Science, 232(1–2): 273–298, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/30/article.pdf.

Ritter:2000:IFC

[56] Eike Ritter, David Pym, and Lincoln Wallen. On the intuitionistic force of classical search. Theoretical Computer Science, 232(1–2):299– 333, February 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
165/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/165/21/31/article.pdf.

Anonymous:2000:Id

[57] Anonymous.  Index.  Theoreti-

cal Computer Science, 232(1–2):335, February 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        165/21/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/165/21/32/article.pdf.

Monti:2000:STL

[58] Angelo Monti and Adriano Peron. Systolic tree ω-languages: the operational and the logical view. Theoretical Computer Science, 233(1–2): 1–18, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/21/article.pdf.

Hashiguchi:2000:NUB

[59] Kosaburo Hashiguchi. New upper bounds to the limitedness of distance automata. Theoretical Computer Science, 233(1–2):19–32, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/22/article.pdf.

Akutsu:2000:ALC

[60] Tatsuya Akutsu and Magnús M. Halldórsson. On the approximation of largest common subtrees and largest common point sets. Theoretical Computer Science, 233(1–2):

33–50, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/23/article.pdf.

Robilliard:2000:UEP

[61] D. Robilliard and D. Simplot. Undecidability of existential properties in picture languages. Theoretical Computer Science, 233(1–2):51–74, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/24/article.pdf.

Caron:2000:CGA

[62] Pascal Caron and Djelloul Ziadi. Characterization of Glushkov automata. Theoretical Computer Science, 233(1– 2):75–90, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/25/article.pdf.

Gardner:2000:CCD

[63] R. J. Gardner, P. Gritzmann, and D. Prangenberg. On the computational complexity of determining polyatomic structures by X-rays. Theoretical Computer Science, 233(1–2): 91–106, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

                                        166/21/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/26/article.pdf.

Cattell:2000:CGM

[64] Kevin Cattell, Michael J. Dinneen, Rodney G. Downey, Michael R. Fellows, and Michael A. Langston. On computing graph minor obstruction sets. Theoretical Computer Science, 233(1–2):107–127, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/27/article.pdf.

Kawahara:2000:SFC

[65] Yasuo Kawahara and Masao Mori. A small final coalgebra theorem. Theoretical Computer Science, 233(1–2): 129–145, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/28/article.pdf.

Cattaneo:2000:ETR

[66] Gianpiero Cattaneo, Enrico Formenti, Giovanni Manzini, and Luciano Margara. Ergodicity, transitivity, and regularity for linear cellular automata over Z_{m}. Theoretical Computer Science, 233(1–2):147– 164, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/29/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/166/21/29/article.pdf.

Bermond:2000:ECC

[67] Jean-Claude Bermond, Luisa Gargano, Stephan Perennes, Adele A. Rescigno, and Ugo Vaccaro. Efficient collective communication in optical networks. Theoretical Computer Science, 233(1–2):165–189, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/30/article.pdf.

Evfimievski:2000:PAU

[68] Alexandre V. Evfimievski. A probabilistic algorithm for updating files over a communication link. Theoretical Computer Science, 233(1–2):191– 199, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/31/article.pdf.

Lungo:2000:LTT

[69] Alberto Del Lungo, Francesco Del Ristoro, and Jean-Guy Penaud. Left ternary trees and non-separable rooted planar maps. Theoretical Computer Science, 233(1–2):201–215, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/32/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/32/article.pdf.

Germain:2000:LRD

[70] Christian Germain and Jean Pallo. Langages rationnels définis avec une concaténation non-associative. Theoretical Computer Science, 233(1–2): 217–231, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/33/article.pdf.

Rabinovich:2000:SFE

[71] Alexander Rabinovich. Star free expressions over the reals. Theoretical Computer Science, 233(1–2):233– 245, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/34/article.pdf.

Aumann:2000:CRT

[72] Yonatan Aumann, Judit Bar-Ilan, and Uriel Feige. On the cost of recomputing: Tight bounds on pebbling with faults. Theoretical Computer Science, 233(1–2):247–261, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        166/21/35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/166/21/35/article.pdf.

Propp:2000:TPI

[73] James Propp. Three-player impartial games. Theoretical Computer Science, 233(1–2):263–278, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/36/article.pdf.

Nishida:2000:SL

[74] T. Y. Nishida and A. Salomaa. On slender 0L languages. Theoretical Computer Science, 233(1–2):279–286, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/37/article.pdf.

Goles:2000:SRC

[75] Eric Goles and Erich Prisner. Source reversal and chip firing on graphs. Theoretical Computer Science, 233(1–2): 287–295, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/38/article.pdf.

Feder:2000:SPA

[76] Tomás Feder, Nimrod Megiddo, and Serge A. Plotkin. A sublinear parallel algorithm for stable matching. Theoretical Computer Science, 233(1–2): 297–308, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/39/article.pdf.

Karpinski:2000:ZTA

[77] Marek Karpinski, Alf van der Poorten, and Igor Shparlinski. Zero testing of p-adic and modular polynomials. Theoretical Computer Science, 233(1–2): 309–317, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/40/article.pdf.

Kalyanasundaram:2000:ODA

[78] Bala Kalyanasundaram and Kirk R. Pruhs. An optimal deterministic algorithm for online b-matching. Theoretical Computer Science, 233(1–2): 319–325, February 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/41/article.pdf.

Anonymous:2000:Ie

[79] Anonymous.  Index.  Theoreti-

cal Computer Science, 233(1–2):327, February 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
166/21/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/166/21/42/article.pdf.

Puitg:2000:FMH

[80] François Puitg and Jean-François Dufourd. Formalizing mathematics in higher-order logic: A case study in geometric modelling. Theoretical Computer Science, 234(1–2):1– 57, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/21/article.pdf.

Habib:2000:LBP

[81] Michel Habib, Ross McConnell, Christophe Paul, and Laurent Viennot. Lex-BFS and partition refinement, with applications to transitive orientation, interval graph recognition and consecutive ones testing. Theoretical Computer Science, 234(1– 2):59–84, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/22/article.pdf.

Kralovic:2000:CSP

[82] Rastislav Kráľovič, Peter Ružička, and DanielŠtefankovič. The complexity of shortest path and dilation bounded interval routing. Theoretical Computer Science, 234(1–2): 85–107, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/23/article.pdf.

Weihrauch:2000:CCL

[83] Klaus Weihrauch and Xizhong Zheng. Computability on continuous, lower semi-continuous and upper semicontinuous real functions. Theoretical Computer Science, 234(1–2):109– 133, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        170/21/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/170/21/24/article.pdf.

Pixton:2000:SAF

[84] Dennis Pixton. Splicing in abstract families of languages. Theoretical Computer Science, 234(1–2): 135–166, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        170/21/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/170/21/25/article.pdf.

Salomaa:2000:AFA

[85] Kai Salomaa and Sheng Yu. Alternating finite automata and starfree languages. Theoretical Computer Science, 234(1–2):167–176, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/170/21/26/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/170/21/26/article.
                                        pdf.

Mohri:2000:MAS

[86] Mehryar Mohri. Minimization algorithms for sequential transducers. Theoretical Computer Science, 234(1–2):

177–201, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/27/article.pdf.

Achlioptas:2000:CAR

[87] Dimitris Achlioptas, Marek Chrobak, and John Noga. Competitive analysis of randomized paging algorithms. Theoretical Computer Science, 234(1– 2):203–218, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/28/article.pdf.

Wang:2000:TDW

[88] Cao An Wang and Binhai Zhu. Threedimensional weak visibility: Complexity and applications. Theoretical Computer Science, 234(1–2):219– 232, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/29/article.pdf.

Francon:2000:TMS

[89] Jean Françon and Yves Bertrand. Topological 3D-manifolds: a statistical study of the cells. Theoretical Computer Science, 234(1–2):233– 254, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

                                        170/21/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/170/21/30/article.pdf.

Arques:2000:CRM

[90] Didier Arquès and Alain Giorgetti. Counting rooted maps on a surface. Theoretical Computer Science, 234(1– 2):255–272, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        170/21/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/170/21/31/article.pdf.

Pavan:2000:CDP

[91] A. Pavan and Alan L. Selman. Complete distributional problems, hard languages, and resource-bounded measure. Theoretical Computer Science, 234(1–2):273–286, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/170/21/32/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/170/21/32/article.
                                        pdf.

Esik:2000:PKR

[92] Z. Ésik. A proof of the Krohn-Rhodes Decomposition Theorem. Theoretical Computer Science, 234(1–2): 287–300, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        170/21/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/170/21/33/article.pdf.

Lange:2000:FPA

[93] Tanja Lange and Arne Winterhof. Factoring polynomials over arbitrary finite fields. Theoretical Computer Science, 234(1–2):301– 308, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/34/article.pdf.

Boudol:2000:SCN

[94] Gérard Boudol. On the semantics of the call-by-name CPS transform. Theoretical Computer Science, 234(1–2): 309–321, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/35/article.pdf.

Hemaspaandra:2000:ERC

[95] L. Hemaspaandra, A. Hoene, and M. Ogihara. Erratum to “Reducibility classes of P -selective sets” [Theoret. Comput. Sci 155(2) (1996) 447– 457]. Theoretical Computer Science, 234(1–2):323, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/36/article.pdf. See [3].

DeAgostino:2000:EPC

[96] Sergio De Agostino. Erratum to “P-complete Problems in Data Compression” [Theoret. Comput. Sci.

127 (1994) 181–186]. Theoretical Computer Science, 234(1–2):325– 326, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/37/article.pdf. See [2].

Anonymous:2000:If

[97] Anonymous.  Index.  Theoreti-

cal Computer Science, 234(1–2):399, March 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/38/article.pdf.

Anonymous:2000:Ed

[98] Anonymous. Editorial(s). Theoretical Computer Science, 235(1):1– 2, March 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/22/21/article.pdf.

Bach:2000:TDS

[99] Eric Bach and Marcos Kiwi. Threshold data structures and coding theory. Theoretical Computer Science, 235 (1):3–23, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/22/22/article.pdf.

Blum:2000:SDR

[100] Avrim Blum, Goran Konjevod, R. Ravi, and Santosh Vempala. Semi-definite relaxations for minimum bandwidth and other vertex-ordering problems. Theoretical Computer Science, 235(1): 25–42, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/22/23/article.pdf.

Boyar:2000:MCB

[101] Joan Boyar, René Peralta, and Denis Pochuev. On the multiplicative complexity of Boolean functions over the basis (,,1). Theoretical Computer Science, 235(1):43– 57, March 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/22/24/article.pdf.

Buhrman:2000:NAI

[102] Harry Buhrman, Tao Jiang, Ming Li, and Paul Vitányi. New applications of the incompressibility method: Part II. Theoretical Computer Science, 235 (1):59–70, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/22/25/article.pdf.

Burgisser:2000:CVV

[103] Peter Bürgisser. Cook’s versus Valiant’s hypothesis. Theoretical Computer Science, 235(1):71–88, March 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        167/22/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/22/26/article.pdf.

Codenotti:2000:SSP

[104] Bruno Codenotti, Pavel Pudlák, and Giovanni Resta. Some structural properties of low-rank matrices related to computational complexity. Theoretical Computer Science, 235(1): 89–107, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        167/22/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/22/27/article.pdf.

Edmonds:2000:SD

[105] Jeff Edmonds. Scheduling in the dark. Theoretical Computer Science, 235(1): 109–141, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        167/22/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/22/28/article.pdf.

Levin:2000:SSC

[106] Leonid A. Levin. Self-stabilization of circular arrays of automata. Theoretical Computer Science, 235(1): 143–144, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/22/29/article.pdf.

Rojas:2000:ULI

[107] J. Maurice Rojas. Uncomputably large integral points on algebraic plane curves? Theoretical Computer Science, 235(1):145–162, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/167/22/30/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/167/22/30/article.
pdf.

Silverman:2000:DIP

[108] Joseph H. Silverman. On the distribution of integer points on curves of genus zero. Theoretical Computer Science, 235(1):163– 170, March 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/22/31/article.pdf.

Song:2000:SRL

[109] Fangmin Song, Yongsen Xu, and Yuechen Qian. The self-reduction in lambda calculus. Theoretical Computer Science, 235(1):171–181, March 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/22/32/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/22/32/article.pdf.

Trahtman:2000:AFO

[110] A. N. Trahtman. Algorithms finding the order of local testability of deterministic finite automaton and estimations of the order. Theoretical Computer Science, 235(1):183– 204, March 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        167/22/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/22/33/article.pdf.

Vazirani:2000:RRA

[111] Vijay V. Vazirani. Recent results on approximating the Steiner tree problem and its generalizations. Theoretical Computer Science, 235(1): 205–216, March 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        167/22/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/22/34/article.pdf.

Anonymous:2000:Ee

[112] Anonymous.  Editorial(s).  The-

oretical Computer Science, 235(2): 217, March 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        167/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/21/21/article.pdf.

Era:2000:UBG

[113] Hiroshi Era, Kenjiro Ogawa, and Morimasa Tsuchiya. On upper bound graphs with respect to operations on graphs. Theoretical Computer Science, 235(2):219–223, March 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/21/22/article.pdf.

Hachimori:2000:CCR

[114] Masahiro Hachimori. Constructible complexes and recursive division of posets. Theoretical Computer Science, 235(2):225–237, March 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/167/21/23/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/167/21/23/article.
pdf.

Kashiwabara:2000:ESF

[115] Kenji Kashiwabara. Extremality of submodular functions. Theoretical Computer Science, 235(2):239– 256, March 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/21/24/article.pdf.

Kojima:2000:WDM

[116] Toru Kojima and Kiyoshi Ando. Widediameter and minimum length of fan.

Theoretical Computer Science, 235(2): 257–266, March 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/21/25/article.pdf.

Maehara:2000:CCP

[117] Hiroshi Maehara. Can a convex polyhedron have a developable face-cycle? Theoretical Computer Science, 235(2): 267–270, March 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/21/26/article.pdf.

Nakamigawa:2000:GDF

[118] Tomoki Nakamigawa. A generalization of diagonal flips in a convex polygon. Theoretical Computer Science, 235(2): 271–282, March 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/21/27/article.pdf.

Ozeki:2000:CRC

[119] Michio Ozeki. On covering radii and coset weight distributions of extremal binary self-dual codes of length 40. Theoretical Computer Science, 235(2): 283–308, March 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/28/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/167/21/28/article.pdf.

Sakuma:2000:FCC

[120] Tadashi Sakuma. Forced color classes, intersection graphs and the strong perfect graph conjecture. Theoretical Computer Science, 235(2):309– 324, March 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/21/29/article.pdf.

Sugihara:2000:TDC

[121] Kokichi Sugihara. Three-dimensional convex hull as a fruitful source of diagrams. Theoretical Computer Science, 235(2):325–337, March 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/167/21/30/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/167/21/30/article.
pdf.

Tamura:2000:PMP

[122] Akihisa Tamura. Perfect (0, ±1)- matrices and perfect bidirected graphs. Theoretical Computer Science, 235(2): 339–356, March 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
167/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/167/21/31/article.pdf.

Anonymous:2000:Ig

[123] Anonymous.  Index.  Theoretical

Computer Science, 235(2):357–358, March 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        167/21/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/167/21/32/article.pdf.

Dauchet:2000:TAP

[124] M. Dauchet. Trees in Algebra and Programming. Theoretical Computer Science, 236(1–2): 1, April 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        168/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/168/21/21/article.pdf.

Astesiano:2000:FM

[125] Egidio Astesiano and Gianna Reggio. Formalism and method. Theoretical Computer Science, 236(1–2): 3–34, April 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        168/21/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/168/21/22/article.pdf.

Bouhoula:2000:SPM

[126] Adel Bouhoula, Jean-Pierre Jouannaud, and José Meseguer. Specification and proof in membership equational logic. Theoretical Computer Science, 236(1–2):35– 132, April 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
168/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/168/21/23/article.pdf.

Arts:2000:TTR

[127] Thomas Arts and Jürgen Giesl. Termination of term rewriting using dependency pairs. Theoretical Computer Science, 236(1–2):133– 178, April 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
168/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/168/21/24/article.pdf.

Dessmark:2000:MPC

[128] Anders Dessmark, Andrzej Lingas, and Andrzej Proskurowski. Maximum packing for k-connected partial k-trees in polynomial time. Theoretical Computer Science, 236(1–2): 179–191, April 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
168/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/168/21/25/article.pdf.

Leivant:2000:CAL

[129] Daniel Leivant and Jean-Yves Marion. A characterization of alternating log time by ramified recurrence. Theoretical Computer Science, 236(1– 2):193–208, April 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

                                        168/21/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/168/21/26/article.pdf.

Yamada:2000:LCR

[130] Toshiyuki Yamada, Jürgen Avenhaus, Carlos Loría-Sáenz, and Aart Middeldorp. Logicality of conditional rewrite systems. Theoretical Computer Science, 236(1–2):209– 232, April 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        168/21/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/168/21/27/article.pdf.

Anonymous:2000:Ih

[131] Anonymous.  Index.  Theoret-

ical Computer Science, 236(1–2): 233, April 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        168/21/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/168/21/28/article.pdf.

Courcelle:2000:MSOa

[132] Bruno Courcelle. The monadic secondorder logic of graphs XII: planar graphs and planar maps. Theoretical Computer Science, 237(1–2):1– 32, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        171/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/171/21/21/article.pdf.

Wang:2000:RBR

[133] Yongge Wang. Resource bounded randomness and computational complexity. Theoretical Computer Science, 237 (1–2):33–55, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/22/article.pdf.

Malvestuto:2000:DHP

[134] Francesco Mario Malvestuto and Marina Moscarini. Decomposition of a hypergraph by partial-edge separators. Theoretical Computer Science, 237(1– 2):57–79, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/23/article.pdf.

Birget:2000:RFP

[135] Jean-Camille Birget. Reductions and functors from problems to word problems. Theoretical Computer Science, 237(1–2):81–104, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/24/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/171/21/24/article.
pdf.

Schuurman:2000:PTA

[136] Petra Schuurman and Gerhard J. Woeginger. A polynomial time approximation scheme for the two-stage multiprocessor flow shop problem. Theoretical Computer Science, 237(1–2): 105–122, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/25/article.pdf.

Alimonti:2000:SAC

[137] Paola Alimonti and Viggo Kann. Some APX-completeness results for cubic graphs. Theoretical Computer Science, 237(1–2):123–134, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/26/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/171/21/26/article.
pdf.

Mandache:2000:CPC

[138] Niculae Mandache. On the computational power of context-free PC grammar systems. Theoretical Computer Science, 237(1–2):135– 148, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/27/article.pdf.

vanderWalt:2000:SLR

[139] Andries van der Walt and Sigrid Ewert. A shrinking lemma for random forbidding context languages. Theoretical Computer Science, 237(1–2): 149–158, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/28/article.pdf.

Geffert:2000:AEP

[140] Viliam Geffert, Jyrki Katajainen, and Tomi Pasanen. Asymptotically efficient in-place merging. Theoretical Computer Science, 237(1–2):159– 181, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/29/article.pdf.

Kim:2000:LBP

[141] Changwook Kim and Ivan Hal Sudborough. Leftmove-bounded picture languages. Theoretical Computer Science, 237(1–2):183–195, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/30/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/171/21/30/article.
pdf.

Sengupta:2000:NCB

[142] Rimli Sengupta and H. Venkateswaran. Non-cancellative Boolean circuits: A generalization of monotone Boolean circuits. Theoretical Computer Science, 237(1–2):197–212, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/31/abstract.
html;   http://www.elsevier.nl/

                                        gej-ng/10/41/16/171/21/31/article.
                                        pdf.

Cheng:2000:STP

[143] Siu-Wing Cheng. The Steiner tree problem for terminals on the boundary of a rectilinear polygon. Theoretical Computer Science, 237(1–2): 213–238, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        171/21/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/171/21/32/article.pdf.

Senellart:2000:FPM

[144] Jean Senellart. Fast pattern matching in indexed texts. Theoretical Computer Science, 237(1–2):239– 262, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        171/21/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/171/21/33/article.pdf.

Apostolico:2000:FAT

[145] Alberto Apostolico and Valentin E. Brimkov. Fibonacci arrays and their two-dimensional repetitions. Theoretical Computer Science, 237(1–2): 263–273, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        171/21/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/171/21/34/article.pdf.

Moore:2000:QAQ

[146] Cristopher Moore and James P. Crutchfield. Quantum automata and quantum grammars. Theoretical Computer Science, 237(1–2):275– 306, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/35/article.pdf.

Kurka:2000:RS

[147] Petr Kurka and Alejandro Maass. Realtime subshifts. Theoretical Computer Science, 237(1–2):307–325, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/36/article.pdf.

Poon:2000:SLB

[148] Chung Keung Poon. A space lower bound for st-connectivity on node-named JAGs. Theoretical Computer Science, 237(1–2):327–345, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/37/article.pdf.

Lodaya:2000:SPL

[149] K. Lodaya and P. Weil. Seriesparallel languages and the boundedwidth property. Theoretical Computer Science, 237(1–2):347–380, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/38/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/171/21/38/article.
pdf.

Tohyama:2000:CPD

[150] Hiroaki Tohyama and Akeo Adachi. Complexity of path discovery game problems. Theoretical Computer Science, 237(1–2):381–406, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/39/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/171/21/39/article.
pdf.

Sarkar:2000:DCA

[151] U. K. Sarkar. On the design of a constructive algorithm to solve the multipeg towers of Hanoi problem. Theoretical Computer Science, 237(1–2): 407–421, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/40/article.pdf.

Meduna:2000:TLH

[152] Alexander Meduna. Terminating left-hand sides of scattered context productions M. Nivat. Theoretical Computer Science, 237(1–2):423– 427, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/41/article.pdf.

Simon:2000:BCA

[153] Janos Simon and Shi-Chun Tsai. On the bottleneck counting argument. Theoretical Computer Science, 237(1– 2):429–437, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/42/article.pdf.

Anderson:2000:IR

[154] James A. Anderson. The intersection of retracts of A *. Theoretical Computer Science, 237(1–2):439– 445, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/43/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/43/article.pdf.

Litow:2000:HSR

[155] B. Litow. On Hadamard square roots of unity. Theoretical Computer Science, 237(1–2):447–454, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/171/21/44/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/171/21/44/article.
pdf.

Navarro:2000:IAP

[156] Gonzalo Navarro. Improved approximate pattern matching on hypertext. Theoretical Computer Science, 237(1– 2):455–463, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        171/21/45/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/171/21/45/article.pdf.

Geffert:2000:VIC

[157] Viliam Geffert. A variant of inductive counting. Theoretical Computer Science, 237(1–2):465–475, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        171/21/46/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/171/21/46/article.pdf.

Lokam:2000:RVM

[158] Satyanarayana V. Lokam. On the rigidity of Vandermonde matrices. Theoretical Computer Science, 237(1–2): 477–483, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        171/21/47/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/171/21/47/article.pdf.

Iwama:2000:TBN

[159] Kazuo Iwama, Yahiko Kambayashi, and Kazuya Takaki. Tight bounds on the number of states of DFAs that are equivalent to n-state NFAs. Theoretical Computer Science, 237(1–2):

485–494, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/48/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/48/article.pdf.

Anonymous:2000:Ii

[160] Anonymous.  Index.  Theoreti-

cal Computer Science, 237(1–2):499– 500, April 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
171/21/49/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/171/21/49/article.pdf.

Sabelfeld:2000:TEL

[161] V. Sabelfeld. The tree equivalence of linear recursion schemes. Theoretical Computer Science, 238(1–2): 1–29, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/21/article.pdf.

Panangaden:2000:GIP

[162] Prakash Panangaden and Clark Verbrugge. Generating irregular partitionable data structures. Theoretical Computer Science, 238(1–2):31– 80, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/22/article.pdf.

Roggenbach:2000:TUV

[163] Markus Roggenbach and Mila Majster-Cederbaum. Towards a unified view of bisimulation: a comparative study. Theoretical Computer Science, 238(1– 2):81–130, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        172/21/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/172/21/23/article.pdf.

Codish:2000:TDL

[164] Michael Codish and Vitaly Lagoon. Type dependencies for logic programs using ACI-unification. Theoretical Computer Science, 238(1–2): 131–159, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        172/21/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/172/21/24/article.pdf.

Czaja:2000:PLN

[165] Ludwik Czaja. Process languages and nets. Theoretical Computer Science, 238(1–2):161–181, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/172/21/25/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/172/21/25/article.
                                        pdf.

Kesner:2000:CEN

[166] Delia Kesner. Confluence of extensional and non-extensional λ-calculi with explicit substitutions. Theoretical Computer Science, 238(1–2):

183–220, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/26/article.pdf.

Pacini:2000:LCS

[167] Giuliano Pacini and Maria I. Sessa. Loop checking in SLD-derivations by well-quasi-ordering of goals. Theoretical Computer Science, 238(1–2): 221–246, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/27/article.pdf.

Vagvolgyi:2000:CCG

[168] Sándor Vágvölgyi. Congruential complements of ground term rewrite systems. Theoretical Computer Science, 238(1–2):247–274, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/172/21/28/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/172/21/28/article.
pdf.

Hesselink:2000:FSS

[169] Wim H. Hesselink and Albert Thijs. Fixpoint semantics and simulation. Theoretical Computer Science, 238(1– 2):275–311, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/29/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/172/21/29/article.pdf.

Boreale:2000:CAB

[170] Michele Boreale and Luca Trevisan. A complexity analysis of bisimilarity for value-passing processes. Theoretical Computer Science, 238(1–2): 313–345, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        172/21/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/172/21/30/article.pdf.

Savicky:2000:ROL

[171] P. Savický and S.Žák. A read-once lower bound and a (1,+ k )-hierarchy for branching programs. Theoretical Computer Science, 238(1–2):347– 362, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        172/21/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/172/21/31/article.pdf.

Margaria:2000:GFM

[172] Ines Margaria and Maddalena Zacchi. Generalized filter models. Theoretical Computer Science, 238(1–2): 363–387, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        172/21/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/172/21/32/article.pdf.

Nicola:2000:LBA

[173] Rocco De Nicola and Rosario Pugliese. Linda-based applicative and imperative process algebras. Theoretical Computer Science, 238(1–2):389– 437, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/33/article.pdf.

Katoen:2000:PMA

[174] Joost-Pieter Katoen and Albert Nymeyer. Pattern-matching algorithms based on term rewrite systems. Theoretical Computer Science, 238(1– 2):439–464, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/34/article.pdf.

Ying:2000:WCI

[175] Mingsheng Ying. Weak confluence and τ-inertness. Theoretical Computer Science, 238(1–2):465– 475, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/35/article.pdf.

Arai:2000:NFM

[176] Noriko H. Arai. No feasible monotone interpolation for simple combinatorial reasoning. Theoretical Computer Science, 238(1–2):477– 482, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/36/article.pdf.

Naundorf:2000:SCF

[177] Holger Naundorf. Strictly causal functions have a unique fixed point. Theoretical Computer Science, 238(1–2): 483–488, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/37/article.pdf.

Besnard:2000:WNC

[178] Ph. Besnard and T. Schaub. What is a (non-constructive) non-monotone logical system? Theoretical Computer Science, 238(1–2):489– 494, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/38/article.pdf.

Savicky:2000:DTL

[179] P. Savický and Jiří Sgall. DNF tautologies with a limited number of occurrences of every variable. Theoretical Computer Science, 238(1–2): 495–498, May 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

172/21/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/39/article.pdf.

Anonymous:2000:Ij

[180] Anonymous.  Index.  Theoreti-

cal Computer Science, 238(1–2):499– 500, May 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
172/21/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/172/21/40/article.pdf.

Anonymous:2000:Ef

[181] Anonymous. Editorial(s). Theoretical Computer Science, 239(1):1– 2, May 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/21/21/article.pdf.

Broy:2000:ASR

[182] Manfred Broy. Algebraic specification of reactive systems. Theoretical Computer Science, 239(1):3– 40, May 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/21/22/article.pdf.

Jacobs:2000:OOH

[183] Bart Jacobs. Object-oriented hybrid systems of coalgebras plus monoid actions. Theoretical Computer Science, 239(1):41–95, May 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/21/23/article.pdf.

Ulidowski:2000:FAS

[184] Irek Ulidowski. Finite axiom systems for testing preorder and De Simone process languages. Theoretical Computer Science, 239(1):97– 139, May 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/21/24/article.pdf.

Rao:2000:SCS

[185] M. R. K. Krishna Rao. Some characteristics of strong innermost normalization. Theoretical Computer Science, 239(1):141–164, May 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/173/21/25/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/173/21/25/article.
pdf.

Gribomont:2000:SBV

[186] E. Pascal Gribomont. Simplification of boolean verification conditions. Theoretical Computer Science, 239(1): 165–185, May 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/21/26/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/173/21/26/article.pdf.

Afrati:2000:F

[187] Foto N. Afrati and Phokion G. Kolaitis. Foreword. Theoretical Computer Science, 239(2):187, May 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/21/article.pdf.

Ullman:2000:IIU

[188] Jeffrey D. Ullman. Information integration using logical views. Theoretical Computer Science, 239(2): 189–210, May 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/22/article.pdf.

Chekuri:2000:CQC

[189] Chandra Chekuri and Anand Rajaraman. Conjunctive query containment revisited. Theoretical Computer Science, 239(2):211–229, May 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/23/article.pdf.

Abiteboul:2000:QCW

[190] Serge Abiteboul and Victor Vianu. Queries and computation on the web.

Theoretical Computer Science, 239(2): 231–255, May 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/24/article.pdf.

Flum:2000:GTD

[191] Jörg Flum, Max Kubierschky, and Bertram Ludäscher. Games and total Datalog ≠ queries. Theoretical Computer Science, 239(2):257– 276, May 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/25/article.pdf.

Dong:2000:LPQ

[192] Guozhu Dong, Leonid Libkin, and Limsoon Wong. Local properties of query languages. Theoretical Computer Science, 239(2):277– 308, May 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/26/article.pdf.

Fagin:2000:FIW

[193] Ronald Fagin and Edward L. Wimmers. A formula for incorporating weights into scoring rules. Theoretical Computer Science, 239(2):309– 338, May 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/27/article.pdf.

Anonymous:2000:Ik

[194] Anonymous.  Index.  Theoreti-

cal Computer Science, 239(2):339, May 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
173/23/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/173/23/28/article.pdf.

Metayer:2000:F

[195] Daniel Le Métayer. Foreword. Theoretical Computer Science, 240(1): 1–2, June 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/21/21/article.pdf.

Bonsangue:2000:TSS

[196] M. M. Bonsangue, F. Arbab, J. W. de Bakker, J. J. M. M. Rutten, A. Scutellà, and G. Zavattaro. A transition system semantics for the controldriven coordination language MANI-FOLD. Theoretical Computer Science, 240(1):3–47, June 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/21/22/article.pdf.

Busi:2000:CTS

[197] Nadia Busi, Roberto Gorrieri, and Gianluigi Zavattaro. Comparing three semantics for Linda-like languages. Theoretical Computer Science, 240 (1):49–90, June 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/21/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/174/21/23/article.pdf.

Hedman:2000:CAS

[198] Eric J. Hedman, Joost N. Kok, and Kaisa Sere. Coordinating action systems. Theoretical Computer Science, 240(1):91–115, June 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/21/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/174/21/24/article.pdf.

Jagannathan:2000:CBT

[199] Suresh Jagannathan. Continuationbased transformations for coordination languages. Theoretical Computer Science, 240(1):117–146, June 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/174/21/25/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/174/21/25/article.
                                        pdf.

Amadio:2000:MM

[200] Roberto M. Amadio. On modelling mobility. Theoretical Computer Science, 240(1):147–176, June 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/21/26/article.pdf.

Cardelli:2000:MA

[201] Luca Cardelli and Andrew D. Gordon. Mobile ambients. Theoretical Computer Science, 240(1):177– 213, June 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/21/27/article.pdf.

Nicola:2000:TAC

[202] Rocco De Nicola, GianLuigi Ferrari, Rosario Pugliese, and Betti Venneri. Types for access control. Theoretical Computer Science, 240(1):215– 254, June 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/21/28/article.pdf.

Anonymous:2000:Eg

[203] Anonymous.  Editorial(s).  The-

oretical Computer Science, 240(2): 255, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/23/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/23/21/article.pdf.

Cai:2000:RHC

[204] Jin-Yi Cai and D. Sivakumar. Resolution of Hartmanis’ conjecture for NL-hard sparse sets. Theoretical Computer Science, 240(2):257– 269, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/23/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/174/23/22/article.pdf.

Berry:2000:IET

[205] Vincent Berry and Olivier Gascuel. Inferring evolutionary trees with strong combinatorial evidence. Theoretical Computer Science, 240(2):271– 298, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/23/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/174/23/23/article.pdf.

Chen:2000:STT

[206] Lin Chen. A selected tour of the theory of identification matrices. Theoretical Computer Science, 240(2): 299–318, June 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/23/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/174/23/24/article.pdf.

Reischuk:2000:CLF

[207] Rüdiger Reischuk. Can large fanin circuits perform reliable computations in the presence of faults? Theoretical Computer Science, 240(2):

319–335, June 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/23/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/23/25/article.pdf.

Kobayashi:2000:RLG

[208] Yuji Kobayashi and Friedrich Otto. Repetitiveness of languages generated by morphisms. Theoretical Computer Science, 240(2):337– 378, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/23/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/23/26/article.pdf.

Eades:2000:SAS

[209] Peter Eades and Xuemin Lin. Spring algorithms and symmetry. Theoretical Computer Science, 240(2):379– 405, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/23/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/23/27/article.pdf.

Kashem:2000:AGV

[210] Md. Abul Kashem, Xiao Zhou, and Takao Nishizeki. Algorithms for generalized vertex-rankings of partial k-trees. Theoretical Computer Science, 240(2):407–427, June 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/174/23/28/abstract.

                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/174/23/28/article.
                                        pdf.

Peng:2000:ENS

[211] Sheng-Lung Peng, Chin-Wen Ho, Tsan sheng Hsu, Ming-Tat Ko, and Chuan Yi Tang. Edge and node searching problems on trees. Theoretical Computer Science, 240(2): 429–446, June 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/23/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/174/23/29/article.pdf.

Poulakidas:2000:ICF

[212] A. S. Poulakidas, A. Srinivasan, Ö. Eğecioğlu, O. Ibarra, and T. Yang. Image compression for fast waveletbased subregion retrieval. Theoretical Computer Science, 240(2):447– 469, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/23/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/174/23/30/article.pdf.

Hofmeister:2000:COS

[213] Thomas Hofmeister, Matthias Krause, and Hans U. Simon. Contrast-optimal k out of n secret sharing schemes in visual cryptography. Theoretical Computer Science, 240(2):471– 485, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        174/23/31/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/174/23/31/article.pdf.

Annexstein:2000:MTR

[214] Fred S. Annexstein, Kenneth A. Berman, Tsan-Sheng Hsu, and Ram Swaminathan. A multi-tree routing scheme using acyclic orientations. Theoretical Computer Science, 240(2): 487–494, June 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/23/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/23/32/article.pdf.

Anonymous:2000:Il

[215] Anonymous.  Index.  Theoreti-

cal Computer Science, 240(2):495– 496, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/23/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/23/33/article.pdf.

Anonymous:2000:Im

[216] Anonymous.  Index.  Theoreti-

cal Computer Science, 240(2):497– 503, June 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
174/23/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/174/23/34/article.pdf.

Anonymous:2000:Eh

[217] Anonymous. Editorial(s). Theoretical Computer Science, 241(1–2):

1–2, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/21/article.pdf.

Vitanyi:2000:DEP

[218] Paul Vitányi. A discipline of evolutionary programming. Theoretical Computer Science, 241(1–2):3– 23, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/22/article.pdf.

Long:2000:IBA

[219] Philip M. Long. Improved bounds about on-line learning of smoothfunctions of a single variable. Theoretical Computer Science, 241(1–2): 25–35, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/23/article.pdf.

Takimoto:2000:LEE

[220] Eiji Takimoto, Yoshifumi Sakai, and Akira Maruoka. The learnability of exclusive-or expansions based on monotone DNF formulas. Theoretical Computer Science, 241(1–2):37– 50, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

175/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/24/article.pdf.

Arvind:2000:ELT

[221] V. Arvind and N. V. Vinodchandran. Exact learning via teaching assistants. Theoretical Computer Science, 241(1– 2):51–81, June 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/25/article.pdf.

Nakamura:2000:QLB

[222] Atsuyoshi Nakamura. Query learning of bounded-width OBDDs. Theoretical Computer Science, 241(1–2): 83–114, June 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/26/article.pdf.

Case:2000:VBL

[223] John Case, Sanjay Jain, and Frank Stephan. Vacillatory and BC learning on noisy data. Theoretical Computer Science, 241(1–2):115–141, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/27/article.pdf.

Jain:2000:LLF

[224] Sanjay Jain, Efim Kinber, Steffen Lange, Rolf Wiehagen, and Thomas Zeugmann. Learning languages and functions by erasing. Theoretical Computer Science, 241(1–2):143– 189, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/28/article.pdf.

Shinohara:2000:IIU

[225] Takeshi Shinohara and Hiroki Arimura. Inductive inference of unbounded unions of pattern languages from positive data. Theoretical Computer Science, 241(1–2):191–209, June 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/175/21/29/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/175/21/29/article.
pdf.

Rao:2000:SCP

[226] M. R. K. Krishna Rao. Some classes of Prolog programs inferable from positive data. Theoretical Computer Science, 241(1–2):211– 234, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/30/article.pdf.

Anonymous:2000:In

[227] Anonymous.  Index.  Theoreti-

cal Computer Science, 241(1–2):235, June 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
175/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/175/21/31/article.pdf.

Kraus:2000:ADT

[228] Sarit Kraus and Tatjana Plotkin. Algorithms of distributed task allocation for cooperative agents. Theoretical Computer Science, 242(1–2): 1–27, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/21/article.pdf.

Hwang:2000:PAA

[229] Hsien-Kuei Hwang, Bo-Yin Yang, and Yeong-Nan Yeh. Presorting algorithms: An average-case point of view. Theoretical Computer Science, 242(1– 2):29–40, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/22/article.pdf.

Meer:2000:CPR

[230] Klaus Meer. Counting problems over the reals. Theoretical Computer Science, 242(1–2):41–58, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/176/21/23/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/176/21/23/article.
                                        pdf.

Leung:2000:SPL

[231] Hing Leung and Detlef Wotschke. On the size of parsers and LR(k)- grammars. Theoretical Computer Science, 242(1–2):59–69, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/176/21/24/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/176/21/24/article.
                                        pdf.

Kucera:2000:EDS

[232] Antonín Kučera. Effective decomposability of sequential behaviours. Theoretical Computer Science, 242(1– 2):71–89, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        176/21/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/176/21/25/article.pdf.

Hubert:2000:SE

[233] P. Hubert. Suites équilibrées. Theoretical Computer Science, 242(1–2): 91–108, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        176/21/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/176/21/26/article.pdf.

Merlini:2000:STR

[234] Donatella Merlini, Renzo Sprugnoli, and M. Cecilia Verri. Strip tiling and regular grammars. Theoretical Computer Science, 242(1–2):109– 124, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/27/article.pdf.

Shankar:2000:NAL

[235] Priti Shankar, Amitranjan Gantait, A. R. Yuvaraj, and Maya Madhavan. A new algorithm for linear regular tree pattern matching. Theoretical Computer Science, 242(1–2): 125–142, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/28/article.pdf.

Hubalek:2000:VIP

[236] Friedrich Hubalek. On the variance of the internal path length of generalized digital trees — the Mellin convolution approach. Theoretical Computer Science, 242(1–2): 143–168, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/29/article.pdf.

Escobar-Molano:2000:CCD

[237] Martha L. Escobar-Molano and Shahram Ghandeharizadeh. On the complexity of coordinated display of multimedia objects. Theoretical Computer Science, 242(1–2):169– 197, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        176/21/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/176/21/30/article.pdf.

Arvind:2000:CCG

[238] V. Arvind and N. V. Vinodchandran. The counting complexity of group-definable languages. Theoretical Computer Science, 242(1–2):199– 218, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        176/21/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/176/21/31/article.pdf.

Calude:2000:FNA

[239] Cristian S. Calude, Elena Calude, and Bakhadyr Khoussainov. Finite nondeterministic automata: Simulation and minimality. Theoretical Computer Science, 242(1–2):219– 235, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        176/21/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/176/21/32/article.pdf.

Khuller:2000:FTC

[240] Samir Khuller, Robert Pless, and Yoram J. Sussmann. Fault tolerant K-center problems. Theoretical Computer Science, 242(1–2):237– 245, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/33/article.pdf.

Birget:2000:PCP

[241] J.-C. Birget, S. Margolis, J. Meakin, and P. Weil. PSPACE-complete problems for subgroups of free groups and inverse finite automata. Theoretical Computer Science, 242(1–2): 247–281, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/34/article.pdf.

Vandeth:2000:SWW

[242] Drew Vandeth. Sturmian words and words with a critical exponent. Theoretical Computer Science, 242(1–2): 283–300, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/35/article.pdf.

Mundhenk:2000:HI

[243] Martin Mundhenk. On hard instances. Theoretical Computer Science, 242(1– 2):301–311, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/36/article.pdf.

Martinez:2000:CMF

[244] Conrado Martínez and Salvador Roura. On the competitiveness of the move-to-front rule. Theoretical Computer Science, 242(1–2):313– 325, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/37/article.pdf.

Ilie:2000:LWC

[245] Lucian Ilie. On lengths of words in context-free languages. Theoretical Computer Science, 242(1–2):327– 359, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/38/article.pdf.

Caron:2000:FLT

[246] Pascal Caron. Families of locally testable languages. Theoretical Computer Science, 242(1–2):361– 376, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/39/article.pdf.

Wang:2000:TBS

[247] Biing-Feng Wang. Tight bounds on the solutions of multidimensional divideand-conquer maximin recurrences. Theoretical Computer Science, 242(1– 2):377–401, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/40/article.pdf.

Makanin:2000:PSP

[248] Gennady S. Makanin and Tatiana A. Makanina. Parametrisation of solutions of parametric equation in free monoid. Theoretical Computer Science, 242(1– 2):403–475, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/41/article.pdf.

Gyenizse:2000:PLL

[249] Pál Gyenizse and Sándor Vágvölgyi. A property of left-linear rewrite systems preserving recognizability. Theoretical Computer Science, 242(1–2): 477–498, July 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/42/article.pdf.

Anonymous:2000:Io

[250] Anonymous.  Index.  Theoreti-

cal Computer Science, 242(1–2):499– 500, July 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
176/21/43/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/176/21/43/article.pdf.

Winkowski:2000:PTP

[251] Józef Winkowski. Processes of timed Petri nets. Theoretical Computer Science, 243(1–2):1–34, July 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/21/article.pdf.

DePrisco:2000:RPA

[252] Roberto De Prisco, Butler Lampson, and Nancy Lynch. Revisiting the PAXOS algorithm. Theoretical Computer Science, 243(1–2):35– 91, July 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/22/article.pdf.

Padawitz:2000:STS

[253] Peter Padawitz. Swinging types=functions+relations+transition systems. Theoretical Computer Science, 243(1–2):93–165, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/177/21/23/abstract.
html;   http://www.elsevier.nl/

gej-ng/10/41/16/177/21/23/article.
pdf.

Degtyarev:2000:DCS

[254] Anatoli Degtyarev, Yuri Gurevich, Paliath Narendran, Margus Veanes, and Andrei Voronkov. Decidability and complexity of simultaneous rigid E-unification with one variable and related results. Theoretical Computer Science, 243(1–2):167– 184, July 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/24/article.pdf.

Arai:2000:TCF

[255] Noriko H. Arai. Tractability of Cutfree Gentzen-type propositional calculus with permutation inference II. Theoretical Computer Science, 243(1– 2):185–197, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/25/article.pdf.

Seldin:2000:GSS

[256] Jonathan P. Seldin. A Gentzenstyle sequent calculus of constructions with expansion rules. Theoretical Computer Science, 243(1–2): 199–215, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/26/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/177/21/26/article.pdf.

Moran:2000:SEN

[257] Shlomo Moran and Sagi Snir. Simple and efficient network decomposition and synchronization. Theoretical Computer Science, 243(1–2): 217–241, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        177/21/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/177/21/27/article.pdf.

Bazzi:2000:PQ

[258] Rida A. Bazzi. Planar quorums. Theoretical Computer Science, 243(1–2): 243–268, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        177/21/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/177/21/28/article.pdf.

Lakshmanan:2000:APS

[259] K. B. Lakshmanan, Daniel J. Rosenkrantz, and S. S. Ravi. Alarm placement in systems with fault propagation. Theoretical Computer Science, 243(1–2):269–288, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/177/21/29/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/177/21/29/article.
                                        pdf.

Kirousis:2000:PCP

[260] Lefteris M. Kirousis, Evangelos Kranakis, Danny Krizanc, and Andrzej Pelc. Power consumption in packet radio networks. Theoretical Computer Science, 243(1–2):289–305, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/177/21/30/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/177/21/30/article.
pdf.

Joung:2000:TDA

[261] Yuh-Jzer Joung. Two decentralized algorithms for strong interaction fairness for systems with unbounded speed variability. Theoretical Computer Science, 243(1–2): 307–338, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/31/article.pdf.

Meier:2000:CLP

[262] David Meier and Beverly Sanders. Composing leads-to properties. Theoretical Computer Science, 243(1–2): 339–361, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/32/article.pdf.

Garay:2000:SDS

[263] Juan A. Garay, Rosario Gennaro, Charanjit Jutla, and Tal Rabin. Secure distributed storage and retrieval. Theoretical Computer Science, 243(1–2): 363–389, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/33/article.pdf.

Fokkink:2000:LPP

[264] Wan Fokkink. Language preorder as a precongruence. Theoretical Computer Science, 243(1–2):391– 408, July 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/34/article.pdf.

Aura:2000:CST

[265] Tuomas Aura and Johan Lilius. A causal semantics for time Petri nets. Theoretical Computer Science, 243(1– 2):409–447, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/35/article.pdf.

Zhou:2000:GHC

[266] Qing Zhou. Grzegorczyk’s hierarchy of computable analysis. Theoretical Computer Science, 243(1–2): 449–466, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/36/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/177/21/36/article.pdf.

Rabinovich:2000:SMC

[267] A. Rabinovich. Symbolic model checking for μ-calculus requires exponential time. Theoretical Computer Science, 243(1–2):467–475, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/177/21/37/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/177/21/37/article.
pdf.

Massacci:2000:PCA

[268] Fabio Massacci. The proof complexity of analytic and clausal tableaux. Theoretical Computer Science, 243(1– 2):477–487, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
177/21/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/177/21/38/article.pdf.

Geser:2000:NNT

[269] Alfons Geser. On normalizing, nonterminating one-rule string rewriting systems. Theoretical Computer Science, 243(1–2):489–498, July 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/177/21/39/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/177/21/39/article.
pdf.

Anonymous:2000:Ip

[270] Anonymous.  Index.  Theoret-

ical Computer Science, 243(1–2): 499, July 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        177/21/40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/177/21/40/article.pdf.

Apel:2000:CIT

[271] Joachim Apel. Computational ideal theory in finitely generated extension rings. Theoretical Computer Science, 244(1–2):1–33, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        178/20/20/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/20/article.pdf.

Park:2000:RCT

[272] Jung-Heum Park and Kyung-Yong Chwa. Recursive circulants and their embeddings among hypercubes. Theoretical Computer Science, 244(1–2): 35–62, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        178/20/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/21/article.pdf.

Courcelle:2000:MSOb

[273] Bruno Courcelle. The monadic secondorder logic of graphs XIII: Graph drawings with edge crossings. Theoretical Computer Science, 244(1–2):

63–94, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/22/article.pdf.

Ekin:2000:CLA

[274] Oya Ekin, Peter L. Hammer, and Alexander Kogan. Convexity and logical analysis of data. Theoretical Computer Science, 244(1–2):95– 116, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/23/article.pdf.

Honkala:2000:DPS

[275] Juha Honkala. On D0L power series. Theoretical Computer Science, 244(1– 2):117–134, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/24/article.pdf.

Ott:2000:SMG

[276] Matthias Ott and Frank Stephan. Structural measures for games and process control in the branch learning model. Theoretical Computer Science, 244(1–2):135–165, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/25/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/25/article.pdf.

Bodlaender:2000:HPP

[277] Hans L. Bodlaender, Michael R. Fellows, Michael T. Hallett, H. Todd Wareham, and Tandy J. Warnow. The hardness of perfect phylogeny, feasible register assignment and other problems on thin colored graphs. Theoretical Computer Science, 244(1–2): 167–188, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        178/20/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/26/article.pdf.

Jonsson:2000:BCS

[278] Peter Jonsson. Boolean constraint satisfaction: complexity results for optimization problems with arbitrary weights. Theoretical Computer Science, 244(1–2):189–203, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        178/20/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/27/article.pdf.

Hemaspaandra:2000:SST

[279] Lane A. Hemaspaandra and Jörg Rothe. A second step towards complexity-theoretic analogs of Rice’s Theorem. Theoretical Computer Science, 244(1–2):205–217, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/178/20/28/abstract.

html;   http://www.elsevier.nl/
gej-ng/10/41/16/178/20/28/article.
pdf.

Cattaneo:2000:ITC

[280] Gianpiero Cattaneo, Michele Finelli, and Luciano Margara. Investigating topological chaos by elementary cellular automata dynamics. Theoretical Computer Science, 244(1–2): 219–241, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/29/article.pdf.

Roka:2000:FSS

[281] Zsuzsanna Róka. The firing squad synchronization problem on Cayley graphs. Theoretical Computer Science, 244(1–2):243–256, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/30/article.pdf.

Hemaspaandra:2000:CEO

[282] Lane A. Hemaspaandra and Jörg Rothe. Characterizing the existence of one-way permutations. Theoretical Computer Science, 244(1–2):257– 261, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/31/article.pdf.

Kern:2000:ARF

[283] Eric Kern and Maurice Mignotte. Applications of the representation of finite fields by matrices. Theoretical Computer Science, 244(1–2):263– 265, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        178/20/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/32/article.pdf.

Honkala:2000:SSH

[284] Juha Honkala. A short solution for the HDT0L sequence equivalence problem. Theoretical Computer Science, 244(1– 2):267–270, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        178/20/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/33/article.pdf.

Rigo:2000:GAS

[285] Michel Rigo. Generalization of automatic sequences for numeration systems on a regular language. Theoretical Computer Science, 244(1–2): 271–281, August 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        178/20/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/178/20/34/article.pdf.

vanMelkebeek:2000:ZOL

[286] Dieter van Melkebeek. The zeroone law holds for BPP. Theoretical Computer Science, 244(1–2):283– 288, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/35/article.pdf.

Kelemen:2000:CLA

[287] Jozef Kelemen, Alica Kelemenová, Carlos Martín-Vide, and Victor Mitrana. Colonies with limited activation of components. Theoretical Computer Science, 244(1–2):289– 298, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/36/article.pdf.

Anonymous:2000:Iq

[288] Anonymous.  Index.  Theoreti-

cal Computer Science, 244(1–2):299, August 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
178/20/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/178/20/37/article.pdf.

Ito:2000:F

[289] Masami Ito and Chrystopher L. Nehaniv. Foreword. Theoretical Computer Science, 245(1):1, August 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
179/20/20/00003416.pdf;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/20/20/abstract.html.

Brzozowski:2000:DIT

[290] J. A. Brzozowski. Delay-insensitivity and ternary simulation. Theoretical Computer Science, 245(1):3–25, August 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        179/20/21/00003417.pdf;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/20/21/abstract.html.

Domosi:2000:CSA

[291] Pál Dömösi and Chrystopher L. Nehaniv. On complete systems of automata. Theoretical Computer Science, 245(1):27–54, August 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        179/20/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/20/22/article.pdf.

Goguen:2000:HA

[292] Joseph Goguen and Grant Malcolm. A hidden agenda. Theoretical Computer Science, 245(1):55– 101, August 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        179/20/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/20/23/article.pdf.

Imreh:2000:ECP

[293] B. Imreh. On the equivalence of the cube-product and the general product of automata. Theoretical Computer Science, 245(1):103–113, August 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
179/20/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/179/20/24/article.pdf.

Ito:2000:SSD

[294] Masami Ito, Lila Kari, and Gabriel Thierrin. Shuffle and scattered deletion closure of languages. Theoretical Computer Science, 245(1):115– 133, August 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
179/20/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/179/20/25/article.pdf.

Maciel:2000:PSD

[295] Alexis Maciel, Pierre Péladeau, and Denis Thérien. Programs over semigroups of dot-depth one. Theoretical Computer Science, 245(1):135– 148, August 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
179/20/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/179/20/26/article.pdf.

Chlebus:2000:APA

[296] Bogdan S. Chlebus, Artur Czumaj, Leszek Ga̧sieniec, Mirosław Kowaluk, and Wojciech Plandowski. Algorithms for the parallel alternating direction access machine. Theoretical Computer Science, 245(2):151– 173, August 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        179/22/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/22/21/article.pdf.

Oesterdiekhoff:2000:PCN

[297] Brigitte Oesterdiekhoff. Periodic comparator networks. Theoretical Computer Science, 245(2):175–202, August 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        179/22/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/22/22/article.pdf.

Fertin:2000:SMB

[298] Guillaume Fertin. On the structure of minimum broadcast digraphs. Theoretical Computer Science, 245(2): 203–216, August 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        179/22/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/22/23/article.pdf.

Gavoille:2000:SIR

[299] Cyril Gavoille. A survey on interval routing. Theoretical Computer Science, 245(2):217–253, August 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        179/22/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/179/22/24/article.pdf.

Ruzicka:2000:CMD

[300] Peter Ružička and DanielŠtefankovič. On the complexity of multi-dimensional interval routing schemes. Theoretical Computer Science, 245(2):255– 280, August 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
179/22/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/179/22/25/article.pdf.

Shahrokhi:2000:NLB

[301] Farhad Shahrokhi, Ondrej Sýkora, László A. Székely, and Imrich Vrt’o. A new lower bound for the bipartite crossing number with applications. Theoretical Computer Science, 245(2): 281–294, August 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
179/22/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/179/22/26/article.pdf.

Anonymous:2000:Ir

[302] Anonymous.  Index.  Theoreti-

cal Computer Science, 245(2):295, August 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
179/22/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/179/22/27/article.pdf.

Drewes:2000:TBP

[303] Frank Drewes. Tree-based picture generation. Theoretical Computer Science, 246(1–2):1–51, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/180/21/21/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/180/21/21/article.
pdf.

Nielsen:2000:FSB

[304] Frank Nielsen. Fast stabbing of boxes in high dimensions. Theoretical Computer Science, 246(1–2):53– 72, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
180/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/180/21/22/article.pdf.

Barneva:2000:TDT

[305] Reneta P. Barneva, Valentin E. Brimkov, and Philippe Nehlig. Thin discrete triangular meshes. Theoretical Computer Science, 246(1–2):73– 105, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
180/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/180/21/23/article.pdf.

Fujito:2000:AMF

[306] Toshihiro Fujito. Approximating minimum feedback vertex sets in hypergraphs. Theoretical Computer Science, 246(1–2):107–116, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/180/21/24/abstract.
html;   http://www.elsevier.nl/

gej-ng/10/41/16/180/21/24/article.
pdf.

Durand-Lose:2000:RST

[307] Jérôme O. Durand-Lose. Reversible space-time simulation of cellular automata. Theoretical Computer Science, 246(1–2):117–129, September 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
180/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/180/21/25/article.pdf.

Nishida:2000:GPE

[308] T. Y. Nishida and S. Seki. Grouped partial ET0L systems and parallel multiple context-free grammars. Theoretical Computer Science, 246(1–2):131– 150, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
180/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/180/21/26/article.pdf.

Arques:2000:SCS

[309] Didier Arquès and Jean-François Béraud. Le Schéma de carte et ses applications. (french) [The sketch of a map and its applications]. Theoretical Computer Science, 246(1–2): 151–194, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
180/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/180/21/27/article.pdf.

Sander:2000:CFL

[310] J. W. Sander and R. Tijdeman. The complexity of functions on lattices. Theoretical Computer Science, 246(1–2):195–225, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/180/21/28/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/180/21/28/article.
                                        pdf.

Lin:2000:SEH

[311] Yih-Kai Lin and Kuo-Liang Chung. A space-efficient Huffman decoding algorithm and its parallelism. Theoretical Computer Science, 246(1–2):227– 238, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        180/21/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/180/21/29/article.pdf.

Liu:2000:DSP

[312] Zhen Liu. Dynamic scheduling of parallel computations. Theoretical Computer Science, 246(1–2):239–252, September 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        180/21/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/180/21/30/article.pdf.

Garcia:2000:RLL

[313] Pedro García and José Ruiz. Right and left locally testable languages. Theoretical Computer Science, 246(1–2):

253–264, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
180/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/180/21/31/article.pdf.

Dyer:2000:PTC

[314] Martin Dyer and Catherine Greenhill. Polynomial-time counting and sampling of two-rowed contingency tables. Theoretical Computer Science, 246(1–2):265–278, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/180/21/32/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/180/21/32/article.
pdf.

Meduna:2000:GPT

[315] Alexander Meduna. Generative power of three-nonterminal scattered context grammars. Theoretical Computer Science, 246(1–2):279– 284, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
180/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/180/21/33/article.pdf.

Ding:2000:SSC

[316] Cunsheng Ding, David R. Kohel, and San Ling. Secret-sharing with a class of ternary codes. Theoretical Computer Science, 246(1–2):285– 298, September 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        180/21/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/180/21/34/article.pdf.

Anonymous:2000:Is

[317] Anonymous.  Index.  Theoreti-

cal Computer Science, 246(1–2):299, September 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        180/21/35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/180/21/35/article.pdf.

Droste:2000:ACA

[318] Manfred Droste, Paul Gastin, and Dietrich Kuske. Asynchronous cellular automata for pomsets. Theoretical Computer Science, 247(1–2): 1–38, September 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        182/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/182/21/21/article.pdf.

Ehrhard:2000:PSH

[319] Thomas Ehrhard. Parallel and serial hypercoherences. Theoretical Computer Science, 247(1–2):39–81, September 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        182/21/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/182/21/22/article.pdf.

Mackie:2000:INL

[320] Ian Mackie. Interaction nets for linear logic. Theoretical Computer Science, 247(1–2):83–140, September 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/23/article.pdf.

Oriat:2000:DEM

[321] Catherine Oriat. Detecting equivalence of modular specifications with categorical diagrams. Theoretical Computer Science, 247(1–2):141–190, September 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/24/article.pdf.

Ognjanovic:2000:SFO

[322] Zoran Ognjanovic and Miodrag Raškovic. Some first-order probability logics. Theoretical Computer Science, 247(1–2):191–212, September 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/25/article.pdf.

Dinitz:2000:TDC

[323] Yefim Dinitz, Tamar Eilam, Shlomo Moran, and Shmuel Zaks. On the total k-diameter of connection networks. Theoretical Computer Science, 247(1–2):213–228, September 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/26/article.pdf.

Blanck:2000:DRT

[324] Jens Blanck. Domain representations of topological spaces. Theoretical Computer Science, 247(1–2):229– 255, September 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/27/article.pdf.

Mili:2000:SDB

[325] Rym Mili, Jules Desharnais, Marc Frappier, and Ali Mili. Semantic distance between specifications. Theoretical Computer Science, 247(1–2):257– 276, September 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/28/article.pdf.

Broda:2000:PTC

[326] Sabine Broda and Luís Damas. On principal types of combinators. Theoretical Computer Science, 247(1–2): 277–290, September 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

182/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/29/article.pdf.

Hesselink:2000:GNF

[327] Wim H. Hesselink. A generalization of Naundorf’s fixpoint theorem. Theoretical Computer Science, 247(1–2): 291–296, September 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/30/article.pdf.

Anonymous:2000:It

[328] Anonymous.  Index.  Theoreti-

cal Computer Science, 247(1–2):299, September 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
182/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/182/21/31/article.pdf.

Consel:2000:F

[329] Charles Consel. Foreword. Theoretical Computer Science, 248(1–2):1– 2, October 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
183/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/183/21/21/article.pdf.

Hornof:2000:ABT

[330] Luke Hornof and Jacques Noyé. Accurate binding-time analysis for imperative languages: flow, context, and return sensitivity. Theoretical Computer Science, 248(1–2):3– 27, October 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
183/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/183/21/22/article.pdf.

Melski:2000:ICS

[331] David Melski and Thomas Reps. Interconvertibility of a class of set constraints and context-free-language reachability. Theoretical Computer Science, 248(1–2):29–98, October 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
183/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/183/21/23/article.pdf.

Heldal:2000:EPE

[332] Rogardt Heldal and John Hughes. Extending a partial evaluator which supports separate compilation. Theoretical Computer Science, 248(1–2): 99–145, October 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
183/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/183/21/24/article.pdf.

Grant:2000:DEA

[333] Brian Grant, Markus Mock, Matthai Philipose, Craig Chambers, and Susan J. Eggers. DyC: an expressive annotation-directed dynamic compiler for C. Theoretical Computer Science, 248(1–2):147–199, October 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
183/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/183/21/25/article.pdf.

Muller:2000:APA

[334] Gilles Muller, Renaud Marlet, and Eugen-Nicolae Volanschi. Accurate program analyses for successful specialization of legacy system software. Theoretical Computer Science, 248(1– 2):201–210, October 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
183/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/183/21/26/article.pdf.

Taha:2000:MMS

[335] Walid Taha and Tim Sheard. MetaML and multi-stage programming with explicit annotations. Theoretical Computer Science, 248(1–2):211–242, October 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
183/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/183/21/27/article.pdf.

Danvy:2000:LDT

[336] Olivier Danvy and Ulrik P. Schultz. Lambda-dropping: transforming recursive equations into programs with block structure. Theoretical Computer Science, 248(1–2):243–287, October 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        183/21/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/183/21/28/article.pdf.

Anonymous:2000:Iu

[337] Anonymous.  Index.  Theoreti-

cal Computer Science, 248(1–2):289, October 6, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        183/21/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/183/21/29/article.pdf.

Keimel:2000:F

[338] Klaus Keimel, Michael Mislove, and Constantine Tsinakis. Foreword. Theoretical Computer Science, 249(1):1, October 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        184/20/20/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/184/20/20/article.pdf.

Rutten:2000:UCT

[339] J. J. M. M. Rutten. Universal coalgebra: a theory of systems. Theoretical Computer Science, 249(1): 3–80, October 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        184/20/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/184/20/21/article.pdf.

Berline:2000:CFF

[340] Chantal Berline. From computation to foundations via functions and application: The λ-calculus and its webbed models. Theoretical Computer Science, 249(1):81–161, October 17, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/20/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/20/22/article.pdf.

Oles:2000:ALT

[341] Frank J. Oles. An application of lattice theory to knowledge representation. Theoretical Computer Science, 249(1): 163–196, October 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/20/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/20/23/article.pdf.

Salibra:2000:AML

[342] Antonino Salibra. On the algebraic models of lambda calculus. Theoretical Computer Science, 249(1):197– 240, October 17, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/20/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/20/24/article.pdf.

Iliopoulos:2000:P

[343] Costas S. Iliopoulos. Preface. Theoretical Computer Science, 249(2): 241, October 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/23/20/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/23/20/article.pdf.

Dunne:2000:CTM

[344] Paul E. Dunne, Alan Gibbons, and Michele Zito. Complexitytheoretic models of phase transitions in search problems. Theoretical Computer Science, 249(2):243– 263, October 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/23/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/23/21/article.pdf.

Fraenkel:2000:RRQ

[345] Aviezri S. Fraenkel. Recent results and questions in combinatorial game complexities. Theoretical Computer Science, 249(2):265– 288, October 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/23/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/23/22/article.pdf.

Franek:2000:RSS

[346] František Franěk, Ayşe Karaman, and W. F. Smyth. Repetitions in Sturmian strings. Theoretical Computer Science, 249(2):289–303, October 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

184/23/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/23/23/article.pdf.

Holub:2000:ASM

[347] Jan Holub and Bořivoj Melichar. Approximate string matching using factor automata. Theoretical Computer Science, 249(2):305–311, October 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/23/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/23/24/article.pdf.

Mouchard:2000:NFQ

[348] Laurent Mouchard. Normal forms of quasiperiodic strings. Theoretical Computer Science, 249(2):313– 324, October 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/23/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/23/25/article.pdf.

Ravindran:2000:DED

[349] S. Ravindran, A. M. Gibbons, and M. S. Paterson. Dense edge-disjoint embedding of complete binary trees in interconnection networks. Theoretical Computer Science, 249(2):325– 342, October 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
184/23/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/184/23/26/article.pdf.

Smyth:2000:RPC

[350] W. F. Smyth. Repetitive perhaps, but certainly not boring. Theoretical Computer Science, 249(2):343– 355, October 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        184/23/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/184/23/27/article.pdf.

Zhou:2000:BBG

[351] Sanming Zhou. Bounding the bandwidths for graphs. Theoretical Computer Science, 249(2):357–368, October 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        184/23/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/184/23/28/article.pdf.

Anonymous:2000:Iv

[352] Anonymous.  Index.  Theoretical

Computer Science, 249(2):369, October 28, 2000. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        184/23/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/184/23/29/article.pdf.

Koch:2001:EAC

[353] Ina Koch. Enumerating all connected maximal common subgraphs in two graphs. Theoretical Computer Science, 250(1–2):1–30, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/22/article.pdf.

Jedrzejowicz:2001:SLP

[354] Joanna Jȩdrzejowicz and Andrzej Szepietowski. Shuffle languages are in P. Theoretical Computer Science, 250 (1–2):31–53, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/23/article.pdf.

Okawa:2001:HCR

[355] Satoshi Okawa and Sadaki Hirose. Homomorphic characterizations of recursively enumerable languages with very small language classes. Theoretical Computer Science, 250(1–2): 55–69, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/24/article.pdf.

Prieur:2001:HDC

[356] Christophe Prieur. How to decide continuity of rational functions on infinite words. Theoretical Computer Science, 250(1–2):71– 82, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/25/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/186/20/25/article.pdf.

Caporaso:2001:PDC

[357] S. Caporaso, M. Zito, and N. Galesi. A predicative and decidable characterization of the polynomial classes of languages. Theoretical Computer Science, 250(1–2):83–99, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        186/20/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/186/20/26/article.pdf.

Merkle:2001:SPB

[358] Wolfgang Merkle. Structural properties of bounded relations with an application to NP optimization problems. Theoretical Computer Science, 250(1– 2):101–124, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        186/20/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/186/20/27/article.pdf.

Blair:2001:PAM

[359] Jean R. S. Blair, Pinar Heggernes, and Jan Arne Telle. A practical algorithm for making filled graphs minimal. Theoretical Computer Science, 250(1–2): 125–141, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        186/20/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/186/20/28/article.pdf.

Ateniese:2001:ECV

[360] Giuseppe Ateniese, Carlo Blundo, Alfredo De Santis, and Douglas R. Stinson. Extended capabilities for visual cryptography. Theoretical Computer Science, 250(1–2):143–161, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/29/article.pdf.

Mihalache:2001:LTA

[361] Valeria Mihalache and Arto Salomaa. Language-theoretic aspects of DNA complematarity. Theoretical Computer Science, 250(1–2):163– 178, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/30/article.pdf.

Bar-Ilan:2001:GSC

[362] Judit Bar-Ilan, Guy Kortsarz, and David Peleg. Generalized submodular cover problems and applications. Theoretical Computer Science, 250(1–2): 179–200, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/31/article.pdf.

Franciosa:2001:SDB

[363] Paolo Giulio Franciosa, Daniele Frigioni, and Roberto Giaccio. Semidynamic breadth-first search in digraphs. Theoretical Computer Science, 250(1–2):201–217, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/186/20/32/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/186/20/32/article.
pdf.

Fulop:2001:RGT

[364] Z. Fülöp and S. Vágvölgyi. Restricted ground tree transducers. Theoretical Computer Science, 250(1–2): 219–233, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/33/article.pdf.

Brodal:2001:CNB

[365] Gerth Støting Brodal and M. Cristina Pinotti. Comparator networks for binary heap construction. Theoretical Computer Science, 250(1–2):235– 245, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/34/article.pdf.

Kim:2001:DGO

[366] Changwook Kim. Double Greibach operator grammars. Theoretical Computer Science, 250(1–2):247–264, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/35/article.pdf.

Labhalla:2001:EMR

[367] S. Labhalla, H. Lombardi, and E. Moutai. Espaces métriques rationnellement présentés et complexité, le cas de l’espace des fonctions réelles uniformément continues sur un intervalle compact. Theoretical Computer Science, 250(1–2):265– 332, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/36/article.pdf.

Knuutila:2001:RDA

[368] Timo Knuutila. Re-describing an algorithm by Hopcroft. Theoretical Computer Science, 250(1–2):333– 363, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/37/article.pdf.

Campbell:2001:AS

[369] Colin M. Campbell, Edmund F. Robertson, Nikola Ruškuc, and Richard M. Thomas. Automatic semigroups. Theoretical Computer Science, 250(1–2):365–391, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/186/20/38/abstract.

                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/186/20/38/article.
                                        pdf.

Anonymous:2001:Ia

[370] Anonymous.  Index.  Theoretical

Computer Science, 250(1–2):393–394, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        186/20/39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/186/20/39/article.pdf.

Anonymous:2001:Ib

[371] Anonymous.  Index.  Theoretical

Computer Science, 250(1–2):395–401, January 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        186/20/40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/186/20/40/article.pdf.

Ausiello:2001:Y

[372] Giorgio Ausiello, Donald Sannella, and Michael Mislove. 25 Years. Theoretical Computer Science, 250(1–2): xv–xvi, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        186/20/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/186/20/21/article.pdf.

Nivat:2001:Y

[373] Maurice Nivat. 25 Years. Theoretical Computer Science, 250(1–2): xi–xiii, January 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
186/20/20/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/186/20/20/article.pdf.

Senizergues:2001:DRC

[374] Géraud Sénizergues. L(A) = L(B)? decidability results from complete formal systems. Theoretical Computer Science, 251(1–2):1–166, January 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../20/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
187/20/20/article.pdf.

Anonymous:2001:Ic

[375] Anonymous. Index. Theoretical Computer Science, 251(1–2):167, January 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../21/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
187/20/21/article.pdf.

vandenHerik:2001:F

[376] H. Jaap van den Herik and Hiroyuki Iida. Foreword. Theoretical Computer Science, 252(1–2):1–3, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../20/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
188/20/20/article.pdf.

Fraenkel:2001:NHG

[377] Aviezri S. Fraenkel and Dmitri Zusman. A new heap game. Theoretical Computer Science, 252(1–2):5– 12, February 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../21/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/188/20/21/article.pdf.

Fraenkel:2001:ICI

[378] Aviezri S. Fraenkel and Ofer Rahat. Infinite cyclic impartial games. Theoretical Computer Science, 252(1–2): 13–22, February 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../22/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/188/20/22/article.pdf.

Spight:2001:ETM

[379] William L. Spight. Extended thermography for multiple kos in go. Theoretical Computer Science, 252(1–2):23– 43, February 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../23/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/188/20/23/article.pdf.

Willmott:2001:AAP

[380] Steven Willmott, Julian Richardson, Alan Bundy, and John Levine. Applying adversarial planning techniques to Go. Theoretical Computer Science, 252(1–2):45–82, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../24/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        188/20/24/article.pdf.

Gao:2001:SAD

[381] Xinbo Gao, Hiroyuki Iida, Jos W. H. M. Uiterwijk, and H. Jaap van den Herik. Strategies anticipating a difference in search depth using opponentmodel search. Theoretical Computer Science, 252(1–2):83–104, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../25/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
188/20/25/article.pdf.

Beal:2001:TDL

[382] Donald F. Beal and Martin C. Smith. Temporal difference learning applied to game playing and the results of application to shogi. Theoretical Computer Science, 252(1–2):105–119, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../26/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
188/20/26/article.pdf.

Breuker:2001:SGP

[383] Dennis M. Breuker, H. Jaap van den Herik, Jos W. H. M. Uiterwijk, and L. Victor Allis. A solution to the GHI problem for best-first search. Theoretical Computer Science, 252(1–2):121– 149, February 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../27/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/188/20/27/article.pdf.

Junghanns:2001:SIS

[384] Andreas Junghanns and Jonathan Schaeffer. Sokoban: improving the search with relevance cuts. Theoretical Computer Science, 252(1–2):151– 175, February 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../28/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/188/20/28/article.pdf.

Bjornsson:2001:MCP

[385] Yngvi Björnsson and Tony A. Marsland. Multi-cut α, β-pruning in gametree search. Theoretical Computer Science, 252(1–2):177–196, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../29/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
188/20/29/article.pdf.

Pijls:2001:GTA

[386] Wim Pijls and Arie de Bruin. Game tree algorithms and solution trees. Theoretical Computer Science, 252(1–2):197–215, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../30/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
188/20/30/article.pdf.

Frank:2001:TEI

[387] Ian Frank and David Basin. A theoretical and empirical investigation of search in imperfect information games. Theoretical Computer Science, 252(1–2):217–256, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../31/abstract.html;   http://

www.elsevier.nl/gej-ng/10/41/16/
188/20/31/article.pdf.

Anonymous:2001:Id

[388] Anonymous. Index. Theoretical Computer Science, 252(1–2):257, February 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../32/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
188/20/32/article.pdf.

Bertran:2001:P

[389] Miquel Bertran and Teodor Rus. Preface. Theoretical Computer Science, 253 (1):1, February 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../20/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/189/20/20/article.pdf.

Broy:2001:RT

[390] Manfred Broy. Refinement of time. Theoretical Computer Science, 253(1): 3–26, February 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../21/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/189/20/21/article.pdf.

Bjorner:2001:DVR

[391] Nikolaj S. Bjørner, Zohar Manna, Henny B. Sipma, and Tomás E. Uribe. Deductive verification of real-time systems using STeP. Theoretical Computer Science, 253(1):27–60, February 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../22/abstract.html;   http://

www.elsevier.nl/gej-ng/10/41/16/
189/20/22/article.pdf.

Dierks:2001:PAN

[392] Henning Dierks. PLC-automata: a new class of implementable realtime automata. Theoretical Computer Science, 253(1):61–93, February 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../23/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
189/20/23/article.pdf.

Campos:2001:VLR

[393] Sérgio Vale Aguiar Campos and Edmund Clarke. The Verus language: representing time efficiently with BDDs. Theoretical Computer Science, 253(1):95–118, February 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../24/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
189/20/24/article.pdf.

Liu:2001:VRS

[394] Zhiming Liu and Mathai Joseph. Verification, refinement and scheduling of real-time programs. Theoretical Computer Science, 253(1):119–152, February 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../25/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
189/20/25/article.pdf.

Recalde:2001:STM

[395] Laura Recalde, Enrique Teruel, and Manuel Silva. Structure theory of multi-level deterministically synchronized sequential processes. Theoretical Computer Science, 254(1–2):1– 33, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../20/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/20/article.pdf.

Ingolfsdottir:2001:FAD

[396] Anna Ingólfsdóttir and Andrea Schalk. A fully abstract denotational model for observational precongruence. Theoretical Computer Science, 254(1–2): 35–61, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../21/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/21/article.pdf.

Monin:2001:OMB

[397] François Monin and Marianne Simonot. An ordinal measure based procedure for termination of functions. Theoretical Computer Science, 254(1– 2):63–94, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../22/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/22/article.pdf.

Resende:2001:QFO

[398] Pedro Resende. Quantales, finite observations and strong bisimulation. Theoretical Computer Science, 254(1–2): 95–149, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../23/abstract.html;

                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/191/20/23/article.pdf.

Crolard:2001:SL

[399] Tristan Crolard. Subtractive logic. Theoretical Computer Science, 254(1– 2):151–185, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.nl../24/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/191/20/24/article.
                                        pdf.

Ramesh:2001:ADE

[400] R. Ramesh, I. V. Ramakrishnan, and R. C. Sekar. Automata-driven efficient subterm unification. Theoretical Computer Science, 254(1–2):187– 223, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../25/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/191/20/25/article.pdf.

Springintveld:2001:TTA

[401] Jan Springintveld, Frits Vaandrager, and Pedro R. D’Argenio. Testing timed automata. Theoretical Computer Science, 254(1–2):225–257, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../26/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        191/20/26/article.pdf.

Raffalli:2001:CML

[402] Christophe Raffalli. Completeness, minimal logic and programs extraction. Theoretical Computer Science, 254(1–2):259–271, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../27/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
191/20/27/article.pdf.

Ruggieri:2001:UTL

[403] Salvatore Ruggieri. ∃-universal termination of logic programs. Theoretical Computer Science, 254(1–2):273– 296, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../28/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/191/20/28/article.pdf.

Guerraoui:2001:GAM

[404] Rachid Guerraoui and André Schiper. Genuine atomic multicast in asynchronous distributed systems. Theoretical Computer Science, 254(1–2): 297–316, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../29/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/191/20/29/article.pdf.

Guerrini:2001:PMP

[405] Stefano Guerrini and Andrea Masini. Parsing MELL proof nets. Theoretical Computer Science, 254(1–2):317– 335, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../30/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/191/20/30/article.pdf.

Chen:2001:ACT

[406] Xiao Jun Chen and Rocco De Nicola. Algebraic characterizations of trace and decorated trace equivalences over tree-like structures. Theoretical Computer Science, 254(1–2):337– 361, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../31/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/31/article.pdf.

VandenBussche:2001:SNR

[407] Jan Van den Bussche. Simulation of the nested relational algebra by the flat relational algebra, with an application to the complexity of evaluating powerset algebra expressions. Theoretical Computer Science, 254(1–2):363– 377, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../32/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/32/article.pdf.

Jenner:2001:FAS

[408] Lars Jenner and Walter Vogler. Fast asynchronous systems in dense time. Theoretical Computer Science, 254(1–2):379–422, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../33/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
191/20/33/article.pdf.

Vandeurzen:2001:ELC

[409] Luc Vandeurzen, Marc Gyssens, and Dirk Van Gucht. On the expressiveness of linear-constraint query languages for spatial databases. Theoretical Computer Science, 254(1–2):423– 463, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../34/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/34/article.pdf.

Sagonas:2001:LFO

[410] Konstantinos Sagonas, Terrance Swift, and David S. Warren. The limits of fixed-order computation. Theoretical Computer Science, 254(1–2):465– 499, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../35/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/35/article.pdf.

Katoen:2001:MST

[411] Joost-Pieter Katoen, Christel Baier, and Diego Latella. Metric semantics for true concurrent real time. Theoretical Computer Science, 254(1–2): 501–542, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../36/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/191/20/36/article.pdf.

Frias:2001:CRC

[412] Marcelo F. Frias and Roger D. Maddux. Completeness of a relational calculus for program schemes. Theoretical Computer Science, 254(1–2):543– 556, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../37/abstract.html;

                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/191/20/37/article.pdf.

Thomsen:2001:FAD

[413] B. Thomsen and S. Abramsky. A fully abstract denotational semantics for the calculus of higher-order communicating systems. Theoretical Computer Science, 254(1–2):557–589, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../38/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        191/20/38/article.pdf.

Spruit:2001:RDU

[414] Paul Spruit, Roel Wieringa, and John-Jules Meyer. Regular database update logics. Theoretical Computer Science, 254(1–2):591–661, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../39/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        191/20/39/article.pdf.

Rondogiannis:2001:SNT

[415] P. Rondogiannis. Stratified negation in temporal logic programming and the cycle-sum test. Theoretical Computer Science, 254(1–2):663–676, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../40/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        191/20/40/article.pdf.

Probst:2001:HNJ

[416] Dieter Probst and Thomas Studer. How to normalize the Jay. Theoretical Computer Science, 254(1–2):677– 681, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../41/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/191/20/41/article.pdf.

Levene:2001:GNI

[417] Mark Levene and George Loizou. Guaranteeing no interaction between functional dependencies and tree-like inclusion dependencies. Theoretical Computer Science, 254(1–2):683– 690, March 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../42/abstract.html;
http://www.elsevier.nl/gej-ng/
10/41/16/191/20/42/article.pdf.

Bernardo:2001:CTE

[418] Marco Bernardo and Roberto Gorrieri. Corrigendum to “A tutorial on EMPA: a theory of concurrent processes with nondeterminism, priorities, probabilities and time” — [Theoret. Comput. Sci. 202 (1998) 1–54]. Theoretical Computer Science, 254(1–2):691–694, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl../
43/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/191/
20/43/article.pdf;   http://www.
elsevier.nl/PII/S0304397597001278.

See [4].

Anonymous:2001:AIVa

[419] Anonymous. Author index volume 254 (2001). Theoretical Computer Science, 254(1–2):699–700, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../44/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
191/20/44/article.pdf.

Stirling:2001:DDE

[420] Colin Stirling. Decidability of DPDA equivalence. Theoretical Computer Science, 255(1–2):1–31, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../21/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/21/article.pdf.

Erlebach:2001:CPC

[421] Thomas Erlebach and Klaus Jansen. The complexity of path coloring and call scheduling. Theoretical Computer Science, 255(1–2):33–50, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../22/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/22/article.pdf.

Arikati:2001:AAM

[422] Srinivasa R. Arikati, Anders Dessmark, Andrzej Lingas, and Madhav V. Marathe. Approximation algorithms for maximum two-dimensional pattern matching. Theoretical Computer Science, 255(1–2):51–62, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../23/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/23/article.pdf.

Duntsch:2001:RAA

[423] Ivo Düntsch, Hui Wang, and Steve Mc-Closkey. A relation-algebraic approach to the region connection calculus. Theoretical Computer Science, 255(1–2): 63–83, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../24/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/24/article.pdf.

Yu:2001:PPL

[424] S. S. Yu. Post-plus languages. Theoretical Computer Science, 255(1–2): 85–105, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../25/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/25/article.pdf.

Jain:2001:BBN

[425] Sanjay Jain. Branch and bound on the network model. Theoretical Computer Science, 255(1–2):107–123, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../26/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/26/article.pdf.

Calude:2001:RER

[426] Cristian S. Calude, Peter H. Hertling, Bakhadyr Khoussainov, and Yongge Wang. Recursively enumerable reals and Chaitin Ω numbers. Theoretical Computer Science, 255(1–2):125– 149, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../27/abstract.html;

                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/197/21/27/article.pdf.

Neraud:2001:CFD

[427] Jean Néraud and Carla Selmi. On codes with a finite deciphering delay: constructing uncompletable words. Theoretical Computer Science, 255(1–2):151–162, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../28/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/28/article.pdf.

Watjen:2001:PCL

[428] Dietmar Wätjen. Parallel communicating limited and uniformly limited 0L systems. Theoretical Computer Science, 255(1–2):163–191, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../29/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/29/article.pdf.

Halava:2001:MPD

[429] Vesa Halava, Mika Hirvensalo, and Ronald de Wolf. Marked PCP is decidable. Theoretical Computer Science, 255(1–2):193–204, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../30/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/30/article.pdf.

Arvind:2001:PRB

[430] V. Arvind and Johannes Köbler. On pseudorandomness and resourcebounded measure. Theoretical Computer Science, 255(1–2):205–221, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../31/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/31/article.pdf.

Finkel:2001:LFL

[431] Olivier Finkel. Locally finite languages. Theoretical Computer Science, 255(1–2):223–261, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../32/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/32/article.pdf.

Calamoneri:2001:OTD

[432] Tiziana Calamoneri and Annalisa Massini. Optimal three-dimensional layout of interconnection networks. Theoretical Computer Science, 255 (1–2):263–279, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl../
33/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/197/
21/33/article.pdf.

Feder:2001:FLC

[433] Tomás Feder. Fanout limitations on constraint systems. Theoretical Computer Science, 255(1–2):281– 293, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../34/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/34/article.pdf.

Apolloni:2001:GDF

[434] B. Apolloni and D. Malchiodi. Gaining degrees of freedom in subsymbolic learning. Theoretical Computer Science, 255(1–2):295–321, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../35/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/35/article.pdf.

Lopez:2001:SIE

[435] Luis-Miguel Lopez and Philippe Narbel. Substitutions and interval exchange transformations of rotation class. Theoretical Computer Science, 255(1–2):323–344, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../36/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/36/article.pdf.

Penaud:2001:CSH

[436] Jean-Guy Penaud, Elisa Pergola, Renzo Pinzani, and Olivier Roques. Chemins de Schröder et hiérarchies aléatoires. (French). [schröder paths and aleatory hierarchies]. Theoretical Computer Science, 255(1–2):345– 361, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../37/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/197/21/37/article.pdf.

Justin:2001:FPS

[437] Jacques Justin and Giuseppe Pirillo. Fractional powers in Sturmian words. Theoretical Computer Science, 255(1–2):363–376, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../38/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/38/article.pdf.

Rahonis:2001:AST

[438] George Rahonis. Alphabetic and synchronized tree transducers. Theoretical Computer Science, 255(1–2):377– 399, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../39/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/39/article.pdf.

Koskinen:2001:NIK

[439] Jukka A. Koskinen. Non-injective knapsack public-key cryptosystems. Theoretical Computer Science, 255 (1–2):401–422, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl../
40/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/197/
21/40/article.pdf.

Fischer:2001:GST

[440] Paul F. Fischer, Franco P. Preparata, and John E. Savage. Generalized scans and tridiagonal systems. Theoretical Computer Science, 255(1–2):423– 436, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../41/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/41/article.pdf.

Picouleau:2001:RDT

[441] C. Picouleau. Reconstruction of domino tiling from its two orthogonal projections. Theoretical Computer Science, 255(1–2):437–447, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../42/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/42/article.pdf.

Jin:2001:TFP

[442] Guohua Jin, Zhiyuan Li, and Fujie Chen. A theoretical foundation for program transformations to reduce cache thrashing due to true data sharing. Theoretical Computer Science, 255(1–2):449–481, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../43/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/43/article.pdf.

Mantaci:2001:CET

[443] Sabrina Mantaci and Antonio Restivo. Codes and equations on trees. Theoretical Computer Science, 255(1–2):483– 509, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../44/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/197/21/44/article.pdf.

Csuhaj-Varju:2001:CFP

[444] Erzsébet Csuhaj-Varjú and György Vaszil. On context-free parallel communicating grammar systems: synchronization, communication, and normal forms. Theoretical Computer Science, 255(1–2):511–538, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../45/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/45/article.pdf.

Droubay:2001:EWS

[445] Xavier Droubay, Jacques Justin, and Giuseppe Pirillo. Episturmian words and some constructions of de Luca and Rauzy. Theoretical Computer Science, 255(1–2):539–553, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../46/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/46/article.pdf.

Xue:2001:COP

[446] Guoliang Xue. A cost optimal parallel algorithm for computing force field in N-body simulations on a CREW PRAM. Theoretical Computer Science, 255(1–2):555–568, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../47/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/47/article.pdf.

Almeida:2001:SHR

[447] Jorge Almeida and Pedro V. Silva. SC-hyperdecidability of R. Theoretical Computer Science, 255(1–2):569– 591, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../48/abstract.html;

                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/197/21/48/article.pdf.

Cucker:2001:WWC

[448] Felipe Cucker. On weak and weighted computations over the real closure of Q. Theoretical Computer Science, 255(1–2):593–600, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../49/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/49/article.pdf.

Incitti:2001:GFC

[449] Roberto Incitti. The growth function of context-free languages. Theoretical Computer Science, 255(1–2):601– 605, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../50/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/197/21/50/article.pdf.

Fournier:2001:SNC

[450] Hervé Fournier. Sparse NP-complete problems over the reals with addition. Theoretical Computer Science, 255(1–2):607–610, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../51/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/51/article.pdf.

Sokratova:2001:MLR

[451] Olga Sokratova. The Mal’cev Lemma and rewriting on semirings. Theoretical Computer Science, 255(1–2):611– 614, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../52/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/52/article.pdf.

Nikolskaia:2001:SOR

[452] M. Nikolskaia and L. Nikolskaia. Size of OBDD representation of 2-level redundancies functions. Theoretical Computer Science, 255(1–2):615– 625, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../53/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/53/article.pdf.

Nandy:2001:DMW

[453] Subhas C. Nandy, Tomohiro Harayama, and Tetsuo Asano. Dynamically maintaining the widest k-dense corridor. Theoretical Computer Science, 255(1–2):627–639, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../54/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/54/article.pdf.

Kumar:2001:USL

[454] S. Ravi Kumar and D. Sivakumar. On the unique shortest lattice vector problem. Theoretical Computer Science, 255(1–2):641–648, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../55/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/55/article.pdf.

Chen:2001:IFT

[455] Yu-Wei Chen and Kuo-Liang Chung. Improved fault-tolerant sorting algorithm in hypercubes. Theoretical Computer Science, 255(1–2):649– 658, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../56/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/197/21/56/article.pdf.

Chen:2001:DCP

[456] Chiuyuan Chen and Kaiping Wu. Disproving a conjecture on planar visibility graphs. Theoretical Computer Science, 255(1–2):659–665, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../57/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/57/article.pdf.

Honkala:2001:PSC

[457] Juha Honkala. On Parikh slender context-free languages. Theoretical Computer Science, 255(1–2):667– 677, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.nl../58/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/41/16/197/21/58/article.pdf.

Barcucci:2001:SLR

[458] E. Barcucci and S. Rinaldi. Some linear recurrences and their combinatorial interpretation by means of regular languages. Theoretical Computer Science, 255(1–2):679–686, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../59/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
197/21/59/article.pdf.

Blondel:2001:DSM

[459] Vincent D. Blondel, Olivier Bournez, Pascal Koiran, Christos H. Papadimitriou, and John N. Tsitsiklis. Deciding stability and mortality of piecewise affine dynamical systems. Theoretical Computer Science, 255(1–2):687– 696, March 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../60/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/197/21/60/article.pdf.

Trahtman:2001:EOE

[460] A. N. Trahtman. Erratum to “Optimal estimation on the order of local testability of finite automata” — [Theoret. Comput. Sci. 231 (2000) 59–74]. Theoretical Computer Science, 255(1–2):697, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl../
61/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/197/
21/61/article.pdf;   http://www.
elsevier.nl/PII/S0304397599000171.

See [27].

Anonymous:2001:AIVb

[461] Anonymous. Author index volume 255 (2001). Theoretical Computer Science, 255(1–2):699–700, March 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        nl../62/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        197/21/62/article.pdf.

Bouajjani:2001:P

[462] Ahmed Bouajjani. Preface. Theoretical Computer Science, 256(1–2):1–2, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../21/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        198/21/21/article.pdf.

Hirshfeld:2001:PAM

[463] Yoram Hirshfeld and Faron Moller. Pushdown automata, multiset automata, and Petri nets. Theoretical Computer Science, 256(1–2):3–21, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../22/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        198/21/22/article.pdf.

Jancar:2001:NRC

[464] Petr Jančar. Nonprimitive recursive complexity and undecidability for Petri net equivalences. Theoretical Computer Science, 256(1–2):23–30, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../23/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        198/21/23/article.pdf.

Mayr:2001:DMC

[465] Richard Mayr. Decidability of model checking with the temporal logic EF. Theoretical Computer Science, 256(1– 2):31–62, April 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.nl../24/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/41/16/198/21/24/article.pdf.

Finkel:2001:WST

[466] A. Finkel and Ph. Schnoebelen. Wellstructured transition systems everywhere! Theoretical Computer Science, 256(1–2):63–92, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl../
25/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/198/
21/25/article.pdf.

Kesten:2001:SMC

[467] Y. Kesten, O. Maler, M. Marcus, A. Pnueli, and E. Shahar. Symbolic model checking with rich assertional languages. Theoretical Computer Science, 256(1–2):93–112, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../26/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
198/21/26/article.pdf.

Lesens:2001:AVP

[468] David Lesens, Nicolas Halbwachs, and Pascal Raymond. Automatic verification of parameterized networks of processes. Theoretical Computer Science, 256(1–2):113–144, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
nl../27/abstract.html;   http://
www.elsevier.nl/gej-ng/10/41/16/
198/21/27/article.pdf.

Abdulla:2001:ECS

[469] Parosh Aziz Abdulla and Bengt Jonsson. Ensuring completeness of symbolic verification methods for infinitestate systems. Theoretical Computer Science, 256(1–2):145–167, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../28/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        198/21/28/article.pdf.

Anonymous:2001:AIVc

[470] Anonymous. Author index volume 256 (2001). Theoretical Computer Science, 256(1–2):169, April 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        nl../29/abstract.html;   http://
                                        www.elsevier.nl/gej-ng/10/41/16/
                                        198/21/29/article.pdf.

Ressayre:2001:WA

[471] J.-P. Ressayre. Weak arithmetics. Theoretical Computer Science, 257(1– 2):1–15, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        199/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/199/21/21/article.pdf.

Richard:2001:WWA

[472] Denis Richard. What are weak arithmetics? Theoretical Computer Science, 257(1–2):17–29, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-

ng/10/41/16/199/21/22/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/199/21/22/article.
pdf.

Anonymous:2001:C

[473] Anonymous. Conference. Theoretical Computer Science, 257(1–2):31– 49, April 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
199/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/199/21/23/article.pdf.

Cegielski:2001:DTN

[474] Patrick Cegielski and Denis Richard. Decidability of the theory of the natural integers with the Cantor pairing function and the successor. Theoretical Computer Science, 257(1–2): 51–77, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
199/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/199/21/24/article.pdf.

Dimitracopoulos:2001:EEM

[475] C. Dimitracopoulos. On end extensions of models of subsystems of Peano arithmetic. Theoretical Computer Science, 257(1–2):79– 84, April 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
199/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/199/21/25/article.pdf.

Duparc:2001:CSF

[476] J. Duparc, O. Finkel, and J.-P. Ressayre. Computer science and the fine structure of Borel sets. Theoretical Computer Science, 257(1–2): 85–105, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        199/21/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/199/21/26/article.pdf.

Esbelin:2001:CMF

[477] Henri-Alex Esbelin. Counting modulo finite semigroups. Theoretical Computer Science, 257(1–2):107– 114, April 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        199/21/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/199/21/27/article.pdf.

Korec:2001:LAS

[478] Ivan Korec. A list of arithmetical structures complete with respect to the first-order definability. Theoretical Computer Science, 257(1–2): 115–151, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        199/21/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/199/21/28/article.pdf.

Margenstern:2001:QUU

[479] Maurice Margenstern. On quasiunilateral universal Turing machines. Theoretical Computer Science, 257(1– 2):153–166, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
199/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/199/21/29/article.pdf.

Matiyasevich:2001:SAR

[480] Yuri Matiyasevich. Some arithmetical restatements of the Four Color Conjecture. Theoretical Computer Science, 257(1–2):167–183, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/199/21/30/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/199/21/30/article.
pdf.

Sudac:2001:PNT

[481] Olivier Sudac. The prime number theorem is PRA-provable. Theoretical Computer Science, 257(1–2): 185–239, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
199/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/199/21/31/article.pdf.

Teytaud:2001:DHP

[482] Olivier Teytaud. Decidability of the halting problem for Matiyasevich deterministic machines. Theoretical Computer Science, 257(1–2): 241–251, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
199/21/32/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/199/21/32/article.pdf.

Anonymous:2001:AIVd

[483] Anonymous. Author index volume 257 (2001). Theoretical Computer Science, 257(1–2):253, April 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        199/21/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/199/21/33/article.pdf.

vanBreugel:2001:IMS

[484] Franck van Breugel. An introduction to metric semantics: operational and denotational models for programming and specification languages. Theoretical Computer Science, 258(1–2): 1–98, May 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        200/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/200/21/21/article.pdf.

Broy:2001:ASP

[485] Manfred Broy and GheorgheŞtefănescu. The algebra of stream processing functions. Theoretical Computer Science, 258(1–2):99–129, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/200/21/22/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/200/21/22/article.
                                        pdf.

Engelfriet:2001:SIP

[486] Joost Engelfriet and Tjalling Gelsema. Structural inclusion in the pi-calculus with replication. Theoretical Computer Science, 258(1–2):131–168, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/200/21/23/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/200/21/23/article.
pdf.

Mano:2001:UNF

[487] Ken Mano and Mizuhito Ogawa. Unique normal form property of compatible term rewriting systems: a new proof of Chew’s theorem. Theoretical Computer Science, 258(1–2): 169–208, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/200/21/24/article.pdf.

Tzouvaras:2001:OTL

[488] Athanassios Tzouvaras. Objects and their lambda calculus. Theoretical Computer Science, 258(1–2):209– 232, May 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/200/21/25/article.pdf.

Hains:2001:CDS

[489] Gaétan Hains, Frédéric Loulergue, and John Mullins. Concrete data structures and functional parallel programming. Theoretical Computer Science, 258(1– 2):233–267, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/200/21/26/article.pdf.

Delzanno:2001:PCL

[490] Giorgio Delzanno and Maurizio Martelli. Proofs as computations in linear logic. Theoretical Computer Science, 258(1– 2):269–297, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/200/21/27/article.pdf.

Birkedal:2001:CBR

[491] Lars Birkedal and Mads Tofte. A constraint-based region inference algorithm. Theoretical Computer Science, 258(1–2):299–392, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/200/21/28/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/200/21/28/article.
pdf.

Corradini:2001:TUT

[492] Flavio Corradini and Dino Di Cola. On testing urgency through laziness over processes with durational actions. Theoretical Computer Science, 258(1– 2):393–407, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/200/21/29/article.pdf.

Jancar:2001:DBL

[493] Petr Jančar, Antonín Kučera, and Richard Mayr. Deciding bisimulationlike equivalences with finite-state processes. Theoretical Computer Science, 258(1–2):409–433, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/200/21/30/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/200/21/30/article.
pdf.

Gramlich:2001:ISC

[494] Bernhard Gramlich. On interreduction of semi-complete term rewriting systems. Theoretical Computer Science, 258(1–2):435–451, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/200/21/31/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/200/21/31/article.
pdf.

Seynhaeve:2001:GSU

[495] Franck Seynhaeve, Sophie Tison, Marc Tommasi, and Ralf Treinen. Grid structures and undecidable constraint theories. Theoretical Computer Science, 258(1–2):453– 490, May 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/32/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/200/21/32/article.pdf.

Emerson:2001:MCC

[496] E. Allen Emerson, Charanjit S. Jutla, and A. Prasad Sistla. On model checking for the μ-calculus and its fragments. Theoretical Computer Science, 258(1– 2):491–522, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        200/21/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/200/21/33/article.pdf.

Brass:2001:ESD

[497] Stefan Brass, Jürgen Dix, Ilkka Niemelä, and Teodor C. Przymusinski. On the equivalence of the static and disjunctive well-founded semantics and its computation. Theoretical Computer Science, 258(1–2):523– 553, May 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        200/21/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/200/21/34/article.pdf.

Shekhar:2001:OPD

[498] Shashi Shekhar and Weili Wu. Optimal placement of data replicas in distributed database with majority voting protocol. Theoretical Computer Science, 258(1–2):555– 571, May 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        200/21/35/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/200/21/35/article.pdf.

Darondeau:2001:PNR

[499] Philippe Darondeau. On the Petri net realization of context-free graphs. Theoretical Computer Science, 258(1– 2):573–598, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/200/21/36/article.pdf.

Anonymous:2001:AIVe

[500] Anonymous. Author index volume 258 (2001). Theoretical Computer Science, 258(1–2):599, May 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
200/21/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/200/21/37/article.pdf.

Schmitt:2001:TGA

[501] Lothar M. Schmitt. Theory of genetic algorithms. Theoretical Computer Science, 259(1–2):1–61, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/21/article.pdf.

Bonizzoni:2001:CMS

[502] Paola Bonizzoni and Gianluca Della Vedova. The complexity of multiple sequence alignment with SPscore that is a metric. Theoretical Computer Science, 259(1–2):63– 79, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/22/article.pdf.

Chrobak:2001:RPS

[503] Marek Chrobak and Christoph Dürr. Reconstructing polyatomic structures from discrete X-rays: NP-completeness proof for three atoms. Theoretical Computer Science, 259(1–2):81– 98, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/23/article.pdf.

Margenstern:2001:NPT

[504] Maurice Margenstern and Kenichi Morita. NP problems are tractable in the space of cellular automata in the hyperbolic plane. Theoretical Computer Science, 259(1–2):99– 128, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/24/article.pdf.

Kobayashi:2001:TOS

[505] Kojiro Kobayashi. On time optimal solutions of the firing squad synchronization problem for twodimensional paths. Theoretical Computer Science, 259(1–2):129–143, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/25/article.pdf.

Carpi:2001:WSF

[506] Arturo Carpi and Aldo de Luca. Words and special factors. Theoretical Computer Science, 259(1–2): 145–182, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/26/article.pdf.

Klimann:2001:NTA

[507] Ines Klimann. New types of automata to solve fixed point problems. Theoretical Computer Science, 259(1–2): 183–197, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/27/article.pdf.

Trow:2001:DPS

[508] Paul Trow. Determining presentations of sofic shifts. Theoretical Computer Science, 259(1–2):199– 216, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/28/article.pdf.

Serna:2001:PAS

[509] Maria Serna and Fatos Xhafa. On the parallel approximability of a subclass of quadratic programming. Theoretical Computer Science, 259(1–2): 217–231, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/29/article.pdf.

Hoang:2001:FHH

[510] Chính T. Hoàng and R. Sritharan. Finding houses and holes in graphs. Theoretical Computer Science, 259(1– 2):233–244, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/30/article.pdf.

Col:2001:AQA

[511] Marie-Andrée Jacob-Da Col. Applications quasi-affines et pavages du plan discret. Theoretical Computer Science, 259(1–2):245–269, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/202/21/31/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/202/21/31/article.
                                        pdf.

Dubacq:2001:KCC

[512] J.-C. Dubacq, B. Durand, and E. Formenti. Kolmogorov complexity and cellular automata classification. Theoretical Computer Science, 259(1–2):

271–285, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/32/article.pdf.

Frigioni:2001:DAC

[513] Daniele Frigioni, Alberto Marchetti-Spaccamela, and Umberto Nanni. Dynamic algorithms for classes of constraint satisfaction problems. Theoretical Computer Science, 259(1–2): 287–305, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/33/article.pdf.

Goerdt:2001:GCT

[514] Andreas Goerdt. The giant component threshold for random regular graphs with edge faults. Theoretical Computer Science, 259(1–2): 307–321, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/34/article.pdf.

Hong:2001:ASN

[515] Dawei Hong and Jean-Camille Birget. Approximation of some NPhard optimization problems by finite machines, in probability. Theoretical Computer Science, 259(1–2): 323–339, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/35/article.pdf.

Bonizzoni:2001:NPS

[516] Paola Bonizzoni and Ross M. Mc-Connell. Nesting of prime substructures in k-ary relations. Theoretical Computer Science, 259(1–2):341– 357, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/36/article.pdf.

Marco:2001:CMW

[517] Gianluca De Marco, Luisa Gargano, and Ugo Vaccaro. Concurrent multicast in weighted networks. Theoretical Computer Science, 259(1–2): 359–377, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/37/article.pdf.

Barbe:2001:DIS

[518] A. Barbé and G. Skordev. Decimationinvariant sequences and their automaticity. Theoretical Computer Science, 259(1–2):379–403, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/202/21/38/abstract.
                                        html;   http://www.elsevier.nl/

gej-ng/10/41/16/202/21/38/article.
pdf.

Fernau:2001:HMC

[519] H. Fernau, M. Holzer, and R. Freund. Hybrid modes in cooperating distributed grammar systems: internal versus external hybridization. Theoretical Computer Science, 259(1–2): 405–426, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/39/article.pdf.

Kaufmann:2001:RLI

[520] Susanne Kaufmann and Frank Stephan. Robust learning with infinite additional information. Theoretical Computer Science, 259(1–2):427–454, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/40/article.pdf.

Case:2001:CGP

[521] John Case, Keh-Jiann Chen, and Sanjay Jain. Costs of general purpose learning. Theoretical Computer Science, 259(1–2):455–473, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/202/21/41/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/202/21/41/article.
pdf.

Xu:2001:ASS

[522] Jianliang Xu, Tsunehiro Yoshinaga, Katsushi Inoue, Yue Wang, and Akira Ito. Alternation for sublogarithmic space-bounded alternating pushdown automata. Theoretical Computer Science, 259(1–2):475– 492, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/42/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/42/article.pdf.

Bonis:2001:EAC

[523] Annalisa De Bonis, Luisa Gargano, and Ugo Vaccaro. Efficient algorithms for chemical threshold testing problems. Theoretical Computer Science, 259(1– 2):493–511, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/43/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/43/article.pdf.

Lin:2001:SGR

[524] Guo-Hui Lin and Guoliang Xue. Signed genome rearrangement by reversals and transpositions: models and approximations. Theoretical Computer Science, 259(1–2):513– 531, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        202/21/44/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/202/21/44/article.pdf.

Bloom:2001:LWT

[525] Stephen L. Bloom and Christian Choffrut. Long words: the theory of concatenation and omega-power. Theoretical Computer Science, 259(1–2): 533–548, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/45/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/45/article.pdf.

Guijarro:2001:MTD

[526] David Guijarro, Víctor Lavín, and Vijay Raghavan. Monotone term decision lists. Theoretical Computer Science, 259(1–2):549–575, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/202/21/46/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/202/21/46/article.
pdf.

Aldaz:2001:NMO

[527] M. Aldaz, G. Matera, J. L. Montaña, and L. M. Pardo. A new method to obtain lower bounds for polynomial evaluation. Theoretical Computer Science, 259(1–2):577–596, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/47/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/47/article.pdf.

Arvind:2001:NNC

[528] V. Arvind and Jacobo Torán. A nonadaptive NC checker for permutation group intersection. Theoretical Computer Science, 259(1–2):597– 611, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/48/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/48/article.pdf.

Grigoriev:2001:LLB

[529] Dima Grigoriev. Linear lower bound on degrees of Positivstellensatz calculus proofs for the parity. Theoretical Computer Science, 259(1–2): 613–622, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/49/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/49/article.pdf.

Kirrinnis:2001:FAS

[530] Peter Kirrinnis. Fast algorithms for the Sylvester equation AX − XB^{T} = C. Theoretical Computer Science, 259(1– 2):623–638, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/50/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/50/article.pdf.

Heuer:2001:TPP

[531] Gerald A. Heuer. Three-part partition games on rectangles. Theoretical Computer Science, 259(1–2): 639–661, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

202/21/51/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/51/article.pdf.

Sgall:2001:SDG

[532] Jiří Sgall. Solution of David Gale’s lion and man problem. Theoretical Computer Science, 259(1–2):663– 670, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/52/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/52/article.pdf.

Stiglic:2001:CDC

[533] Anton Stiglic. Computations with a deck of cards. Theoretical Computer Science, 259(1–2):671–678, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/53/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/53/article.pdf.

Fuster-Sabater:2001:EAG

[534] A. Fúster-Sabater and L. J. García-Villalba. An efficient algorithm to generate binary sequences for cryptographic purposes. Theoretical Computer Science, 259(1–2):679– 688, May 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
202/21/54/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/202/21/54/article.pdf.

Honkala:2001:WCD

[535] Juha Honkala and Arto Salomaa. Watson–Crick D0L systems with regular triggers. Theoretical Computer Science, 259(1–2):689–698, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/202/21/55/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/202/21/55/article.
                                        pdf.

Anonymous:2001:AIVf

[536] Anonymous. Author index volume 259 (2001). Theoretical Computer Science, 259(1–2):699–700, May 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/202/21/56/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/202/21/56/article.
                                        pdf.

Jacobs:2001:F

[537] Bart Jacobs, Larry Moss, Horst Reichel, and Jan Rutten. Foreword. Theoretical Computer Science, 260(1– 2):1, June 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        201/21/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/201/21/21/article.pdf.

Cirstea:2001:SCS

[538] Corina Cîrstea. Semantic constructions for the specification of objects. Theoretical Computer Science, 260 (1–2):3–25, June 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/201/21/22/article.pdf.

Corradini:2001:CPS

[539] Andrea Corradini, Martin Große-Rhode, and Reiko Heckel. A coalgebraic presentation of structured transition systems. Theoretical Computer Science, 260(1–2):27– 55, June 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/201/21/23/article.pdf.

Gumm:2001:EIC

[540] H. Peter Gumm. Equational and implicational classes of coalgebras. Theoretical Computer Science, 260(1–2): 57–69, June 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/201/21/24/article.pdf.

Gumm:2001:CCC

[541] H. Peter Gumm and Tobias Schröder. Covarieties and complete covarieties. Theoretical Computer Science, 260(1– 2):71–86, June 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/25/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/201/21/25/article.pdf.

Johnstone:2001:SCC

[542] Peter Johnstone, John Power, Toru Tsujishita, Hiroshi Watanabe, and James Worrell. On the structure of categories of coalgebras. Theoretical Computer Science, 260(1–2):87– 117, June 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        201/21/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/201/21/26/article.pdf.

Kurz:2001:SCM

[543] Alexander Kurz. Specifying coalgebras with modal logic. Theoretical Computer Science, 260(1–2):119– 138, June 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        201/21/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/201/21/27/article.pdf.

Moss:2001:PC

[544] Lawrence S. Moss. Parametric corecursion. Theoretical Computer Science, 260(1–2):139–163, June 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/201/21/28/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/201/21/28/article.
                                        pdf.

Pardo:2001:FRP

[545] Alberto Pardo. Fusion of recursive programs with computational effects. Theoretical Computer Science, 260(1– 2):165–207, June 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/201/21/29/article.pdf.

Ro:2001:MLT

[546] Martin Rö. From modal logic to terminal coalgebras. Theoretical Computer Science, 260(1–2):209– 228, June 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/201/21/30/article.pdf.

Rosu:2001:EAC

[547] Grigore Roşu. Equational axiomatizability for coalgebra. Theoretical Computer Science, 260(1–2):229– 247, June 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/201/21/31/article.pdf.

Anonymous:2001:AIVg

[548] Anonymous. Author index volume 260 (2001). Theoretical Computer Science, 260(1–2):249, June 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
201/21/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/201/21/32/article.pdf.

Anonymous:2001:MIV

[549] Anonymous. Master index volumes 251-260. Theoretical Computer Science, 260(1–2):251–259, June 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/201/21/33/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/201/21/33/article.
pdf.

Li:2001:F

[550] Ming Li.  Foreword.  Theo-

retical Computer Science, 261(1): 1, June 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/21/21/article.pdf.

Jain:2001:LRE

[551] Sanjay Jain, Steffen Lange, and Jochen Nessel. On the learnability of recursively enumerable languages from good examples. Theoretical Computer Science, 261(1):3– 29, June 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/21/22/article.pdf.

Case:2001:SNT

[552] John Case, Sanjay Jain, and Arun Sharma. Synthesizing noise-tolerant language learners. Theoretical Computer Science, 261(1):31–56, June 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/203/21/23/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/203/21/23/article.
pdf.

Vovk:2001:PTB

[553] V. Vovk. Probability theory for the Brier game. Theoretical Computer Science, 261(1):57–79, June 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/203/21/24/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/203/21/24/article.
pdf.

Gurvits:2001:NSS

[554] Leonid Gurvits. A note on a scalesensitive dimension of linear bounded functionals in Banach spaces. Theoretical Computer Science, 261(1):81– 90, June 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/21/25/article.pdf.

Ambainis:2001:HPT

[555] Andris Ambainis, Kalvis Apsītis, Rūsiņš Freivalds, and Carl H. Smith. Hierarchies of probabilistic and team FIN-learning. Theoretical Computer Science, 261(1):91–117, June 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/203/21/26/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/203/21/26/article.
pdf.

Erlebach:2001:LOV

[556] Thomas Erlebach, Peter Rossmanith, Hans Stadtherr, Angelika Steger, and Thomas Zeugmann. Learning one-variable pattern languages very efficiently on average, in parallel, and by asking queries. Theoretical Computer Science, 261(1):119– 156, June 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/21/27/article.pdf.

Maass:2001:RTN

[557] Wolfgang Maass. On the relevance of time in neural computation and learning. Theoretical Computer Science, 261 (1):157–178, June 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/21/28/article.pdf.

Takimoto:2001:PNW

[558] Eiji Takimoto, Akira Maruoka, and Volodya Vovk. Predicting nearly as well as the best pruning of a decision tree through dynamic programming scheme. Theoretical Computer Science, 261(1):179– 209, June 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/21/29/article.pdf.

Hsu:2001:F

[559] Wen-Lian Hsu and Ming-Yang Kao. Foreword. Theoretical Computer Science, 261(2):211, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/203/25/21/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/203/25/21/article.
pdf.

Chung:2001:DLP

[560] Fan Chung and Ronald Graham. Dynamic location problems with limited look-ahead. Theoretical Computer Science, 261(2):213–226, June 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/25/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/25/22/article.pdf.

Iwama:2001:ERR

[561] Kazuo Iwama, Eiji Miyano, Satoshi Tajima, and Hisao Tamaki. Efficient randomized routing algorithms on the two-dimensional mesh of buses. Theoretical Computer Science, 261(2): 227–239, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        203/25/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/203/25/23/article.pdf.

Kim:2001:MMS

[562] Dongsoo S. Kim and Ding-Zhu Du. Multirate multicast switching networks. Theoretical Computer Science, 261(2):241–251, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/203/25/24/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/203/25/24/article.
                                        pdf.

Kiwi:2001:MMB

[563] Marcos Kiwi, Daniel A. Spielman, and Shang-Hua Teng. Min-max-boundary domain decomposition. Theoretical Computer Science, 261(2): 253–266, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        203/25/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/203/25/25/article.pdf.

Lopez-Ortiz:2001:USS

[564] Alejandro López-Ortiz and Sven Schuierer. The ultimate strategy to search on m rays? Theoretical Computer Science, 261(2):267– 295, June 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        203/25/26/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/203/25/26/article.pdf.

Lu:2001:ECS

[565] Chi-Jen Lu. An exact characterization of symmetric functions in qAC^{0}[2]. Theoretical Computer Science, 261(2): 297–303, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/25/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/25/27/article.pdf.

Reith:2001:BLB

[566] Steffen Reith and Klaus W. Wagner. On boolean lowness and boolean highness. Theoretical Computer Science, 261(2):305–321, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/203/25/28/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/203/25/28/article.
pdf.

Tsai:2001:LBW

[567] Kuo-Hui Tsai, Da-Wei Wang, and Frank Hwang. Lower bounds for wide-sense nonblocking Clos network. Theoretical Computer Science, 261(2): 323–328, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
203/25/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/203/25/29/article.pdf.

Anonymous:2001:AIVh

[568] Anonymous. Author index volume 261 (2001). Theoretical Computer Science, 261(2):333–334, June 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/203/25/30/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/203/25/30/article.
                                        pdf.

Schnitger:2001:PVI

[569] Juraj Hromkovičand Georg Schnitger. On the power of Las Vegas II: Two-way finite automata. Theoretical Computer Science, 262(1–2): 1–24, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        204/21/21/00003627.pdf;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/204/21/21/abstract.html.

Holub:2001:LGC

[570] Štěpán Holub. Local and global cyclicity in free semigroups. Theoretical Computer Science, 262(1–2): 25–36, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        204/21/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/204/21/22/article.pdf.

Lee:2001:AEC

[571] David Lee and Henryk Woźniakowski. Approximate evaluations of characteristic polynomials of Boolean functions. Theoretical Computer Science, 262(1–2):37–68, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/23/article.pdf.

Bertoni:2001:ADB

[572] Alberto Bertoni and Marco Carpentieri. Analogies and differences between quantum and stochastic automata. Theoretical Computer Science, 262(1–2):69–81, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/24/article.pdf.

Chen:2001:AST

[573] Donghui Chen, Ding-Zhu Du, Xiao-Dong Hu, Guo-Hui Lin, Lusheng Wang, and Guoliang Xue. Approximations for Steiner trees with minimum number of Steiner points. Theoretical Computer Science, 262(1–2): 83–99, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/25/article.pdf.

Kao:2001:MRE

[574] Ming-Yang Kao and Jie Wang. Minimizing roundoff errors of prefix sums via dynamic construction of Huffman trees. Theoretical Computer Science, 262(1–2):101– 115, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        204/21/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/204/21/26/article.pdf.

Kobayashi:2001:FPP

[575] Satoshi Kobayashi, Victor Mitrana, Gheorghe Păun, and Grzegorz Rozenberg. Formal properties of PAmatching. Theoretical Computer Science, 262(1–2):117–131, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/204/21/27/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/204/21/27/article.
                                        pdf.

Kalorkoti:2001:PQB

[576] K. Kalorkoti and D. H. Tulley. Priority queues with binary priorities. Theoretical Computer Science, 262(1– 2):133–144, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        204/21/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/204/21/28/article.pdf.

Aspvall:2001:AGB

[577] Bengt Aspvall, Magnús M. Halldórsson, and Fredrik Manne. Approximations for the general block distribution of a matrix. Theoretical Computer Science, 262(1–2):145– 160, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/

204/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/29/article.pdf.

DasGupta:2001:PTA

[578] Bhaskar DasGupta and Eduardo D. Sontag. A polynomial-time algorithm for checking equivalence under certain semiring congruences motivated by the state-space isomorphism problem for hybrid systems. Theoretical Computer Science, 262(1–2):161– 189, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/30/article.pdf.

Jenkinson:2001:SCT

[579] Oliver Jenkinson. Strong cocycle triviality for Z^{2} subshifts. Theoretical Computer Science, 262(1–2): 191–213, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/31/article.pdf.

Hsieh:2001:LFF

[580] Sun-Yuan Hsieh, Gen-Huey Chen, and Chin-Wen Ho. Longest fault-free paths in star graphs with vertex faults. Theoretical Computer Science, 262(1–2): 215–227, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/32/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/204/21/32/article.pdf.

Danyi:2001:CHC

[581] G. Dányi and Z. Fülöp. The component hierarchy of chain-free cooperating distributed regular tree grammars. Theoretical Computer Science, 262(1– 2):229–240, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/33/article.pdf.

Fenner:2001:HPH

[582] Stephen Fenner, Steven Homer, Randall Pruim, and Marcus Schaefer. Hyper-polynomial hierarchies and the polynomial jump. Theoretical Computer Science, 262(1–2): 241–256, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/34/article.pdf.

Lu:2001:WSB

[583] Bing Lu, Jun Gu, Xiaodong Hu, and Eugene Shragowitz. Wire segmenting for buffer insertion based on RSTP-MSP. Theoretical Computer Science, 262(1–2):257–267, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/204/21/35/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/204/21/35/article.
pdf.

Duval:2001:RPI

[584] J.-P. Duval, F. Mignosi, and A. Restivo. Recurrence and periodicity in infinite words from local periods. Theoretical Computer Science, 262(1–2): 269–284, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/36/article.pdf.

Kindler:2001:ANF

[585] Ekkart Kindler and Hagen Völzer. Algebraic nets with flexible arcs. Theoretical Computer Science, 262(1–2): 285–310, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/37/article.pdf.

Ohlebusch:2001:ICT

[586] Enno Ohlebusch. Implementing conditional term rewriting by graph rewriting. Theoretical Computer Science, 262 (1–2):311–331, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/38/article.pdf.

Lam:2001:FMS

[587] Nguyen Huong Lam. Finite maximal solid codes. Theoretical Computer Science, 262(1–2):333–347, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/204/21/39/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/204/21/39/article.
pdf.

Fiebig:2001:MCS

[588] Doris Fiebig, Ulf-Rainer Fiebig, and Nataša Jonoska. Multiplicities of covers for sofic shifts. Theoretical Computer Science, 262(1–2):349– 375, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/40/article.pdf.

Drewes:2001:TBG

[589] Frank Drewes. Tree-based generation of languages of fractals. Theoretical Computer Science, 262(1–2): 377–414, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/41/article.pdf.

Park:2001:PAR

[590] Heejin Park and Kunsoo Park. Parallel algorithms for red-black trees. Theoretical Computer Science, 262(1– 2):415–435, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/42/article.pdf.

Seiden:2001:PMS

[591] Steven S. Seiden. Preemptive multiprocessor scheduling with rejection. Theoretical Computer Science, 262(1– 2):437–458, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/43/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/43/article.pdf.

Cheng:2001:OSS

[592] Siu-Wing Cheng and Yin-Feng Xu. Onbeta-skeleton as a subgraph of the minimum weight triangulation. Theoretical Computer Science, 262(1–2): 459–471, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/44/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/44/article.pdf.

Spreen:2001:RVN

[593] Dieter Spreen. Representations versus numberings: on the relationship of two computability notions. Theoretical Computer Science, 262(1–2): 473–499, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/45/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/45/article.pdf.

Borchert:2001:DO

[594] Bernd Borchert and Riccardo Silvestri. Dot operators. Theoretical Computer Science, 262(1–2):501– 523, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/46/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/46/article.pdf.

Latapy:2001:SSS

[595] Matthieu Latapy, Roberto Mantaci, Michel Morvan, and Ha Duong Phan. Structure of some sand piles model. Theoretical Computer Science, 262(1– 2):525–556, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/47/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/47/article.pdf.

Sim:2001:APS

[596] Jeong Seop Sim, Costas S. Iliopoulos, Kunsoo Park, and W. F. Smyth. Approximate periods of strings. Theoretical Computer Science, 262(1–2): 557–568, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/48/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/48/article.pdf.

Czumaj:2001:EWS

[597] Artur Czumaj, Ian Finch, Leszek Ga̧sieniec, Alan Gibbons, Paul Leng, Wojciech Rytter, and Michele Zito. Efficient Web searching using temporal factors. Theoretical Computer Science, 262(1–2):569–582, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/204/21/49/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/204/21/49/article.
pdf.

Kobayashi:2001:TDC

[598] Yuji Kobayashi, Masashi Katsura, and Kayoko Shikishima-Tsuji. Termination and derivational complexity of confluent one-rule string-rewriting systems. Theoretical Computer Science, 262(1– 2):583–632, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/50/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/50/article.pdf.

Takenaga:2001:TSB

[599] Yasuhiko Takenaga, Kouji Nakajima, and Shuzo Yajima. Tree-shellability of Boolean functions. Theoretical Computer Science, 262(1–2):633– 647, July 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/51/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/204/21/51/article.pdf.

Janssen:2001:AAC

[600] Jeannette Janssen and Lata Narayanan. Approximation algorithms for channel assignment with constraints. Theoretical Computer Science, 262(1–2): 649–667, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
204/21/52/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/204/21/52/article.pdf.

Finkel:2001:TPO

[601] Olivier Finkel. Topological properties of omega context-free languages. Theoretical Computer Science, 262(1– 2):669–697, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        204/21/53/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/204/21/53/article.pdf.

Anonymous:2001:AIVi

[602] Anonymous. Author index volume 262 (2001). Theoretical Computer Science, 262(1–2):699–700, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/204/21/54/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/204/21/54/article.
                                        pdf.

Chang:2001:P

[603] Gerard Chang, Michel Deza, Yannis Manoussakis, and Jean-Marc Steyaert. Preface. Theoretical Computer Science, 263(1–2):1, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/27/00003694.pdf;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/27/abstract.html.

Yeh:2001:WCD

[604] Hong-Gwa Yeh and Gerard J. Chang. Weighted connected k-domination and weighted k-dominating clique in distance-hereditary graphs. Theoretical Computer Science, 263(1–2): 3–8, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/28/article.pdf.

Fukuda:2001:NAO

[605] Komei Fukuda, Alain Prodon, and Tadashi Sakuma. Notes on acyclic orientations and the shelling lemma. Theoretical Computer Science, 263(1– 2):9–16, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/29/article.pdf.

Nielsen:2001:PCO

[606] Frank Nielsen. On point covers of c-oriented polygons. Theoretical Computer Science, 263(1–2):17– 29, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/30/article.pdf.

Cara:2001:NIM

[607] Philippe Cara, Serge Lehman, and Dimitrii V. Pasechnik. On the number of inductively minimal geometries. Theoretical Computer Science, 263(1– 2):31–35, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/31/article.pdf.

Pasechnik:2001:CHB

[608] Dmitrii V. Pasechnik. On computing Hilbert bases via the Elliot-MacMahon algorithm. Theoretical Computer Science, 263(1–2):37– 46, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/32/article.pdf.

Euler:2001:CBT

[609] Reinhardt Euler. Characterizing bipartite Toeplitz graphs. Theoretical Computer Science, 263(1–2):47– 58, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/33/article.pdf.

Yegnanarayanan:2001:GCP

[610] V. Yegnanarayanan. Graph colourings and partitions. Theoretical Computer Science, 263(1–2):59– 74, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/34/article.pdf.

Manoussakis:2001:RNT

[611] Yannis Manoussakis and Zsolt Tuza. Ramsey numbers for tournaments. Theoretical Computer Science, 263(1– 2):75–85, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/35/article.pdf.

Karolyi:2001:GGR

[612] Gyula Károlyi and Vera Rosta. Generalized and geometric Ramsey numbers for cycles. Theoretical Computer Science, 263(1–2):87– 98, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/36/article.pdf.

Damiand:2001:SPG

[613] Guillaume Damiand, Michel Habib, and Christophe Paul. A simple paradigm for graph recognition: application to cographs and distance hereditary graphs. Theoretical Computer Science, 263(1–2):99– 111, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/37/article.pdf.

Lu:2001:FSB

[614] Bor-Kuan Lu, Fang-Rong Hsu, and Chuan Yi Tang. Finding the shortest boundary guard of a simple polygon. Theoretical Computer Science, 263(1– 2):113–121, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/38/article.pdf.

Belaga:2001:MAT

[615] E. G. Belaga. Mod 3 arithmetic on triangulated Riemann surfaces. Theoretical Computer Science, 263(1–2): 123–137, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/39/article.pdf.

Tong:2001:CGB

[616] Li-Da Tong, Frank K. Hwang, and Gerard J. Chang. Channel graphs of bit permutation networks. Theoretical Computer Science, 263(1–2): 139–143, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/40/article.pdf.

Hwang:2001:UAS

[617] Hsien-Kuei Hwang. Uniform asymptotics of some Abel sums arising in coding theory. Theoretical Computer Science, 263(1–2):145– 158, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/41/article.pdf.

Angel:2001:LRQ

[618] Eric Angel and Vassilis Zissimopoulos. On the landscape ruggedness of the quadratic assignment problem. Theoretical Computer Science, 263(1– 2):159–172, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/42/article.pdf.

Harbane:2001:ERA

[619] Rabah Harbane and Marie-Claude Heydemann. Efficient reconfiguration algorithms of de Bruijn and Kautz networks into linear arrays. Theoretical Computer Science, 263(1–2): 173–189, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/43/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/43/article.pdf.

Quilliot:2001:CGO

[620] A. Quilliot. Convexity and global optimization: a theoretical link. Theoretical Computer Science, 263(1–2): 191–204, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/44/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/44/article.pdf.

Negami:2001:NRT

[621] Seiya Negami. Note on Ramsey theorems for spatial graphs. Theoretical Computer Science, 263(1–2):205– 210, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/45/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/45/article.pdf.

Hwang:2001:CSD

[622] F. K. Hwang. A complementary survey on double-loop networks. Theoretical Computer Science, 263(1–2): 211–229, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/46/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/46/article.pdf.

Martin:2001:SCA

[623] Bruno Martin. A simulation of cellular automata on hexagons by cellular automata on rings. Theoretical Computer Science, 263(1–2):231– 234, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/47/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/47/article.pdf.

Calegari:2001:COA

[624] Patrice Calégari, Frédéric Guidec, Pierre Kuonen, and Frank Nielsen. Combinatorial optimization algorithms for radio network planning. Theoretical Computer Science, 263(1–2):

235–245, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/48/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/48/article.pdf.

Favaron:2001:NUR

[625] O. Favaron and Y. Redouane. Neighborhood unions and regularity in graphs. Theoretical Computer Science, 263(1–2):247–254, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/212/27/49/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/212/27/49/article.
pdf.

Bazgan:2001:PVT

[626] Cristina Bazgan, Amel Harkat-Benhamdine, Hao Li, and Mariusz Woźniak. Partitioning vertices of 1-tough graphs into paths. Theoretical Computer Science, 263(1–2):255–261, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/212/27/50/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/212/27/50/article.
pdf.

Indermitte:2001:VDP

[627] C. Indermitte, Th. M. Liebling, M. Troyanov, and H. Clémençon. Voronoi diagrams on piecewise flat surfaces and an application to biological growth. Theoretical Computer Science, 263(1–2):263–274, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/212/27/51/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/212/27/51/article.
                                        pdf.

Yeh:2001:FDC

[628] Roger K. Yeh. Finite-dimensional T-colorings of graphs. Theoretical Computer Science, 263(1–2):275– 281, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/52/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/52/article.pdf.

Golin:2001:CAG

[629] Mordecai J. Golin. A combinatorial approach to Golomb forests. Theoretical Computer Science, 263(1–2): 283–304, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/53/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/53/article.pdf.

Matsui:2001:NCC

[630] Yasuko Matsui and Tomomi Matsui. NP-completeness for calculating power indices of weighted majority games. Theoretical Computer Science, 263(1– 2):305–310, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        212/27/54/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/212/27/54/article.pdf.

Ozeki:2001:CRP

[631] Michio Ozeki. On the covering radius problem for ternary self-dual codes. Theoretical Computer Science, 263(1– 2):311–332, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/55/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/55/article.pdf.

Li:2001:ITC

[632] Hao Li and Jianping Li. Independent triangles covering given vertices of a graph. Theoretical Computer Science, 263(1–2):333– 344, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/56/00003723.pdf;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/56/abstract.html.

Maehara:2001:WDP

[633] H. Maehara and N. Tokushige. When does a planar bipartite framework admit a continuous deformation? Theoretical Computer Science, 263(1–2): 345–354, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/57/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/57/article.pdf.

Cheng:2001:ISP

[634] Szu-En Cheng and Ko-Wei Lih. An improvement on a spernerity proof of Horrocks. Theoretical Computer Science, 263(1–2):355– 377, July 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
212/27/58/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/212/27/58/article.pdf.

Anonymous:2001:AIVj

[635] Anonymous. Author index volume 263 (2001). Theoretical Computer Science, 263(1–2):379–380, July 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/212/27/59/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/212/27/59/article.
pdf.

Freivalds:2001:F

[636] Rūsiņš Freivalds, Juraj Hromkovič, George Păun, and Walter Unger. Foreword. Theoretical Computer Science, 264(1):1–2, August 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/21/article.pdf.

Kobayashi:2001:MSS

[637] Satoshi Kobayashi and Yasubumi Sakakibara. Multiple splicing systems and the universal computability. Theoretical Computer Science, 264 (1):3–23, August 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/22/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/205/23/22/article.pdf.

Manca:2001:LSR

[638] Vincenzo Manca. Logical string rewriting. Theoretical Computer Science, 264 (1):25–51, August 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/23/article.pdf.

Barriere:2001:GCR

[639] Lali Barrière, Johanne Cohen, and Margarida Mitjana. Gossiping in chordal rings under the line model. Theoretical Computer Science, 264(1): 53–64, August 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/24/article.pdf.

Bockenhauer:2001:CTW

[640] Hans-Joachim Böckenhauer. Communication in the two-way listen-in vertex-disjoint paths mode. Theoretical Computer Science, 264(1):65– 90, August 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/25/article.pdf.

Goerdt:2001:RRG

[641] Andreas Goerdt. Random regular graphs with edge faults: Expansion through cores. Theoretical Computer Science, 264(1):91– 125, August 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/26/article.pdf.

Ablayev:2001:BVN

[642] Farid Ablayev, Marek Karpinski, and Rustam Mubarakzjanov. On BPP versus NP coNP for ordered readonce branching programs. Theoretical Computer Science, 264(1):127– 137, August 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/27/article.pdf.

Mosca:2001:CQE

[643] Michele Mosca. Counting by quantum eigenvalue estimation. Theoretical Computer Science, 264(1):139– 153, August 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/28/article.pdf.

Ambainis:2001:PII

[644] Andris Ambainis. Probabilistic inductive inference: a survey. Theoretical Computer Science, 264(1): 155–167, August 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
205/23/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/205/23/29/article.pdf.

Dubois:2001:E

[645] O. Dubois, R. Monasson, B. Selman, and R. Zecchina. Editorial. Theoretical Computer Science, 265(1–2): 1, August 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
215/27/27/00004029.pdf;   http:
//www.elsevier.nl/gej-ng/10/41/
16/215/27/27/abstract.html.

Martin:2001:SMM

[646] Olivier C. Martin, Rémi Monasson, and Riccardo Zecchina. Statistical mechanics methods and phase transitions in optimization problems. Theoretical Computer Science, 265(1–2): 3–67, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
215/27/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/215/27/28/article.pdf.

Talagrand:2001:RRM

[647] Michel Talagrand. Rigorous results for mean field models for spin glasses. Theoretical Computer Science, 265(1– 2):69–77, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
215/27/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/215/27/29/article.pdf.

Mertens:2001:PAN

[648] Stephan Mertens. A physicist’s approach to number partitioning. Theoretical Computer Science, 265(1–2): 79–108, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        215/27/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/215/27/30/article.pdf.

Achlioptas:2001:RRR

[649] Dimitris Achlioptas, Lefteris M. Kirousis, Evangelos Kranakis, and Danny Krizanc. Rigorous results for random (2 + p)-SAT. Theoretical Computer Science, 265(1–2):109– 129, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        215/27/31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/215/27/31/article.pdf.

delaVega:2001:RSR

[650] W. Fernandez de la Vega. Random 2-SAT: results and problems. Theoretical Computer Science, 265(1–2): 131–146, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        215/27/32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/215/27/32/article.pdf.

Franco:2001:RRT

[651] John Franco. Results related to threshold phenomena research in satisfiability: lower bounds. Theoretical Computer Science, 265(1–2):147– 157, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
215/27/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/215/27/33/article.pdf.

Achlioptas:2001:LBR

[652] Dimitris Achlioptas. Lower bounds for random 3-SAT via differential equations. Theoretical Computer Science, 265(1–2):159–185, August 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
215/27/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/215/27/34/article.pdf.

Dubois:2001:UBS

[653] Olivier Dubois. Upper bounds on the satisfiability threshold. Theoretical Computer Science, 265(1–2): 187–197, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
215/27/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/215/27/35/article.pdf.

Hartmann:2001:SMP

[654] Alexander K. Hartmann and Martin Weigt. Statistical mechanics perspective on the phase transition in vertex covering of finiteconnectivity random graphs. Theoretical Computer Science, 265(1–2): 199–225, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        215/27/36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/215/27/36/article.pdf.

Culberson:2001:FDG

[655] Joseph Culberson and Ian Gent. Frozen development in graph coloring. Theoretical Computer Science, 265(1–2): 227–264, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        215/27/37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/215/27/37/article.pdf.

Smith:2001:CAP

[656] Barbara M. Smith. Constructing an asymptotic phase transition in random binary constraint satisfaction problems. Theoretical Computer Science, 265(1–2):265–283, August 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        215/27/38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/215/27/38/article.pdf.

Engel:2001:CLA

[657] Andreas Engel. Complexity of learning in artificial neural networks. Theoretical Computer Science, 265(1–2): 285–306, August 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        215/27/39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/215/27/39/article.pdf.

Anonymous:2001:AIVk

[658] Anonymous. Author index volume 265 (2001). Theoretical Computer Science, 265(1–2):307, August 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
215/27/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/215/27/40/article.pdf.

Schurmann:2001:PRH

[659] Carsten Schürmann, Joëlle Despeyroux, and Frank Pfenning. Primitive recursion for higher-order abstract syntax. Theoretical Computer Science, 266(1–2):1–57, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/216/27/27/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/216/27/27/article.
pdf.

Matsushita:2001:APU

[660] Tatsuru Matsushita and Colin Runciman. The accepting power of unary string logic programs. Theoretical Computer Science, 266(1–2):59– 79, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/28/article.pdf.

Aguilera:2001:RNC

[661] G. Aguilera, I. P. de Guzmán, M. Ojeda-Aciego, and A. Valverde.

Reductions for non-clausal theorem proving. Theoretical Computer Science, 266(1–2):81–112, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/216/27/29/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/216/27/29/article.
pdf.

Schulz:2001:CCS

[662] Klaus U. Schulz and Stephan Kepser. Combination of constraint systems II: Rational amalgamation. Theoretical Computer Science, 266(1–2):113– 157, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/30/article.pdf.

David:2001:ABP

[663] René David. On the asymptotic behaviour of primitive recursive algorithms. Theoretical Computer Science, 266(1–2):159–193, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/31/article.pdf.

Lermer:2001:RSM

[664] Karl Lermer and Paul Strooper. Refinement and state machine abstraction. Theoretical Computer Science, 266(1–2):195–235, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/216/27/32/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/216/27/32/article.
pdf.

Boreale:2001:DTR

[665] Michele Boreale, Rocco De Nicola, and Rosario Pugliese. Divergence in testing and readiness semantics. Theoretical Computer Science, 266(1–2): 237–248, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/33/article.pdf.

Hashimoto:2001:TCC

[666] Masatomo Hashimoto and Atsushi Ohori. A typed context calculus. Theoretical Computer Science, 266(1–2): 249–272, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/34/article.pdf.

Aspinall:2001:SDT

[667] David Aspinall and Adriana Compagnoni. Subtyping dependent types. Theoretical Computer Science, 266 (1–2):273–309, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/216/27/35/abstract.
html;   http://www.elsevier.nl/

gej-ng/10/41/16/216/27/35/article.
pdf.

Koubarakis:2001:TDL

[668] Manolis Koubarakis. Tractable disjunctions of linear constraints: basic results and applications to temporal reasoning. Theoretical Computer Science, 266(1–2):311–339, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/36/article.pdf.

Loader:2001:FPD

[669] Ralph Loader. Finitary PCF is not decidable. Theoretical Computer Science, 266(1–2):341–364, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/216/27/37/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/216/27/37/article.
pdf.

Wijesekera:2001:NFS

[670] Duminda Wijesekera, M. Ganesh, Jaideep Srivastava, and Anil Nerode. Normal forms and syntactic completeness proofs for functional independencies. Theoretical Computer Science, 266(1–2):365–405, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/216/27/38/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/216/27/38/article.
pdf.

Munoz:2001:PTS

[671] César Muñoz. Proof-term synthesis on dependent-type systems via explicit substitutions. Theoretical Computer Science, 266(1–2):407–440, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/39/article.pdf.

Shen:2001:LCL

[672] Yi-Dong Shen, Li-Yan Yuan, and Jia-Huai You. Loop checks for logic programs with functions. Theoretical Computer Science, 266(1–2):441– 461, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/40/article.pdf.

Bauderon:2001:NRG

[673] Michel Bauderon and Hélène Jacquet. Node rewriting in graphs and hypergraphs: a categorical framework. Theoretical Computer Science, 266(1–2): 463–487, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/41/article.pdf.

Yamasaki:2001:SCP

[674] Susumu Yamasaki and Yoshinori Kurose. A sound and complete procedure for a general logic program in non-floundering derivations with respect to the 3-valued stable model semantics. Theoretical Computer Science, 266(1–2):489–512, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/42/article.pdf.

McIver:2001:PCP

[675] A. K. McIver and Carroll Morgan. Partial correctness for probabilistic demonic programs. Theoretical Computer Science, 266(1–2):513– 541, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/43/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/43/article.pdf.

Pucella:2001:EPF

[676] Riccardo Pucella and Prakash Panangaden. On the expressive power of firstorder boolean functions in PCF. Theoretical Computer Science, 266(1–2): 543–567, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/44/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/44/article.pdf.

Fournet:2001:BJC

[677] Cédric Fournet and Cosimo Laneve. Bisimulations in the join-calculus. Theoretical Computer Science, 266(1–2):

569–603, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/45/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/45/article.pdf.

Srba:2001:BPA

[678] Jiří Srba. Basic process algebra with deadlocking states. Theoretical Computer Science, 266(1–2):605– 630, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/46/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/46/article.pdf.

Groote:2001:PCU

[679] Jan Friso Groote and Jos van Wamel. The parallel composition of uniform processes with data. Theoretical Computer Science, 266(1–2):631–652, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/47/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/47/article.pdf.

Abraham:2001:SSE

[680] Uri Abraham, Shlomi Dolev, Ted Herman, and Irit Koll. Selfstabilizing l-exclusion. Theoretical Computer Science, 266(1–2):653–692, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/48/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/216/27/48/article.pdf.

Riely:2001:DPL

[681] James Riely and Matthew Hennessy. Distributed processes and location failures. Theoretical Computer Science, 266(1–2):693–735, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/216/27/49/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/216/27/49/article.
                                        pdf.

Khasidashvili:2001:LPR

[682] Zurab Khasidashvili. On the longest perpetual reductions in orthogonal expression reduction systems. Theoretical Computer Science, 266(1–2):737– 772, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        216/27/50/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/216/27/50/article.pdf.

Barthe:2001:IPP

[683] Gilles Barthe, John Hatcliff, and Morten Heine B. Sørensen. An induction principle for pure type systems. Theoretical Computer Science, 266(1–2):773–818, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.elsevier.nl/gej-
                                        ng/10/41/16/216/27/51/abstract.
                                        html;   http://www.elsevier.nl/
                                        gej-ng/10/41/16/216/27/51/article.
                                        pdf.

Schlechta:2001:EUM

[684] Karl Schlechta and Jürgen Dix. Explaining updates by minimal sums. Theoretical Computer Science, 266 (1–2):819–838, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/216/27/52/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/216/27/52/article.
pdf.

Ying:2001:REH

[685] Mingsheng Ying and Martin Wirsing. Recursive equations in higherorder process calculi. Theoretical Computer Science, 266(1–2):839–852, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/53/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/53/article.pdf.

Goldblatt:2001:WCA

[686] Robert Goldblatt. What is the coalgebraic analogue of Birkhoff’s variety theorem? Theoretical Computer Science, 266(1–2):853–886, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/54/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/54/article.pdf.

Auletta:2001:SLW

[687] Vincenzo Auletta, Ioannis Caragiannis, Luisa Gargano, Christos Kaklamanis, and Pino Persiano. Sparse and limited wavelength conversion in all-optical tree networks. Theoretical Computer Science, 266(1–2):887– 934, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/55/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/55/article.pdf.

Majster-Cederbaum:2001:USP

[688] M. E. Majster-Cederbaum. Underspecification for a simple process algebra of recursive processes. Theoretical Computer Science, 266(1–2): 935–950, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/56/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/56/article.pdf.

Turner:2001:TIS

[689] Raymond Turner. Type inference for set theory. Theoretical Computer Science, 266(1–2):951–974, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/57/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/57/article.pdf.

Joly:2001:CTP

[690] Thierry Joly. Constant time parallel computations in λ-calculus. Theoretical Computer Science, 266(1–2): 975–985, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/58/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/58/article.pdf.

Krantz:2001:EHC

[691] Thomas Krantz and Virgile Mogbil. Encoding Hamiltonian circuits into multiplicative linear logic. Theoretical Computer Science, 266(1–2):987– 996, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/59/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/59/article.pdf.

Spreen:2001:CFP

[692] Dieter Spreen. Corrigendum to “On functions preserving levels of approximation: a refined model construction for various lambda calculi” [Theoret. Comput. Sci. 212 (1999) 261–303]. Theoretical Computer Science, 266(1–2): 997–998, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
216/27/60/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/216/27/60/article.pdf;   http:
//www.elsevier.nl/PII/S0304397598001443.

See [5].

Anonymous:2001:AIVl

[693] Anonymous. Author index volume 266 (2001). Theoretical Computer Science, 266(1–2):999–1000, September 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

                                        216/27/61/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/216/27/61/article.pdf.

Champarnaud:2001:F

[694] Jean-Marc Champarnaud, Denis Maurel, and Djelloul Ziadi. Foreword. Theoretical Computer Science, 267(1– 2):1, September 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        218/27/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/218/27/27/article.pdf.

Campeanu:2001:MCA

[695] C. Câmpeanu, N. Sântean, and S. Yu. Minimal cover-automata for finite languages. Theoretical Computer Science, 267(1–2):3–16, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        218/27/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/218/27/28/article.pdf.

Champarnaud:2001:SCC

[696] J.-M. Champarnaud. Subset construction complexity for homogeneous automata, position automata and ZPC-structures. Theoretical Computer Science, 267(1–2):17–34, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        218/27/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/218/27/29/article.pdf.

Albert:2001:NFA

[697] Jürgen Albert, Dora Giammarresi, and Derick Wood. Normal form algorithms for extended context-free grammars. Theoretical Computer Science, 267(1– 2):35–47, September 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/30/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/30/article.pdf.

Blum:2001:PLL

[698] Norbert Blum. On parsing LLlanguages. Theoretical Computer Science, 267(1–2):49–59, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/31/article.pdf.

Goeman:2001:PCS

[699] Heiko Goeman. On parsing and condensing substrings of LR languages in linear time. Theoretical Computer Science, 267(1–2):61– 82, September 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/32/article.pdf.

Nykanen:2001:UAT

[700] Matti Nykänen. Using acceptors as transducers. Theoretical Computer Science, 267(1–2):83–104, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/33/article.pdf.

Duchamp:2001:DDL

[701] G. Duchamp, M. Flouret, É. Laugerotte, and J.-G. Luque. Direct and dual laws for automata with multiplicities. Theoretical Computer Science, 267(1–2): 105–120, September 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/34/article.pdf.

Maurel:2001:SPT

[702] Denis Maurel and Brigitte Le Pévédic. The syntactic prediction with token automata: application to HandiAS system. Theoretical Computer Science, 267(1–2):121–129, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/35/article.pdf.

Poibeau:2001:PNL

[703] T. Poibeau. Parsing natural language idioms with bidirectional finitestate machines. Theoretical Computer Science, 267(1–2):131–140, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/36/article.pdf.

LHer:2001:PSF

[704] Dominique L’Her, Philippe Le Parc, and Lionel Marcé. Proving sequential function chart programs using timed automata. Theoretical Computer Science, 267(1–2):141–155, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/37/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/37/article.pdf.

Anonymous:2001:AIVm

[705] Anonymous. Author index volume 267 (2001). Theoretical Computer Science, 267(1–2):157, September 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
218/27/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/218/27/38/article.pdf.

Leonardi:2001:P

[706] Stefano Leonardi and Alberto Marchetti-Spaccamela. Preface. Theoretical Computer Science, 268(1):1, October 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/27/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/27/27/article.pdf.

Ambuhl:2001:NLB

[707] Christoph Ambühl, Bernd Gärtner, and Bernhard von Stengel. A new lower bound for the list update problem in the partial cost model. Theoretical Computer Science, 268(1):3– 16, October 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        217/27/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/217/27/28/article.pdf.

Azar:2001:LS

[708] Yossi Azar and Oded Regev. Online bin-stretching. Theoretical Computer Science, 268(1):17–41, October 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        217/27/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/217/27/29/article.pdf.

Bartal:2001:PMO

[709] Yair Bartal, Moses Charikar, and Piotr Indyk. On page migration and other relaxed task systems. Theoretical Computer Science, 268(1): 43–66, October 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        217/27/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/217/27/30/article.pdf.

Bischof:2001:LSP

[710] Stefan Bischof and Ernst W. Mayr. On-line scheduling of parallel jobs with runtime restrictions. Theoretical Computer Science, 268(1):67– 90, October 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/27/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/27/31/article.pdf.

Feuerstein:2001:LSS

[711] Esteban Feuerstein and Leen Stougie. On-line single-server dial-a-ride problems. Theoretical Computer Science, 268(1):91–105, October 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/217/27/32/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/217/27/32/article.
pdf.

Glazek:2001:OAP

[712] Włodzimierz Głazek. Online algorithms for page replication in rings. Theoretical Computer Science, 268(1): 107–117, October 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/27/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/27/33/article.pdf.

Kimbrel:2001:OPF

[713] Tracy Kimbrel. Online paging and file caching with expiration times. Theoretical Computer Science, 268(1): 119–131, October 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

                                        217/27/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/217/27/34/article.pdf.

Noga:2001:OOA

[714] John Noga and Steven S. Seiden. An optimal online algorithm for scheduling two machines with release times. Theoretical Computer Science, 268(1): 133–143, October 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        217/27/35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/217/27/35/article.pdf.

Riedel:2001:ORS

[715] Marco Riedel. Online request server matching. Theoretical Computer Science, 268(1):145–160, October 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        217/27/36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/217/27/36/article.pdf.

Fleischer:2001:BP

[716] Rudolf Fleischer. On the Bahncard problem. Theoretical Computer Science, 268(1):161–174, October 6, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        217/27/37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/217/27/37/article.pdf.

Wiehagen:2001:F

[717] R. Wiehagen and T. Zeugmann. Foreword. Theoretical Computer Science, 268(2):175–177, October 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/217/31/27/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/217/31/27/article.
pdf.

Rao:2001:PTL

[718] M. R. K. Krishna Rao and A. Sattar. Polynomial-time learnability of logic programs with local variables from entailment. Theoretical Computer Science, 268(2):179– 198, October 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/31/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/31/28/article.pdf.

Fischlin:2001:CLP

[719] Marc Fischlin. Cryptographic limitations on parallelizing membership and equivalence queries with applications to random-self-reductions. Theoretical Computer Science, 268(2):199– 219, October 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/31/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/31/29/article.pdf.

Stephan:2001:LAS

[720] Frank Stephan and Yuri Ventsov. Learning algebraic structures from text. Theoretical Computer Science, 268(2):221–273, October 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/217/31/30/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/217/31/30/article.
pdf.

Meyer:2001:ACP

[721] Léa Meyer. Aspects of complexity of probabilistic learning under monotonicity constraints. Theoretical Computer Science, 268(2):275– 322, October 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/31/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/31/31/article.pdf.

Case:2001:PLM

[722] John Case, Sanjay Jain, Susanne Kaufmann, Arun Sharma, and Frank Stephan. Predictive learning models for concept drift. Theoretical Computer Science, 268(2):323–349, October 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/31/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/31/32/article.pdf.

Hirowatari:2001:CIC

[723] Eiju Hirowatari and Setsuo Arikawa. A comparison of identification criteria for inductive inference of recursive real-valued functions. Theoretical Computer Science, 268(2):351– 366, October 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/31/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/31/33/article.pdf.

Apsitis:2001:CPE

[724] Kalvis Apsītis, Rūsiņš Freivalds, Raimonds Simanovskis, and Juris Smotrovs. Closedness properties in ex-identification. Theoretical Computer Science, 268(2):367–393, October 17, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/31/34/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/31/34/article.pdf.

Anonymous:2001:AIVn

[725] Anonymous. Author index volume 268 (2001). Theoretical Computer Science, 268(2):395, October 17, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
217/31/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/217/31/35/article.pdf.

Safavi-Naini:2001:BAG

[726] Rei Safavi-Naini and Huaxiong Wang. Broadcast authentication for group communication. Theoretical Computer Science, 269(1–2):1–21, October 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/27/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/219/27/27/article.pdf.

Kerth:2001:CSM

[727] Rainer Kerth. On the construction of stable models of untyped λ-calculus. Theoretical Computer Science, 269(1– 2):23–46, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        219/27/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/219/27/28/article.pdf.

Corradini:2001:SDA

[728] Flavio Corradini, GianLuigi Ferrari, and Marco Pistore. On the semantics of durational actions. Theoretical Computer Science, 269(1–2): 47–82, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        219/27/29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/219/27/29/article.pdf.

Quaglia:2001:ESP

[729] Paola Quaglia. Explicit substitutions for pi-congruences. Theoretical Computer Science, 269(1–2):83– 134, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        219/27/30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/219/27/30/article.pdf.

deMoor:2001:HOM

[730] Oege de Moor and Ganesh Sittampalam. Higher-order matching for program transformation. Theoretical Computer Science, 269(1–2):135– 162, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/31/article.pdf.

Riany:2001:TPS

[731] Yaron Riany, Nir Shavit, and Dan Touitou. Towards a practical snapshot algorithm. Theoretical Computer Science, 269(1–2):163–201, October 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/32/article.pdf.

Bergstra:2001:NRI

[732] Jan A. Bergstra and Alban Ponse. Nonregular iterators in process algebra. Theoretical Computer Science, 269(1– 2):203–229, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/33/article.pdf.

Tini:2001:ASE

[733] Simone Tini. An axiomatic semantics for Esterel. Theoretical Computer Science, 269(1–2):231–282, October 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/

                                        219/27/34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/219/27/34/article.pdf.

Finkel:2001:WHO

[734] Olivier Finkel. Wadge hierarchy of omega context-free languages. Theoretical Computer Science, 269(1–2): 283–315, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        219/27/35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/219/27/35/article.pdf.

Barthe:2001:WNI

[735] Gilles Barthe, John Hatcliff, and Morten Heine Sørensen. Weak normalization implies strong normalization in a class of non-dependent pure type systems. Theoretical Computer Science, 269(1–2):317–361, October 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        219/27/36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/219/27/36/article.pdf.

Lago:2001:DFO

[736] Joaquín Mateos Lago and Mario Rodríguez Artalejo. A declarative framework for object-oriented programming with genetic inheritance. Theoretical Computer Science, 269(1–2):363–417, October 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        219/27/37/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/219/27/37/article.pdf.

Tasic:2001:OCC

[737] Dragan Mašulovićand Boža Tasić. Operators on classes of coalgebras. Theoretical Computer Science, 269(1–2): 419–431, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/38/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/38/article.pdf.

Lepper:2001:DLO

[738] Ingo Lepper. Derivation lengths and order types of Knuth–Bendix orders. Theoretical Computer Science, 269(1– 2):433–450, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/39/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/39/article.pdf.

Duntsch:2001:CSD

[739] Ivo Düntsch and Szabolcs Mikulás. Cylindric structures and dependencies in relational databases. Theoretical Computer Science, 269(1–2): 451–468, October 28, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/40/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/40/article.pdf.

Rigo:2001:NSR

[740] Michel Rigo. Numeration systems on a regular language: arithmetic operations, recognizability and formal power series. Theoretical Computer Science, 269(1–2):469–498, October 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/41/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/41/article.pdf.

Anonymous:2001:AIVo

[741] Anonymous. Author index volume 269 (2001). Theoretical Computer Science, 269(1–2):499, October 28, 2001. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
219/27/42/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/219/27/42/article.pdf.

Matz:2002:DDM

[742] Oliver Matz. Dot-depth, monadic quantifier alternation, and first-order closure over grids and pictures. Theoretical Computer Science, 270(1–2): 1–70, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/27/abstract.html.

Pelc:2002:SGE

[743] Andrzej Pelc. Searching games with errors—fifty years of coping with liars. Theoretical Computer Science, 270(1– 2):71–109, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/28/abstract.html.

Hwang:2002:SPR

[744] F. K. Hwang, Y. C. Yao, and B. Dasgupta. Some permutation routing algorithms for low-dimensional hypercubes. Theoretical Computer Science, 270(1– 2):111–124, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/29/abstract.html.

vanderAalst:2002:IWA

[745] W. M. P. van der Aalst and T. Basten. Inheritance of workflows: an approach to tackling problems related to change. Theoretical Computer Science, 270(1– 2):125–203, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/30/abstract.html.

Bourgade:2002:CSL

[746] Ménard Bourgade. Calculs sur les structures de langage dénombrable. Theoretical Computer Science, 270(1– 2):205–222, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/31/abstract.html.

Cockett:2002:RCC

[747] J. R. B. Cockett and Stephen Lack. Restriction categories I: categories of partial maps. Theoretical Computer Science, 270(1–2):223– 259, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/32/abstract.html.

Lenhart:2002:DPM

[748] William Lenhart and Giuseppe Liotta. The drawability problem for minimum weight triangulations. Theoretical Computer Science, 270(1–2): 261–286, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/33/abstract.html.

Case:2002:CSH

[749] John Case, Sanjay Jain, and Mandayam Suraj. Control structures in hypothesis spaces: the influence on learning. Theoretical Computer Science, 270(1–2):287–308, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/34/abstract.html.

Herley:2002:DPB

[750] Kieran T. Herley, Andrea Pietracaprina, and Geppino Pucci. Deterministic parallel backtrack search. Theoretical Computer Science, 270(1–2): 309–324, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/35/abstract.html.

Lam:2002:LLB

[751] Tak-Wah Lam, Hing-Fung Ting, Kar-Keung To, and Wai-Ha Wong. On-line load balancing of temporary tasks revisited. Theoretical Computer Science, 270(1–2):325–340, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/36/abstract.html.

Becchetti:2002:DEA

[752] Luca Becchetti, Paola Bertolazzi, Carlo Gaibisso, and Giorgio Gambosi. On the design of efficient ATM routing schemes. Theoretical Computer Science, 270(1–2):341– 359, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/37/abstract.html.

Caragiannis:2002:ECB

[753] Ioannis Caragiannis, Christos Kaklamanis, and Pino Persiano. Edge coloring of bipartite graphs with constraints. Theoretical Computer Science, 270(1– 2):361–399, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/38/abstract.html.

Blanchet-Sadri:2002:PWT

[754] F. Blanchet-Sadri and Robert A. Hegstrom. Partial words and a theorem of Fine and Wilf revisited. Theoretical Computer Science, 270(1–2): 401–419, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/39/abstract.html.

Fujimura:2002:TMP

[755] Kikuo Fujimura. Time-minimal paths amidst moving obstacles in three dimensions. Theoretical Computer Science, 270(1–2):421–440, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/40/abstract.html.

Nebel:2002:SST

[756] Markus E. Nebel. The stack-size of tries: a combinatorial study. Theoretical Computer Science, 270(1–2): 441–461, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/41/abstract.html.

Cherubini:2002:ALD

[757] Alessandra Cherubini, Stefano Crespi Reghizzi, and Pierluigi San Pietro. Associative language descriptions. Theoretical Computer Science, 270(1–2): 463–491, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/42/abstract.html.

Eiter:2002:DLR

[758] Thomas Eiter, Toshihide Ibaraki, and Kazuhisa Makino. Decision lists and related Boolean functions. Theoretical Computer Science, 270(1–2): 493–524, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/43/abstract.html.

Mairesse:2002:ABH

[759] Jean Mairesse and Laurent Vuillon. Asymptotic behavior in a heap model with two pieces. Theoretical Computer Science, 270(1–2):525– 560, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/44/abstract.html.

Murat:2002:POP

[760] Cécile Murat and Vangelis Th. Paschos. A priori optimization for the probabilistic maximum independent set problem. Theoretical Computer Science, 270(1–2):561– 590, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/45/abstract.html.

Bergman:2002:CCS

[761] Clifford Bergman and Giora Slutzki. Computational complexity of some problems involving congruences on algebras. Theoretical Computer Science, 270(1–2):591–608, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/gej-ng/
10/41/16/223/27/46/abstract.html.

Mehta:2002:DTA

[762] Dinesh Mehta and Vijay Raghavan. Decision tree approximations of Boolean functions. Theoretical Computer Science, 270(1–2):609– 623, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/47/abstract.html.

Durand:2002:CRH

[763] Arnaud Durand, Miki Hermann, and Laurent Juban. On the complexity of recognizing the Hilbert basis of a linear Diophantine system. Theoretical Computer Science, 270(1–2): 625–642, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/48/abstract.html.

Pergola:2002:AAF

[764] E. Pergola, R. Pinzani, and S. Rinaldi. Approximating algebraic functions by means of rational ones. Theoretical Computer Science, 270(1–2): 643–657, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/49/abstract.html.

Shallit:2002:TSI

[765] Jeffrey Shallit and Ming wei Wang. On two-sided infinite fixed points of morphisms. Theoretical Computer Science, 270(1–2):659–675, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/50/abstract.html.

Kucera:2002:WBB

[766] Antonín Kučera and Richard Mayr. Weak bisimilarity between finite-state systems and BPA or normed BPP is decidable in polynomial time. Theoretical Computer Science, 270(1–2): 677–700, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/51/abstract.html.

Dassow:2002:OLG

[767] Jürgen Dassow, Victor Mitrana, and Arto Salomaa. Operations and language generating devices suggested by the genome evolution. Theoretical Computer Science, 270(1–2): 701–738, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/52/abstract.html.

Zhang:2002:PRT

[768] Lan Zhang, Katsushi Inoue, Akira Ito, and Yue Wang. Probabilistic rebound Turing machines. Theoretical Computer Science, 270(1–2): 739–760, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/53/abstract.html.

Gottlob:2002:CLC

[769] Georg Gottlob, Nicola Leone, and Francesco Scarcello. Computing LOGCFL certificates. Theoretical Computer Science, 270(1–2):761–777, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/54/abstract.html.

Martin-Vide:2002:MSC

[770] Carlos Martín-Vide, Gheorghe Păun, and Grzegorz Rozenberg. Membrane systems with carriers. Theoretical Computer Science, 270(1–2): 779–796, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/55/abstract.html.

Iwamoto:2002:CFC

[771] Chuzo Iwamoto, Tomonobu Hatsuyama, Kenichi Morita, and Katsunobu Imai. Constructible functions in cellular automata and their applications to hierarchy results. Theoretical Computer Science, 270(1–2): 797–809, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/56/abstract.html.

Busch:2002:TCI

[772] Costas Busch, Neophytos Demetriou, Maurice Herlihy, and Marios Mavronicolas. Threshold counters with increments and decrements. Theoretical Computer Science, 270(1–2): 811–826, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/57/abstract.html.

Goles:2002:SLC

[773] Eric Goles, Michel Morvan, and Ha Duong Phan. The structure of a linear chip firing game and related models. Theoretical Computer Science, 270(1–2):827–841, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/58/abstract.html.

Stoye:2002:SFD

[774] Jens Stoye and Dan Gusfield. Simple and flexible detection of contiguous repeats using a suffix tree. Theoretical Computer Science, 270(1–2): 843–856, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/59/abstract.html.

Sander:2002:RCF

[775] J. W. Sander and R. Tijdeman. The rectangle complexity of functions on two-dimensional lattices. Theoretical Computer Science, 270(1–2):857– 863, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/60/abstract.html.

Pluhar:2002:ARG

[776] András Pluhár. The accelerated kin-a-row game. Theoretical Computer Science, 270(1–2):865–875, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/61/abstract.html.

Cicalese:2002:LAO

[777] Ferdinando Cicalese, Daniele Mundici, and Ugo Vaccaro. Least adaptive optimal search with unreliable tests. Theoretical Computer Science, 270(1–2): 877–893, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/62/abstract.html.

Flake:2002:RHP

[778] Gary William Flake and Eric B. Baum. Rush hour is PSPACEcomplete, or “Why you should generously tip parking lot attendants”. Theoretical Computer Science, 270(1–2): 895–911, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/63/abstract.html.

Drmota:2002:VHB

[779] Michael Drmota. The variance of the height of binary search trees. Theoretical Computer Science, 270(1–2): 913–919, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/64/abstract.html.

Akiyama:2002:CNS

[780] Shigeki Akiyama and Attila Pethő. On canonical number systems. Theoretical Computer Science, 270(1–2): 921–933, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/65/abstract.html.

Fiebig:2002:CFC

[781] Doris Fiebig and Ulf-Rainer Fiebig. Compact factors of countable state Markov shifts. Theoretical Computer Science, 270(1–2):935–946, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/66/abstract.html.

Becher:2002:ECA

[782] Verónica Becher and Santiago Figueira. An example of a computable absolutely normal number. Theoretical Computer Science, 270(1–2):947– 958, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/67/abstract.html.

Ewert:2002:PLR

[783] Sigrid Ewert and Andries van der Walt. A pumping lemma for random permitting context languages. Theoretical Computer Science, 270(1–2): 959–967, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/68/abstract.html.

Kim:2002:TUR

[784] Changwook Kim. Two undecidability results for chain code picture languages. Theoretical Computer Science, 270(1–2):969–976, January 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/69/abstract.html.

Anonymous:2002:AIa

[785] Anonymous. Author index. Theoretical Computer Science, 270(1–2): 977–979, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/223/27/70/abstract.html.

Anonymous:2002:MIa

[786] Anonymous. Master index. Theoretical Computer Science, 270(1–2): 981–992, January 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/223/27/71/abstract.html.

Durand:2002:F

[787] Bruno Durand. Foreword. Theoretical Computer Science, 271(1– 2):1, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/220/27/27/abstract.html.

Calude:2002:CCR

[788] Cristian S. Calude. A characterization of c.e. random reals. Theoretical Computer Science, 271(1–2): 3–14, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/220/27/28/abstract.html.

Muchnik:2002:KEC

[789] Andrej A. Muchnik and Semen Ye. Positselsky. Kolmogorov entropy in the context of computability theory. Theoretical Computer Science, 271(1– 2):15–35, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/220/27/29/abstract.html.

Durand:2002:CBC

[790] Bruno Durand and Sylvain Porrot. Comparison between the complexity of a function and the complexity of its graph. Theoretical Computer Science, 271(1–2):37– 46, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/30/abstract.html.

Durand:2002:DCC

[791] Bruno Durand, Alexander Shen, and Nikolai Vereshchagin. Descriptive complexity of computable sequences. Theoretical Computer Science, 271(1–2): 47–58, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/31/abstract.html.

Vereshchagin:2002:KCC

[792] Nikolai K. Vereshchagin. Kolmogorov complexity conditional to large integers. Theoretical Computer Science, 271(1–2):59–67, January 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/32/abstract.html.

Chernov:2002:USL

[793] Alexei Chernov, Andrej Muchnik, Andrei Romashchenko, Alexander Shen, and Nikolai Vereshchagin. Upper semilattice of binary strings with the relation “x is simple conditional to y”. Theoretical Computer Science, 271(1– 2):69–95, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/33/abstract.html.

Muchnik:2002:CCC

[794] Andrej A. Muchnik. Conditional complexity and codes. Theoretical Computer Science, 271(1–2):97– 109, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/220/27/34/abstract.html.

Romashchenko:2002:CIK

[795] A. Romashchenko, A. Shen, and N. Vereshchagin. Combinatorial interpretation of Kolmogorov complexity. Theoretical Computer Science, 271(1– 2):111–123, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/220/27/35/abstract.html.

Shen:2002:LOK

[796] Alexander Shen and Nikolai Vereshchagin. Logical operations and Kolmogorov complexity. Theoretical Computer Science, 271(1–2):125–129, January 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/220/27/36/abstract.html.

Vereshchagin:2002:IML

[797] Nikolai K. Vereshchagin and Michael V. Vyugin. Independent minimum length programs to translate between given strings. Theoretical Computer Science, 271(1–2):131–143, January 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/220/27/37/abstract.html.

Vyugin:2002:IDC

[798] Mikhail V. Vyugin. Information distance and conditional complexities. Theoretical Computer Science, 271(1– 2):145–150, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/38/abstract.html.

Grigorieff:2002:KCN

[799] Serge Grigorieff and Jean-Yves Marion. Kolmogorov complexity and non-determinism. Theoretical Computer Science, 271(1–2):151–180, January 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/39/abstract.html.

Kalnishkan:2002:GLR

[800] Yuri Kalnishkan. General linear relations between different types of predictive complexity. Theoretical Computer Science, 271(1–2):181– 200, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/40/abstract.html.

Anonymous:2002:AIb

[801] Anonymous. Author index. Theoretical Computer Science, 271(1–2): 201, January 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/220/27/41/abstract.html.

Dezani-Ciancaglini:2002:P

[802] Mariangiola Dezani-Ciancaglini, Mitsu Okada, and Masako Takahashi. Preface. Theoretical Computer Science, 272 (1–2):1–2, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/27/abstract.html.

vanBakel:2002:ITT

[803] Steffen van Bakel, Franco Barbanera, Mariangiola Dezani-Ciancaglini, and Fer-Jan de Vries. Intersection types for λ-trees. Theoretical Computer Science, 272(1–2):3–40, February 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/28/abstract.html.

Blanqui:2002:IDT

[804] Frédéric Blanqui, Jean-Pierre Jouannaud, and Mitsuhiro Okada. Inductivedata-type systems. Theoretical Computer Science, 272(1–2):41–68, February 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/29/abstract.html. See

corrigendum [2229].

Coppo:2002:STN

[805] Mario Coppo, Ferruccio Damiani, and Paola Giannini. Strictness, totality, and non-standard-type inference. Theoretical Computer Science, 272(1–2): 69–112, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/30/abstract.html.

Hasegawa:2002:TAA

[806] Ryu Hasegawa. Two applications of analytic functors. Theoretical Computer Science, 272(1–2):113–175, February 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/31/abstract.html.

Hayashi:2002:TAP

[807] Susumu Hayashi, Ryosuke Sumitomo, and Ken ichiro Shii. Towards the animation of proofs — testing proofs by examples. Theoretical Computer Science, 272(1–2):177– 195, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/32/abstract.html.

Ishihara:2002:CIU

[808] Hajime Ishihara and Toshihiko Kurata. Completeness of intersection and union type assignment systems for call-by-value λ-models. Theoretical Computer Science, 272(1–2):197– 221, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/33/abstract.html.

Kameyama:2002:SNN

[809] Yukiyoshi Kameyama and Masahiko Sato. Strong normalizability of the non-deterministic catch/throw calculi. Theoretical Computer Science, 272(1– 2):223–245, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/34/abstract.html.

Ker:2002:IGM

[810] Andrew D. Ker, Hanno Nickau, and C.-H. Luke Ong. Innocent game models of untyped λ-calculus. Theoretical Computer Science, 272(1–2):

247–292, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/222/27/35/abstract.html.

Oostdijk:2002:PCC

[811] Martijn Oostdijk and Herman Geuvers. Proof by computation in the Coq system. Theoretical Computer Science, 272(1–2):293–314, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/41/16/222/27/36/abstract.html.

Uustalu:2002:LGF

[812] Tarmo Uustalu and Varmo Vene. Least and greatest fixed points in intuitionistic natural deduction. Theoretical Computer Science, 272(1–2): 315–339, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/222/27/37/abstract.html.

Yokouchi:2002:CTA

[813] Hirofumi Yokouchi. Completeness of type assignment systems with intersection, union, and type quantifiers. Theoretical Computer Science, 272(1–2): 341–398, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/222/27/38/abstract.html.

Anonymous:2002:AIc

[814] Anonymous. Author index. Theoretical Computer Science, 272(1–2): 399, February 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/222/27/39/abstract.html.

Neraud:2002:P

[815] J. Neraud.  Preface.  Theoreti-

cal Computer Science, 273(1–2):1– 3, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/27/abstract.html.

Brown:2002:ASS

[816] Tom C. Brown. Applications of standard Sturmian words to elementary number theory. Theoretical Computer Science, 273(1–2):5– 9, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/28/abstract.html.

DAlessandro:2002:CPT

[817] Flavio D’Alessandro. A combinatorial problem on Trapezoidal words. Theoretical Computer Science, 273(1–2): 11–33, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/29/abstract.html.

Heinis:2002:FBI

[818] Alex Heinis. The P (n)/n-function for bi-infinite words. Theoretical Computer Science, 273(1–2):35–46, February 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/30/abstract.html.

Berstel:2002:SFU

[819] Jean Berstel and Luc Boasson. Shuffle factorization is unique. Theoretical Computer Science, 273(1–2): 47–67, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/226/27/31/abstract.html.

Choffrut:2002:CFS

[820] Christian Choffrut, Juhani Karhumäki, and Nicolas Ollinger. The commutation of finite sets: a challenging problem. Theoretical Computer Science, 273(1– 2):69–79, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/226/27/32/abstract.html.

Karhumaki:2002:MFW

[821] Juhani Karhumäki and Ján Maňuch. Multiple factorizations of words and defect effect. Theoretical Computer Science, 273(1–2):81–97, February 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/226/27/33/abstract.html.

Mignosi:2002:WFF

[822] F. Mignosi, A. Restivo, and M. Sciortino. Words and forbidden factors. Theoretical Computer Science, 273(1–2): 99–117, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/226/27/34/abstract.html.

Richomme:2002:SRP

[823] G. Richomme and F. Wlazinski. Some results on k-power-free morphisms. Theoretical Computer Science, 273 (1–2):119–142, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/gej-ng/
10/41/16/226/27/35/abstract.html.

deLuca:2002:SCR

[824] Aldo de Luca. Some combinatorial results on Bernoulli sets and codes. Theoretical Computer Science, 273(1–2): 143–165, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/36/abstract.html.

Guesnet:2002:MCF

[825] Yannick Guesnet. On maximal codes with a finite interpreting delay. Theoretical Computer Science, 273(1–2): 167–183, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/37/abstract.html.

Neraud:2002:LCS

[826] Jean Néraud and Carla Selmi. Locally complete sets and finite decomposable codes. Theoretical Computer Science, 273(1–2):185–196, February 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/38/abstract.html.

Berthe:2002:BPM

[827] Valérie Berthé and Robert Tijdeman. Balance properties of multidimensional words. Theoretical Computer Science, 273(1–2):197–224, February 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/39/abstract.html.

Esik:2002:ASS

[828] Zoltán Ésik. Axiomatizing the subsumption and subword preorders on finite and infinite partial words. Theoretical Computer Science, 273(1–2): 225–248, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/40/abstract.html.

Mauduit:2002:FIP

[829] Christian Mauduit. Finite and infinite pseudorandom binary words. Theoretical Computer Science, 273(1–2): 249–261, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/41/abstract.html.

Malyshev:2002:QEW

[830] V. Malyshev. Quantum evolution of words. Theoretical Computer Science, 273(1–2):263–269, February 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/42/abstract.html.

Bigotte:2002:LWS

[831] M. Bigotte, G. Jacob, N. E. Oussous, and M. Petitot. Lyndon words and shuffle algebras for generating the coloured multiple zeta values relations tables. Theoretical Computer Science, 273(1–2):271– 282, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/43/abstract.html.

Duchamp:2002:SGL

[832] G. Duchamp, É. Laugerotte, and J.-G. Luque. On the support of graph Lie algebras. Theoretical Computer Science, 273(1–2):283– 294, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/44/abstract.html.

Encheva:2002:FCA

[833] Sylvia Encheva and Gérard Cohen. Frameproof codes against limited coalitions of pirates. Theoretical Computer Science, 273(1–2):295– 304, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/45/abstract.html.

Anonymous:2002:AId

[834] Anonymous. Author index. Theoretical Computer Science, 273(1–2): 305, February 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/226/27/46/abstract.html.

Sangiorgi:2002:E

[835] Davide Sangiorgi and Robert de Simone. Editorial. Theoretical Computer Science, 274(1–2):1, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/41/16/227/27/27/abstract.html.

deBoer:2002:HLD

[836] F. S. de Boer. A Hoare logic for dynamic networks of asynchronously communicating deterministic processes. Theoretical Computer Science, 274(1– 2):3–41, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/227/27/28/abstract.html.

Hermanns:2002:PAP

[837] Holger Hermanns, Ulrich Herzog, and Joost-Pieter Katoen. Process algebra for performance evaluation. Theoretical Computer Science, 274(1–2): 43–87, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/227/27/29/abstract.html.

Lugiez:2002:RVP

[838] D. Lugiez and Ph. Schnoebelen. The regular viewpoint on PA-processes. Theoretical Computer Science, 274(1– 2):89–115, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/227/27/30/abstract.html.

Madhusudan:2002:BTC

[839] P. Madhusudan and P. S. Thiagarajan. Branching time controllers for discrete event systems. Theoretical Computer Science, 274(1–2): 117–149, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/227/27/31/abstract.html.

Schobbens:2002:ART

[840] P.-Y. Schobbens, J.-F. Raskin, and T. A. Henzinger. Axioms for realtime logics. Theoretical Computer Science, 274(1–2):151–182, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/gej-ng/
10/41/16/227/27/32/abstract.html.

Sewell:2002:RRB

[841] Peter Sewell. From rewrite rules to bisimulation congruences. Theoretical Computer Science, 274(1–2): 183–230, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/227/27/33/abstract.html.

Yoshida:2002:MSR

[842] Nobuko Yoshida. Minimality and separation results on asynchronous mobile processes — representability theorems by concurrent combinators. Theoretical Computer Science, 274(1–2): 231–276, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/227/27/34/abstract.html.

Anonymous:2002:AIVa

[843] Anonymous. Author index volume 274 (2002). Theoretical Computer Science, 274(1–2):277, March 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/227/27/35/abstract.html.

Ying:2002:BIT

[844] Mingsheng Ying. Bisimulation indexes and their applications. Theoretical Computer Science, 275(1–2): 1–68, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/27/abstract.html.

Burdick:2002:NID

[845] Bruce S. Burdick. A note on iterated duals of certain topological spaces. Theoretical Computer Science, 275(1– 2):69–77, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/28/abstract.html.

Manes:2002:TMT

[846] Ernest G. Manes. Taut Monads and T0-spaces. Theoretical Computer Science, 275(1–2):79–109, March 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/29/abstract.html.

Rabinovich:2002:FVI

[847] A. Rabinovich. Finite variability interpretation of monadic logic of order.

Theoretical Computer Science, 275(1– 2):111–125, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/30/abstract.html.

Busi:2002:AIP

[848] Nadia Busi. Analysis issues in Petri nets with inhibitor arcs. Theoretical Computer Science, 275(1–2):127– 177, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/31/abstract.html.

Abiteboul:2002:CTH

[849] Serge Abiteboul, Sophie Cluet, and Tova Milo. Correspondence and translation for heterogeneous data. Theoretical Computer Science, 275(1–2): 179–213, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/32/abstract.html.

Duggan:2002:OCS

[850] Dominic Duggan and John Ophel. Open and closed scopes for constrained genericity. Theoretical Computer Science, 275(1–2):215–258, March 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/33/abstract.html.

Degano:2002:CSC

[851] Pierpaolo Degano, Fabio Gadducci, and Corrado Priami. A causal semantics for CCS via rewriting logic. Theoretical Computer Science, 275(1–2):

259–282, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/34/abstract.html.

Stewart:2002:PSA

[852] Iain A. Stewart. Program schemes, arrays, Lindström quantifiers and zeroone laws. Theoretical Computer Science, 275(1–2):283–310, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/gej-ng/10/41/16/246/27/35/abstract.
                                        html.

Walukiewicz:2002:MSO

[853] Igor Walukiewicz. Monadic secondorder logic on tree-like structures. Theoretical Computer Science, 275(1–2): 311–346, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/36/abstract.html.

Beauquier:2002:DVR

[854] Danièle Beauquier and Anatol Slissenko. Decidable verification for reducible timed automata specified in a first order logic with time. Theoretical Computer Science, 275(1–2): 347–388, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/37/abstract.html.

Sessa:2002:ARS

[855] Maria I. Sessa. Approximate reasoning by similarity-based SLD resolution. Theoretical Computer Science, 275(1– 2):389–426, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/38/abstract.html.

Adamek:2002:ADT

[856] J. Adámek, M. Hébert, and J. Rosický. On abstract data types presented by multiequations. Theoretical Computer Science, 275(1–2):427–462, March 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/39/abstract.html.

Vorobyov:2002:ETC

[857] Sergei Vorobyov. ∀∃^{5}-equational theory of context unification is undecidable. Theoretical Computer Science, 275(1– 2):463–479, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/40/abstract.html.

Ying:2002:AMP

[858] Mingsheng Ying. Additive models of probabilistic processes. Theoretical Computer Science, 275(1–2): 481–519, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/41/abstract.html.

Cortesi:2002:OAS

[859] Agostino Cortesi, Agostino Dovier, Elisa Quintarelli, and Letizia Tanca. Operational and abstract semantics of the query language G-Log. Theoretical Computer Science, 275(1–2): 521–560, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/42/abstract.html.

Zimmermann:2002:PRN

[860] Ernst Zimmermann. Peirce’s Rule in Natural Deduction. Theoretical Computer Science, 275(1–2):561– 574, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/43/abstract.html.

Cardone:2002:CCP

[861] Felice Cardone. A coinductive completeness proof for the equivalence of recursive types. Theoretical Computer Science, 275(1–2): 575–587, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/44/abstract.html.

Vogler:2002:EAS

[862] Walter Vogler. Efficiency of asynchronous systems, read arcs, and the MUTEX-problem. Theoretical Computer Science, 275(1–2):589– 631, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/246/27/45/abstract.html.

Neven:2002:QAF

[863] Frank Neven and Thomas Schwentick. Query automata over finite trees. Theoretical Computer Science, 275(1–2): 633–674, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/46/abstract.html.

Bouhoula:2002:OPR

[864] Adel Bouhoula and Michaël Rusinowitch. Observational proofs by rewriting. Theoretical Computer Science, 275(1– 2):675–698, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/47/abstract.html.

Anonymous:2002:AIVb

[865] Anonymous. Author index volume 275 (2002). Theoretical Computer Science, 275(1–2):699, March 28, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/246/27/48/abstract.html.

Cori:2002:PIS

[866] Robert Cori, Dominique Rossin, and Bruno Salvy. Polynomial ideals for sandpiles and their Gröbner bases. Theoretical Computer Science, 276(1– 2):1–15, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/27/abstract.html.

Bouchitte:2002:LAP

[867] Vincent Bouchitté and Ioan Todinca. Listing all potential maximal cliques of a graph. Theoretical Computer Science, 276(1–2):17–32, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/28/abstract.html.

Calinescu:2002:STP

[868] Gruia Călinescu and Peng-Jun Wan. Splittable traffic partition in WDM/ SONET rings to minimize SONET ADMs. Theoretical Computer Science, 276(1–2):33–50, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/29/abstract.html.

Droste:2002:AEA

[869] Stefan Droste, Thomas Jansen, and Ingo Wegener. On the analysis of the (1 + 1) evolutionary algorithm. Theoretical Computer Science, 276(1– 2):51–81, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/30/abstract.html.

Settle:2002:SSF

[870] Amber Settle and Janos Simon. Smaller solutions for the firing squad. Theoretical Computer Science, 276(1– 2):83–109, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/31/abstract.html.

Ma:2002:CSB

[871] Bin Ma, Lusheng Wang, and Kaizhong Zhang. Computing similarity between RNA structures. Theoretical Computer Science, 276(1–2):111– 132, April 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/32/abstract.html.

Maitra:2002:CSB

[872] Subhamoy Maitra and Palash Sarkar. Cryptographically significant Boolean functions with five valued Walsh spectra. Theoretical Computer Science, 276 (1–2):133–146, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/33/abstract.html.

Nishimura:2002:CCU

[873] Harumichi Nishimura and Masanao Ozawa. Computational complexity of uniform quantum circuit families and quantum Turing machines. Theoretical Computer Science, 276(1–2): 147–181, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/34/abstract.html. Com-

municated by O. Watanabe.

Halava:2002:BGP

[874] Vesa Halava, Tero Harju, and Mika Hirvensalo. Binary (generalized) Post Correspondence Problem. Theoretical Computer Science, 276(1–2): 183–204, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/35/abstract.html.

Csuhaj-Varju:2002:PCG

[875] Erzsébet Csuhaj-Varjú and György Vaszil. Parallel communicating grammar systems with bounded resources. Theoretical Computer Science, 276(1– 2):205–219, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/36/abstract.html.

Fulop:2002:IRT

[876] Z. Fülöp and A. Terlutte. Iterated relabeling tree transducers. Theoretical Computer Science, 276(1–2): 221–244, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/37/abstract.html.

Salomaa:2002:DES

[877] Kai Salomaa and Sheng Yu. Decidability of EDT0L structural equivalence. Theoretical Computer Science, 276(1– 2):245–259, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/38/abstract.html.

Manlove:2002:HVS

[878] David F. Manlove, Robert W. Irving, Kazuo Iwama, Shuichi Miyazaki, and Yasufumi Morita. Hard variants of stable marriage. Theoretical Computer Science, 276(1–2): 261–279, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/39/abstract.html.

Justin:2002:EWE

[879] Jacques Justin and Giuseppe Pirillo. Episturmian words and episturmian morphisms. Theoretical Computer Science, 276(1–2):281–313, April 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/40/abstract.html.

Cicerone:2002:SDL

[880] Serafino Cicerone, Gabriele Di Stefano, and Michele Flammini. Static and dynamic low-congested interval routing schemes. Theoretical Computer Science, 276(1–2):315– 354, April 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/41/abstract.html.

Konstantinidis:2002:EDP

[881] Stavros Konstantinidis and Amber O’Hearn. Error-detecting properties of languages. Theoretical Computer Science, 276(1–2):355–375, April 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/42/abstract.html.

Fernau:2002:SGA

[882] Henning Fernau and Ralf Stiebe. Sequential grammars and automata with valences. Theoretical Computer Science, 276(1–2):377–405, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/gej-ng/
10/41/16/247/27/43/abstract.html.

Vyugin:2002:NDS

[883] V. V. V’yugin. Note: Does snooping help? Theoretical Computer Science, 276(1–2):407–415, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/41/16/247/27/44/abstract.html.

Chen:2002:NNC

[884] Chiuyuan Chen. Note: A necessary condition for a graph to be the visibility graph of a simple polygon. Theoretical Computer Science, 276(1– 2):417–424, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/45/abstract.html.

Malajovich:2002:NLB

[885] Gregorio Malajovich. Note: Lower bounds for some decision problems over C. Theoretical Computer Science, 276 (1–2):425–434, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/46/abstract.html.

Robson:2002:NCB

[886] J. M. Robson. Note: Constant bounds on the moments of the height of binary search trees. Theoretical Computer Science, 276(1–2):435– 444, April 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/41/
                                        16/247/27/47/abstract.html.

Prieur:2002:NHD

[887] Christophe Prieur. Note: How to decide continuity of rational functions on infinite words. Theoretical Computer Science, 276(1–2):445– 447, April 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/48/abstract.html.

Wang:2002:NCT

[888] Yongge Wang. Note: A comparison of two approaches to pseudorandomness. Theoretical Computer Science, 276(1– 2):449–459, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/gej-ng/10/41/
16/247/27/49/abstract.html.

Anonymous:2002:AIVc

[889] Anonymous. Author index volume 276 (2002). Theoretical Computer Science, 276(1–2):465–466, April 6, 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/gej-ng/
10/41/16/247/27/50/abstract.html.

VanHentenryck:2002:E

[890] Pascal Van Hentenryck. Editorial. Theoretical Computer Science, 277(1– 2):1–2, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bagnara:2002:SSR

[891] Roberto Bagnara, Patricia M. Hill, and Enea Zaffanella. Set-sharing is redundant for pair-sharing. Theoretical Computer Science, 277(1–2):3–46, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cousot:2002:CDH

[892] Patrick Cousot. Constructive design of a hierarchy of semantics of a transition system by abstract interpretation. Theoretical Computer Science, 277(1– 2):47–103, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frey:2002:SSI

[893] Alexandre Frey. Satisfying subtype inequalities in polynomial space. Theoretical Computer Science, 277(1–2): 105–117, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ramalingam:2002:SER

[894] G. Ramalingam. On sparse evaluation representations. Theoretical Computer Science, 277(1–2):119–147, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Scozzari:2002:LOG

[895] Francesca Scozzari. Logical optimality of groundness analysis. Theoretical Computer Science, 277(1–2): 149–184, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yi:2002:CEE

[896] Kwangkeun Yi and Sukyoung Ryu. A cost-effective estimation of uncaught exceptions in Standard ML programs. Theoretical Computer Science, 277(1– 2):185–217, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIVd

[897] Anonymous. Author index volume 277 (2002). Theoretical Computer Science, 277(1–2):223, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brookes:2002:F

[898] Stephen Brookes and Michael Mislove. Foreword. Theoretical Computer Science, 278(1–2):1–2, April 2002. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freyd:2002:CL

[899] Peter Freyd. Cartesian logic. Theoretical Computer Science, 278(1–2): 3–21, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brooks:2002:D

[900] Stephen Brooks and Michael Mislove. Dedication. Theoretical Computer Science, 278(1–2):23, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fitting:2002:FSL

[901] Melvin Fitting. Fixpoint semantics for logic programming a survey. Theoretical Computer Science, 278(1–2): 25–51, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hennessy:2002:FAD

[902] Matthew Hennessy. A fully abstract denotational semantics for the π-calculus. Theoretical Computer Science, 278(1– 2):53–89, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bucciarelli:2002:RDB

[903] Antonio Bucciarelli and Pasquale Malacaria. Relative definability of boolean functions via hypergraphs.

Theoretical Computer Science, 278(1– 2):91–110, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiech:2002:PLC

[904] Adrian Fiech and David A. Schmidt. Polymorphic lambda calculus and subtyping. Theoretical Computer Science, 278(1–2):111–140, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flagg:2002:EIC

[905] Robert C. Flagg and Philipp Sünderhauf. The essence of ideal completion in quantitative form. Theoretical Computer Science, 278(1–2):141–158, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kunzi:2002:YCQ

[906] H. P. Künzi and M. P. Schellekens. On the Yoneda completion of a quasimetric space. Theoretical Computer Science, 278(1–2):159–194, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gastin:2002:RTD

[907] Paul Gastin and Dan Teodosiu. Resource traces: a domain for processes sharing exclusive resources. Theoretical Computer Science, 278(1–2): 195–221, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gupta:2002:TCC

[908] V. Gupta, R. Jagadeesan, and V. A. Saraswat. Truly concurrent constraint programming. Theoretical Computer Science, 278(1–2):223–255, April 2002.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mulry:2002:LRC

[909] P. S. Mulry. Lifting results for categories of algebras. Theoretical Computer Science, 278(1–2):257–269, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Naumann:2002:SDR

[910] David A. Naumann. Soundness of data refinement for a higher-order imperative language. Theoretical Computer Science, 278(1–2):271–301, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Power:2002:PCC

[911] John Power. Premonoidal categories as categories with algebraic structure. Theoretical Computer Science, 278(1– 2):303–321, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Power:2002:FOD

[912] John Power and Giuseppe Rosolini. Fixpoint operators for domain equations. Theoretical Computer Science, 278(1–2):323–333, April 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIVe

[913] Anonymous. Author index volume 278 (2002). Theoretical Computer Science, 278(1–2):335, April 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chesneaux:2002:FRN

[914] Jean-Marie Chesneaux, Christiane Frougny, and Jean-Michel Muller. Foreword: Real numbers. Theoretical Computer Science, 279(1–2):1–2, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Etiemble:2002:CAH

[915] Daniel Etiemble. Computer arithmetic and hardware: “off the shelf” microprocessors versus “custom hardware”. Theoretical Computer Science, 279(1– 2):3–27, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Turner:2002:RPS

[916] Peter R. Turner. Residue polynomial systems. Theoretical Computer Science, 279(1–2):29–49, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lester:2002:CAF

[917] David Lester, Scott Chambers, and Heoi Lee Lu. A constructive algorithm for finding the exact roots of polynomials with computable real coefficients. Theoretical Computer Science, 279(1– 2):51–64, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heckmann:2002:CLF

[918] Reinhold Heckmann. Contractivity of linear fractional transformations. Theoretical Computer Science, 279(1–2): 65–82, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hertling:2002:LBR

[919] Peter Hertling. A lower bound for range enclosure in interval arithmetic. Theoretical Computer Science, 279(1– 2):83–95, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIVf

[920] Anonymous. Author index volume 279 (2002). Theoretical Computer Science, 279(1–2):97, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacobs:2002:F

[921] Bart Jacobs and Jan Rutten. Foreword. Theoretical Computer Science, 280(1–2):1, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wolter:2002:CPA

[922] Uwe Wolter. CSP, partial automata, and coalgebras. Theoretical Computer Science, 280(1–2):3–34, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cirstea:2002:CEA

[923] Corina Cîrstea. A coalgebraic equational approach to specifying observational structures. Theoretical Computer Science, 280(1–2):35–68, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kurz:2002:IMC

[924] Alexander Kurz and Rolf Hennicker. On institutions for modular coalgebraic specifications. Theoretical Computer Science, 280(1–2):69–103, May 2002.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pavlovi:2002:CFC

[925] D. Pavlovi and V. Pratt. The continuum as a final coalgebra. Theoretical Computer Science, 280(1–2): 105–122, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lasota:2002:CMS

[926] Sławomir Lasota. Coalgebra morphisms subsume open maps. Theoretical Computer Science, 280(1–2): 123–135, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Power:2002:CMC

[927] John Power and Hiroshi Watanabe. Combining a monad and a comonad. Theoretical Computer Science, 280(1– 2):137–162, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corradini:2002:CSB

[928] Andrea Corradini, Reiko Heckel, and Ugo Montanari. Compositional SOS and beyond: a coalgebraic view of open systems. Theoretical Computer Science, 280(1–2):163–192, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIe

[929] Anonymous. Author index. Theoretical Computer Science, 280(1–2):193, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:MIb

[930] Anonymous. Master index. Theoretical Computer Science, 280(1–2): 195–200, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ausiello:2002:P

[931] G. Ausiello. Preface. Theoretical Computer Science, 281(1–2):1, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Curien:2002:BBS

[932] Pierre-Louis Curien. Une brève biographie scientifique de Maurice Nivat. (French) [A brief scientific biography of Maurice Nivat]. Theoretical Computer Science, 281(1–2):3–23, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2002:IEM

[933] Grzegorz Rozenberg and Arto Salomaa. ICALP, EATCS and Maurice Nivat. Theoretical Computer Science, 281 (1–2):25–30, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:2002:NPT

[934] André Arnold. Nivat’s processes and their synchronization. Theoretical Computer Science, 281(1–2):31–36, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Banderier:2002:BAC

[935] Cyril Banderier and Philippe Flajolet. Basic analytic combinatorics of directed lattice paths. Theoretical Computer Science, 281(1–2):37–80, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:2002:GT

[936] Danièle Beauquier and Jean-Claude Fournier. Groups and tilings. Theoretical Computer Science, 281(1–2): 81–97, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berstel:2002:CRI

[937] Jean Berstel and Laurent Vuillon. Coding rotations on intervals. Theoretical Computer Science, 281(1–2): 99–107, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boudol:2002:NCP

[938] Gérard Boudol and Ilaria Castellani. Noninterference for concurrent programs and thread systems. Theoretical Computer Science, 281(1–2): 109–130, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruni:2002:DCC

[939] Roberto Bruni and Ugo Montanari. Dynamic connectors for concurrency. Theoretical Computer Science, 281(1– 2):131–176, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:2002:EFO

[940] Bruno Courcelle and Teodor Knapik. The evaluation of first-order substitution is monadic second-order compatible. Theoretical Computer Science, 281(1–2):177–206, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cousineau:2002:TPE

[941] Guy Cousineau. Tilings as a programming exercise. Theoretical Computer Science, 281(1–2):207–217, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dauchet:2002:RNL

[942] Max Dauchet, Sophie Tison, and Marc Tommasi. Réduction de la non-linéarité des morphismes d’arbres. (French) [Recognizable tree-languages and non-linear morphisms]. Theoretical Computer Science, 281(1–2): 219–233, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lungo:2002:RPM

[943] Alberto Del Lungo. Reconstructing permutation matrices from diagonal sums. Theoretical Computer Science, 281(1–2):235–249, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delorme:2002:RPL

[944] Marianne Delorme and Jacques Mazoyer. Reconnaissance parallèle des langages rationnels sur automates cellulaires plans. (French) [Parallel recognition of rational languages on planar cellular automata]. Theoretical Computer Science, 281(1–2):251–289, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2002:CCP

[945] Josep Díaz, Maria Serna, and Dimitrios M. Thilikos. Counting Hcolorings of partial k-trees. Theoretical Computer Science, 281(1–2): 291–309, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durand:2002:LPF

[946] Bruno Durand. De la logique aux pavages. (French) [On the logic of tiling]. Theoretical Computer Science, 281(1–2):311–324, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:2002:GAT

[947] A. Ehrenfeucht, T. Harju, and G. Rozenberg. Gene assembly through cyclic graph decomposition. Theoretical Computer Science, 281(1–2): 325–349, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferrari:2002:ACS

[948] Luca Ferrari, Elisa Pergola, Renzo Pinzani, and Simone Rinaldi. An algebraic characterization of the set of succession rules. Theoretical Computer Science, 281(1–2):351–367, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gastin:2002:TCS

[949] Paul Gastin and Michael Mislove. A truly concurrent semantics for a process algebra using resource pomsets. Theoretical Computer Science, 281(1– 2):369–421, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grigorieff:2002:MDR

[950] Serge Grigorieff. Modelization of deterministic rational relations. Theoretical Computer Science, 281(1–2): 423–453, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gritzmann:2002:AID

[951] Peter Gritzmann and Sven de Vries. On the algorithmic inversion of the discrete Radon transform. Theoretical Computer Science, 281(1–2):455–469, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okada:2002:USP

[952] Mitsuhiro Okada. A uniform semantic proof for cut-elimination and completeness of various first and higher order logics. Theoretical Computer Science, 281(1–2):471–498, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rus:2002:ULP

[953] Teodor Rus. A unified language processing methodology. Theoretical Computer Science, 281(1–2):499–536, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2002:UTW

[954] Arto Salomaa. Uni-transitional Watson–Crick D0L systems. Theoretical Computer Science, 281(1–2): 537–553, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Senizergues:2002:BSD

[955] Géraud Sénizergues. L(A)=L(B)? A simplified decidability proof. Theoretical Computer Science, 281(1–2): 555–608, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huet:2002:SYG

[956] Gérard Huet. Śrī Yantra geometry. Theoretical Computer Science, 281(1– 2):609–628, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIf

[957] Anonymous. Author index. Theoretical Computer Science, 281(1–2): 629–630, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Katoen:2002:F

[958] Joost-Pieter Katoen. Foreword. Theoretical Computer Science, 282(1):1–3, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bravetti:2002:TIG

[959] Mario Bravetti and Roberto Gorrieri. The theory of interactive generalized semi-Markov processes. Theoretical Computer Science, 282(1):5–32, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonsson:2002:TPP

[960] Bengt Jonsson and Wang Yi. Testing preorders for probabilistic processes can be characterized by simulations. Theoretical Computer Science, 282(1):33–51, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kolano:2002:PAR

[961] Paul Z. Kolano. Proof assistance for real-time systems using an interactive theorem prover. Theoretical Computer Science, 282(1):53–99, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kwiatkowska:2002:AVR

[962] Marta Kwiatkowska, Gethin Norman, Roberto Segala, and Jeremy Sproston. Automatic verification of real-time systems with discrete probability distributions. Theoretical Computer Science, 282(1):101–150, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lermer:2002:FMR

[963] Karl Lermer and Colin Fidge. A formal model of real-time program compilation. Theoretical Computer Science, 282(1):151–190, May 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McIver:2002:QPL

[964] A. K. McIver. Quantitative program logic and expected time bounds in probabilistic distributed algorithms. Theoretical Computer Science, 282(1): 191–219, May 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lodi:2002:F

[965] Elena Lodi, Linda Pagli, and Nicola Santoro. Foreword. Theoretical Computer Science, 282(2):221–222, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luccio:2002:ANS

[966] F. Luccio. Algorithms, nymphs, and shepherds. Theoretical Computer Science, 282(2):223–229, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peleg:2002:LMC

[967] David Peleg. Local majorities, coalitions and monopolies in graphs: a review. Theoretical Computer Science, 282(2):231–257, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boldi:2002:MJ

[968] Paolo Boldi, Massimo Santini, and Sebastiano Vigna. Measuring with jugs. Theoretical Computer Science, 282(2): 259–270, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fraenkel:2002:ANS

[969] Aviezri S. Fraenkel. Arrays, numeration systems and Frankenstein games. Theoretical Computer Science, 282(2): 271–284, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pedrotti:2002:PST

[970] Alberto Pedrotti. Playing by searching: two strategies against a linearly bounded liar. Theoretical Computer Science, 282(2):285–302, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Itai:2002:PCP

[971] Alon Itai, Michael Rodeh, and Hadas Shachnai. The passport control problem or how to keep a dynamic service system load balanced? Theoretical Computer Science, 282(2):303–318, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daurat:2002:CP

[972] Alain Daurat, Yan Gérard, and Maurice Nivat. The chords’ problem. Theoretical Computer Science, 282(2): 319–336, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Merlini:2002:SLT

[973] Donatella Merlini, Renzo Sprugnoli, and M. Cecilia Verri. A strip-like tiling algorithm. Theoretical Computer Science, 282(2):337–352, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becker:2002:SAC

[974] Ronald Becker, Bruno Simeone, and Yen-I Chiang. A shifting algorithm for continuous tree partitioning. Theoretical Computer Science, 282(2):353–380, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seiden:2002:MCM

[975] Steve Seiden. A manifesto for the computational method. Theoretical Computer Science, 282(2):381–395, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIg

[976] Anonymous. Author index. Theoretical Computer Science, 282(2):397, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertrand:2002:P

[977] Gilles Bertrand and Rémy Malgouyres. Preface. Theoretical Computer Science, 283(1):1, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kong:2002:TAR

[978] T. Y. Kong. Topological adjacency relations on Z^{n}. Theoretical Computer Science, 283(1):3–28, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ayala:2002:WLF

[979] R. Ayala, E. Domínguez, A. R. Francés, and A. Quintero. Weak lighting functions and strong 26-surfaces. Theoretical Computer Science, 283(1):29–66, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malgouyres:2002:CCR

[980] Rémy Malgouyres and Malika More. On the computational complexity of reachability in 2D binary images and some basic problems of 2D digital topology. Theoretical Computer Science, 283(1):67–108, June 2002. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fourey:2002:INT

[981] Sébastien Fourey and Rémy Malgouyres. Intersection number and topology preservation within digital surfaces. Theoretical Computer Science, 283(1):109–150, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brimkov:2002:GPL

[982] Valentin E. Brimkov and Reneta P. Barneva. Graceful planes and lines. Theoretical Computer Science, 283(1): 151–170, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gerard:2002:PGC

[983] Yan Gérard. Periodic graphs and connectivity of the rational digital hyperplanes. Theoretical Computer Science, 283(1):171–182, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Col:2002:ALC

[984] Marie-Andrée Jacob-Da Col. About local configurations in arithmetic planes. Theoretical Computer Science, 283(1): 183–201, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Devillers:2002:RVD

[985] Olivier Devillers and Pierre-Marie Gandoin. Rounding Voronoi diagram. Theoretical Computer Science, 283(1): 203–221, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuba:2002:RCD

[986] Attila Kuba and Emese Balogh. Reconstruction of convex 2D discrete sets in polynomial time. Theoretical Computer Science, 283(1):223–242, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tajine:2002:TPH

[987] Mohamed Tajine and Christian Ronse. Topological properties of Hausdorff discretization, and comparison to other discretization schemes. Theoretical Computer Science, 283(1):243–268, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gorrieri:2002:E

[988] Roberto Gorrieri. Editorial. Theoretical Computer Science, 283(2):269–270, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodei:2002:PAP

[989] Chiara Bodei, Pierpaolo Degano, Riccardo Focardi, and Corrado Priami. Primitives for authentication in process algebras. Theoretical Computer Science, 283(2):271–304, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Denney:2002:CJC

[990] Ewen Denney and Thomas Jensen. Correctness of Java card method lookup via logical relations. Theoretical Computer Science, 283(2):305–331, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guttman:2002:ATS

[991] Joshua D. Guttman and F. Javier Thayer. Authentication tests and the structure of bundles. Theoretical Computer Science, 283(2):333–380, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nielson:2002:VFU

[992] Flemming Nielson, Hanne Riis Nielson, and René Rydhof Hansen. Validating firewalls using flow logics. Theoretical Computer Science, 283(2):381–418, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shmatikov:2002:FSA

[993] Vitaly Shmatikov and John C. Mitchell. Finite-state analysis of two contract signing protocols. Theoretical Computer Science, 283(2):419–450, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIh

[994] Anonymous. Author index. Theoretical Computer Science, 283(2):451, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fischer:2002:F

[995] Paul Fischer, Hans Simon, and Carl Smith. Foreword. Theoretical Computer Science, 284(1):1–2, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Antos:2002:LBR

[996] András Antos. Lower bounds for the rate of convergence in nonparametric pattern recognition. Theoretical Computer Science, 284(1):3–24, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fine:2002:QCL

[997] Shai Fine, Ran Gilad-Bachrach, and Eli Shamir. Query by committee, linear separation and random walks. Theoretical Computer Science, 284(1): 25–51, June 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartlett:2002:HRN

[998] Peter L. Bartlett and Shai Ben-David. Hardness results for neural network approximation problems. Theoretical Computer Science, 284(1):53–66, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duffy:2002:GAL

[999] Nigel Duffy and David Helmbold. A geometric approach to leveraging weak learners. Theoretical Computer Science, 284(1):67–108, June 2002. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Helmbold:2002:DIA

[1000] D. P. Helmbold, S. Panizza, and M. K. Warmuth. Direct and indirect algorithms for on-line learning of disjunctions. Theoretical Computer Science, 284(1):109–142, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2002:MCC

[1001] Sanjay Jain and Arun Sharma. Mind change complexity of learning logic programs. Theoretical Computer Science, 284(1):143–160, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ott:2002:ACT

[1002] Matthias Ott and Frank Stephan.

Avoiding coding tricks by hyperrobust learning. Theoretical Computer Science, 284(1):161–180, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pinter:2002:RCE

[1003] Márta Pintér. On the rate of convergence of error estimates for the partitioning classification rule. Theoretical Computer Science, 284(1):181–196, June 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ko:2002:F

[1004] Ker l Ko, Anil Nerode, and Klaus Weihrauch. Foreword. Theoretical Computer Science, 284(2):197, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blaser:2002:UCC

[1005] Markus Bläser. Uniform computational complexity of the derivatives of C-functions. Theoretical Computer Science, 284(2):199–206, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanck:2002:DRP

[1006] J. Blanck, V. Stoltenberg-Hansen, and J. V. Tucker. Domain representations of partial functions, with applications to spatial objects and constructive volume geometry. Theoretical Computer Science, 284(2):207–240, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brattka:2002:TPR

[1007] Vasco Brattka and Peter Hertling. Topological properties of real number representations. Theoretical Computer Science, 284(2):241–257, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bridges:2002:KSC

[1008] Douglas S. Bridges and Nicholas Dudley Ward. Kernels of seminorms in constructive analysis. Theoretical Computer Science, 284(2):259–267, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calude:2002:CNS

[1009] Cristian S. Calude. Chaitin numbers, Solovay machines, and Gödel incompleteness. Theoretical Computer Science, 284(2):269–277, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cenzer:2002:ECS

[1010] Douglas Cenzer and Jeffrey B. Remmel. Effectively closed sets and graphs of computable real functions. Theoretical Computer Science, 284(2):279–318, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edalat:2002:FCS

[1011] Abbas Edalat and André Lieutier. Foundation of a computable solid modelling. Theoretical Computer Science, 284(2):319–345, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemmerling:2002:EMS

[1012] Armin Hemmerling. Effective metric spaces and representations of the reals. Theoretical Computer Science, 284 (2):347–372, July 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Konecny:2002:RFC

[1013] Michal Konečný. Real functions computable by finite automata using affine representations. Theoretical Computer Science, 284(2):373–396, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krupski:2002:ESA

[1014] Vladimir N. Krupski. Effective simultaneous approximability of reals. Theoretical Computer Science, 284(2): 397–417, July 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mori:2002:CWF

[1015] Takakazu Mori. On the computability of Walsh functions. Theoretical Computer Science, 284(2):419–436, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Normann:2002:ERN

[1016] Dag Normann. Exact real number computations relative to hereditarily total functionals. Theoretical Computer Science, 284(2):437–453, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Staiger:2002:KCR

[1017] Ludwig Staiger. The Kolmogorov complexity of real numbers. Theoretical Computer Science, 284(2):455–466, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsuiki:2002:RNC

[1018] Hideki Tsuiki. Real number computation through Gray code embedding. Theoretical Computer Science, 284(2): 467–485, July 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yoshikawa:2002:IFC

[1019] Atsushi Yoshikawa. Interpolation functor and computability. Theoretical Computer Science, 284(2):487–498, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zheng:2002:CPR

[1020] Xizhong Zheng. The closure properties on real numbers under limits and computable operators. Theoretical Computer Science, 284(2):499–518, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schroder:2002:EA

[1021] Matthias Schröder. Extended admissibility. Theoretical Computer Science, 284(2):519–538, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Downey:2002:PCE

[1022] Rodney G. Downey and Geoffrey L. LaForte. Presentations of computably enumerable reals. Theoretical Computer Science, 284(2):539–555, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIVg

[1023] Anonymous. Author index volume 284 (2002). Theoretical Computer Science, 284(2):557–558, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bongiovanni:2002:F

[1024] Giancarlo Bongiovanni, Giorgio Gambosi, and Rossella Petreschi. Foreword. Theoretical Computer Science, 285(1):1, July 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bockenhauer:2002:TNS

[1025] Hans-Joachim Böckenhauer, Juraj Hromkovi, Ralf Klasing, Sebastian Seibert, and Walter Unger. Towards the notion of stability of approximation for hard optimization tasks and the traveling salesman problem. Theoretical Computer Science, 285(1):3–24, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cantone:2002:QEM

[1026] D. Cantone and G. Cincotti. Quick-Heapsort, an efficient mix of classical sorting algorithms. Theoretical Computer Science, 285(1):25–42, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Damaschke:2002:OSB

[1027] Peter Damaschke. Online strategies for backups. Theoretical Computer Science, 285(1):43–53, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dinur:2002:ASW

[1028] Irit Dinur. Approximating SVP to within almost-polynomial factors is NP-hard. Theoretical Computer Science, 285(1):55–71, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Merlini:2002:MBS

[1029] Donatella Merlini, Renzo Sprugnoli, and M. Cecilia Verri. Modified binary searching for static tables. Theoretical Computer Science, 285(1):73–88, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seibert:2002:HPS

[1030] Sebastian Seibert and Walter Unger. The hardness of placing street names in a Manhattan type map. Theoretical Computer Science, 285(1):89–99, July 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Steinhofel:2002:CSA

[1031] K. Steinhöfel, A. Albrecht, and C. K. Wong. The convergence of stochastic algorithms solving flow shop scheduling. Theoretical Computer Science, 285 (1):101–117, July 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marti-Oliet:2002:P

[1032] Narciso Martí-Oliet and José Meseguer. Preface. Theoretical Computer Science, 285(2):119–120, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marti-Oliet:2002:RLR

[1033] Narciso Martí-Oliet and José Meseguer.

Rewriting logic: roadmap and bibliography. Theoretical Computer Science, 285(2):121–154, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borovansky:2002:ERL

[1034] Peter Borovanský, Claude Kirchner, Hélène Kirchner, and Pierre-Etienne Moreau. ELAN from a rewriting logic point of view. Theoretical Computer Science, 285(2):155–185, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clavel:2002:MSP

[1035] M. Clavel, F. Durán, S. Eker, P. Lincoln, N. Martí-Oliet, J. Meseguer, and J. F. Quesada. Maude: specification and programming in rewriting logic. Theoretical Computer Science, 285(2): 187–243, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clavel:2002:RCR

[1036] Manuel Clavel and José Meseguer. Reflection in conditional rewriting logic. Theoretical Computer Science, 285(2): 245–288, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaconescu:2002:LFC

[1037] Răzvan Diaconescu and Kokichi Futatsugi. Logical foundations of cafeOBJ. Theoretical Computer Science, 285(2): 289–318, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gadducci:2002:CLR

[1038] Fabio Gadducci and Ugo Montanari. Comparing logics for rewriting: rewriting logic, action calculi and tile logic. Theoretical Computer Science, 285(2): 319–358, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Olveczky:2002:SRT

[1039] Peter Csaba Ölveczky and José Meseguer. Specification of real-time and hybrid systems in rewriting logic. Theoretical Computer Science, 285(2): 359–405, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pita:2002:MSO

[1040] Isabel Pita and Narciso Martí-Oliet. A Maude specification of an objectoriented model for telecommunication networks. Theoretical Computer Science, 285(2):407–439, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Talcott:2002:ATR

[1041] Carolyn Talcott. Actor theories in rewriting logic. Theoretical Computer Science, 285(2):441–485, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Viry:2002:ERR

[1042] Patrick Viry. Equational rules for rewriting logic. Theoretical Computer Science, 285(2):487–517, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wirsing:2002:FAO

[1043] Martin Wirsing and Alexander Knapp.

A formal approach to object-oriented software engineering. Theoretical Computer Science, 285(2):519–560, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIi

[1044] Anonymous. Author index. Theoretical Computer Science, 285(2):563– 564, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ruzicka:2002:P

[1045] Peter Ružička. Preface. Theoretical Computer Science, 286(1):1, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Afrati:2002:ED

[1046] Foto Afrati, Irène Guessarian, and Michel de Rougemont. The expressiveness of DAC. Theoretical Computer Science, 286(1):3–32, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vogler:2002:POS

[1047] Walter Vogler. Partial order semantics and read arcs. Theoretical Computer Science, 286(1):33–63, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khasidashvili:2002:RCF

[1048] Zurab Khasidashvili and John Glauert.

Relating conflict-free stable transition and event models via redex families. Theoretical Computer Science, 286(1):

65–95, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holzer:2002:MHF

[1049] Markus Holzer. Multi-head finite automata: data-independent versus datadependent computations. Theoretical Computer Science, 286(1):97–116, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:2002:DBL

[1050] Christian Choffrut and Giovanni Pighizzini. Distances between languages and reflexivity of relations. Theoretical Computer Science, 286(1): 117–138, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krznaric:2002:OAC

[1051] Drago Krznaric and Christos Levcopoulos. Optimal algorithms for complete linkage clustering in d dimensions. Theoretical Computer Science, 286(1): 139–149, August 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:EBa

[1052] Anonymous. Editorial board. Theoretical Computer Science, 286(1):v–ix, August 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiadeiro:2002:P

[1053] J. L. Fiadeiro. Preface. Theoretical Computer Science, 286(2):151, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Astesiano:2002:CCA

[1054] Egidio Astesiano, Michel Bidoit, Hélène Kirchner, Bernd Krieg-Brückner, Peter D. Mosses, Donald Sannella, and Andrzej Tarlecki. CASL: the common algebraic specification language. Theoretical Computer Science, 286(2):153– 196, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borzyszkowski:2002:LSS

[1055] Tomasz Borzyszkowski. Logical systems for structured specifications. Theoretical Computer Science, 286(2):197– 245, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruni:2002:NFA

[1056] Roberto Bruni, Fabio Gadducci, and Ugo Montanari. Normal forms for algebras of connections. Theoretical Computer Science, 286(2):247–292, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corradini:2002:FSM

[1057] Andrea Corradini and Fabio Gadducci. A functorial semantics for multi-algebras and partial algebras, with applications to syntax. Theoretical Computer Science, 286(2):293– 322, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Konikowska:2002:RSD

[1058] Beata Konikowska. Rasiowa-Sikorski deduction systems in computer science applications. Theoretical Computer Science, 286(2):323–366, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mossakowski:2002:RCO

[1059] Till Mossakowski. Relating CASL with other specification languages: the institution level. Theoretical Computer Science, 286(2):367–475, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIj

[1060] Anonymous. Author index. Theoretical Computer Science, 286(2):477, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2002:P

[1061] G. Rozenberg, A. E. Eiben, and J. N. Kok. Preface. Theoretical Computer Science, 287(1):1–2, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amos:2002:TTD

[1062] Martyn Amos, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Topics in the theory of DNA computing. Theoretical Computer Science, 287 (1):3–38, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brenneman:2002:SDB

[1063] Arwen Brenneman and Anne Condon. Strand design for biomolecular computation. Theoretical Computer Science, 287(1):39–58, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hagiya:2002:CAS

[1064] Masami Hagiya, John A. Rose, Ken Komiya, and Kensaku Sakamoto. Complexity analysis of the SAT engine: DNA algorithms as probabilistic algorithms. Theoretical Computer Science, 287(1):59–71, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paun:2002:GMC

[1065] Gheorghe Păun and Grzegorz Rozenberg. A guide to membrane computing. Theoretical Computer Science, 287(1): 73–100, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beyer:2002:HAE

[1066] Hans-Georg Beyer, Hans-Paul Schwefel, and Ingo Wegener. How to analyse evolutionary algorithms. Theoretical Computer Science, 287(1):101– 130, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:2002:ORS

[1067] Stefan Droste, Thomas Jansen, and Ingo Wegener. Optimization with randomized search heuristics the (A)NFL theorem, realistic scenarios, and difficult functions. Theoretical Computer Science, 287(1):131–144, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muhlenbein:2002:ECW

[1068] H. Mühlenbein and Th. Mahnig. Evolutionary computation and Wright’s equation. Theoretical Computer Science, 287(1):145–165, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Naudts:2002:GPD

[1069] B. Naudts and L. Schoofs. GA performance distributions and randomly generated binary constraint satisfaction problems. Theoretical Computer Science, 287(1):167–185, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oja:2002:ULN

[1070] Erkki Oja. Unsupervised learning in neural computation. Theoretical Computer Science, 287(1):187–207, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gammerman:2002:PAC

[1071] Alex Gammerman and Volodya Vovk. Prediction algorithms and confidence measures based on algorithmic randomness theory. Theoretical Computer Science, 287(1):209–217, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heskes:2002:AAN

[1072] Tom Heskes, Bart Bakker, and Bert Kappen. Approximate algorithms for neural-Bayesian approaches. Theoretical Computer Science, 287(1):219– 238, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Legenstein:2002:NCP

[1073] Robert A. Legenstein and Wolfgang Maass. Neural circuits for pattern recognition with small total wire length. Theoretical Computer Science, 287(1):239–249, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Natschlager:2002:SNI

[1074] Thomas Natschläger and Wolfgang Maass. Spiking neurons and the induction of finite state machines. Theoretical Computer Science, 287(1):251– 265, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hirvensalo:2002:CQI

[1075] Mika Hirvensalo. Computing with quanta impacts of quantum theory on computation. Theoretical Computer Science, 287(1):267–298, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambainis:2002:TWF

[1076] Andris Ambainis and John Watrous.

Two-way finite automata with quantum and classical states. Theoretical Computer Science, 287(1):299– 311, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terhal:2002:DQE

[1077] Barbara M. Terhal. Detecting quantum entanglement. Theoretical Computer Science, 287(1):313–335, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deWolf:2002:QCC

[1078] Ronald de Wolf. Quantum communication and complexity. Theoretical Computer Science, 287(1):337– 353, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:EBb

[1079] Anonymous. Editorial board. Theoretical Computer Science, 287(1):v– ix, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nesetril:2002:P

[1080] Jaroslav Nešetřil. Preface. Theoretical Computer Science, 287(2):355– 357, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akhavi:2002:RLT

[1081] Ali Akhavi. Random lattices, threshold phenomena and efficient reduction algorithms. Theoretical Computer Science, 287(2):359–385, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anderson:2002:SPD

[1082] Eric Anderson, Marek Chrobak, John Noga, Ji í Sgall, and Gerhard J. Woeginger. Solution of a problem in DNA computing. Theoretical Computer Science, 287(2):387–391, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anderson:2002:LUW

[1083] Eric J. Anderson, Kirsten Hildrum, Anna R. Karlin, April Rasala, and Michael Saks. On list update and work function algorithms. Theoretical Computer Science, 287(2):393– 418, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Azar:2002:LTT

[1084] Yossi Azar, Oded Regev, Ji í Sgall, and Gerhard J. Woeginger. Off-line temporary tasks assignment. Theoretical Computer Science, 287(2):419– 428, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becchetti:2002:AAR

[1085] Luca Becchetti, Miriam Di Ianni, and Alberto Marchetti-Spaccamela. Approximation algorithms for routing and call scheduling in all-optical chains and rings. Theoretical Computer Science, 287(2):429–448, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diks:2002:IIB

[1086] Krzysztof Diks, Evangelos Kranakis, Danny Krizanc, and Andrzej Pelc. The impact of information on broadcasting time in linear radio networks. Theoretical Computer Science, 287(2):449– 471, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drori:2002:FES

[1087] Limor Drori and David Peleg. Faster exact solutions for some NP-hard problems. Theoretical Computer Science, 287(2):473–499, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finocchiaro:2002:CDP

[1088] Daniele V. Finocchiaro and Marco Pellegrini. On computing the diameter of a point set in high dimensional Euclidean space. Theoretical Computer Science, 287(2):501–514, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flajolet:2002:RIR

[1089] P. Flajolet, K. Hatzis, S. Nikoletseas, and P. Spirakis. On the robustness of interconnections in random graphs: a symbolic approach. Theoretical Computer Science, 287(2):515– 534, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frankel:2002:ASD

[1090] Yair Frankel, Philip MacKenzie, and Moti Yung. Adaptively secure distributed public-key systems. Theoretical Computer Science, 287(2):535– 561, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gaillard:2002:SMR

[1091] Arlette Gaillard, Heinz Groeflin, Alan J. Hoffman, and William R. Pulleyblank. On the submodular matrix representation of a digraph. Theoretical Computer Science, 287(2):563– 570, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laber:2002:SSD

[1092] Eduardo Sany Laber, Ruy Luiz Milidiú, and Artur Alves Pessoa. A strategy for searching with different access costs. Theoretical Computer Science, 287(2):571–584, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nesetril:2002:DD

[1093] Jaroslav Nešetřil and Claude Tardif.

Density via duality. Theoretical Computer Science, 287(2):585–591, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nicodeme:2002:MS

[1094] Pierre Nicodème, Bruno Salvy, and Philippe Flajolet. Motif statistics. Theoretical Computer Science, 287(2):593– 617, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIk

[1095] Anonymous. Author index. Theoretical Computer Science, 287(2):619– 620, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moser:2002:F

[1096] G. Moser. Foreword. Theoretical Computer Science, 288(1):1, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beckmann:2002:NE

[1097] Arnold Beckmann. Notations for exponentiation. Theoretical Computer Science, 288(1):3–19, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buhrman:2002:CMD

[1098] Harry Buhrman and Ronald de Wolf. Complexity measures and decision tree complexity: a survey. Theoretical Computer Science, 288(1):21–43, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carbone:2002:SSF

[1099] A. Carbone. Streams and strings in formal proofs. Theoretical Computer Science, 288(1):45–83, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crescenzi:2002:HDC

[1100] P. Crescenzi and G. Rossi. On the Hamming distance of constraint satisfaction problems. Theoretical Computer Science, 288(1):85–100, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eiter:2002:CDD

[1101] Thomas Eiter and Helmut Veith. On the complexity of data disjunctions. Theoretical Computer Science, 288(1): 101–128, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gradel:2002:GFP

[1102] Erich Grädel. Guarded fixed point logics and the monadic theory of countable trees. Theoretical Computer Science, 288(1):129–152, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Libkin:2002:LBI

[1103] Leonid Libkin and Limsoon Wong.

Lower bounds for invariant queries in logics with counting. Theoretical Computer Science, 288(1):153–180, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sadowski:2002:OPP

[1104] Zenon Sadowski. On an optimal propositional proof system and the structure of easy subsets of TAUT. Theoretical Computer Science, 288(1):181– 193, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:EBc

[1105] Anonymous. Editorial board. Theoretical Computer Science, 288(1):v– ix, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:2002:P

[1106] Osamu Watanabe and Arun Sharma.

Preface. Theoretical Computer Science, 288(2):195–196, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balcazar:2002:CDD

[1107] José L. Balcázar, Jorge Castro, David Guijarro, and Hans-Ulrich Simon. The consistency dimension and distribution-dependent learning from queries. Theoretical Computer Science, 288(2):197–215, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Takimoto:2002:PNW

[1108] Eiji Takimoto and Manfred K. Warmuth. Predicting nearly as well as the best pruning of a planar decision graph. Theoretical Computer Science, 288(2):217–235, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldman:2002:LUP

[1109] Sally A. Goldman and Stephen S. Kwek. On learning unions of pattern languages and tree patterns in the mistake bound model. Theoretical Computer Science, 288(2):237– 254, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bshouty:2002:PLN

[1110] Nader H. Bshouty, Nadav Eiron, and Eyal Kushilevitz. PAC learning with nasty noise. Theoretical Computer Science, 288(2):255–275, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:2002:PIL

[1111] Steffen Lange and Gunter Grieser. On the power of incremental learning. Theoretical Computer Science, 288(2):277– 307, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stephan:2002:LCA

[1112] Frank Stephan and Thomas Zeugmann. Learning classes of approximations to non-recursive functions. Theoretical Computer Science, 288(2):309– 341, September 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIl

[1113] Anonymous. Author index. Theoretical Computer Science, 288(2):343, September 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durchholz:2002:GCM

[1114] Reiner Durchholz. A generic causal model for place latency. Theoretical Computer Science, 289(1):1–49, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grabner:2002:SAB

[1115] Peter J. Grabner and Helmut Prodinger.

Sorting algorithms for broadcast communications: mathematical analysis. Theoretical Computer Science, 289(1): 51–67, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dantsin:2002:DAS

[1116] Evgeny Dantsin, Andreas Goerdt, Edward A. Hirsch, Ravi Kannan, Jon Kleinberg, Christos Papadimitriou, Prabhakar Raghavan, and Uwe Schöning. A deterministic (2 − 2/(k + 1))^{n} algorithm for k-SAT based on local search. Theoretical Computer Science, 289(1):69–83, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eilam:2002:CCN

[1117] T. Eilam, S. Moran, and S. Zaks. The complexity of the characterization of networks supporting shortest-path interval routing. Theoretical Computer Science, 289(1):85–104, October 2002.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbe:2002:SDI

[1118] A. Barbé and F. von Haeseler. Symmetries of decimation invariant sequences and digit sets. Theoretical Computer Science, 289(1):105–136, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Champarnaud:2002:CDP

[1119] J.-M. Champarnaud and D. Ziadi. Canonical derivatives, partial derivatives and finite automaton constructions. Theoretical Computer Science, 289(1):137–163, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2002:CMV

[1120] Oscar H. Ibarra, Jianwen Su, Zhe Dang, Tevfik Bultan, and Richard A. Kemmerer. Counter machines and verification problems. Theoretical Computer Science, 289(1):165–189, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2002:ADT

[1121] Oscar H. Ibarra and Jianwen Su. Augmenting the discrete timed automaton with other data structures. Theoretical Computer Science, 289(1):191– 204, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atkinson:2002:STO

[1122] M. D. Atkinson, M. M. Murphy, and N. Ruškuc. Sorting with two ordered stacks in series. Theoretical Computer Science, 289(1):205–223, October 2002.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:2002:DTF

[1123] Marie-Pierre Béal and Olivier Carton. Determinization of transducers over finite and infinite words. Theoretical Computer Science, 289(1):225– 251, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klein:2002:DTM

[1124] Andreas Klein and Martin Kutrib. Deterministic Turing machines in the range between real-time and lineartime. Theoretical Computer Science, 289(1):253–275, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2002:RAC

[1125] Wei Chen, Koichi Wada, and Kimio Kawaguchi. Robust algorithms for constructing strongly convex hulls in parallel. Theoretical Computer Science, 289 (1):277–295, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet-Sadri:2002:CPW

[1126] F. Blanchet-Sadri and D. K. Luhmann.

Conjugacy on partial words. Theoretical Computer Science, 289(1):297– 312, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chung:2002:LBP

[1127] Kuo-Liang Chung, Wen-Ming Yan, and Jung-Gen Wu. Load-balanced parallel banded-system solvers. Theoretical Computer Science, 289(1):313– 334, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bein:2002:SPP

[1128] Wolfgang W. Bein, Marek Chrobak, and Lawrence L. Larmore. The 3-server problem in the plane. Theoretical Computer Science, 289(1):335–354, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Auletta:2002:RPC

[1129] Vincenzo Auletta, Ioannis Caragiannis, Christos Kaklamanis, and Pino Persiano. Randomized path coloring on binary trees. Theoretical Computer Science, 289(1):355–399, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brit:2002:PDS

[1130] Hagit Brit, Shlomo Moran, and Gadi Taubenfeld. Public data structures: counters as a special case. Theoretical Computer Science, 289(1):401– 423, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernau:2002:ELS

[1131] Henning Fernau. Even linear simple matrix languages: formal language properties and grammatical inference. Theoretical Computer Science, 289(1): 425–456, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Valiant:2002:EM

[1132] Leslie G. Valiant. Expressiveness of matchgates. Theoretical Computer Science, 289(1):457–471, October 2002.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See corrigendum [1471].

Feige:2002:DSS

[1133] Uriel Feige and Giora Rayzman. On the drift of short schedules. Theoretical Computer Science, 289(1):473– 484, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sosik:2002:UCW

[1134] Petr Sosík. Universal computation with Watson–Crick D0L systems. Theoretical Computer Science, 289(1):485– 501, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fleischner:2002:PTR

[1135] Herbert Fleischner, Oliver Kullmann, and Stefan Szeider. Polynomial-time recognition of minimal unsatisfiable formulas with fixed clause-variable difference. Theoretical Computer Science, 289(1):503–516, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eriksen:2002:ASR

[1136] Niklas Eriksen. (1+ϵ)-approximation of sorting by reversals and transpositions. Theoretical Computer Science, 289(1): 517–529, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crvenkovic:2002:ACR

[1137] Siniša Crvenković and Igor Dolinka.

On axioms for commutative regular equations without addition. Theoretical Computer Science, 289(1):531– 551, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2002:INN

[1138] Liming Cai, David Juedes, and Iyad Kanj. The inapproximability of non-NP-hard optimization problems. Theoretical Computer Science, 289(1):553– 571, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blondel:2002:PPC

[1139] Vincent D. Blondel, Julien Cassaigne, and Codrin Nichitiu. On the presence of periodic configurations in Turing machines and in counter machines. Theoretical Computer Science, 289(1): 573–590, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baumgartner:2002:PDL

[1140] Robert Baumgartner and Georg Gottlob. Propositional default logics made easier: computational complexity of model checking. Theoretical Computer Science, 289(1):591–627, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:2002:PAE

[1141] Dirk V. Arnold and Hans-Georg Beyer. Performance analysis of evolution strategies with multi-recombination in high-dimensional R^{N}-search spaces disturbed by noise. Theoretical Computer Science, 289(1):629–647, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knessl:2002:HBS

[1142] Charles Knessl and Wojciech Szpankowski. The height of a binary search tree: the limiting distribution perspective. Theoretical Computer Science, 289(1):649–703, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:2002:CPT

[1143] Juhani Karhumäki and Ion Petre. Conway’s problem for three-word sets. Theoretical Computer Science, 289(1):705– 725, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Silva:2002:ESA

[1144] Pedro V. Silva and Benjamin Steinberg. Extensions and submonoids of automatic monoids. Theoretical Computer Science, 289(1):727–754, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:2002:LPC

[1145] Antonio Restivo and Pedro V. Silva.

On the lattice of prefix codes. Theoretical Computer Science, 289(1):755– 782, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wan:2002:ACL

[1146] Honghui Wan and John C. Wootton.

Algorithms for computing lengths of chains in integral partition lattices. Theoretical Computer Science, 289(1): 783–800, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciobanu:2002:MI

[1147] Gabriel Ciobanu and Mihai Rotaru. Molecular interaction. Theoretical Computer Science, 289(1):801–827, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Slany:2002:EPS

[1148] Wolfgang Slany. Endgame problems of Sim-like graph Ramsey avoidance games are PSPACE-complete. Theoretical Computer Science, 289(1):829– 843, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Damaschke:2002:TSN

[1149] Peter Damaschke. Two short notes on the on-line travelling salesman: handling times and lookahead. Theoretical Computer Science, 289(1):845– 852, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:2002:ADS

[1150] Kosaburo Hashiguchi. Algorithms for determining the smallest number of nonterminals (states) sufficient for generating (accepting) a regular language R with R_{1} ⊆ R ⊆ R_{2} for given regular languages R_{1}, R_{2}. Theoretical Computer Science, 289(1):853–859, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hitchcock:2002:MEH

[1151] John M. Hitchcock. MAX3SAT is exponentially hard to approximate if NP has positive dimension. Theoretical Computer Science, 289(1):861– 869, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ageev:2002:MGL

[1152] M. Ageev. Martin’s game: a lower bound for the number of sets. Theoretical Computer Science, 289(1):871– 876, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:EBd

[1153] Anonymous. Editorial board. Theoretical Computer Science, 289(1):v– ix, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Du:2002:F

[1154] Dingzhu Du, Peter Eades, and Xuemin Lin. Foreword. Theoretical Computer Science, 289(2):877, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aurenhammer:2002:AUT

[1155] Franz Aurenhammer, Naoki Katoh, Hiromichi Kojima, Makoto Ohsaki, and Yinfeng Xu. Approximating uniform triangular meshes in polygons. Theoretical Computer Science, 289(2):879– 895, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Battista:2002:EPP

[1156] Giuseppe Di Battista, Giuseppe Liotta, Anna Lubiw, and Sue Whitesides. Embedding problems for paths with direction constrained edges. Theoretical Computer Science, 289(2):897– 917, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hernando:2002:LTP

[1157] Carmen Hernando, Michael E. Houle, and Ferran Hurtado. On local transformation of polygons with visibility properties. Theoretical Computer Science, 289(2):919–937, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fujita:2002:TDL

[1158] Satoshi Fujita and Takeshi Hada. Twodimensional on-line bin packing problem with rotatable items. Theoretical Computer Science, 289(2):939– 952, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halldorsson:2002:OIS

[1159] Magnús M. Halldórsson, Kazuo Iwama, Shuichi Miyazaki, and Shiro Taketomi. Online independent sets. Theoretical Computer Science, 289(2):953– 962, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamasaki:2002:OWP

[1160] Tomohiro Yamasaki, Hirotada Kobayashi, Yuuki Tokunaga, and Hiroshi Imai. One-way probabilistic reversible and quantum one-counter automata. Theoretical Computer Science, 289(2): 963–976, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ono:2002:LAD

[1161] Hirotaka Ono, Kazuhisa Makino, and Toshihide Ibaraki. Logical analysis of data with decomposable structures. Theoretical Computer Science, 289(2): 977–995, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khot:2002:PCF

[1162] Subhash Khot and Venkatesh Raman.

Parameterized complexity of finding subgraphs with hereditary properties. Theoretical Computer Science, 289(2): 997–1008, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reznik:2002:SRT

[1163] Yuriy A. Reznik. Some results on tries with adaptive branching. Theoretical Computer Science, 289(2):1009– 1026, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:AIm

[1164] Anonymous. Author index. Theoretical Computer Science, 289(2):1027– 1029, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mavronicolas:2002:TRC

[1165] Marios Mavronicolas and Nikos Papadakis. Trade-off results for connection management. Theoretical Computer Science, 290(1):1–57, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stomp:2002:CSP

[1166] Frank Stomp. Correctness of substringpreprocessing in Boyer–Moore’s pattern matching algorithm. Theoretical Computer Science, 290(1):59–78, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caucal:2002:ITG

[1167] Didier Caucal. On infinite transition graphs having a decidable monadic theory. Theoretical Computer Science, 290 (1):79–115, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernardo:2002:PMS

[1168] Marco Bernardo and Mario Bravetti. Performance measure sensitive congruences for Markovian process algebras. Theoretical Computer Science, 290(1): 117–160, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laurent:2002:PPN

[1169] Olivier Laurent. Polarized proof-nets and μ-calculus. Theoretical Computer Science, 290(1):161–188, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Valentini:2002:CCC

[1170] Silvio Valentini. A cartesian closed category in Martin-Löf’s intuitionistic type theory. Theoretical Computer Science, 290(1):189–219, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mukund:2002:BTS

[1171] Madhavan Mukund, K. Narayan Kumar, and Milind Sohoni. Bounded time-stamping in message-passing systems. Theoretical Computer Science, 290(1):221–239, October 2002.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abdulla:2002:MCS

[1172] Parosh Aziz Abdulla and Bengt Jonsson. Model checking of systems with many identical timed processes. Theoretical Computer Science, 290(1):241– 264, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lucanu:2002:RMR

[1173] Dorel Lucanu. Relaxed models for rewriting logic. Theoretical Computer Science, 290(1):265–289, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tinelli:2002:UND

[1174] Cesare Tinelli and Christophe Ringeissen. Unions of non-disjoint theories and combinations of satisfiability procedures. Theoretical Computer Science, 290(1):291–353, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bravetti:2002:DTG

[1175] M. Bravetti and A. Aldini. Discrete time generative-reactive probabilistic processes with different advancing speeds. Theoretical Computer Science, 290(1):355–406, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbanera:2002:FCM

[1176] Franco Barbanera and Stefano Berardi.

A full continuous model of polymorphism. Theoretical Computer Science, 290(1):407–428, October 2002.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nakazawa:2002:CSN

[1177] Koji Nakazawa. Confluency and strong normalizability of call-by-value μ-calculus. Theoretical Computer Science, 290(1):429–463, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maggiolo-Schettini:2002:CSS

[1178] Andrea Maggiolo-Schettini, Adriano Peron, and Simone Tini. A comparison of statecharts step semantics. Theoretical Computer Science, 290(1): 465–498, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chesnevar:2002:RDN

[1179] Carlos Iván Chesñevar, Jürgen Dix, Frieder Stolzenburg, and G. R. Guillermo Ricardo Simari. Relating defeasible and normal logic programming through transformation properties. Theoretical Computer Science, 290 (1):499–529, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cousot:2002:PAI

[1180] Patrick Cousot and Radhia Cousot. Parsing as abstract interpretation of grammar semantics. Theoretical Computer Science, 290(1):531–544, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2002:RTR

[1181] Zhaohui Zhu, Bin Li, Xi’an Xiao, Shifu Chen, and Wujia Zhu. A representation theorem for recovering contraction relations satisfying wci. Theoretical Computer Science, 290(1):545–564, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deAlfaro:2002:HD

[1182] Luca de Alfaro and Arjun Kapur. Hybrid diagrams. Theoretical Computer Science, 290(1):565–597, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alessi:2002:CCD

[1183] Fabio Alessi, Paolo Baldan, and Furio Honsell. A category of compositional domain-models for separable stone spaces. Theoretical Computer Science, 290(1):599–635, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernardinello:2002:AMO

[1184] Luca Bernardinello, Carlo Ferigato, and Lucia Pomello. An algebraic model of observable properties in distributed systems. Theoretical Computer Science, 290(1):637–668, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calders:2002:AFI

[1185] T. Calders and J. Paredaens. Axiomatization of frequent itemsets. Theoretical Computer Science, 290(1):669– 693, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amadio:2002:SRP

[1186] Roberto M. Amadio, Denis Lugiez, and Vincent Vanackère. On the symbolic reduction of processes with cryptographic functions. Theoretical Computer Science, 290(1):695–740, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laneve:2002:TSJ

[1187] Cosimo Laneve. A type system for JVM threads. Theoretical Computer Science, 290(1):741–778, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2002:UMC

[1188] Yuxi Fu and Zhenrong Yang. Understanding the mismatch combinator in chi calculus. Theoretical Computer Science, 290(1):779–830, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laurence:2002:ECF

[1189] Michael R. Laurence, Sebastian Danicic, Mark Harman, Rob Hierons, and John Howroyd. Equivalence of conservative, free, linear program schemas is decidable. Theoretical Computer Science, 290(1):831–862, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gori:2002:AIF

[1190] Roberta Gori. An abstract interpretation framework to reason on finite failure and other properties of finite and infinite computations. Theoretical Computer Science, 290(1):863– 936, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ronkko:2002:HAS

[1191] Mauno Rönkkö, Anders P. Ravn, and Kaisa Sere. Hybrid action systems.

Theoretical Computer Science, 290(1): 937–973, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanBakel:2002:NAS

[1192] Steffen van Bakel and Maribel Fernández. Normalization, approximation, and semantics for combinator systems. Theoretical Computer Science, 290(1):975–1019, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pichler:2002:EVI

[1193] Reinhard Pichler. Explicit versus implicit representations of subsets of the Herbrand universe. Theoretical Computer Science, 290(1):1021–1056, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martinelli:2002:ASP

[1194] Fabio Martinelli. Analysis of security protocols as open systems. Theoretical Computer Science, 290(1):1057– 1106, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Coquand:2002:SPM

[1195] Thierry Coquand. A syntactical proof of the marriage lemma. Theoretical Computer Science, 290(1):1107– 1113, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:EBe

[1196] Anonymous. Editorial board. Theoretical Computer Science, 290(1):v– ix, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Motet:2002:DCO

[1197] G. Motet and J.-C. Geffroy. Dependable computing: an overview. Theoretical Computer Science, 290(2):1115– 1126, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cao:2002:CMC

[1198] Guohong Cao and Mukesh Singhal. Checkpointing with mutable checkpoints. Theoretical Computer Science, 290(2):1127–1148, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caruso:2002:FDG

[1199] Antonio Caruso, Stefano Chessa, Piero Maestrini, and Paolo Santi. Faultdiagnosis of grid structures. Theoretical Computer Science, 290(2):1149– 1174, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Colby:2002:ATP

[1200] Christopher Colby, Karl Crary, Robert Harper, Peter Lee, and Frank Pfenning. Automated techniques for provably safe mobile code. Theoretical Computer Science, 290(2):1175– 1199, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prasetya:2002:FFT

[1201] I. S. W. B. Prasetya and S. D. Swierstra. Factorizing fault tolerance. Theoretical Computer Science, 290(2):1201– 1222, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Walter:2002:CFE

[1202] C. J. Walter and N. Suri. The customizable fault/error model for dependable distributed systems. Theoretical Computer Science, 290(2):1223– 1251, October 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bampis:2002:AAP

[1203] Evripidis Bampis, Rodolphe Giroudeau, and Jean-Claude König. An approximation algorithm for the precedence constrained scheduling problem with hierarchical communications. Theoretical Computer Science, 290(3):1883– 1895, November 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutten:2002:P

[1204] Shay Kutten and Paul Spirakis. Preface. Theoretical Computer Science, 291 (1):1, November 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bermond:2002:DVP

[1205] Jean-Claude Bermond, Nausica Marlin, David Peleg, and Stéphane Perennes. Directed virtual path layouts in ATM networks. Theoretical Computer Science, 291(1):3–28, November 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flocchini:2002:SDD

[1206] Paola Flocchini, Bernard Mans, and Nicola Santoro. Sense of direction in distributed computing. Theoretical Computer Science, 291(1):29– 53, November 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herlihy:2002:CWF

[1207] Maurice Herlihy and Sergio Rajsbaum. A classification of wait-free loop agreement tasks. Theoretical Computer Science, 291(1):55–77, November 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pedone:2002:OAB

[1208] Fernando Pedone and André Schiper. Optimistic atomic broadcast: a pragmatic viewpoint. Theoretical Computer Science, 291(1):79–101, November 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sussman:2002:BPE

[1209] Jeremy Sussman and Keith Marzullo. The Bancomat problem: an example of resource allocation in a partitionable asynchronous system. Theoretical Computer Science, 291(1):103– 131, November 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2002:EBf

[1210] Anonymous. Editorial board. Theoretical Computer Science, 291(1):v– ix, November 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kornerup:2002:PRN

[1211] Peter Kornerup, Jean-Claude Bajard, Christiane Frougny, and Jean-Michel Muller. Preface: Real numbers and computers. Theoretical Computer Science, 291(2):133–134, November 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hanrot:2002:DRF

[1212] G. Hanrot, J. Rivat, G. Tenenbaum, and P. Zimmermann. Density results on floating-point invertible numbers. Theoretical Computer Science, 291(2): 135–141, November 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daumas:2002:ASN

[1213] Marc Daumas and Philippe Langlois.

Additive symmetries: the non-negative case. Theoretical Computer Science, 291(2):143–157, November 2002. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matula:2002:BFM

[1214] David W. Matula and Lee D. McFearin.

A p × p bit fraction model of binary floating point division and extremal rounding cases. Theoretical Computer Science, 291(2):159–182, November 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ziv:2002:SGM

[1215] Abraham Ziv and Laurent Fournier.

Solving the generalized mask constraint for test generation of binary floating point add operation. Theoretical Computer Science, 291(2):183– 201, November 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lester:2002:UPV

[1216] David Lester and Paul Gowland. Using PVS to validate the algorithms of an exact arithmetic. Theoretical Computer Science, 291(2):203–218, November 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tinelli:2003:P

[1217] Cesare Tinelli and Teodor Rus. Preface. Theoretical Computer Science, 291 (3):219–221, January 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adi:2003:NLE

[1218] Kamel Adi, Mourad Debbabi, and Mohamed Mejri. A new logic for electronic commerce protocols. Theoretical Computer Science, 291(3):223–283, January 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Focardi:2003:CTA

[1219] Riccardo Focardi, Roberto Gorrieri, and Fabio Martinelli. A comparison of three authentication properties. Theoretical Computer Science, 291(3):285– 327, January 6, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacobs:2003:CMS

[1220] Bart Jacobs and Erik Poll. Coalgebras and monads in the semantics of Java. Theoretical Computer Science, 291(3):329–349, January 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

VanWyk:2003:SLA

[1221] Eric Van Wyk. Specification languages in algebraic compilers. Theoretical Computer Science, 291(3):351– 385, January 6, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIa

[1222] Anonymous. Author index. Theoretical Computer Science, 291(3):387– 388, January 6, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seebold:2003:F

[1223] Patrice Séébold. Foreword. Theoretical Computer Science, 292(1):1, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:PJB

[1224] Anonymous. Publications of Jean Berstel. Theoretical Computer Science, 292 (1):3–7, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:2003:PC

[1225] Jean-Paul Allouche, Michael Baake, Julien Cassaigne, and David Damanik. Palindrome complexity. Theoretical Computer Science, 292(1):9–31, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Austinat:2003:SPR

[1226] Holger Austinat, Volker Diekert, and Ulrich Hertrampf. A structural property of regular frequency computations.

Theoretical Computer Science, 292(1): 33–43, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:2003:STE

[1227] Marie-Pierre Béal, Olivier Carton, Christophe Prieur, and Jacques Sakarovitch. Squaring transducers: an efficient procedure for deciding functionality and sequentiality. Theoretical Computer Science, 292(1):45–63, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:2003:PTA

[1228] Danièle Beauquier. On probabilistic timed automata. Theoretical Computer Science, 292(1):65–84, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2003:VRS

[1229] Srečko Brlek and Christophe Reutenauer. On a valuation of rational subsets of Z^{k} Dédié à Jean Berstel. Theoretical Computer Science, 292(1):85– 96, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruyere:2003:CD

[1230] Véronique Bruyère. Cumulative defect. Theoretical Computer Science, 292(1): 97–109, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:2003:SWR

[1231] Arturo Carpi and Aldo de Luca. Semiperiodic words and root-conjugacy. Theoretical Computer Science, 292(1):

111–130, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:2003:MST

[1232] Christian Choffrut. Minimizing subsequential transducers: a survey. Theoretical Computer Science, 292(1):131– 143, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chretienne:2003:PSC

[1233] Philippe Chrétienne and Francis Sourd.

PERT scheduling with convex cost functions. Theoretical Computer Science, 292(1):145–164, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cori:2003:DTT

[1234] Robert Cori and Gilles Schaeffer. Description trees and Tutte formulas. Theoretical Computer Science, 292(1): 165–183, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:2003:RSI

[1235] Maxime Crochemore. Reducing space for index implementation. Theoretical Computer Science, 292(1):185–197, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:2003:FSG

[1236] Andrzej Ehrenfeucht, Tero Harju, Ion Petre, David M. Prescott, and Grzegorz Rozenberg. Formal systems for gene assembly in ciliates. Theoretical Computer Science, 292(1):199–219, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frougny:2003:LDS

[1237] Christiane Frougny. On-line digit set conversion in real base. Theoretical Computer Science, 292(1):221– 235, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:2003:DTB

[1238] Juhani Karhumäki, Ján Maňuch, and Wojciech Plandowski. A defect theorem for bi-infinite words. Theoretical Computer Science, 292(1):237– 243, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mignosi:2003:FWT

[1239] Filippo Mignosi, Antonio Restivo, and Pedro V. Silva. On Fine and Wilf’s theorem for bidimensional words. Theoretical Computer Science, 292(1):245– 262, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2003:CSF

[1240] Arto Salomaa. Composition sequences for functions over a finite domain. Theoretical Computer Science, 292(1):263– 281, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seebold:2003:SGT

[1241] Patrice Séébold. On some generalizations of the Thue–Morse morphism. Theoretical Computer Science, 292(1): 283–298, January 10, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thomas:2003:UNR

[1242] Wolfgang Thomas. Uniform and nonuniform recognizability. Theoretical Computer Science, 292(1):299–316, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pin:2003:ATC

[1243] Jean-Eric Pin. Algebraic tools for the concatenation product. Theoretical Computer Science, 292(1):317– 342, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBa

[1244] Anonymous. Editorial board. Theoretical Computer Science, 292(1):v– ix, January 10, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arikawa:2003:P

[1245] S. Arikawa, K. Furukawa, S. Morishita, and H. Motoda. Preface. Theoretical Computer Science, 292(2):343– 344, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hajek:2003:FLD

[1246] Petr Hájek and Martin Hole a. Formal logics of discovery and hypothesis formation by machine. Theoretical Computer Science, 292(2):345– 357, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:2003:VIL

[1247] Steffen Lange and Gunter Grieser. Variants of iterative learning. Theoretical Computer Science, 292(2):359– 376, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hayashi:2003:UCP

[1248] Yosuke Hayashi, Satoshi Matsumoto, Ayumi Shinohara, and Masayuki Takeda. Uniform characterizations of polynomial-query learnabilities. Theoretical Computer Science, 292(2):377– 385, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kudoh:2003:DAD

[1249] Yoshimitsu Kudoh, Makoto Haraguchi, and Yoshiaki Okubo. Data abstractions for decision tree induction. Theoretical Computer Science, 292(2):387– 416, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gama:2003:IB

[1250] João Gama. Iterative Bayes. Theoretical Computer Science, 292(2):417– 430, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kitagawa:2003:SPA

[1251] Genshiro Kitagawa, Tomoyuki Higuchi, and Fumiyo N. Kondo. Smoothness prior approach to explore mean structure in large-scale time series. Theoretical Computer Science, 292(2):431– 446, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Takimoto:2003:TDT

[1252] Eiji Takimoto and Akira Maruoka.

Top-down decision tree learning as information based boosting. Theoretical Computer Science, 292(2):447– 464, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hirao:2003:PAF

[1253] Masahiro Hirao, Hiromasa Hoshino, Ayumi Shinohara, Masayuki Takeda, and Setsuo Arikawa. A practical algorithm to find the best subsequence patterns. Theoretical Computer Science, 292(2):465–479, January 27, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akutsu:2003:SGA

[1254] Tatsuya Akutsu, Satoru Miyano, and Satoru Kuhara. A simple greedy algorithm for finding functional relations: efficient implementation and average case analysis. Theoretical Computer Science, 292(2):481–495, January 27, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Takeda:2003:DIP

[1255] Masayuki Takeda, Tomoko Fukuda, Ichiro Nanri, Mayumi Yamasaki, and Koichi Tamari. Discovering instances of poetic allusion from anthologies of classical Japanese poems. Theoretical Computer Science, 292(2):497–524, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Takeda:2003:DCE

[1256] Masayuki Takeda, Tetsuya Matsumoto, Tomoko Fukuda, and Ichiro Nanri. Discovering characteristic expressions in literary works. Theoretical Computer Science, 292(2):525– 546, January 27, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ohwada:2003:IIV

[1257] Hayato Ohwada and Fumio Mizoguchi. Integrating information visualization and retrieval for WWW information discovery. Theoretical Computer Science, 292(2):547–571, January 27, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berthiaume:2003:F

[1258] André Berthiaume and John D. Rogers. Foreword. Theoretical Computer Science, 292(3):573, January 31, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watrous:2003:PCR

[1259] John Watrous. PSPACE has constantround quantum interactive proof systems. Theoretical Computer Science, 292(3):575–588, January 31, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Horodecki:2003:RTB

[1260] Pawel Horodecki, John A. Smolin, Barbara M. Terhal, and Ashish V. Thapliyal. Rank two bipartite bound entangled states do not exist. Theoretical Computer Science, 292(3):589– 596, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fortnow:2003:OCT

[1261] Lance Fortnow. One complexity theorist’s view of quantum computing. Theoretical Computer Science, 292(3):597– 610, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:2003:FES

[1262] A. Ehrenfeucht and G. Rozenberg.

Forbidding–enforcing systems. Theoretical Computer Science, 292(3):611– 638, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jung:2003:EDP

[1263] Hermann Jung, Maria Serna, and Paul Spirakis. An efficient deterministic parallel algorithm for two processors precedence constraint scheduling. Theoretical Computer Science, 292(3):639– 652, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2003:DCB

[1264] G. Rozenberg and H. Spaink. DNA computing by blocking. Theoretical Computer Science, 292(3):653–665, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2003:DCR

[1265] Wei-Mei Chen and Gen-Huey Chen.

Divide-and-conquer recurrences associated with generalized heaps, optimal merge, and related structures. Theoretical Computer Science, 292(3):667– 677, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2003:TAR

[1266] Xuemin Lin and Peter Eades. Towards area requirements for drawing hierarchically planar graphs. Theoretical Computer Science, 292(3):679– 695, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zheng:2003:CAN

[1267] Yuliang Zheng and Xian-Mo Zhang. Connections among nonlinearity, avalanche and correlation immunity. Theoretical Computer Science, 292(3):697– 710, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moreira:2003:UDN

[1268] Andrés Moreira. Universality and decidability of number-conserving cellular automata. Theoretical Computer Science, 292(3):711–721, January 31, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dai:2003:SKZ

[1269] Jack Jie Dai. A stronger Kolmogorov zero-one law for resource-bounded measure. Theoretical Computer Science, 292(3):723–732, January 31, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIb

[1270] Anonymous. Author index. Theoretical Computer Science, 292(3):733– 735, January 31, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gaubert:2003:F

[1271] Stephane Gaubert, Jean-Jacques Loiseau, Jean Mairesse, Maurice Nivat, and Jean-Eric Pin. Foreword. Theoretical Computer Science, 293(1):1– 2, February 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cuninghame-Green:2003:EM

[1272] R. A. Cuninghame-Green and P. Butkovic.

The equation A ⊗ x = B ⊗ y over (max,+). Theoretical Computer Science, 293(1):3–12, February 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanderWoude:2003:CSE

[1273] Jacob van der Woude and Subiono.

Conditions for the structural existence of an eigenvalue of a bipartite (min,max,+)-system. Theoretical Computer Science, 293(1):13–24, February 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernhard:2003:MFV

[1274] Pierre Bernhard. Minimax—or feared value—L^{1}/L^{∞} control. Theoretical Computer Science, 293(1):25–44, February 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zimmermann:2003:DOM

[1275] Karel Zimmermann. Disjunctive optimization, max-separable problems and extremal algebras. Theoretical Computer Science, 293(1):45–54, February 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

dAlessandro:2003:FPP

[1276] Flavio d’Alessandro and Jacques Sakarovitch. The finite power property in free groups. Theoretical Computer Science, 293(1):55–82, February 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kobayashi:2003:SPR

[1277] Nami Kobayashi. Some properties of recognizable Z-subsets. Theoretical Computer Science, 293(1):83–113, February 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klimann:2003:SPI

[1278] Ines Klimann. A solution to the problem of (A, B)-invariance for series. Theoretical Computer Science, 293(1):115– 139, February 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gunawardena:2003:MPA

[1279] Jeremy Gunawardena. From maxplus algebra to nonexpansive mappings: a nonlinear theory for discrete event systems. Theoretical Computer Science, 293(1):141–167, February 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aceto:2003:MPA

[1280] Luca Aceto, Zoltán Ésik, and Anna Ingólfsdóttir. The max-plus algebra of the natural numbers has no finite equational basis. Theoretical Computer Science, 293(1):169–188, February 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Comet:2003:AMP

[1281] J.-P. Comet. Application of Max-Plus algebra to biological sequence comparisons. Theoretical Computer Science, 293(1):189–217, February 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ducourthial:2003:SSP

[1282] Bertrand Ducourthial and Sébastien Tixeuil. Self-stabilization with path algebra. Theoretical Computer Science, 293(1):219–236, February 3, 2003. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBb

[1283] Anonymous. Editorial board. Theoretical Computer Science, 293(1):v– ix, February 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nijholt:2003:E

[1284] A. Nijholt, G. Scollo, and D. Heylen.

Editorial. Theoretical Computer Science, 293(2):237–241, February 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Joshi:2003:NSW

[1285] Aravind K. Joshi. A note on the strong and weak generative powers of formal systems. Theoretical Computer Science, 293(2):243–259, February 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kolb:2003:ODA

[1286] Hans-Peter Kolb, Jens Michaelis, Uwe Mönnich, and Frank Morawietz. An operational and denotational approach to non-context-freeness. Theoretical Computer Science, 293(2):261–289, February 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rogers:2003:WTG

[1287] James Rogers. wMSO theories as grammar formalisms. Theoretical Computer Science, 293(2):291–320, February 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchier:2003:DCB

[1288] Denys Duchier. Dominance constraints with Boolean connectives: a model-eliminative treatment. Theoretical Computer Science, 293(2):321– 343, February 6, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stabler:2003:SSW

[1289] Edward P. Stabler and Edward L. Keenan. Structural similarity within and among languages. Theoretical Computer Science, 293(2):345–363, February 6, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schneider:2003:TPA

[1290] Karl-Michael Schneider. Tabular parsing and algebraic transformations. Theoretical Computer Science, 293(2): 365–389, February 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boullier:2003:CRC

[1291] P. Boullier. Counting with range concatenation grammars. Theoretical Computer Science, 293(2):391–416, February 6, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asveld:2003:AAF

[1292] Peter R. J. Asveld. Algebraic aspects of families of fuzzy languages. Theoretical Computer Science, 293(2):417– 445, February 6, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cantone:2003:CDF

[1293] D. Cantone, A. Formisano, E. G.

Omodeo, and C. G. Zarba. Compiling dyadic first-order specifications into map algebra. Theoretical Computer Science, 293(2):447–475, February 6, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fidge:2003:F

[1294] Colin Fidge. Foreword. Theoretical Computer Science, 293(3):477, February 9, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Torre:2003:FAT

[1295] Salvatore La Torre and Margherita Napoli. Finite automata on timed ωtrees. Theoretical Computer Science, 293(3):479–505, February 9, 2003. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McIver:2003:ACE

[1296] Annabelle McIver and Carroll Morgan. Almost-certain eventualities and abstract probabilities in the quantitative temporal logic qTL. Theoretical Computer Science, 293(3):507– 534, February 9, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saunders:2003:ISP

[1297] Shane Saunders and Tadao Takaoka. Improved shortest path algorithms for nearly acyclic graphs. Theoretical Computer Science, 293(3):535–556, February 9, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schachte:2003:PGI

[1298] Peter Schachte. Precise goalindependent abstract interpretation of constraint logic programs. Theoretical Computer Science, 293(3):557– 577, February 9, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIc

[1299] Anonymous. Author index. Theoretical Computer Science, 293(3):583– 584, February 9, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:2003:P

[1300] Ji í Adámek, Martín Escardó, and Martin Hofmann. Preface. Theoretical Computer Science, 294(1–2): 1, February 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:2003:FCC

[1301] Ji í Adámek. On final coalgebras of continuous functors. Theoretical Computer Science, 294(1–2):3–29, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bucalo:2003:ENL

[1302] Anna Bucalo, Carsten Führmann, and Alex Simpson. An equational notion of lifting monad. Theoretical Computer Science, 294(1–2):31–60, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cockett:2003:RCI

[1303] J. R. B. Cockett and Stephen Lack.

Restriction categories II: partial map classification. Theoretical Computer Science, 294(1–2):61–102, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiorentini:2003:CWP

[1304] Camillo Fiorentini and Silvio Ghilardi. Combining word problems through rewriting in categories with products. Theoretical Computer Science, 294(1–2):103–149, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hildebrandt:2003:TCM

[1305] Thomas T. Hildebrandt. Towards categorical models for fairness: fully abstract presheaf semantics of SCCS with finite delay. Theoretical Computer Science, 294(1–2):151–181, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hyland:2003:GOM

[1306] Martin Hyland and Andrea Schalk.

Glueing and orthogonality for models of linear logic. Theoretical Computer Science, 294(1–2):183–231, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moss:2003:RCS

[1307] Lawrence S. Moss. Recursion and corecursion have the same equational logic. Theoretical Computer Science, 294(1–2):233–267, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murawski:2003:ESJ

[1308] A. S. Murawski and C.-H. L. Ong. Exhausting strategies, joker games and full completeness for IMLL with Unit. Theoretical Computer Science, 294(1–2):269–305, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsuiki:2003:DTS

[1309] Hideki Tsuiki. A domain-theoretic semantics of lax generic functions. Theoretical Computer Science, 294(1–2): 307–331, February 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBc

[1310] Anonymous. Editorial board. Theoretical Computer Science, 294(1–2):v– ix, February 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girard:2003:P

[1311] Jean-Yves Girard, Mitsuhiro Okada, and Andre Scedrov. Preface. Theoretical Computer Science, 294(3):333, February 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abrusci:2003:TSP

[1312] V. Michele Abrusci. Towards a semantics of proofs for non-commutative logic: multiplicatives and additives. Theoretical Computer Science, 294(3): 335–351, February 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danos:2003:CIC

[1313] Vincent Danos, Jean-Baptiste Joinet, and Harold Schellinx. Computational isomorphisms in classical logic. Theoretical Computer Science, 294(3):353– 378, February 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guerrini:2003:CSP

[1314] Stefano Guerrini, Simone Martini, and Andrea Masini. Coherence for sharing proof-nets. Theoretical Computer Science, 294(3):379–409, February 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McDowell:2003:ETS

[1315] Raymond McDowell, Dale Miller, and Catuscia Palamidessi. Encoding transition systems in sequent calculus. Theoretical Computer Science, 294(3):411– 437, February 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pratt:2003:CSS

[1316] Vaughan Pratt. Chu spaces as a semantic bridge between linear logic and mathematics. Theoretical Computer Science, 294(3):439–471, February 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Retore:2003:HPN

[1317] Christian Retoré. Handsome proofnets: perfect matchings and cographs. Theoretical Computer Science, 294(3): 473–488, February 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

TortoradeFalco:2003:ALL

[1318] Lorenzo Tortora de Falco. Additives of linear logic and normalization—Part I: A (restricted) Church–Rosser property. Theoretical Computer Science, 294(3): 489–524, February 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kanovich:2003:PSL

[1319] Max I. Kanovich, Mitsuhiro Okada, and Andre Scedrov. Phase semantics for light linear logic. Theoretical Computer Science, 294(3):525–549, February 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagayama:2003:GTC

[1320] Misao Nagayama and Mitsuhiro Okada. A graph-theoretic characterization theorem for multiplicative fragment of non-commutative linear logic. Theoretical Computer Science, 294(3): 551–573, February 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AId

[1321] Anonymous. Author index. Theoretical Computer Science, 294(3):575– 576, February 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sgall:2003:F

[1322] Jiří Sgall and Ale Pultr. Foreword.

Theoretical Computer Science, 295(1– 3):1, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambainis:2003:ERA

[1323] Andris Ambainis and Arnolds Ķikusts.

Exact results for accepting probabilities of quantum automata. Theoretical Computer Science, 295(1–3):3– 25, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atserias:2003:IBW

[1324] Albert Atserias. Improved bounds on the Weak Pigeonhole Principle and infinitely many primes from weaker axioms. Theoretical Computer Science, 295(1–3):27–39, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barrett:2003:RPS

[1325] Chris Barrett, Harry B. Hunt III, Madhav V. Marathe, S. S. Ravi, Daniel J. Rosenkrantz, and Richard E. Stearns. Reachability problems for sequential dynamical systems with threshold functions. Theoretical Computer Science, 295(1–3):41–64, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blaser:2003:CBP

[1326] Markus Bläser. The complexity of bivariate power series arithmetic. Theoretical Computer Science, 295(1–3):65– 83, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouajjani:2003:AVR

[1327] Ahmed Bouajjani, Peter Habermehl, and Richard Mayr. Automatic verification of recursive procedures with one integer parameter. Theoretical Computer Science, 295(1–3):85–106, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cautis:2003:PMM

[1328] Sabin Cautis, Filippo Mignosi, Jeffrey Shallit, Ming wei Wang, and Soroosh Yazdani. Periodicity, morphisms, and matrices. Theoretical Computer Science, 295(1–3):107–121, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crescenzo:2003:SOS

[1329] Giovanni Di Crescenzo. Sharing one secret vs. sharing many secrets. Theoretical Computer Science, 295(1–3):123– 140, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duri:2003:CCI

[1330] Pavol Ďuri and Ján Maňuch. On the computational complexity of infinite words. Theoretical Computer Science, 295(1–3):141–151, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fournier:2003:QRP

[1331] Hervé Fournier. Quantifier rank for parity of embedded finite models. Theoretical Computer Science, 295(1–3): 153–169, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geffert:2003:SHT

[1332] Viliam Geffert. Space hierarchy theorem revised. Theoretical Computer Science, 295(1–3):171–187, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geffert:2003:CTW

[1333] Viliam Geffert, Carlo Mereghetti, and Giovanni Pighizzini. Converting twoway nondeterministic unary automata into simpler automata. Theoretical Computer Science, 295(1–3):189–203, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoang:2003:CCM

[1334] Thanh Minh Hoang and Thomas Thierauf. The complexity of the characteristic and the minimal polynomial. Theoretical Computer Science, 295(1–3):205–222, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2003:SFA

[1335] Jarkko Kari. Synchronizing finite automata on Eulerian digraphs. Theoretical Computer Science, 295(1–3):223– 232, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klein:2003:FOW

[1336] Andreas Klein and Martin Kutrib. Fast one-way cellular automata. Theoretical Computer Science, 295(1–3):233– 250, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koo:2003:LST

[1337] Chiu-Yuen Koo, Tak-Wah Lam, Tsuen-Wan Ngan, Kunihiko Sadakane, and Kar-Keung To. On-line scheduling with tight deadlines. Theoretical Computer Science, 295(1–3):251–261, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kral:2003:MHB

[1338] Daniel Král’, Jan Kratochvíl, and Heinz-Jürgen Voss. Mixed hypergraphs with bounded degree: edge-coloring of mixed multigraphs. Theoretical Computer Science, 295(1–3):263–278, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krumke:2003:NOT

[1339] Sven O. Krumke, Willem E. de Paepe, Diana Poensgen, and Leen Stougie. News from the online traveling repairman. Theoretical Computer Science, 295(1–3):279–294, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [2227].

Piterman:2003:BA

[1340] Nir Piterman and Moshe Y. Vardi. From bidirectionality to alternation. Theoretical Computer Science, 295(1– 3):295–321, February 24, 2003. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pudlak:2003:RSD

[1341] Pavel Pudlák. On reducibility and symmetry of disjoint NP pairs. Theoretical Computer Science, 295(1–3):323– 339, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santocanale:2003:EDL

[1342] Luigi Santocanale. On the equational definition of the least prefixed point. Theoretical Computer Science, 295(1–3):341–370, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sutner:2003:SPA

[1343] K. Sutner. The size of power automata. Theoretical Computer Science, 295(1–3):371–386, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thimm:2003:AST

[1344] Martin Thimm. On the approximability of the Steiner tree problem. Theoretical Computer Science, 295(1–3): 387–402, February 24, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIe

[1345] Anonymous. Author index. Theoretical Computer Science, 295(1–3):403– 404, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBd

[1346] Anonymous. Editorial board. Theoretical Computer Science, 295(1–3):v– ix, February 24, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2003:P

[1347] Jie Wang. Preface. Theoretical Computer Science, 296(1):1–2, March 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aichholzer:2003:TCT

[1348] Oswin Aichholzer, Franz Aurenhammer, Ferran Hurtado, and Hannes Krasser. Towards compatible triangulations. Theoretical Computer Science, 296(1):3–13, March 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2003:TZP

[1349] Jin-Yi Cai and Eric Bach. On testing for zero polynomials by a set of points with bounded precision. Theoretical Computer Science, 296(1):15– 25, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chan:2003:LSM

[1350] Wun-Tat Chan, Tak-Wah Lam, Hing-Fung Ting, and Prudence W. H. Wong. On-line stream merging in a general setting. Theoretical Computer Science, 296(1):27–46, March 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheong:2003:CFN

[1351] Otfried Cheong, Chan-Su Shin, and Antoine Vigneron. Computing farthest neighbors on a convex polytope. Theoretical Computer Science, 296(1): 47–58, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dang:2003:GDT

[1352] Zhe Dang, Oscar H. Ibarra, and Richard A. Kemmerer. Generalized discrete timed automata: decidable approximations for safety verification. Theoretical Computer Science, 296(1): 59–74, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duncan:2003:PTA

[1353] Rob Duncan, Jianbo Qian, Antoine Vigneron, and Binhai Zhu. Polynomial time algorithms for three-label point labeling. Theoretical Computer Science, 296(1):75–87, March 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kang:2003:LBM

[1354] Liying Kang, Hong Qiao, Erfang Shan, and Dingzhu Du. Lower bounds on the minus domination and ksubdomination numbers. Theoretical Computer Science, 296(1):89–98, March 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsu:2003:PTC

[1355] Wen-Lian Hsu and Ross M. McConnell.

PC trees and circular-ones arrangements. Theoretical Computer Science, 296(1):99–116, March 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koucky:2003:LSC

[1356] Michal Koucký. Log-space constructible universal traversal sequences for cycles of length O(n^{4.03}). Theoretical Computer Science, 296(1):117– 144, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2003:GWS

[1357] Xiang-Yang Li. Generating well-shaped d-dimensional Delaunay Meshes. Theoretical Computer Science, 296(1):145– 165, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nardelli:2003:FMV

[1358] Enrico Nardelli, Guido Proietti, and Peter Widmayer. Finding the most vital node of a shortest path. Theoretical Computer Science, 296(1):167– 177, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manthey:2003:NAW

[1359] Bodo Manthey. Non-approximability of weighted multiple sequence alignment. Theoretical Computer Science, 296(1): 179–192, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBe

[1360] Anonymous. Editorial board. Theoretical Computer Science, 296(1):v– ix, March 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Margenstern:2003:F

[1361] Maurice Margenstern. Foreword. Theoretical Computer Science, 296(2):193– 194, March 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caucal:2003:TGT

[1362] Didier Caucal. On the transition graphs of Turing machines. Theoretical Computer Science, 296(2):195– 223, March 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernau:2003:NCP

[1363] Henning Fernau. Nonterminal complexity of programmed grammars. Theoretical Computer Science, 296(2):225– 251, March 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Besozzi:2003:GPS

[1364] D. Besozzi, G. Mauri, G. Păun, and C. Zandron. Gemmating P systems: collapsing hierarchies. Theoretical Computer Science, 296(2):253– 267, March 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frisco:2003:DCU

[1365] Pierluigi Frisco. Direct constructions of universal extended H systems. Theoretical Computer Science, 296(2):269– 293, March 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin-Vide:2003:TPS

[1366] Carlos Martín-Vide, Gheorghe Păun, Juan Pazos, and Alfonso Rodríguez-Patón. Tissue P systems. Theoretical Computer Science, 296(2):295– 326, March 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herrmann:2003:UCA

[1367] Francine Herrmann and Maurice Margenstern. A universal cellular automaton in the hyperbolic plane. Theoretical Computer Science, 296(2):327– 364, March 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sutner:2003:CAI

[1368] K. Sutner. Cellular automata and intermediate degrees. Theoretical Computer Science, 296(2):365–375, March 8, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

VandenBussche:2003:F

[1369] Jan Van den Bussche and Victor Vianu. Foreword. Theoretical Computer Science, 296(3):377, March 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Libkin:2003:EPS

[1370] Leonid Libkin. Expressive power of SQL. Theoretical Computer Science, 296(3):379–404, March 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arenas:2003:SAI

[1371] Marcelo Arenas, Leopoldo Bertossi, Jan Chomicki, Xin He, Vijay Raghavan, and Jeremy Spinrad. Scalar aggregation in inconsistent databases. Theoretical Computer Science, 296(3):405– 434, March 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edmonds:2003:MES

[1372] Jeff Edmonds, Jarek Gryz, Dongming Liang, and Renée J. Miller. Mining for empty spaces in large data sets. Theoretical Computer Science, 296(3):435– 452, March 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grahne:2003:ARO

[1373] Gösta Grahne and Alex Thomo. Algebraic rewritings for optimizing regular path queries. Theoretical Computer Science, 296(3):453–471, March 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hesse:2003:DCT

[1374] William Hesse. The dynamic complexity of transitive closure is in DynTC^{0}. Theoretical Computer Science, 296(3): 473–485, March 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Poon:2003:DOR

[1375] Chung Keung Poon. Dynamic orthogonal range queries in OLAP. Theoretical Computer Science, 296(3):487– 510, March 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sinha:2003:AOD

[1376] Rakesh K. Sinha, Randeep Bhatia, and Chung-Min Chen. Asymptotically optimal declustering schemes for 2-dim range queries. Theoretical Computer Science, 296(3):511–534, March 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIf

[1377] Anonymous. Author index. Theoretical Computer Science, 296(3):535– 536, March 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Viola:2003:P

[1378] A. Viola. Preface. Theoretical Computer Science, 297(1–3):1, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akhavi:2003:OLA

[1379] Ali Akhavi. The optimal LLL algorithm is still polynomial in fixed dimension. Theoretical Computer Science, 297(1–3):3–23, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berrizbeitia:2003:PTN

[1380] Pedro Berrizbeitia, Mauricio Odremán, and Juan Tena Ayuso. Primality test for numbers M with a large power of 5 dividing M^{4} − 1. Theoretical Computer Science, 297(1–3):25–36, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carton:2003:UBA

[1381] Olivier Carton and Max Michel. Unambiguous Büchi automata. Theoretical Computer Science, 297(1–3):37– 81, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cicerone:2003:FDA

[1382] Serafino Cicerone, Gabriele Di Stefano, Daniele Frigioni, and Umberto Nanni. A fully dynamic algorithm for distributed shortest paths. Theoretical Computer Science, 297(1–3):83– 102, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2003:OPO

[1383] Myra B. Cohen and Charles J. Colbourn. Optimal and pessimal orderings of Steiner triple systems in disk arrays. Theoretical Computer Science, 297(1– 3):103–117, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corteel:2003:PPC

[1384] Sylvie Corteel, Mario Valencia-Pabon, Danièle Gardy, Dominique Barth, and Alain Denise. The permutation-path coloring problem on trees. Theoretical Computer Science, 297(1–3):119– 143, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deFigueiredo:2003:DEC

[1385] Celina M. H. de Figueiredo, João Meidanis, Célia Picinin de Mello, and Carmen Ortiz. Decompositions for the edge colouring of reduced indifference graphs. Theoretical Computer Science, 297(1–3):145–155, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernandez:2003:OEI

[1386] Maribel Fernández and Ian Mackie.

Operational equivalence for interaction nets. Theoretical Computer Science, 297(1–3):157–181, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernandez-Baca:2003:DMP

[1387] David Fernández-Baca. Decomposable multi-parameter matroid optimization problems. Theoretical Computer Science, 297(1–3):183–198, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vonzurGathen:2003:SR

[1388] Joachim von zur Gathen and Thomas Lücking. Subresultants revisited. Theoretical Computer Science, 297(1–3): 199–239, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goerdt:2003:AED

[1389] Andreas Goerdt and Mike Molloy. Analysis of edge deletion processes on faulty random regular graphs. Theoretical Computer Science, 297(1–3):241– 260, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grabner:2003:CGD

[1390] Peter J. Grabner, Arnold Knopfmacher, and Helmut Prodinger. Combinatorics of geometrically distributed random variables: run statistics. Theoretical Computer Science, 297(1–3): 261–270, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutierrez:2003:EFS

[1391] Claudio Gutiérrez. Equations in free semigroups with involution and their relation to equations in free groups. Theoretical Computer Science, 297(1– 3):271–280, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kabanets:2003:AWI

[1392] Valentine Kabanets. Almost k-wise independence and hard Boolean functions. Theoretical Computer Science, 297(1–3):281–295, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laroussinie:2003:ECQ

[1393] F. Laroussinie, Ph. Schnoebelen, and M. Turuani. On the expressivity and complexity of quantitative branchingtime temporal logics. Theoretical Computer Science, 297(1–3):297–315, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Louchard:2003:GCM

[1394] Guy Louchard and John W. Turner.

Generalized covariances of multidimensional Brownian excursion local times. Theoretical Computer Science, 297(1–3):317–336, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mayr:2003:UPU

[1395] Richard Mayr. Undecidable problems in unreliable computations. Theoretical Computer Science, 297(1–3):337– 354, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Miyazawa:2003:CP

[1396] F. K. Miyazawa and Y. Wakabayashi.

Cube packing. Theoretical Computer Science, 297(1–3):355–366, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moura:2003:RIS

[1397] Lucia Moura. Rank inequalities and separation algorithms for packing designs and sparse triple systems. Theoretical Computer Science, 297(1–3): 367–384, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Opatrny:2003:UMH

[1398] Jaroslav Opatrny. Uniform multi-hop all-to-all optical routings in rings. Theoretical Computer Science, 297(1–3): 385–397, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stevens:2003:AOS

[1399] Brett Stevens. The anti-Oberwolfach solution: pancyclic 2-factorizations of complete graphs. Theoretical Computer Science, 297(1–3):399–424, March 17, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Taylor:2003:EBB

[1400] Stephen Taylor and Marianne Durand. Emerging behavior as binary search trees are symmetrically updated. Theoretical Computer Science, 297(1–3): 425–445, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vallee:2003:DAC

[1401] Brigitte Vallée. Dynamical analysis of a class of Euclidean algorithms. Theoretical Computer Science, 297(1–3):447– 486, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zito:2003:SMM

[1402] Michele Zito. Small maximal matchings in random graphs. Theoretical Computer Science, 297(1–3):487– 507, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIg

[1403] Anonymous. Author index. Theoretical Computer Science, 297(1–3):509– 510, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBf

[1404] Anonymous. Editorial board. Theoretical Computer Science, 297(1–3): v–ix, March 17, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shinohara:2003:F

[1405] Takeshi Shinohara, Carl H. Smith, and Thomas Zeugmann. Foreword. Theoretical Computer Science, 298(1): 1–4, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamamoto:2003:HFB

[1406] Akihiro Yamamoto. Hypothesis finding based on upward refinement of residue hypotheses. Theoretical Computer Science, 298(1):5–19, April 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sakamoto:2003:LEF

[1407] Hiroshi Sakamoto, Kouichi Hirata, and Hiroki Arimura. Learning elementary formal systems with queries. Theoretical Computer Science, 298(1):21– 50, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:2003:AEF

[1408] Steffen Lange, Gunter Grieser, and Klaus P. Jantke. Advanced elementary formal systems. Theoretical Computer Science, 298(1):51–70, April 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lange:2003:DLR

[1409] Steffen Lange and Jochen Nessel. Decision lists over regular patterns. Theoretical Computer Science, 298(1):71– 87, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mukouchi:2003:RLL

[1410] Yasuhito Mukouchi and Masako Sato. Refutable language learning with a neighbor system. Theoretical Computer Science, 298(1):89–110, April 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2003:LFR

[1411] Sanjay Jain, Efim Kinber, Rolf Wiehagen, and Thomas Zeugmann. On learning of functions refutably. Theoretical Computer Science, 298(1):111– 143, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Merkle:2003:RLR

[1412] Wolfgang Merkle and Frank Stephan. Refuting learning revisited. Theoretical Computer Science, 298(1):145– 177, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yokomori:2003:PTI

[1413] Takashi Yokomori. Polynomial-time identification of very simple grammars from positive data. Theoretical Computer Science, 298(1):179–206, April 4, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [2228].

Okamoto:2003:EDC

[1414] Seishi Okamoto and Nobuhiro Yugami.

Effects of domain characteristics on instance-based learning algorithms. Theoretical Computer Science, 298(1): 207–233, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akutsu:2003:IGN

[1415] Tatsuya Akutsu, Satoru Kuhara, Osamu Maruyama, and Satoru Miyano. Identification of genetic networks by strategic gene disruptions and gene overexpressions under a boolean model. Theoretical Computer Science, 298(1): 235–251, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kida:2003:CSU

[1416] Takuya Kida, Tetsuya Matsumoto, Yusuke Shibata, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Collage system: a unifying framework for compressed pattern matching. Theoretical Computer Science, 298(1):253– 272, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBg

[1417] Anonymous. Editorial board. Theoretical Computer Science, 298(1):v– ix, April 4, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wolfram:2003:P

[1418] David Wolfram. Preface. Theoretical Computer Science, 298(2):273, April 8, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dezani-Ciancaglini:2003:ILC

[1419] Mariangiola Dezani-Ciancaglini, Paula Severi, and Fer-Jan de Vries. Infinitary lambda calculus and discrimination of Berarducci trees. Theoretical Computer Science, 298(2):275–302, April 8, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Downey:2003:UHL

[1420] Rod Downey and Lance Fortnow. Uniformly hard languages. Theoretical Computer Science, 298(2):303–315, April 8, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fellows:2003:PCS

[1421] Michael R. Fellows and Catherine Mc-Cartin. On the parametric complexity of schedules to minimize tardy tasks. Theoretical Computer Science, 298(2): 317–324, April 8, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khoussainov:2003:ALS

[1422] Bakhadyr Khoussainov. On algebraic and logical specifications of classes of regular languages. Theoretical Computer Science, 298(2):325–346, April 8, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krishnan:2003:ASS

[1423] Padmanabhan Krishnan. Automatic synthesis of a subclass of schedulers in timed systems. Theoretical Computer Science, 298(2):347–363, April 8, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2003:LPL

[1424] Eric Martin, Arun Sharma, and Frank Stephan. Learning power and language expressiveness. Theoretical Computer Science, 298(2):365–383, April 8, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honsell:2003:P

[1425] Furio Honsell and Marino Miculan.

Preface. Theoretical Computer Science, 298(3):385–386, April 11, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abadi:2003:STA

[1426] Martín Abadi and Bruno Blanchet. Secrecy types for asymmetric communication. Theoretical Computer Science, 298(3):387–415, April 11, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aceto:2003:ETT

[1427] Luca Aceto, Zoltán Ésik, and Anna Ingólfsdóttir. Equational theories of tropical semirings. Theoretical Computer Science, 298(3):417–469, April 11, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bidoit:2003:OLC

[1428] Michel Bidoit, Rolf Hennicker, and Alexander Kurz. Observational logic, constructor-based logic, and their duality. Theoretical Computer Science, 298 (3):471–510, April 11, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bojanczyk:2003:FGP

[1429] Mikołaj Bojańczyk. The finite graph problem for two-way alternating automata. Theoretical Computer Science, 298(3):511–528, April 11, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busi:2003:EDC

[1430] Nadia Busi and Gianluigi Zavattaro. Expired data collection in shared dataspaces. Theoretical Computer Science, 298(3):529–556, April 11, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calcagno:2003:PLE

[1431] Cristiano Calcagno, Peter O’Hearn, and Richard Bornat. Program logic and equivalence in the presence of garbage collection. Theoretical Computer Science, 298(3):557–581, April 11, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klein:2003:VBV

[1432] Gerwin Klein and Tobias Nipkow. Verified bytecode verifiers. Theoretical Computer Science, 298(3):583– 626, April 11, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIh

[1433] Anonymous. Author index. Theoretical Computer Science, 298(3):627– 628, April 11, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:2003:MSO

[1434] Bruno Courcelle. The monadic secondorder logic of graphs XIV: uniformly sparse graphs and edge set quantifications. Theoretical Computer Science, 299(1–3):1–36, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Magnoni:2003:PLP

[1435] Letizia Magnoni, Massimo Mirolli, Franco Montagna, and Giulia Simi. PAC learning of probability distributions over a discrete domain. Theoretical Computer Science, 299(1–3): 37–63, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aracena:2003:CPR

[1436] Julio Aracena and Eric Goles. Complexity of perceptron recognition for a class of geometric patterns. Theoretical Computer Science, 299(1–3): 65–79, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kiwi:2003:ATW

[1437] M. Kiwi. Algebraic testing and weight distributions of codes. Theoretical Computer Science, 299(1–3):81– 106, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hwang:2003:SML

[1438] F. K. Hwang. A survey on multi-loop networks. Theoretical Computer Science, 299(1–3):107–121, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epifanio:2003:CBW

[1439] Chiara Epifanio, Michel Koskas, and Filippo Mignosi. On a conjecture on bidimensional words. Theoretical Computer Science, 299(1–3):123– 150, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wood:2003:OTD

[1440] David R. Wood. Optimal threedimensional orthogonal graph drawing in the general position model. Theoretical Computer Science, 299(1–3):151– 178, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartha:2003:SEC

[1441] Miklós Bartha and Miklós Krész. Structuring the elementary components of graphs having a perfect internal matching. Theoretical Computer Science, 299(1–3):179–210, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2003:AAN

[1442] Zhi-Zhong Chen, Tao Jiang, Guohui Lin, Jianjun Wen, Dong Xu, Jinbo Xu, and Ying Xu. Approximation algorithms for NMR spectral peak assignment. Theoretical Computer Science, 299(1–3):211–229, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeza-Yates:2003:FAS

[1443] R. Baeza-Yates, J. Gabarró, and X. Messeguer. Fringe analysis of synchronized parallel insertion algorithms in 2–3 Trees. Theoretical Computer Science, 299(1–3):231–271, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nandy:2003:ENN

[1444] Subhas C. Nandy, Sandip Das, and Partha P. Goswami. An efficient k nearest neighbors searching algorithm for a query line. Theoretical Computer Science, 299(1–3):273–288, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malkhi:2003:DFR

[1445] Dahlia Malkhi, Yishay Mansour, and Michael K. Reiter. Diffusion without false rumors: on propagating updates in a Byzantine environment. Theoretical Computer Science, 299(1–3):289– 306, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holzer:2003:AEA

[1446] Markus Holzer and Pierre McKenzie. Alternating and empty alternating auxiliary stack automata. Theoretical Computer Science, 299(1–3):307– 326, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkel:2003:OCF

[1447] Olivier Finkel. On omega context free languages which are Borel sets of infinite rank. Theoretical Computer Science, 299(1–3):327–346, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuske:2003:TLT

[1448] Dietrich Kuske. Towards a language theory for infinite N-free pomsets. Theoretical Computer Science, 299(1–3): 347–386, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lemstrom:2003:APM

[1449] Kjell Lemström and Lauri Hella. Approximate pattern matching and transitive closure logics. Theoretical Computer Science, 299(1–3):387–412, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dang:2003:PLV

[1450] Zhe Dang, Pierluigi San Pietro, and Richard A. Kemmerer. Presburger liveness verification of discrete timed automata. Theoretical Computer Science, 299(1–3):413–438, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2003:LBL

[1451] Leah Epstein and Rob van Stee. Lower bounds for on-line single-machine scheduling. Theoretical Computer Science, 299(1–3):439–450, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rusinowitch:2003:PIF

[1452] Michaël Rusinowitch and Mathieu Turuani. Protocol insecurity with a finite number of sessions and composed keys is NP-complete. Theoretical Computer Science, 299(1–3):451–475, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiorenzi:2003:CAS

[1453] Francesca Fiorenzi. Cellular automata and strongly irreducible shifts of finite type. Theoretical Computer Science, 299(1–3):477–493, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krishna:2003:BUP

[1454] Shankara Narayanan Krishna and Raghavan Rama. Breaking DES using P systems. Theoretical Computer Science, 299(1–3):495–508, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watson:2003:NRG

[1455] Bruce W. Watson. A new regular grammar pattern matching algorithm. Theoretical Computer Science, 299(1–3): 509–521, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durand:2003:NCC

[1456] Bruno Durand, Enrico Formenti, and Zsuzsanna Róka. Number-conserving cellular automata I: decidability. Theoretical Computer Science, 299(1–3): 523–535, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engebretsen:2003:TOL

[1457] Lars Engebretsen and Jonas Holmerin. Towards optimal lower bounds for clique and chromatic number. Theoretical Computer Science, 299(1–3):537– 584, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giorgetti:2003:ASP

[1458] Alain Giorgetti. An asymptotic study for path reversal. Theoretical Computer Science, 299(1–3):585–602, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wild:2003:ICI

[1459] Marcel Wild. Idempotent and coidempotent stack filters and min–max operators. Theoretical Computer Science, 299(1–3):603–631, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernau:2003:HMC

[1460] H. Fernau, M. Holzer, and R. Freund. Hybrid modes in cooperating distributed grammar systems: combining the t-mode with the modes ≤ k and = k. Theoretical Computer Science, 299(1–3):633–662, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2003:CPL

[1461] Alexander Okhotin. On the closure properties of linear conjunctive languages. Theoretical Computer Science, 299(1–3):663–685, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2003:EST

[1462] Oscar H. Ibarra and Zhe Dang. Eliminating the storage tape in reachability constructions. Theoretical Computer Science, 299(1–3):687–706, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buchsbaum:2003:FCN

[1463] Adam L. Buchsbaum, Raffaele Giancarlo, and Jeffery R. Westbrook. On finding common neighborhoods in massive graphs. Theoretical Computer Science, 299(1–3):707–718, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gerber:2003:AVP

[1464] Michael U. Gerber and Daniel Kobler.

Algorithms for vertex-partitioning problems on graphs with fixed cliquewidth. Theoretical Computer Science, 299(1–3):719–734, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heam:2003:SCR

[1465] P.-C. Héam. Some complexity results for polynomial rational expressions. Theoretical Computer Science, 299(1– 3):735–741, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2003:RCD

[1466] Y. J. Liu. Regular component decomposition of regular languages. Theoretical Computer Science, 299(1–3):743– 749, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clementi:2003:MBR

[1467] Andrea E. F. Clementi, Miriam Di Ianni, and Riccardo Silvestri. The minimum broadcast range assignment problem on linear multi-hop wireless networks. Theoretical Computer Science, 299(1–3):751–761, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rytter:2003:MSC

[1468] Wojciech Rytter. On maximal suffixes and constant-space linear-time versions of KMP algorithm. Theoretical Computer Science, 299(1–3):763– 774, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcia:2003:BLT

[1469] Pedro García, José Ruiz, and Manuel Vazquez de Parga. Bilateral locally testable languages. Theoretical Computer Science, 299(1–3):775–783, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:2003:PGS

[1470] Erzsébet Csuhaj-Varjú, Gheorghe Păun, and György Vaszil. PC grammar systems with five context-free components generate all recursively enumerable languages. Theoretical Computer Science, 299(1–3):785–794, April 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Valiant:2003:CEM

[1471] Leslie G. Valiant. Corrigendum to “Expressiveness of matchgates”: [Theoret. Comput. Sci. 289(1) (2002) 457–471].

Theoretical Computer Science, 299(1– 3):795, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1132].

Anonymous:2003:AIi

[1472] Anonymous. Author index. Theoretical Computer Science, 299(1–3):797– 799, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBh

[1473] Anonymous. Editorial board. Theoretical Computer Science, 299(1–3): v–ix, April 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aczel:2003:ITC

[1474] Peter Aczel, Ji í Adámek, Stefan Milius, and Ji í Velebil. Infinite trees and completely iterative theories: a coalgebraic view. Theoretical Computer Science, 300(1–3):1–45, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cattani:2003:PMC

[1475] Gian Luca Cattani and Glynn Winskel. Presheaf models for CCS-like languages. Theoretical Computer Science, 300(1–3):47–89, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkelstein:2003:NFD

[1476] Stacy E. Finkelstein, Peter Freyd, and James Lipton. A new framework for declarative programming. Theoretical Computer Science, 300(1–3): 91–160, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [1781].

Mohamed:2003:NTM

[1477] Otmane Aït Mohamed, Xiaoyu Song, and Eduard Cerny. On the nontermination of M-based abstract state enumeration. Theoretical Computer Science, 300(1–3):161–179, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wabenhorst:2003:ITI

[1478] Axel Wabenhorst. Induction in the Timed Interval Calculus. Theoretical Computer Science, 300(1–3):181– 207, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vagvolgyi:2003:IFG

[1479] Sándor Vágvölgyi. Intersection of finitely generated congruences over term algebra. Theoretical Computer Science, 300(1–3):209–234, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demri:2003:PSC

[1480] Stéphane Demri. A polynomial space construction of tree-like models for logics with local chains of modal connectives. Theoretical Computer Science, 300(1–3):235–258, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Devillers:2003:GPR

[1481] Raymond Devillers, Hanna Klaudel, and Robert-C. Riemann. General parameterised refinement and recursion for the M-net calculus. Theoretical Computer Science, 300(1–3):259– 300, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ayari:2003:DPI

[1482] Abdelwaheb Ayari, David Basin, and Felix Klaedtke. Decision procedures for inductive Boolean functions based on alternating automata. Theoretical Computer Science, 300(1–3):301– 329, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rabinovich:2003:ACT

[1483] Alexander Rabinovich. Automata over continuous time. Theoretical Computer Science, 300(1–3):331–363, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carlucci:2003:NPT

[1484] Lorenzo Carlucci. A new prooftheoretic proof of the independence of Kirby–Paris’ Hydra Theorem. Theoretical Computer Science, 300(1–3): 365–378, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gordon:2003:TCA

[1485] Andrew D. Gordon and Alan Jeffrey.

Typing correspondence assertions for communication protocols. Theoretical Computer Science, 300(1–3):379– 409, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aceto:2003:PRT

[1486] Luca Aceto, Patricia Bouyer, Augusto Burgueño, and Kim G. Larsen. The power of reachability testing for timed automata. Theoretical Computer Science, 300(1–3):411–475, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

David:2003:DRP

[1487] R. David. Decidability results for primitive recursive algorithms. Theoretical Computer Science, 300(1–3): 477–504, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIj

[1488] Anonymous. Author index. Theoretical Computer Science, 300(1–3):505, May 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:MI

[1489] Anonymous. Master index. Theoretical Computer Science, 300(1–3): 507–519, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBi

[1490] Anonymous. Editorial board. Theoretical Computer Science, 300(1–3): v–ix, May 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moschovakis:2003:PRA

[1491] Yiannis N. Moschovakis. On primitive recursive algorithms and the greatest common divisor function. Theoretical Computer Science, 301(1–3): 1–30, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kelarev:2003:CAD

[1492] A. V. Kelarev and O. V. Sokratova.

On congruences of automata defined by directed graphs. Theoretical Computer Science, 301(1–3):31–43, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sauerhoff:2003:ABF

[1493] Martin Sauerhoff. Approximation of boolean functions by combinatorial rectangles. Theoretical Computer Science, 301(1–3):45–78, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Katsura:2003:UPM

[1494] Masashi Katsura, Yuji Kobayashi, and Friedrich Otto. Undecidable properties of monoids with word problem solvable in linear time. Part II. Cross sections and homological and homotopical finiteness conditions. Theoretical Computer Science, 301(1–3):79– 101, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sgarbas:2003:OID

[1495] Kyriakos N. Sgarbas, Nikos D. Fakotakis, and George K. Kokkinakis. Optimal insertion in deterministic DAWGs. Theoretical Computer Science, 301(1– 3):103–117, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gerbessiotis:2003:AIP

[1496] Alexandros V. Gerbessiotis and Constantinos J. Siniolakis. Architecture independent parallel selection with applications to parallel priority queues. Theoretical Computer Science, 301(1–3):

119–142, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nock:2003:CCA

[1497] Richard Nock. Complexity in the case against accuracy estimation. Theoretical Computer Science, 301(1–3):143– 165, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terrier:2003:TDC

[1498] Véronique Terrier. Two-dimensional cellular automata and deterministic online tessalation automata. Theoretical Computer Science, 301(1–3):167– 186, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2003:WCD

[1499] Arto Salomaa and Petr Sosík. Watson– Crick D0L systems: the power of one transition. Theoretical Computer Science, 301(1–3):187–200, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferretti:2003:TVR

[1500] Claudio Ferretti, Giancarlo Mauri, Gheorghe Păun, and Claudio Zandron. On three variants of rewriting P systems. Theoretical Computer Science, 301(1–3):201–215, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkel:2003:AOC

[1501] Olivier Finkel. Ambiguity in omega context free languages. Theoretical Computer Science, 301(1–3):217– 270, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boliac:2003:IDF

[1502] R. Boliac and V. Lozin. Independent domination in finitely defined classes of graphs. Theoretical Computer Science, 301(1–3):271–284, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peter:2003:ADE

[1503] Manfred Peter. The asymptotic distribution of elements in automatic sequences. Theoretical Computer Science, 301(1–3):285–312, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanderWalt:2003:PRC

[1504] Andries van der Walt and Sigrid Ewert.

A property of random context picture grammars. Theoretical Computer Science, 301(1–3):313–320, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ellis:2003:ES

[1505] John Ellis, Frank Ruskey, Joe Sawada, and Jamie Simpson. Euclidean strings. Theoretical Computer Science, 301(1– 3):321–340, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Formenti:2003:SAC

[1506] Enrico Formenti. On the sensitivity of additive cellular automata in Besicovitch topologies. Theoretical Computer Science, 301(1–3):341–354, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flocchini:2003:CAN

[1507] Paola Flocchini, Alessandro Roncato, and Nicola Santoro. Computing on anonymous networks with sense of direction. Theoretical Computer Science, 301(1–3):355–379, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2003:EMS

[1508] Chin Lung Lu, Sheng-Lung Peng, and Chuan Yi Tang. Efficient minus and signed domination in graphs. Theoretical Computer Science, 301(1–3): 381–397, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lam:2003:CCF

[1509] Nguyen Huong Lam. Completing comma-free codes. Theoretical Computer Science, 301(1–3):399–415, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolinka:2003:MFY

[1510] Igor Dolinka. The multiplicative fragment of the Yanov equational theory. Theoretical Computer Science, 301(1– 3):417–425, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Georgiadis:2003:AOP

[1511] Leonidas Georgiadis. Arborescence optimization problems solvable by Edmonds’ algorithm. Theoretical Computer Science, 301(1–3):427–437, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barge:2003:AOP

[1512] Marcy Barge, Beverly Diamond, and Charles Holton. Asymptotic orbits of primitive substitutions. Theoretical Computer Science, 301(1–3):439– 450, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iwama:2003:FNW

[1513] Kazuo Iwama, Akihiro Matsuura, and Mike Paterson. A family of NFAs which need 2^{n}-deterministic states. Theoretical Computer Science, 301(1–3):451– 462, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Le:2003:SCL

[1514] Van Bang Le and Bert Randerath. On stable cutsets in line graphs. Theoretical Computer Science, 301(1–3): 463–475, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sawada:2003:FAG

[1515] Joe Sawada. A fast algorithm to generate necklaces with fixed content. Theoretical Computer Science, 301(1–3): 477–489, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harper:2003:BHG

[1516] L. H. Harper. On the bandwidth of a Hamming graph. Theoretical Computer Science, 301(1–3):491–498, May 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIk

[1517] Anonymous. Author index. Theoretical Computer Science, 301(1–3):499– 500, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBj

[1518] Anonymous. Editorial board. Theoretical Computer Science, 301(1–3): v–ix, May 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Richomme:2003:CEM

[1519] G. Richomme. Conjugacy and episturmian morphisms. Theoretical Computer Science, 302(1–3):1–34, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Damaschke:2003:NOS

[1520] Peter Damaschke. Nearly optimal strategies for special cases of on-line capital investment. Theoretical Computer Science, 302(1–3):35–44, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Panholzer:2003:AMQ

[1521] Alois Panholzer. Analysis of multiple quickselect variants. Theoretical Computer Science, 302(1–3):45–91, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dang:2003:PTA

[1522] Zhe Dang. Pushdown timed automata: a binary reachability characterization and safety verification. Theoretical Computer Science, 302(1–3):

93–121, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mycka:2003:RIL

[1523] Jerzy Mycka. -recursion and infinite limits. Theoretical Computer Science, 302(1–3):123–133, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vagvolgyi:2003:TRR

[1524] Sándor Vágvölgyi. Term rewriting restricted to ground terms. Theoretical Computer Science, 302(1–3):135– 165, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2003:NDP

[1525] S. Brlek and A. Ladouceur. A note on differentiable palindromes. Theoretical Computer Science, 302(1–3):167– 178, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeNicola:2003:NRE

[1526] Rocco De Nicola and Anna Labella.

Nondeterministic regular expressions as solutions of equational systems. Theoretical Computer Science, 302(1–3): 179–189, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2003:OSP

[1527] Jingchao Chen. Optimizing stable inplace merging. Theoretical Computer Science, 302(1–3):191–210, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rytter:2003:ALZ

[1528] Wojciech Rytter. Application of Lempel–Ziv factorization to the approximation of grammar-based compression. Theoretical Computer Science, 302(1–3):211–222, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2003:GSAa

[1529] Dong Kyue Kim, Yoo Ah Kim, and Kunsoo Park. Generalizations of suffix arrays to multi-dimensional matrices. Theoretical Computer Science, 302 (1–3):223–238, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:2003:ASS

[1530] Klaus Jansen. Approximate strong separation with application in fractional graph coloring and preemptive scheduling. Theoretical Computer Science, 302 (1–3):239–256, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karakostas:2003:CIF

[1531] George Karakostas, Richard J. Lipton, and Anastasios Viglas. On the complexity of intersecting finite state automata and versus. Theoretical Computer Science, 302(1–3):257– 274, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiang:2003:UEW

[1532] Zhuhan Jiang, Olivier de Vel, and Bruce Litow. Unification and extension of weighted finite automata applicable to image compression. Theoretical Computer Science, 302(1–3):

275–294, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domosi:2003:FSF

[1533] Pál Dömösi, Chrystopher L. Nehaniv, and John L. Rhodes. Finite semigroups, feedback, and the Letichevsky criteria on non-empty words in finite automata. Theoretical Computer Science, 302(1–3):295–317, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fishkin:2003:MTM

[1534] Aleksei V. Fishkin and Guochuan Zhang. On maximizing the throughput of multiprocessor tasks. Theoretical Computer Science, 302(1–3):319– 335, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clementi:2003:DBR

[1535] Andrea E. F. Clementi, Angelo Monti, and Riccardo Silvestri. Distributed broadcast in radio networks of unknown topology. Theoretical Computer Science, 302(1–3):337–364, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2003:RPA

[1536] Alexander Okhotin. A recognition and parsing algorithm for arbitrary conjunctive grammars. Theoretical Computer Science, 302(1–3):365–399, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2003:GSAb

[1537] Dong Kyue Kim, Yoo Ah Kim, and Kunsoo Park. Generalizations of suffix arrays to multi-dimensional matrices. Theoretical Computer Science, 302 (1–3):401–416, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Creignou:2003:GSP

[1538] Nadia Creignou and Hervé Daudé. Generalized satisfiability problems: minimal elements and phase transitions. Theoretical Computer Science, 302(1–3):417–430, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertoni:2003:NOS

[1539] Alberto Bertoni, Christian Choffrut, Massimiliano Goldwurm, and Violetta Lonati. On the number of occurrences of a symbol in words of regular languages. Theoretical Computer Science, 302(1–3):431–456, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2003:PIS

[1540] Lane A. Hemaspaandra and Harald Hempel. P-immune sets with holes lack self-reducibility properties. Theoretical Computer Science, 302(1–3): 457–466, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tantau:2003:QCM

[1541] Till Tantau. Query complexity of membership comparable sets. Theoretical Computer Science, 302(1–3):467– 474, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biedl:2003:PRU

[1542] Therese Biedl, Jonathan F. Buss, Erik D. Demaine, Martin L. Demaine, Mohammadtaghi Hajiaghayi, and Tomá Vina. Palindrome recognition using a multidimensional tape. Theoretical Computer Science, 302(1– 3):475–480, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honkala:2003:DRW

[1543] Juha Honkala. Decidability results for Watson–Crick D0L systems with nonregular triggers. Theoretical Computer Science, 302(1–3):481–488, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czumaj:2003:PTA

[1544] Artur Czumaj, Leszek Ga̧sieniec, Daya Ram Gaur, Ramesh Krishnamurti, Wojciech Rytter, and Michele Zito. On polynomial-time approximation algorithms for the variable length scheduling problem. Theoretical Computer Science, 302(1–3):489– 495, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:2003:EFA

[1545] Marek Chrobak. Errata to: “Finite Automata and Unary Languages”: [Theoret. Comput. Sci. 47 (1986) 149–158]. Theoretical Computer Science, 302(1– 3):497–498, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1].

Anonymous:2003:AIl

[1546] Anonymous. Author index. Theoretical Computer Science, 302(1–3):499– 500, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBk

[1547] Anonymous. Editorial board. Theoretical Computer Science, 302(1–3): v–ix, June 13, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:2003:F

[1548] Danièle Beauquier and Yuri Matijassevich. Foreword. Theoretical Computer Science, 303(1):1, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:2003:BS

[1549] Danièle Beauquier, Dimitri Grigoriev, and Yuri Matiyasevich. Biography of A. O. Slissenko. Theoretical Computer Science, 303(1):3–5, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:2003:GSC

[1550] A. Arnold, A. Vincent, and I. Walukiewicz. Games for synthesis of controllers with partial observation. Theoretical Computer Science, 303(1):7–34, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carbone:2003:FLS

[1551] A. Carbone and M. Gromov. Functional labels and syntactic entropy on DNA strings and proteins. Theoretical Computer Science, 303(1):35– 51, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cegielski:2003:AIN

[1552] Patrick Cégielski, François Heroult, and Denis Richard. On the amplitude of intervals of natural numbers whose every element has a common prime divisor with at least an extremity. Theoretical Computer Science, 303 (1):53–62, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dekhtyar:2003:FCC

[1553] Michael Dekhtyar, Alexander Dikovsky, and Mars Valiev. On feasible cases of checking multi-agent systems behavior. Theoretical Computer Science, 303(1): 63–81, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grigoriev:2003:APS

[1554] Dima Grigoriev and Edward A. Hirsch.

Algebraic proof systems over formulas. Theoretical Computer Science, 303(1): 83–102, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guessarian:2003:TLV

[1555] Irène Guessarian, Eugénie Foustoucos, Theodore Andronikos, and Foto Afrati. On temporal logic versus datalog. Theoretical Computer Science, 303(1):103– 133, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kolpakov:2003:FAR

[1556] Roman Kolpakov and Gregory Kucherov. Finding approximate repetitions under Hamming distance. Theoretical Computer Science, 303(1):135–156, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Makowsky:2003:TWM

[1557] J. A. Makowsky and J. P. Mariño. Tree-width and the monadic quantifier hierarchy. Theoretical Computer Science, 303(1):157–170, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maksimova:2003:CSP

[1558] L. Maksimova. Complexity of some problems in positive and related calculi. Theoretical Computer Science, 303(1): 171–185, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mints:2003:TPE

[1559] G. Mints. A termination proof for epsilon substitution using partial derivations. Theoretical Computer Science, 303(1):187–213, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Niwinski:2003:GPD

[1560] Damian Niwiński and Igor Walukiewicz. A gap property of deterministic tree languages. Theoretical Computer Science, 303(1):215–231, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Razborov:2003:RLB

[1561] Alexander A. Razborov. Resolution lower bounds for the weak functional pigeonhole principle. Theoretical Computer Science, 303(1):233–243, June 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Soloviev:2003:CEG

[1562] Sergei Soloviev and Vladimir Orevkov.

On categorical equivalence of Gentzenstyle derivations in IMLL. Theoretical Computer Science, 303(1):245– 260, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBl

[1563] Anonymous. Editorial board. Theoretical Computer Science, 303(1):v– ix, June 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durand:2003:F

[1564] B. Durand and L. Vuillon. Foreword.

Theoretical Computer Science, 303(2– 3):265, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Propp:2003:GDS

[1565] James Propp. Generalized dominoshuffling. Theoretical Computer Science, 303(2–3):267–301, July 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pak:2003:TIN

[1566] Igor Pak. Tile invariants: new horizons.

Theoretical Computer Science, 303(2– 3):303–331, July 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fournier:2003:CPM

[1567] J. C. Fournier. Combinatorics of perfect matchings in plane bipartite graphs and application to tilings. Theoretical Computer Science, 303(2–3): 333–351, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thiant:2003:AFD

[1568] Nicolas Thiant. An O(n log n)- algorithm for finding a domino tiling of a plane picture whose number of holes is bounded. Theoretical Computer Science, 303(2–3):353–374, July 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Desreux:2003:AGE

[1569] Sébastien Desreux. An algorithm to generate exactly once every tiling with lozenges of a domain. Theoretical Computer Science, 303(2–3):375– 408, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lam:2003:FNS

[1570] Fumei Lam and Lior Pachter. Forcing numbers of stop signs. Theoretical Computer Science, 303(2–3):409– 416, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:2003:CUP

[1571] Danièle Beauquier and Maurice Nivat. A codicity undecidable problem in the plane. Theoretical Computer Science, 303(2–3):417–430, July 15, 2003.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delgado-Friedrichs:2003:DSA

[1572] Olaf Delgado-Friedrichs. Data structures and algorithms for tilings I. Theoretical Computer Science, 303(2–3): 431–445, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fitzkee:2003:WMT

[1573] Thomas L. Fitzkee, Kevin G. Hockett, and E. Arthur Robinson, Jr. A weakly mixing tiling dynamical system with a smooth model. Theoretical Computer Science, 303(2–3):447– 462, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lenz:2003:HSS

[1574] Daniel Lenz. Hierarchical structures in Sturmian dynamical systems. Theoretical Computer Science, 303(2–3): 463–490, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frougny:2003:AMP

[1575] Christiane Frougny, Jean-Pierre Gazeau, and Rudolf Krejcar. Additive and multiplicative properties of point sets based on beta-integers. Theoretical Computer Science, 303(2–3):491–516, July 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lorys:2003:NAA

[1576] Krzysztof Lorys and Katarzyna E.

Paluch. New approximation algorithm for RTILE problem. Theoretical Computer Science, 303(2–3):517– 537, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eloranta:2003:BEV

[1577] Kari Eloranta. The bounded eightvertex model. Theoretical Computer Science, 303(2–3):539–552, July 15, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIm

[1578] Anonymous. Author index. Theoretical Computer Science, 303(2–3):553– 554, July 15, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muchnik:2003:APS

[1579] An. Muchnik, A. Semenov, and M. Ushakov. Almost periodic sequences. Theoretical Computer Science, 304(1–3):1–33, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brunetti:2003:ARC

[1580] Sara Brunetti and Alain Daurat. An algorithm reconstructing convex lattice sets. Theoretical Computer Science, 304(1–3):35–57, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Louchard:2003:ARS

[1581] Guy Louchard and Helmut Prodinger. Ascending runs of sequences of geometrically distributed random variables: a probabilistic analysis. Theoretical Computer Science, 304(1–3): 59–86, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Na:2003:TST

[1582] Joong Chae Na, Alberto Apostolico, Costas S. Iliopoulos, and Kunsoo Park. Truncated suffix trees and their application to data compression. Theoretical Computer Science, 304(1–3): 87–101, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpentieri:2003:SQT

[1583] Marco Carpentieri. On the simulation of quantum Turing machines. Theoretical Computer Science, 304(1–3): 103–128, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liskiewicz:2003:CCS

[1584] Maciej Liśkiewicz, Mitsunori Ogihara, and Seinosuke Toda. The complexity of counting self-avoiding walks in subgraphs of two-dimensional grids and hypercubes. Theoretical Computer Science, 304(1–3):129–156, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kucera:2003:CBC

[1585] Antonín Kučera. The complexity of bisimilarity-checking for one-counter processes. Theoretical Computer Science, 304(1–3):157–183, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lee:2003:SOD

[1586] Jia Lee, Katsunobu Imai, and Kenichi Morita. Simulation of one-dimensional cellular automata by uniquely parallel parsable grammars. Theoretical Computer Science, 304(1–3):185– 200, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sheng:2003:RAT

[1587] Li Sheng. 2-role assignments on triangulated graphs. Theoretical Computer Science, 304(1–3):201–214, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feigenbaum:2003:HRM

[1588] Joan Feigenbaum, Arvind Krishnamurthy, Rahul Sami, and Scott Shenker. Hardness results for multicast cost sharing. Theoretical Computer Science, 304(1–3):215–236, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2003:PWC

[1589] Guangting Chen and Guoliang Xue. A PTAS for weight constrained Steiner trees in series–parallel graphs. Theoretical Computer Science, 304(1–3): 237–247, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertet:2003:WOE

[1590] Karell Bertet, Jens Gustedt, and Michel Morvan. Weak-order extensions of an order. Theoretical Computer Science, 304(1–3):249–268, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Formenti:2003:NCC

[1591] Enrico Formenti and Aristide Grange. Number conserving cellular automata II: dynamics. Theoretical Computer Science, 304(1–3):269–290, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:2003:RBE

[1592] Kosaburo Hashiguchi, Yoshito Wada, and Shuji Jimbo. Regular binoid expressions and regular binoid languages. Theoretical Computer Science, 304(1– 3):291–313, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fulop:2003:SPT

[1593] Zoltán Fülöp and Zsolt Gazdag. Shape preserving top-down tree transducers. Theoretical Computer Science, 304(1– 3):315–339, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crescenzi:2003:TSL

[1594] Pilu Crescenzi, Alberto Del Lungo, Roberto Grossi, Elena Lodi, Linda Pagli, and Gianluca Rossi. Text sparsification via local maxima. Theoretical Computer Science, 304(1–3):341– 364, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cao:2003:SPF

[1595] Wei-Ting Cao and Zhi-Ying Wen.

Some properties of the factors of Sturmian sequences. Theoretical Computer Science, 304(1–3):365–385, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fukuyama:2003:NGPa

[1596] Masahiko Fukuyama. A Nim game played on graphs. Theoretical Computer Science, 304(1–3):387–399, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fukuyama:2003:NGPb

[1597] Masahiko Fukuyama. A Nim game played on graphs II. Theoretical Computer Science, 304(1–3):401–419, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hachaichi:2003:DCA

[1598] Yassine Hachaïchi. A descriptive complexity approach to the linear hierarchy. Theoretical Computer Science, 304 (1–3):421–429, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hitchcock:2003:FDL

[1599] John M. Hitchcock. Fractal dimension and logarithmic loss unpredictability. Theoretical Computer Science, 304(1– 3):431–441, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Popov:2003:APP

[1600] V. Y. Popov. The approximate period problem for DNA alphabet. Theoretical Computer Science, 304(1–3): 443–447, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vagvolgyi:2003:GTT

[1601] Sándor Vágvölgyi. On ground tree transformations and congruences induced by tree automata. Theoretical Computer Science, 304(1–3):449– 459, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2003:SLC

[1602] Xiaoyan Cheng, Xiufeng Du, Manki Min, Hung Q. Ngo, Lu Ruan, Jianhua Sun, and Weili Wu. Super linkconnectivity of iterated line digraphs. Theoretical Computer Science, 304(1– 3):461–469, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demaine:2003:UEC

[1603] Erik D. Demaine, Alejandro López-Ortiz, and J. Ian Munro. On universally easy classes for NP-complete problems. Theoretical Computer Science, 304(1–3):471–476, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gasarch:2003:WDR

[1604] William Gasarch, Evan Golub, and Aravind Srinivasan. When does a random Robin Hood win? Theoretical Computer Science, 304(1–3):477– 484, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dartnell:2003:CCA

[1605] Pablo Dartnell, Alejandro Maass, and Fernando Schwartz. Combinatorial constructions associated to the dynamics of one-sided cellular automata. Theoretical Computer Science, 304(1–3): 485–497, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIn

[1606] Anonymous. Author index. Theoretical Computer Science, 304(1–3):499– 500, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBm

[1607] Anonymous. Editorial board. Theoretical Computer Science, 304(1–3): v–ix, July 28, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kopperman:2003:F

[1608] Ralph Kopperman, Michael B. Smyth, and Dieter Spreen. Foreword. Theoretical Computer Science, 305(1–3): 1–2, August 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amenta:2003:CTA

[1609] Nina Amenta, Thomas J. Peters, and Alexander C. Russell. Computational topology: ambient isotopic approximation of 2-manifolds. Theoretical Computer Science, 305(1–3):3–15, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brattka:2003:RQM

[1610] Vasco Brattka. Recursive quasi-metric spaces. Theoretical Computer Science, 305(1–3):17–42, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brattka:2003:CSM

[1611] Vasco Brattka and Gero Presser. Computability on subsets of metric spaces.

Theoretical Computer Science, 305(1– 3):43–76, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Coquand:2003:RSC

[1612] Thierry Coquand and Guo-Qiang Zhang. A representation of stably compact spaces, and patch topology. Theoretical Computer Science, 305(1–3): 77–84, August 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Curi:2003:CMP

[1613] Giovanni Curi. Constructive metrisability in point-free topology. Theoretical Computer Science, 305(1–3):85– 109, August 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galton:2003:GTV

[1614] Antony Galton. A generalized topological view of motion in discrete space. Theoretical Computer Science, 305(1– 3):111–134, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hardie:2003:NHQ

[1615] K. A. Hardie, S. Salbany, J. J. C.

Vermeulen, and P. J. Witbooi. A non-Hausdorff quaternion multiplication. Theoretical Computer Science, 305(1–3):135–158, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heckmann:2003:NTV

[1616] Reinhold Heckmann. A non-topological view of dcpos as convergence spaces.

Theoretical Computer Science, 305(1– 3):159–186, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hitzler:2003:GMU

[1617] Pascal Hitzler and Anthony Karel Seda. Generalized metrics and uniquely determined logic programs. Theoretical Computer Science, 305(1–3):187– 219, August 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kong:2003:KTP

[1618] T. Yung Kong. The Khalimsky topologies are precisely those simply connected topologies on Z^{n} whose connected sets include all 2^{n}-connected sets but no (3^{n} − 1)-disconnected sets. Theoretical Computer Science, 305(1– 3):221–235, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kummetz:2003:TMT

[1619] Ralph Kummetz and Dietrich Kuske. The topology of Mazurkiewicz traces. Theoretical Computer Science, 305(1– 3):237–258, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lawson:2003:REI

[1620] Jimmie D. Lawson and Bin Lu. Riemann and Edalat integration on domains. Theoretical Computer Science, 305(1–3):259–275, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2003:IMS

[1621] Keye Martin. Ideal models of spaces.

Theoretical Computer Science, 305(1– 3):277–297, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2003:RSC

[1622] Keye Martin. The regular spaces with countably based models. Theoretical Computer Science, 305(1–3):299– 310, August 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Resende:2003:LSL

[1623] Pedro Resende and Steven Vickers. Localic sup-lattices and tropological systems. Theoretical Computer Science, 305(1–3):311–346, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sambin:2003:SPF

[1624] Giovanni Sambin. Some points in formal topology. Theoretical Computer Science, 305(1–3):347–408, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schellekens:2003:CPM

[1625] M. P. Schellekens. A characterization of partial metrizability: domains are quantifiable. Theoretical Computer Science, 305(1–3):409–432, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schuster:2003:UEA

[1626] Peter M. Schuster. Unique existence, approximate solutions, and countable choice. Theoretical Computer Science, 305(1–3):433–455, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Slapal:2003:COD

[1627] J. Šlapal. Closure operations for digital topology. Theoretical Computer Science, 305(1–3):457–471, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vi:2003:CTP

[1628] L. S. Vî and D. S. Bridges. A constructive theory of point-set nearness. Theoretical Computer Science, 305(1– 3):473–489, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Webster:2003:CCO

[1629] Julian Webster. Cell complexes, oriented matroids and digital geometry. Theoretical Computer Science, 305(1– 3):491–502, August 18, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIo

[1630] Anonymous. Author index. Theoretical Computer Science, 305(1–3): 503, August 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBn

[1631] Anonymous. Editorial board. Theoretical Computer Science, 305(1–3):v– ix, August 18, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Farr:2003:GPG

[1632] G. E. Farr. The Go polynomials of a graph. Theoretical Computer Science, 306(1–3):1–18, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daley:2003:CDP

[1633] Mark Daley, Oscar H. Ibarra, and Lila Kari. Closure and decidability properties of some language classes with respect to ciliate bio-operations. Theoretical Computer Science, 306(1–3):19– 38, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watjen:2003:FDT

[1634] Dietmar Wätjen. Function-dependent teams in eco-grammar systems. Theoretical Computer Science, 306(1–3):39– 53, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2003:FST

[1635] Chin Lung Lu, Chuan Yi Tang, and Richard Chia-Tung Lee. The full Steiner tree problem. Theoretical Computer Science, 306(1–3):55–67, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balogh:2003:PCU

[1636] József Balogh, János A. Csirik, Yuval Ishai, and Eyal Kushilevitz. Private computation using a PEZ dispenser. Theoretical Computer Science, 306(1– 3):69–84, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Albert:2003:RCS

[1637] M. H. Albert, M. D. Atkinson, and N. Ruškuc. Regular closed sets of permutations. Theoretical Computer Science, 306(1–3):85–100, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sosik:2003:WCD

[1638] Petr Sosík. Watson–Crick D0L systems: generative power and undecidable problems. Theoretical Computer Science, 306(1–3):101–112, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andrei:2003:NAR

[1639] Ştefan Andrei, Salvador Valerio Cavadini, and Wei-Ngan Chin. A new algorithm for regularizing one-letter context-free grammars. Theoretical Computer Science, 306(1–3):113–122, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2003:CCF

[1640] Liang Zhang, K. P. Shum, and Shou-Li Peng. Completion of codes with finite bi-decoding delays. Theoretical Computer Science, 306(1–3):123– 137, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iwata:2003:NCP

[1641] Tetsu Iwata, Tomonobu Yoshino, and Kaoru Kurosawa. Non-cryptographic primitive for pseudorandom permutation. Theoretical Computer Science, 306(1–3):139–154, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Francez:2003:ACD

[1642] Nissim Francez and Michael Kaminski.

An algebraic characterization of deterministic regular languages over infinite alphabets. Theoretical Computer Science, 306(1–3):155–175, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambos-Spies:2003:ACS

[1643] Klaus Ambos-Spies, Wolfgang Merkle, Jan Reimann, and Sebastiaan A. Terwijn. Almost complete sets. Theoretical Computer Science, 306(1–3):177– 194, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bunimovich:2003:LGC

[1644] L. A. Bunimovich and D. M.

Kreslavskiy. Lorentz gas cellular automata on graphs. Theoretical Computer Science, 306(1–3):195–221, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonis:2003:CGS

[1645] Annalisa De Bonis and Ugo Vaccaro. Constructions of generalized superimposed codes with applications to group testing and conflict resolution in multiple access channels. Theoretical Computer Science, 306(1–3):223– 243, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rosaz:2003:WPC

[1646] Laurent Rosaz. The word problem for congruences is NP-hard. Theoretical Computer Science, 306(1–3):245– 268, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bienvenue:2003:GCE

[1647] Alexis Bienvenüe and Olivier François. Global convergence for evolution strategies in spherical problems: some simple proofs and difficulties. Theoretical Computer Science, 306(1–3):269– 289, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gravier:2003:MDG

[1648] Sylvain Gravier, Mehdi Mhalla, and Eric Tannier. On a modular domination game. Theoretical Computer Science, 306(1–3):291–303, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2003:MWS

[1649] Leah Epstein, Csanád Imreh, and Rob van Stee. More on weighted servers or F is better than L. Theoretical Computer Science, 306(1–3):305–317, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angel:2003:ATB

[1650] Eric Angel, Evripidis Bampis, and Alexander Kononov. On the approximate tradeoff for bicriteria batching and parallel machine scheduling problems. Theoretical Computer Science, 306(1–3):319–338, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Milidiu:2003:CMM

[1651] Ruy Luiz Milidiú, Artur Alves Pessoa, and Eduardo Sany Laber. The complexity of makespan minimization for pipeline transportation. Theoretical Computer Science, 306(1–3):339– 351, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sima:2003:ETC

[1652] Jiří Šíma and Pekka Orponen. Exponential transients in continuoustime Liapunov systems. Theoretical Computer Science, 306(1–3):353–372, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ilie:2003:RNI

[1653] Lucian Ilie and Sheng Yu. Reducing NFAs by invariant equivalences. Theoretical Computer Science, 306(1–3): 373–390, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dias:2003:SMP

[1654] Vânia M. F. Dias, Guilherme D.

da Fonseca, Celina M. H. de Figueiredo, and Jayme L. Szwarcfiter. The stable marriage problem with restricted pairs. Theoretical Computer Science, 306(1–3):391–405, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Evans:2003:CFC

[1655] Patricia A. Evans, Andrew D. Smith, and H. Todd Wareham. On the complexity of finding common approximate substrings. Theoretical Computer Science, 306(1–3):407–430, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halldorsson:2003:ARS

[1656] Magnús M. Halldórsson, Robert W. Irving, Kazuo Iwama, David F. Manlove, Shuichi Miyazaki, Yasufumi Morita, and Sandy Scott. Approximability results for stable marriage problems with ties. Theoretical Computer Science, 306(1–3):431–447, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meduna:2003:FEG

[1657] Alexander Meduna and Martin Švec. Forbidding ET0L grammars. Theoretical Computer Science, 306(1–3):449– 469, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dulucq:2003:RSS

[1658] Serge Dulucq and Laurent Tichit. RNA secondary structure comparison: exact analysis of the Zhang–Shasha tree edit algorithm. Theoretical Computer Science, 306(1–3):471–484, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dalmau:2003:LQF

[1659] Víctor Dalmau and Peter Jeavons. Learnability of quantified formulas. Theoretical Computer Science, 306(1– 3):485–511, September 5, 2003. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ndoundam:2003:ETL

[1660] René Ndoundam and Maurice Tchuente. Exponential transient length generated by a neuronal recurrence equation. Theoretical Computer Science, 306(1–3):513–533, September 5, 2003.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frid:2003:ACS

[1661] A. E. Frid. Arithmetical complexity of symmetric D0L words. Theoretical Computer Science, 306(1–3):535– 542, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:2003:AFP

[1662] Klaus Jansen and Roberto Solis-Oba.

An asymptotic fully polynomial time approximation scheme for bin covering. Theoretical Computer Science, 306(1–3):543–551, September 5, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIp

[1663] Anonymous. Author index. Theoretical Computer Science, 306(1–3):553– 554, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBo

[1664] Anonymous. Editorial board. Theoretical Computer Science, 306(1–3):v– x, September 5, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neraud:2003:P

[1665] J. Néraud. Preface. Theoretical Computer Science, 307(1):1, September 26, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:2003:RRS

[1666] Jean-Paul Allouche and Jeffrey Shallit. The ring of k-regular sequences, II. Theoretical Computer Science, 307 (1):3–29, September 26, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aberkane:2003:WWC

[1667] Ali Aberkane. Words whose complexity satisfies. Theoretical Computer Science, 307(1):31–46, September 26, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamczewski:2003:BFP

[1668] Boris Adamczewski. Balances for fixed points of primitive substitutions. Theoretical Computer Science, 307(1):47– 75, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ananichev:2003:IRW

[1669] D. S. Ananichev, A. Cherubini, and M. V. Volkov. Image reducing words and subgroups of free groups. Theoretical Computer Science, 307(1):77– 92, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ceccherini-Silberstein:2003:ERL

[1670] Tullio Ceccherini-Silberstein, Antonio Machì, and Fabio Scarabotti. On the entropy of regular languages. Theoretical Computer Science, 307(1):93–102, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ceccherini-Silberstein:2003:GSC

[1671] Tullio Ceccherini-Silberstein and Wolfgang Woess. Growth-sensitivity of context-free languages. Theoretical Computer Science, 307(1):103–116, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fory:2003:SCS

[1672] Wit Fory, Tomasz Krawczyk, and James A. Anderson. Semiretracts—a counterexample and some results. Theoretical Computer Science, 307(1):117– 127, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guesnet:2003:MSC

[1673] Yannick Guesnet. On maximal synchronous codes. Theoretical Computer Science, 307(1):129–138, September 26, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2003:IET

[1674] Tero Harju and Dirk Nowotka. On the independence of equations in three variables. Theoretical Computer Science, 307(1):139–172, September 26, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hohlweg:2003:LWP

[1675] Christophe Hohlweg and Christophe Reutenauer. Lyndon words, permutations and trees. Theoretical Computer Science, 307(1):173–178, September 26, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seebold:2003:LFP

[1676] Patrice Séébold. Lyndon factorization of the Prouhet words. Theoretical Computer Science, 307(1):179–197, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Silva:2003:HPT

[1677] Pedro V. Silva. The homomorphism problem for trace monoids. Theoretical Computer Science, 307(1):199–215, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBp

[1678] Anonymous. Editorial board. Theoretical Computer Science, 307(1):v– x, September 26, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barcucci:2003:P

[1679] Elena Barcucci and Alberto Del Lungo. Preface. Theoretical Computer Science, 307(2):219–220, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arques:2003:GLU

[1680] Didier Arquès and Anne Micheli. A generalization of the language of ukasiewicz coding rooted planar hypermaps. Theoretical Computer Science, 307(2):221–239, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonichon:2003:WUR

[1681] Nicolas Bonichon and Mohamed Mosbah. Watermelon uniform random generation with applications. Theoretical Computer Science, 307(2):241– 256, October 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bousquet-Melou:2003:WCQ

[1682] Mireille Bousquet-Mélou and Marko Petkovšek. Walks confined in a quadrant are not always D-finite. Theoretical Computer Science, 307(2):257– 276, October 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bousquet:2003:TBP

[1683] Michel Bousquet, Cedric Chauve, Gilbert Labelle, and Pierre Leroux. Two bijective proofs for the arborescent form of the Good–Lagrange formula and some applications to colored rooted trees and cacti. Theoretical Computer Science, 307(2):277–302, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chandran:2003:GCP

[1684] L. S. Chandran, L. Ibarra, F. Ruskey, and J. Sawada. Generating and characterizing the perfect elimination orderings of a chordal graph. Theoretical Computer Science, 307(2):303– 317, October 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deutsch:2003:BBD

[1685] Emeric Deutsch and Helmut Prodinger.

A bijection between directed columnconvex polyominoes and ordered trees of height at most three. Theoretical Computer Science, 307(2):319– 325, October 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferrari:2003:SBR

[1686] L. Ferrari, E. Grazzini, E. Pergola, and S. Rinaldi. Some bijective results about the area of Schröder paths. Theoretical Computer Science, 307(2):327– 335, October 7, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Labelle:2003:CPP

[1687] G. Labelle, C. Lamathe, and P. Leroux. A classification of plane and planar 2-trees. Theoretical Computer Science, 307(2):337–363, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Noy:2003:GDP

[1688] Marc Noy. Graphs determined by polynomial invariants. Theoretical Computer Science, 307(2):365–384, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Poulalhon:2003:BTP

[1689] Dominique Poulalhon and Gilles Schaeffer. A bijection for triangulations of a polygon with interior points and multiple edges. Theoretical Computer Science, 307(2):385–401, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shapiro:2003:BRG

[1690] Louis W. Shapiro. Bijections and the Riordan group. Theoretical Computer Science, 307(2):403–413, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vajnovszki:2003:LAG

[1691] Vincent Vajnovszki. A loopless algorithm for generating the permutations of a multiset. Theoretical Computer Science, 307(2):415–431, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Voge:2003:NHP

[1692] Markus Vöge and Anthony J. Guttmann.

On the number of hexagonal polyominoes. Theoretical Computer Science, 307(2):433–453, October 7, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2003:F

[1693] Josep Díaz and Tom Payne. Foreword.

Theoretical Computer Science, 307(3): 455, October 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harper:2003:ACA

[1694] L. H. Harper. Accidental combinatorist, an autobiography. Theoretical Computer Science, 307(3):457–472, October 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bezrukov:2003:EIP

[1695] Sergei L. Bezrukov and Robert Elsässer. Edge-isoperimetric problems for cartesian powers of regular graphs. Theoretical Computer Science, 307(3): 473–492, October 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bollobas:2003:US

[1696] Béla Bollobás and Imre Leader. Union of shadows. Theoretical Computer Science, 307(3):493–502, October 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calamoneri:2003:NRE

[1697] Tiziana Calamoneri, Annalisa Massini, and Imrich Vrt’o. New results on edge-bandwidth. Theoretical Computer Science, 307(3):503–513, October 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Canfield:2003:IPS

[1698] E. Rodney Canfield. Integer partitions and the Sperner property. Theoretical Computer Science, 307(3):515– 529, October 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaz:2003:BMM

[1699] J. Díaz, N. Do, M. J. Serna, and N. C. Wormald. Bounds on the max and min bisection of random cubic and random 4-regular graphs. Theoretical Computer Science, 307(3):531–547, October 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elsasser:2003:STS

[1700] R. Elsässer, R. Královi, and B. Monien. Sparse topologies with small spectrum size. Theoretical Computer Science, 307(3):549–565, October 14, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muradian:2003:BMP

[1701] David Muradian. The bandwidth minimization problem for cyclic caterpillars with hair length 1 is NP-complete. Theoretical Computer Science, 307(3):567– 572, October 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIq

[1702] Anonymous. Author index. Theoretical Computer Science, 307(3):573– 574, October 14, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rutten:2003:BDE

[1703] J. J. M. M. Rutten. Behavioural differential equations: a coinductive calculus of streams, automata, and power series. Theoretical Computer Science, 308(1– 3):1–53, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2003:TLP

[1704] Yuxi Fu and Zhenrong Yang. Tau laws for pi calculus. Theoretical Computer Science, 308(1–3):55–130, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gumm:2003:DCC

[1705] H. Peter Gumm, Jesse Hughes, and Tobias Schröder. Distributivity of categories of coalgebras. Theoretical Computer Science, 308(1–3):131– 143, November 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Keller:2003:CPN

[1706] Walter Keller. Clustering for Petri nets. Theoretical Computer Science, 308(1–3):145–197, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Afrati:2003:LDP

[1707] Foto Afrati, Manolis Gergatsoulis, and Francesca Toni. Linearisability on datalog programs. Theoretical Computer Science, 308(1–3):199–226, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nour:2003:SPC

[1708] Karim Nour and Christophe Raffalli. Simple proof of the completeness theorem for second-order classical and intuitionistic logic by reduction to firstorder mono-sorted logic. Theoretical Computer Science, 308(1–3):227– 237, November 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danvy:2003:FOO

[1709] Olivier Danvy and Lasse R. Nielsen. A first-order one-pass CPS transformation. Theoretical Computer Science, 308(1–3):239–257, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krivine:2003:DCQ

[1710] Jean-Louis Krivine. Dependent choice, ‘quote’ and the clock. Theoretical Computer Science, 308(1–3):259–276, November 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Charatonik:2003:MCM

[1711] Witold Charatonik, Silvano Dal Zilio, Andrew D. Gordon, Supratik Mukhopadhyay, and Jean-Marc Talbot. Model checking mobile ambients. Theoretical Computer Science, 308(1–3):277–331, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ker:2003:AIG

[1712] Andrew D. Ker, Hanno Nickau, and C.-H. Luke Ong. Adapting innocent game models for the Böhm tree λtheory. Theoretical Computer Science, 308(1–3):333–366, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2003:FTN

[1713] Yifeng Chen. A fixpoint theory for non-monotonic parallelism. Theoretical Computer Science, 308(1–3):367– 392, November 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mateus:2003:CFR

[1714] P. Mateus, M. Morais, C. Nunes, A. Pacheco, A. Sernadas, and C. Sernadas. Categorical foundations for randomly timed automata. Theoretical Computer Science, 308(1–3):393– 427, November 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Loyer:2003:PSL

[1715] Yann Loyer, Nicolas Spyratos, and Daniel Stamate. Parametrized semantics of logic programs—a unifying framework. Theoretical Computer Science, 308(1–3):429–447, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abraham:2003:SST

[1716] Uri Abraham. Self-stabilizing timestamps. Theoretical Computer Science, 308(1–3):449–515, November 3, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIr

[1717] Anonymous. Author index. Theoretical Computer Science, 308(1–3): 517, November 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBq

[1718] Anonymous. Editorial board. Theoretical Computer Science, 308(1–3):v– x, November 3, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrhard:2003:DLC

[1719] Thomas Ehrhard and Laurent Regnier. The differential lambda-calculus. Theoretical Computer Science, 309(1–3):1– 41, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Selinger:2003:OIF

[1720] Peter Selinger. Order-incompleteness and finite lambda reduction models. Theoretical Computer Science, 309(1– 3):43–63, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galpin:2003:FSE

[1721] Vashti Galpin. A format for semantic equivalence comparison. Theoretical Computer Science, 309(1–3):65– 109, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Forster:2003:BQO

[1722] Thomas Forster. Better-quasi-orderings and coinduction. Theoretical Computer Science, 309(1–3):111– 123, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hermida:2003:PIP

[1723] Claudio Hermida and Paulo Mateus.

Paracategories I: internal paracategories and saturated partial algebras. Theoretical Computer Science, 309(1– 3):125–156, December 2, 2003. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gardiner:2003:PSR

[1724] Paul Gardiner. Power simulation and its relation to traces and failures refinement. Theoretical Computer Science, 309(1–3):157–176, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pattinson:2003:CML

[1725] Dirk Pattinson. Coalgebraic modal logic: soundness, completeness and decidability of local consequence. Theoretical Computer Science, 309(1–3): 177–193, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vagvolgyi:2003:RLH

[1726] Sándor Vágvölgyi. Right-linear halfmonadic term rewrite systems. Theoretical Computer Science, 309(1–3): 195–211, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Strassburger:2003:MCS

[1727] Lutz Straßburger. MELL in the calculus of structures. Theoretical Computer Science, 309(1–3):213–285, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2003:NCI

[1728] Zhaohui Zhu, Xian Xiao, Yong Zhou, and Wujia Zhu. Normal conditions for inference relations and injective models. Theoretical Computer Science, 309(1–3):287–311, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:2003:BTO

[1729] Jan A. Bergstra, Alban Ponse, and Mark B. van der Zwaag. Branching time and orthogonal bisimulation equivalence. Theoretical Computer Science, 309(1–3):313–355, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duran:2003:STI

[1730] Francisco Durán and José Meseguer. Structured theories and institutions. Theoretical Computer Science, 309(1– 3):357–380, December 2, 2003. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kirsten:2003:TTA

[1731] Daniel Kirsten and Jerzy Marcinkowski. Two techniques in the area of the star problem in trace monoids. Theoretical Computer Science, 309(1–3):381– 412, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boigelot:2003:ILT

[1732] Bernard Boigelot. On iterating linear transformations over recognizable sets of integers. Theoretical Computer Science, 309(1–3):413–468, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ghica:2003:RLS

[1733] Dan R. Ghica and Guy McCusker. The regular-language semantics of secondorder idealized Algol. Theoretical Computer Science, 309(1–3):469–502, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lanotte:2003:CTA

[1734] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Simone Tini. Concurrency in timed automata. Theoretical Computer Science, 309(1–3):503– 527, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lohrey:2003:RHL

[1735] Markus Lohrey. Realizability of highlevel message sequence charts: closing the gaps. Theoretical Computer Science, 309(1–3):529–554, December 2, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:AIs

[1736] Anonymous. Author index. Theoretical Computer Science, 309(1–3): 555, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2003:EBr

[1737] Anonymous. Editorial Board. Theoretical Computer Science, 309(1–3):v– x, December 2, 2003. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dai:2004:FSD

[1738] Jack J. Dai, James I. Lathrop, Jack H. Lutz, and Elvira Mayordomo. Finitestate dimension. Theoretical Computer Science, 310(1–3):1–33, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chemillier:2004:SMW

[1739] Marc Chemillier. Synchronization of musical words. Theoretical Computer Science, 310(1–3):35–60, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tse:2004:NBM

[1740] Savio S. H. Tse and Francis C. M. Lau. New bounds for multi-label interval routing. Theoretical Computer Science, 310(1–3):61–77, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Serre:2004:VLL

[1741] Olivier Serre. Vectorial languages and linear temporal logic. Theoretical Computer Science, 310(1–3):79– 116, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choquet-Geniet:2004:MSI

[1742] Annie Choquet-Geniet and Emmanuel Grolleau. Minimal schedulability interval for real-time systems of periodic tasks with offsets. Theoretical Computer Science, 310(1–3):117– 134, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angel:2004:APC

[1743] Eric Angel, Evripidis Bampis, and Laurent Gourvès. Approximating the Pareto curve with local search for the bicriteria TSP(1,2) problem. Theoretical Computer Science, 310(1–3):135– 146, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leung:2004:LPD

[1744] Hing Leung and Viktor Podolskiy. The limitedness problem on distance automata: Hashiguchi’s method revisited. Theoretical Computer Science, 310(1–3):147–158, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lavi:2004:CAI

[1745] Ron Lavi and Noam Nisan. Competitive analysis of incentive compatible on-line auctions. Theoretical Computer Science, 310(1–3):159–180, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schmitt:2004:TGA

[1746] Lothar M. Schmitt. Theory of Genetic Algorithms II: models for genetic operators over the string-tensor representation of populations and convergence to global optima for arbitrary fitness function under scaling. Theoretical Computer Science, 310(1–3):181– 231, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Serbanuta:2004:EPM

[1747] Traian-Florin Şerbănuţă. Extending Parikh matrices. Theoretical Computer Science, 310(1–3):233–246, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jenkinson:2004:CBW

[1748] Oliver Jenkinson and Luca Q. Zamboni. Characterisations of balanced words via orderings. Theoretical Computer Science, 310(1–3):247–271, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chuan:2004:MCC

[1749] Wai-Fong Chuan. Moments of conjugacy classes of binary words. Theoretical Computer Science, 310(1–3):273– 285, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biedl:2004:FHI

[1750] Therese Biedl, Brovňa Brejová, Erik D. Demaine, Angèle M. Hamel, Alejandro López-Ortiz, and Tomáš Vinař. Finding hidden independent sets in interval graphs. Theoretical Computer Science, 310(1–3):287–307, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Groult:2004:SDE

[1751] Richard Groult, Martine Léonard, and Laurent Mouchard. Speeding up the detection of evolutive tandem repeats. Theoretical Computer Science, 310(1– 3):309–328, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandstadt:2004:TSC

[1752] Andreas Brandstädt, Feodor F. Dragan, Hoàng-Oanh Le, and Van Bang Le. Tree spanners on chordal graphs: complexity and algorithms. Theoretical Computer Science, 310(1–3):329– 354, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mityagin:2004:CFL

[1753] Anton Mityagin. On the complexity of finding a local maximum of functions on discrete planar subsets. Theoretical Computer Science, 310(1–3):355– 363, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bjorklund:2004:MDP

[1754] Henrik Björklund, Sven Sandberg, and Sergei Vorobyov. Memoryless determinacy of parity and mean payoff games: a simple proof. Theoretical Computer Science, 310(1–3):365–378, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manabe:2004:AME

[1755] Yoshifumi Manabe and Naka Tajima.

(h, k)-arbiters for h-out-of-k mutual exclusion problem. Theoretical Computer Science, 310(1–3):379–392, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kunc:2004:UTC

[1756] Michal Kunc. Undecidability of the trace coding problem and some decidable cases. Theoretical Computer Science, 310(1–3):393–456, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ahn:2004:CFL

[1757] Hee-Kap Ahn, Siu-Wing Cheng, Otfried Cheong, Mordecai Golin, and René van Oostrum. Competitive facility location: the Voronoi game. Theoretical Computer Science, 310(1–3): 457–467, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cervelle:2004:TRR

[1758] Julien Cervelle and Bruno Durand. Tilings: recursivity and regularity. Theoretical Computer Science, 310(1– 3):469–477, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dukes:2004:CEB

[1759] P. Dukes and Alan C. H. Ling. A combinatorial error bound for t-point-based sampling. Theoretical Computer Science, 310(1–3):479–488, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chalopin:2004:FFF

[1760] Jérémie Chalopin and Hing Leung. On factorization forests of finite height. Theoretical Computer Science, 310(1– 3):489–499, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhou:2004:CAP

[1761] Sanming Zhou. A channel assignment problem for optical networks modelled by Cayley graphs. Theoretical Computer Science, 310(1–3):501– 511, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alekhnovich:2004:MCP

[1762] Michael Alekhnovich. Mutilated chessboard problem is exponentially hard for resolution. Theoretical Computer Science, 310(1–3):513–525, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lopez-Ortiz:2004:LPH

[1763] Alejandro López-Ortiz and Sven Schuierer. On-line parallel heuristics, processor scheduling and robot searching under the competitive framework. Theoretical Computer Science, 310(1– 3):527–537, January 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIa

[1764] Anonymous. Author index. Theoretical Computer Science, 310(1–3):539– 540, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:MIa

[1765] Anonymous. Master index. Theoretical Computer Science, 310(1–3):541– 553, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBa

[1766] Anonymous. Editorial Board. Theoretical Computer Science, 310(1–3): v–x, January 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kfoury:2004:PTI

[1767] A. J. Kfoury and J. B. Wells. Principality and type inference for intersection types using expansion variables. Theoretical Computer Science, 311(1–3):1– 70, January 23, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hermida:2004:PIA

[1768] Claudio Hermida and P. Mateus. Paracategories II: adjunctions, fibrations and examples from probabilistic automata theory. Theoretical Computer Science, 311(1–3):71–103, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Joachimski:2004:CCC

[1769] Felix Joachimski. Confluence of the coinductive λ-calculus. Theoretical Computer Science, 311(1–3):105–119, January 23, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Igarashi:2004:GTS

[1770] Atsushi Igarashi and Naoki Kobayashi. A generic type system for the Pi-calculus. Theoretical Computer Science, 311(1–3):121–163, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiao:2004:LBA

[1771] Li Jiao, To-Yat Cheung, and Weiming Lu. On liveness and boundedness of asymmetric choice nets. Theoretical Computer Science, 311(1–3):165– 197, January 23, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akama:2004:LPC

[1772] Yohji Akama. Limiting partial combinatory algebras. Theoretical Computer Science, 311(1–3):199–220, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dovier:2004:EAC

[1773] Agostino Dovier, Carla Piazza, and Alberto Policriti. An efficient algorithm for computing bisimulation equivalence. Theoretical Computer Science, 311(1–3):221–256, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:2004:TCC

[1774] J. Adámek and H.-E. Porst. On tree coalgebras and coalgebra presentations. Theoretical Computer Science, 311(1– 3):257–283, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doberkat:2004:MFR

[1775] Ernst-Erich Doberkat and Eugenio G.

Omodeo. ER modelling from first relational principles. Theoretical Computer Science, 311(1–3):285–323, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller-Olm:2004:PID

[1776] Markus Müller-Olm. Precise interprocedural dependence analysis of parallel programs. Theoretical Computer Science, 311(1–3):325–388, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bailey:2004:DTP

[1777] James Bailey, Guozhu Dong, and Kotagiri Ramamohanarao. On the decidability of the termination problem of active database systems. Theoretical Computer Science, 311(1–3):389– 437, January 23, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaconescu:2004:IGI

[1778] Răzvan Diaconescu. Interpolation in Grothendieck Institutions. Theoretical Computer Science, 311(1–3):439– 461, January 23, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanGlabbeek:2004:WBF

[1779] Rob van Glabbeek and Ursula Goltz. Well-behaved flow event structures for parallel composition and action refinement. Theoretical Computer Science, 311(1–3):463–478, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alpuente:2004:RST

[1780] María Alpuente, Moreno Falaschi, Ginés Moreno, and Germán Vidal. Rules + strategies for transforming lazy functional logic programs. Theoretical Computer Science, 311(1–3): 479–525, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkelstein:2004:ENF

[1781] Stacy E. Finkelstein, Peter Freyd, and James Lipton. Erratum to: “A new framework for declarative programming”: [Theoret. Comput. Sci. 300 (2003) 91–160]. Theoretical Computer Science, 311(1–3):527, January 23, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1476].

Anonymous:2004:AIb

[1782] Anonymous. Author index. Theoretical Computer Science, 311(1–3): 529, January 23, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBb

[1783] Anonymous. Editorial Board. Theoretical Computer Science, 311(1–3):v– x, January 23, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eidenbenz:2004:P

[1784] Stephan Eidenbenz, Matthew Hennessy, Rafael Morales, Francisco Triguero, Peter Widmayer, and Ricardo Conejo. Preface. Theoretical Computer Science, 312(1):1–2, January 26, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Charikar:2004:FFI

[1785] Moses Charikar, Kevin Chen, and Martin Farach-Colton. Finding frequent items in data streams. Theoretical Computer Science, 312(1):3– 15, January 26, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engebretsen:2004:IRE

[1786] Lars Engebretsen, Jonas Holmerin, and Alexander Russell. Inapproximability results for equations over finite groups. Theoretical Computer Science, 312(1): 17–45, January 26, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pettie:2004:NAA

[1787] Seth Pettie. A new approach to allpairs shortest paths on real-weighted graphs. Theoretical Computer Science, 312(1):47–74, January 26, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Radzik:2004:ITB

[1788] Tomasz Radzik. Improving time bounds on maximum generalised flow computations by contracting the network. Theoretical Computer Science, 312(1):75–97, January 26, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2004:MPP

[1789] Keye Martin, Michael Mislove, and James Worrell. Measuring the probabilistic powerdomain. Theoretical Computer Science, 312(1):99–119, January 26, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ong:2004:GCL

[1790] C.-H. L. Ong and P. Di Gianantonio. Games characterizing Levy– Longo trees. Theoretical Computer Science, 312(1):121–142, January 26, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBc

[1791] Anonymous. Editorial Board. Theoretical Computer Science, 312(1):v– x, January 26, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freund:2004:RRC

[1792] Rudolf Freund, Carlos Martín-Vide, and Gheorghe Păun. From regulated rewriting to computing with membranes: collapsing hierarchies. Theoretical Computer Science, 312(2–3):143– 188, January 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rao:2004:PQP

[1793] Nageswara S. V. Rao. Probabilistic quickest path algorithm. Theoretical Computer Science, 312(2–3):189– 201, January 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terrier:2004:TDC

[1794] Véronique Terrier. Two-dimensional cellular automata and their neighborhoods. Theoretical Computer Science, 312(2–3):203–222, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vialette:2004:CCI

[1795] Stéphane Vialette. On the computational complexity of 2-interval pattern matching problems. Theoretical Computer Science, 312(2–3):223– 249, January 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:2004:ERB

[1796] Kosaburo Hashiguchi, Naoto Sakakibara, and Shuji Jimbo. Equivalence of regular binoid expressions and regular expressions denoting binoid languages over free binoids. Theoretical Computer Science, 312(2–3):251–266, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karuno:2004:ARM

[1797] Yoshiyuki Karuno and Hiroshi Nagamochi. An approximability result of the multi-vehicle scheduling problem on a path with release and handling times. Theoretical Computer Science, 312(2–3):267–280, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yi:2004:GEC

[1798] Zhang Yi. Global exponential convergence of recurrent neural networks with variable delays. Theoretical Computer Science, 312(2–3):281–293, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frisco:2004:CPS

[1799] Pierluigi Frisco. The conformon-P system: a molecular and cell biologyinspired computability model. Theoretical Computer Science, 312(2–3):295– 319, January 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2004:CIR

[1800] Chuan-Kun Wu and Ed Dawson. Correlation immunity and resiliency of symmetric Boolean functions. Theoretical Computer Science, 312(2–3):321– 335, January 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alber:2004:CST

[1801] Jochen Alber, Jens Gramm, Jiong Guo, and Rolf Niedermeier. Computing the similarity of two sequences with nested arc annotations. Theoretical Computer Science, 312(2–3):337–358, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2004:TWF

[1802] Oscar H. Ibarra and Zhe Dang. On two-way FA with monotonic counters and quadratic Diophantine equations. Theoretical Computer Science, 312(2– 3):359–378, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2004:CPS

[1803] Oscar H. Ibarra, Zhe Dang, and Omer Egecioglu. Catalytic P systems, semilinear sets, and vector addition systems. Theoretical Computer Science, 312(2–3):379–399, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Habib:2004:CAD

[1804] M. Habib, L. Nourine, O. Raynaud, and E. Thierry. Computational aspects of the 2-dimension of partially ordered sets. Theoretical Computer Science, 312(2–3):401–431, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gomez:2004:SPV

[1805] Antonio Cano Gómez and Jean-Éric Pin. Shuffle on positive varieties of languages. Theoretical Computer Science, 312(2–3):433–461, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doerr:2004:TRP

[1806] Benjamin Doerr. Typical rounding problems. Theoretical Computer Science, 312(2–3):463–477, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Qiu:2004:CQA

[1807] Daowen Qiu and Mingsheng Ying. Characterizations of quantum automata. Theoretical Computer Science, 312(2–3):479–489, January 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIc

[1808] Anonymous. Author index. Theoretical Computer Science, 312(2–3):491– 492, January 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watson:2004:I

[1809] Bruce W. Watson and Derick Wood. Introduction. Theoretical Computer Science, 313(1):1, February 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergeron:2004:CDB

[1810] Anne Bergeron and Sylvie Hamel. From cascade decompositions to bitvector algorithms. Theoretical Computer Science, 313(1):3–16, February 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boigelot:2004:CSP

[1811] Bernard Boigelot and Louis Latour. Counting the solutions of Presburger equations without enumerating them. Theoretical Computer Science, 313(1): 17–29, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Champarnaud:2004:DRE

[1812] Jean-Marc Champarnaud and Gérard Duchamp. Derivatives of rational expressions and related theorems. Theoretical Computer Science, 313(1):31– 44, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daciuk:2004:FAC

[1813] Jan Daciuk and Gertjan van Noord.

Finite automata for compact representation of tuple dictionaries. Theoretical Computer Science, 313(1):45– 56, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dang:2004:PPT

[1814] Zhe Dang, Tevfik Bultan, Oscar H.

Ibarra, and Richard A. Kemmerer. Past pushdown timed automata and safety verification. Theoretical Computer Science, 313(1):57–71, February 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Farre:2004:BCN

[1815] Jacques Farré and J. Fortes Gálvez.

Bounded-connect noncanonical discriminatingreverse parsers. Theoretical Computer Science, 313(1):73–91, February 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Friburger:2004:FST

[1816] N. Friburger and D. Maurel. Finitestate transducer cascades to extract named entities in texts. Theoretical Computer Science, 313(1):93–104, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gaal:2004:DSF

[1817] Tamás Gaál. Deciding sequentiability of finite-state transducers by finite-state pattern-matching. Theoretical Computer Science, 313(1):105– 117, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geniet:2004:SHS

[1818] Dominique Geniet and Jean-Philippe Dubernard. Scheduling hard sporadic tasks with regular languages and generating functions. Theoretical Computer Science, 313(1):119–132, February 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Katritzke:2004:EPT

[1819] F. Katritzke, W. Merzenich, and M. Thomas. Enhancements of partitioning techniques for image compression using weighted finite automata. Theoretical Computer Science, 313(1): 133–144, February 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kempe:2004:ERI

[1820] André Kempe. Extraction and recoding of input-ϵ-cycles in finite state transducers. Theoretical Computer Science, 313(1):145–158, February 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanZijl:2004:GAS

[1821] Lynette van Zijl. Generalized acceptance, succinctness and supernondeterministic finite automata. Theoretical Computer Science, 313(1):159– 172, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBd

[1822] Anonymous. Editorial Board. Theoretical Computer Science, 313(1):v– ix, February 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abe:2004:F

[1823] N. Abe and R. Khardon. Foreword.

Theoretical Computer Science, 313(2): 173–174, February 17, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angluin:2004:QR

[1824] Dana Angluin. Queries revisited. Theoretical Computer Science, 313(2):175– 194, February 17, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kalnishkan:2004:LFC

[1825] Yuri Kalnishkan, Volodya Vovk, and Michael V. Vyugin. Loss functions, complexities, and the Legendre transformation. Theoretical Computer Science, 313(2):195–207, February 17, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2004:LHS

[1826] Sanjay Jain and Frank Stephan. Learning how to separate. Theoretical Computer Science, 313(2):209–228, February 17, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zilles:2004:SUL

[1827] Sandra Zilles. Separation of uniform learning classes. Theoretical Computer Science, 313(2):229–265, February 17, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Denis:2004:LRL

[1828] François Denis, Aurélien Lemay, and Alain Terlutte. Learning regular languages using RFSAs. Theoretical Computer Science, 313(2):267–294, February 17, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

delaHiguera:2004:ILP

[1829] C. de la Higuera and J. C. Janodet. Inference of ω-languages from prefixes. Theoretical Computer Science, 313(2): 295–312, February 17, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fleischer:2004:P

[1830] Rudolf Fleischer and Richard J. Nowakowski. Preface. Theoretical Computer Science, 313(3):313, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Althofer:2004:IGP

[1831] Ingo Althöfer. Improved game play by multiple computer hints. Theoretical Computer Science, 313(3):315– 324, February 19, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demaine:2004:SC

[1832] Erik D. Demaine, Martin L. Demaine, and Rudolf Fleischer. Solitaire Clobber. Theoretical Computer Science, 313(3):325–338, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doerr:2004:ETG

[1833] Benjamin Doerr. European tenure games. Theoretical Computer Science, 313(3):339–351, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dumitriu:2004:HG

[1834] Ioana Dumitriu and Joel Spencer. A Halfliar’s game. Theoretical Computer Science, 313(3):353–369, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Erdos:2004:NGC

[1835] Peter L. Erdős, Ulrich Faigle, Winfried Hochstättler, and Walter Kern. Note on the game chromatic index of trees. Theoretical Computer Science, 313(3): 371–376, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fekete:2004:TSP

[1836] Sándor P. Fekete, Rudolf Fleischer, Aviezri Fraenkel, and Matthias Schmitt. Traveling salesmen in the presence of competition. Theoretical Computer Science, 313(3):377–392, February 19, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fraenkel:2004:CAC

[1837] Aviezri S. Fraenkel. Complexity, appeal and challenges of combinatorial games. Theoretical Computer Science, 313(3): 393–415, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grossman:2004:POD

[1838] J. P. Grossman. Periodicity in onedimensional peg duotaire. Theoretical Computer Science, 313(3):417–425, February 19, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harary:2004:LSB

[1839] F. Harary, W. Slany, and O. Verbitsky. On the lengths of symmetry breaking– preserving games on graphs. Theoretical Computer Science, 313(3):427– 446, February 19, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holzer:2004:AMA

[1840] Markus Holzer and Stefan Schwoon. Assembling molecules in ATOMIX is hard. Theoretical Computer Science, 313(3):447–462, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Howse:2004:PAP

[1841] S. Howse and R. J. Nowakowski. Periodicity and arithmetic-periodicity in hexadecimal games. Theoretical Computer Science, 313(3):463–472, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kral:2004:ITP

[1842] Daniel Král’, Vladan Majerech, Jiří Sgall, Tomáš Tichý, and Gerhard Woeginger. It is tough to be a plumber. Theoretical Computer Science, 313(3): 473–484, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lorenz:2004:EAM

[1843] U. Lorenz and B. Monien. Error analysis in minimax trees. Theoretical Computer Science, 313(3):485–498, February 19, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Snatzke:2004:NRE

[1844] Raymond Georg Snatzke. New results of exhaustive search in the game Amazons. Theoretical Computer Science, 313(3):499–509, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Winands:2004:ETL

[1845] Mark H. M. Winands, Jos W. H. M.

Uiterwijk, and H. Jaap van den Herik. An effective two-level proof-number search algorithm. Theoretical Computer Science, 313(3):511–525, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wolfe:2004:CNG

[1846] David Wolfe and William Fraser.

Counting the number of games. Theoretical Computer Science, 313(3):527– 532, February 19, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grossman:2004:ARF

[1847] J. P. Grossman. Appendix A: Report on the First International Clobber Tournament. Theoretical Computer Science, 313(3):533–537, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demaine:2004:ABO

[1848] Erik D. Demaine, Rudolf Fleischer, Aviezri S. Fraenkel, and Richard J. Nowakowski. Appendix B: Open problems at the 2002 Dagstuhl Seminar on Algorithmic Combinatorial Game Theory. Theoretical Computer Science, 313(3):539–543, February 19, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AId

[1849] Anonymous. Author index. Theoretical Computer Science, 313(3):545– 546, February 19, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feder:2004:DCH

[1850] Tomás Feder, Florent Madelaine, and Iain A. Stewart. Dichotomies for classes of homomorphism problems involving unary functions. Theoretical Computer Science, 314(1–2):1–43, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deppe:2004:SRU

[1851] Christian Deppe. Strategies for the Renyi–Ulam Game with fixed number of lies. Theoretical Computer Science, 314(1–2):45–55, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchi:2004:OGE

[1852] Enrica Duchi, Jean-Marc Fedou, and Simone Rinaldi. From object grammars to ECO systems. Theoretical Computer Science, 314(1–2):57–95, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sano:2004:BWG

[1853] Shinya Sano, Naoto Miyoshi, and Ryohei Kataoka. m-balanced words: A generalization of balanced words. Theoretical Computer Science, 314(1–2): 97–120, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ravelomanana:2004:FSC

[1854] Vlady Ravelomanana and Loÿs Thimonier. Forbidden subgraphs in connected graphs. Theoretical Computer Science, 314(1–2):121–171, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2004:TDP

[1855] Amihood Amir, Ayelet Butman, Maxime Crochemore, Gad M. Landau, and Mary Schaps. Two-dimensional pattern matching with rotations. Theoretical Computer Science, 314(1–2): 173–187, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet-Sadri:2004:LPB

[1856] F. Blanchet-Sadri and Ajay Chriscoe.

Local periods and binary partial words: an algorithm. Theoretical Computer Science, 314(1–2):189–216, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angiulli:2004:CIC

[1857] Fabrizio Angiulli, Giovambattista Ianni, and Luigi Palopoli. On the complexity of inducing categorical and quantitative association rules. Theoretical Computer Science, 314(1–2):217– 249, February 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yao:2004:SPI

[1858] Jia-Yan Yao. Some properties of Ising automata. Theoretical Computer Science, 314(1–2):251–279, February 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rybalov:2004:PNP

[1859] A. Rybalov. On the P–NP problem over real matrix rings. Theoretical Computer Science, 314(1–2):281– 285, February 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burckel:2004:SCL

[1860] Serge Burckel and Marianne Morillon. Sequential computation of linear Boolean mappings. Theoretical Computer Science, 314(1–2):287–292, February 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gu:2004:RAM

[1861] Jun Gu, Xiao-Dong Hu, Xiaohua Jia, and Mu-Hong Zhang. Routing algorithm for multicast under multi-tree model in optical networks. Theoretical Computer Science, 314(1–2):293– 301, February 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBe

[1862] Anonymous. Editorial Board. Theoretical Computer Science, 314(1–2):v– ix, February 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Noguchi:2004:SSM

[1863] Kenichiro Noguchi. Simple 8-state minimal time solution to the firing squad synchronization problem. Theoretical Computer Science, 314(3):303– 334, April 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dimitri:2004:EEE

[1864] Nicola Dimitri. Efficiency and equilibrium in the electronic mail game; the general case. Theoretical Computer Science, 314(3):335–349, April 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonis:2004:RSS

[1865] Annalisa De Bonis and Alfredo De Santis. Randomness in secret sharing and visual cryptography schemes. Theoretical Computer Science, 314(3):351– 374, April 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2004:WSN

[1866] F. H. Chang, J. Y. Guo, F. K. Hwang, and C. K. Lin. Wide-sense nonblocking for symmetric or asymmetric 3-stage Clos networks under various routing strategies. Theoretical Computer Science, 314(3):375–386, April 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fulop:2004:HTS

[1867] Zoltán Fülöp, Zsolt Gazdag, and Heiko Vogler. Hierarchies of tree series transformations. Theoretical Computer Science, 314(3):387–429, April 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tsai:2004:LAR

[1868] Chang-Hsiung Tsai. Linear array and ring embeddings in conditional faulty hypercubes. Theoretical Computer Science, 314(3):431–443, April 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bordihn:2004:CFP

[1869] Henning Bordihn. Context-freeness of the power of context-free languages is undecidable. Theoretical Computer Science, 314(3):445–449, April 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domaratzki:2004:RRE

[1870] Michael Domaratzki and Alexander Okhotin. Representing recursively enumerable languages by iterated deletion. Theoretical Computer Science, 314(3): 451–457, April 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elmasry:2004:SAT

[1871] Amr Elmasry. On the sequential access theorem and deque conjecture for splay trees. Theoretical Computer Science, 314(3):459–466, April 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIe

[1872] Anonymous. Author index. Theoretical Computer Science, 314(3):467– 468, April 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mislove:2004:MFP

[1873] Michael Mislove. Mathematical Foundations of Programming Semantics: Papers from MFPS 14 and MFPS 16. Theoretical Computer Science, 315(1): 1–2, May 5, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berardi:2004:BCW

[1874] S. Berardi and C. Berline. Building continuous webbed models for system F. Theoretical Computer Science, 315 (1):3–34, May 5, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bauer:2004:ES

[1875] Andrej Bauer, Lars Birkedal, and D. S.

Dana S. Scott. Equilogical spaces. Theoretical Computer Science, 315(1): 35–59, May 5, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calcagno:2004:TLL

[1876] Cristiano Calcagno. Two-level languages for program optimization. Theoretical Computer Science, 315(1):61– 81, May 5, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schalk:2004:PVS

[1877] Andrea Schalk and Valeria de Paiva. Poset-valued sets or how to build models for linear logics. Theoretical Computer Science, 315(1):83–107, May 5, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Konecny:2004:RFI

[1878] Michal Konečný. Real functions incrementally computable by finite automata. Theoretical Computer Science, 315(1):109–133, May 5, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schellekens:2004:CBP

[1879] M. P. Schellekens. The correspondence between partial metrics and semivaluations. Theoretical Computer Science, 315(1):135–149, May 5, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2004:ETM

[1880] Zhe Yang. Encoding types in MLlike languages. Theoretical Computer Science, 315(1):151–190, May 5, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Furedi:2004:MLT

[1881] Z. Füredi and R. P. Kurshan. Minimal length test vectors for multiplefault detection. Theoretical Computer Science, 315(1):191–208, May 5, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lowe:2004:SMI

[1882] Gavin Lowe. Semantic models for information flow. Theoretical Computer Science, 315(1):209–256, May 5, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pym:2004:PWR

[1883] David J. Pym, P. W. Peter W.

O’Hearn, and Hongseok Yang. Possible worlds and resources: the semantics of BI. Theoretical Computer Science, 315(1):257–305, May 5, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBf

[1884] Anonymous. Editorial board. Theoretical Computer Science, 315(1):v– ix, May 5, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Emiris:2004:P

[1885] Ioannis Z. Emiris, Bernard Mourrain, and Victor Y. Pan. Preface. Theoretical Computer Science, 315(2–3): 307–308, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bai:2004:IEC

[1886] Zheng-Jian Bai and R. H. Raymond H.

Chan. Inverse eigenproblem for centrosymmetric and centroskew matrices and their approximation. Theoretical Computer Science, 315(2–3): 309–318, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bini:2004:BBM

[1887] D. A. Bini and L. Gemignani. Bernstein–Bezoutian matrices. Theoretical Computer Science, 315(2–3): 319–333, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bompadre:2004:PES

[1888] A. Bompadre, G. Matera, R. Wachenchauzer, and A. Waissbein. Polynomial equation solving by lifting procedures for ramified fibers. Theoretical Computer Science, 315(2–3):335–369, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burago:2004:ASP

[1889] D. Burago, D. Grigoriev, and A. Slissenko. Approximating shortest path for the skew lines problem in time doubly logarithmic in 1/ϵ. Theoretical Computer Science, 315(2–3):371–404, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Croot:2004:ACC

[1890] Ernie Croot, Ren-Cang Li, and H. J. Hui June Zhu. The abc conjecture and correctly rounded reciprocal square roots. Theoretical Computer Science, 315(2–3):405–417, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vonzurGathen:2004:FAG

[1891] J. Joachim von zur Gathen and Michael Nöcker. Fast arithmetic with general Gauß periods. Theoretical Computer Science, 315(2–3):419–452, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heinig:2004:SAS

[1892] Georg Heinig and K. Karla Rost. Split algorithms for skewsymmetric Toeplitz matrices with arbitrary rank profile. Theoretical Computer Science, 315(2– 3):453–468, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaporin:2004:ACT

[1893] Igor Kaporin. The aggregation and cancellation techniques as a practical tool for faster matrix multiplication. Theoretical Computer Science, 315(2– 3):469–510, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2004:FIT

[1894] Fu-Rong Lin, Wai-Ki Ching, and M. K.

Michael K. Ng. Fast inversion of triangular Toeplitz matrices. Theoretical Computer Science, 315(2–3):511– 523, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malajovich:2004:HPA

[1895] Gregorio Malajovich and J. M. J. Maurice Rojas. High probability analysis of the condition number of sparse polynomial systems. Theoretical Computer Science, 315(2–3):525–555, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Noutsos:2004:MAP

[1896] D. Noutsos, S. Serra Capizzano, and P. Vassalos. Matrix algebra preconditioners for multilevel Toeplitz systems do not insure optimal convergence rate.

Theoretical Computer Science, 315(2– 3):557–579, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pan:2004:IIS

[1897] Victor Y. Pan, M. Marc Van Barel, Xinmao Wang, and Gianni Codevico. Iterative inversion of structured matrices. Theoretical Computer Science, 315 (2–3):581–592, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pardo:2004:DTS

[1898] Luis Miguel Pardo and J. S. Jorge San Martín. Deformation techniques to solve generalised Pham systems. Theoretical Computer Science, 315(2–3): 593–625, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perez-Diaz:2004:PAA

[1899] Sonia Pérez-Díaz, Juana Sendra, and J. R. J. Rafael Sendra. Parametrization of approximate algebraic curves by lines. Theoretical Computer Science, 315(2–3):627–650, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sommese:2004:NFM

[1900] Andrew J. Sommese, Jan Verschelde, and C. W. Charles W. Wampler. Numerical factorization of multivariate complex polynomials. Theoretical Computer Science, 315(2–3):651– 669, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIf

[1901] Anonymous. Author index. Theoretical Computer Science, 315(2–3): 671–672, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Birkedal:2004:P

[1902] Lars Birkedal, Martín Escardó, Achim Jung, and Giuseppe Rosolini. Preface. Theoretical Computer Science, 316(1– 3):1–2, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:2004:CBC

[1903] Jiří Adámek, Stefan Milius, and Jiří Velebil. On coalgebra based on classes. Theoretical Computer Science, 316(1– 3):3–23, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alessi:2004:ITD

[1904] Fabio Alessi, Mariangiola Dezani-Ciancaglini, and Stefania Lusin. Intersection types and domain operators. Theoretical Computer Science, 316(1– 3):25–47, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dezani-Ciancaglini:2004:BIL

[1905] Mariangiola Dezani-Ciancaglini, Silvia Ghilezan, and Silvia Likavec. Behavioural inverse limit λ-models. Theoretical Computer Science, 316(1–3): 49–74, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lawson:2004:IAC

[1906] J. D. Jimmie D. Lawson. Idempotent analysis and continuous semilattices. Theoretical Computer Science, 316(1– 3):75–87, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lawson:2004:PHC

[1907] J. D. Jimmie D. Lawson and Luoshan Xu. Posets having continuous intervals. Theoretical Computer Science, 316(1– 3):89–103, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lawvere:2004:LRA

[1908] F. W. F. William Lawvere. Left and right adjoint operations on spaces and data types. Theoretical Computer Science, 316(1–3):105–111, May 28, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moshier:2004:RBC

[1909] M. A. M. Andrew Moshier. On the relationship between compact regularity and Gentzen’s cut rule. Theoretical Computer Science, 316(1–3):113– 136, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Normann:2004:HTF

[1910] Dag Normann. Hierarchies of total functionals over the reals. Theoretical Computer Science, 316(1–3):137– 151, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nygaard:2004:DTC

[1911] Mikkel Nygaard and Glynn Winskel.

Domain theory for concurrency. Theoretical Computer Science, 316(1–3): 153–190, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reus:2004:SLO

[1912] Bernhard Reus and Thomas Streicher.

Semantics and logic of object calculi. Theoretical Computer Science, 316(1– 3):191–213, May 28, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rummelhoff:2004:PPM

[1913] Ivar Rummelhoff. Polynat in PER models. Theoretical Computer Science, 316(1–3):215–224, May 28, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Townsend:2004:PLP

[1914] C. F. Townsend. Presenting locale pullback via directed complete posets. Theoretical Computer Science, 316(1–3): 225–258, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vickers:2004:ESS

[1915] Steven Vickers. Entailment systems for stably locally compact locales. Theoretical Computer Science, 316(1–3): 259–296, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vickers:2004:UCD

[1916] S. J. Vickers and C. F. Townsend. A universal characterization of the double powerlocale. Theoretical Computer Science, 316(1–3):297–321, May 28, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIg

[1917] Anonymous. Author index. Theoretical Computer Science, 316(1–3): 323, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBg

[1918] Anonymous. Editorial board. Theoretical Computer Science, 316(1–3): v–ix, May 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burgin:2004:TAS

[1919] Mark Burgin and Allen Klinger. Three aspects of super-recursive algorithms and hypercomputation or finding black swans. Theoretical Computer Science, 317(1–3):1–11, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kugel:2004:TTI

[1920] Peter Kugel. Toward a theory of intelligence. Theoretical Computer Science, 317(1–3):13–30, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burgin:2004:ACR

[1921] Mark Burgin. Algorithmic complexity of recursive and inductive algorithms. Theoretical Computer Science, 317(1– 3):31–60, June 4, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wiedermann:2004:CST

[1922] Jiří Wiedermann. Characterizing the super-Turing computing power and efficiency of classical fuzzy Turing machines. Theoretical Computer Science, 317(1–3):61–69, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burgin:2004:EGL

[1923] Mark Burgin and Allen Klinger. Experience, generations, and limits in machine learning. Theoretical Computer Science, 317(1–3):71–91, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kieu:2004:HQA

[1924] T. D. Tien D. Kieu. Hypercomputation with quantum adiabatic processes. Theoretical Computer Science, 317(1– 3):93–104, June 4, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shagrir:2004:STA

[1925] Oron Shagrir. Super-tasks, accelerating Turing machines and uncomputability. Theoretical Computer Science, 317(1– 3):105–114, June 4, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

MacLennan:2004:NCN

[1926] B. J. Bruce J. MacLennan. Natural computation and non-Turing models of computation. Theoretical Computer Science, 317(1–3):115–145, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Campagnolo:2004:CTC

[1927] M. L. Manuel Lameiras Campagnolo. Continuous-time computation with restricted integration capabilities. Theoretical Computer Science, 317(1–3): 147–165, June 4, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bringsjord:2004:MAH

[1928] Selmer Bringsjord and Konstantine Arkoudas. The modal argument for hypercomputing minds. Theoretical Computer Science, 317(1–3):167–190, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wells:2004:HD

[1929] Benjamin Wells. Hypercomputation by definition. Theoretical Computer Science, 317(1–3):191–207, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cleland:2004:CC

[1930] C. E. Carol E. Cleland. The concept of computability. Theoretical Computer Science, 317(1–3):209–225, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kelly:2004:UPI

[1931] K. T. Kevin T. Kelly. Uncomputability: the problem of induction internalized. Theoretical Computer Science, 317(1– 3):227–249, June 4, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Copeland:2004:HPI

[1932] B. J. B. Jack Copeland. Hypercomputation: philosophical issues. Theoretical Computer Science, 317(1–3): 251–267, June 4, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIh

[1933] Anonymous. Author index. Theoretical Computer Science, 317(1–3):269, June 4, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBh

[1934] Anonymous. Editorial board. Theoretical Computer Science, 317(1–3): v–ix, June 4, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marion:2004:E

[1935] Jean-Yves Marion. Editorial. Theoretical Computer Science, 318(1–2):1, June 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aehlig:2004:ANS

[1936] Klaus Aehlig, Ulrich Berger, Martin Hofmann, and Helmut Schwichtenberg. An arithmetic for non-size-increasing polynomial-time computation. Theoretical Computer Science, 318(1–2): 3–27, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baillot:2004:SCS

[1937] Patrick Baillot. Stratified coherence spaces: a denotational semantics for light linear logic. Theoretical Computer Science, 318(1–2):29–55, June 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bellantoni:2004:SNA

[1938] S. Bellantoni and I. Oitavem. Separating NC along the δ axis. Theoretical Computer Science, 318(1–2): 57–78, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benzinger:2004:AHO

[1939] R. Ralph Benzinger. Automated higher-order complexity analysis. Theoretical Computer Science, 318(1–2): 79–103, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danner:2004:MNM

[1940] N. Danner and C. Pollett. Minimization and NP multifunctions. Theoretical Computer Science, 318(1–2): 105–119, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hofmann:2004:RMB

[1941] M. Hofmann and P. J. Scott. Realizability models for BLL-like languages. Theoretical Computer Science, 318(1– 2):121–137, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kristiansen:2004:CCI

[1942] L. Kristiansen and K.-H. Niggl. On the computational complexity of imperative programming languages. Theoretical Computer Science, 318(1–2): 139–161, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lafont:2004:SLL

[1943] Yves Lafont. Soft linear logic and polynomial time. Theoretical Computer Science, 318(1–2):163–180, June 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leivant:2004:IRA

[1944] Daniel Leivant. Intrinsic reasoning about functional programs II: unipolar induction and primitive-recursion. Theoretical Computer Science, 318(1– 2):181–196, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murawski:2004:ISR

[1945] A. S. Murawski and C.-H. L. Ong. On an interpretation of safe recursion in light affine logic. Theoretical Computer Science, 318(1–2):197–223, June 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Royer:2004:CCL

[1946] J. S. (James S.) Royer. On the computational complexity of Longley’s H functional. Theoretical Computer Science, 318(1–2):225–241, June 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBi

[1947] Anonymous. Editorial board. Theoretical Computer Science, 318(1–2): v–ix, June 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Majster-Cederbaum:2004:THV

[1948] Mila Majster-Cederbaum and Frank Salger. Towards the hierarchical verification of reactive systems. Theoretical Computer Science, 318(3):243– 296, June 8, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alur:2004:OPW

[1949] Rajeev Alur, Salvatore La Torre, and George J. Pappas. Optimal paths in weighted timed automata. Theoretical Computer Science, 318(3):297– 322, June 8, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Desharnais:2004:MLM

[1950] Josée Desharnais, Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Metrics for labelled Markov processes. Theoretical Computer Science, 318(3):323–354, June 8, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karner:2004:CMS

[1951] Georg Karner. Continuous monoids and semirings. Theoretical Computer Science, 318(3):355–372, June 8, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Broda:2004:DFB

[1952] Sabine Broda, Luís Damas, Marcelo Finger, and Paulo Silva e Silva. The decidability of a fragment of BB’IW-logic. Theoretical Computer Science, 318(3): 373–408, June 8, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lago:2004:PSD

[1953] Ugo Dal Lago and Simone Martini. Phase semantics and decidability of elementary affine logic. Theoretical Computer Science, 318(3):409–433, June 8, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Banerjee:2004:GIL

[1954] R. N. Banerjee and A. Bujosa. A geometric interpretation of LD-resolution. Theoretical Computer Science, 318(3): 435–470, June 8, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIi

[1955] Anonymous. Author index. Theoretical Computer Science, 318(3):471, June 8, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vuillon:2004:F

[1956] L. Vuillon. Foreword. Theoretical Computer Science, 319(1–3):1, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Korn:2004:TRT

[1957] Michael Korn and Igor Pak. Tilings of rectangles with T-tetrominoes. Theoretical Computer Science, 319(1–3): 3–27, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuo:2004:AGC

[1958] E. H. Eric H. Kuo. Applications of graphical condensation for enumerating matchings and tilings. Theoretical Computer Science, 319(1–3):29– 57, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodini:2004:TTC

[1959] Olivier Bodini and Eric Rémila. Tilings with trichromatic colored-edges triangles. Theoretical Computer Science, 319(1–3):59–70, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Destainville:2004:FNT

[1960] N. Destainville, R. Mosseri, and F. Bailly. A formula for the number of tilings of an octagon by rhombi. Theoretical Computer Science, 319(1–3): 71–81, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Desreux:2004:DTR

[1961] S. Sébastien Desreux, Martin Matamala, Ivan Rapaport, and Eric Rémila. Domino tilings and related models: space of configurations of domains with holes. Theoretical Computer Science, 319(1–3):83–101, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Oger:2004:AMT

[1962] Francis Oger. Algebraic and modeltheoretic properties of tilings. Theoretical Computer Science, 319(1–3): 103–126, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delvenne:2004:QPC

[1963] Jean-Charles Delvenne and V. D. Vincent D. Blondel. Quasi-periodic configurations and undecidable dynamics for tilings, infinite words and Turing machines. Theoretical Computer Science, 319(1–3):127–143, June 10, 2004.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnoux:2004:TDI

[1964] Pierre Arnoux, Valérie Berthé, and Anne Siegel. Two-dimensional iterated morphisms and discrete planes. Theoretical Computer Science, 319(1–3): 145–176, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berthe:2004:LMD

[1965] Valérie Berthé and Robert Tijdeman.

Lattices and multi-dimensional words. Theoretical Computer Science, 319(1– 3):177–202, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brimkov:2004:CDP

[1966] Valentin E. Brimkov and R. P.

Reneta P. Barneva. Connectivity of discrete planes. Theoretical Computer Science, 319(1–3):203–227, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Quas:2004:PLC

[1967] Anthony Quas and Luca Zamboni. Periodicity and local complexity. Theoretical Computer Science, 319(1–3): 229–240, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harriss:2004:CST

[1968] E. O. Harriss and J. S. W. Lamb.

Canonical substitutions tilings of Ammann–Beenker type. Theoretical Computer Science, 319(1–3):241–279, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elkharrat:2004:SGB

[1969] Avi Elkharrat, Christiane Frougny, Jean-Pierre Gazeau, and J.-L. Jean-Louis Verger-Gaugry. Symmetry groups for beta-lattices. Theoretical Computer Science, 319(1–3):281– 305, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grabner:2004:DRL

[1970] P. J. Peter J. Grabner, Clemens Heuberger, and Helmut Prodinger. Distribution results for low-weight binary representations for pairs of integers. Theoretical Computer Science, 319(1– 3):307–331, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fereti:2004:ECP

[1971] Svjetlan Fereti¸’. A q-enumeration of convex polyominoes by the festoon approach. Theoretical Computer Science, 319(1–3):333–356, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Furedi:2004:DGN

[1972] Zoltán Füredi and J.-H. Jeong-Hyun Kang. Distance graph on Z^{n} with ℓ_{1} norm. Theoretical Computer Science, 319(1–3):357–366, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nichitiu:2004:LEP

[1973] Codrin Nichitiu, Christophe Papazian, and Eric Rémila. Leader election in plane cellular automata, only with left–right global convention. Theoretical Computer Science, 319(1–3):367– 384, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Humphreys:2004:CLP

[1974] Katherine Humphreys and Heinrich Niederhausen. Counting lattice paths taking steps in infinitely many directions under special access restrictions. Theoretical Computer Science, 319(1– 3):385–409, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daniel:2004:ICS

[1975] Marc Daniel, Sylvain Gravier, and Julien Moncel. Identifying codes in some subgraphs of the square lattice. Theoretical Computer Science, 319(1– 3):411–421, June 10, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aguilo:2004:SAA

[1976] F. Aguiló, D. Arteaga, and I. Diego.

A symbolical algorithm on additive basis and double-loop networks. Theoretical Computer Science, 319(1–3): 423–439, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Currie:2004:NBW

[1977] J. D. James D. Currie. The number of binary words avoiding abelian fourth powers grows exponentially. Theoretical Computer Science, 319(1–3):441– 446, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frosini:2004:NCT

[1978] A. Frosini and G. Simi. The NPcompleteness of a tomographical problem on bicolored domino tilings. Theoretical Computer Science, 319(1–3): 447–454, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbanchon:2004:UGC

[1979] R. Régis Barbanchon. On unique graph 3-colorability and parsimonious reductions in the plane. Theoretical Computer Science, 319(1–3):455– 482, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIj

[1980] Anonymous. Author index. Theoretical Computer Science, 319(1–3):483– 484, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBj

[1981] Anonymous. Editorial board. Theoretical Computer Science, 319(1–3): v–ix, June 10, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2004:P

[1982] G. Rozenberg. Preface. Theoretical Computer Science, 320(1):1–2, June 12, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arulanandham:2004:FNA

[1983] Joshua J. Arulanandham, Cristian S. Calude, and M. J. Michael J. Dinneen.

A fast natural algorithm for searching. Theoretical Computer Science, 320(1): 3–13, June 12, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biham:2004:QCE

[1984] Eli Biham, Gilles Brassard, Dan Kenigsberg, and T. Tal Mor. Quantum computing without entanglement. Theoretical Computer Science, 320(1): 15–33, June 12, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Condon:2004:CRP

[1985] Anne Condon, Beth Davy, Baharak Rastegari, Shelly Zhao, and Finbarr Tarrant. Classifying RNA pseudoknotted structures. Theoretical Computer Science, 320(1):35–50, June 12, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daley:2004:FLD

[1986] Mark Daley, Lila Kari, and Ian Mc-Quillan. Families of languages defined by ciliate bio-operations. Theoretical Computer Science, 320(1):51– 69, June 12, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diligenti:2004:NCS

[1987] Michelangelo Diligenti, Marco Gori, and Marco Maggini. Neural computation, social networks, and topological spectra. Theoretical Computer Science, 320(1):71–87, June 12, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2004:CCM

[1988] O. H. Oscar H. Ibarra. On the computational complexity of membrane systems. Theoretical Computer Science, 320(1):89–109, June 12, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deAngulo:2004:NLM

[1989] V. R. Vicente Ruiz de Angulo and Carme Torras. Neural learning methods yielding functional invariance. Theoretical Computer Science, 320(1):111– 121, June 12, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Storch:2004:RRR

[1990] Tobias Storch and Ingo Wegener. Real royal road functions for constant population size. Theoretical Computer Science, 320(1):123–134, June 12, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prugel-Bennett:2004:WGA

[1991] Adam Prügel-Bennett. When a genetic algorithm outperforms hill-climbing. Theoretical Computer Science, 320(1): 135–153, June 12, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBk

[1992] Anonymous. Editorial board. Theoretical Computer Science, 320(1):v– ix, June 12, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bezrukov:2004:NSL

[1993] S. Bezrukov, R. Elsässer, B. Monien, R. Preis, and J.-P. Tillich. New spectral lower bounds on the bisection width of graphs. Theoretical Computer Science, 320(2–3):155–174, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elston:2004:GWW

[1994] G. Z. Gillian Z. Elston and Gretchen Ostheimer. On groups whose word problem is solved by a counter automaton. Theoretical Computer Science, 320(2–3):175–185, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:2004:SMO

[1995] Ch. Choffrut and Y. Haddad. Stringmatching with OBDDs. Theoretical Computer Science, 320(2–3):187– 198, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yuan:2004:USM

[1996] J. J. Yuan, Z. H. Liu, C. T. Ng, and T. C. E. Cheng. The unbounded single machine parallel batch scheduling problem with family jobs and release dates to minimize makespan. Theoretical Computer Science, 320(2–3):199– 212, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berthiaume:2004:BFS

[1997] A. André Berthiaume, Todd Bittner, Ljubomir Perković, Amber Settle, and Janos Simon. Bounding the firing synchronization problem on a ring. Theoretical Computer Science, 320(2–3):

213–228, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Powell:2004:NMP

[1998] Olivier Powell. A note on measuring in P. Theoretical Computer Science, 320(2–3):229–246, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Middendorf:2004:CSS

[1999] Martin Middendorf and D. F. David F. Manlove. Combined super-/substring and super-/subsequence problems. Theoretical Computer Science, 320(2– 3):247–267, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blundo:2004:BCU

[2000] Carlo Blundo, Paolo D’Arco, Vanessa Daza, and Carles Padró. Bounds and constructions for unconditionally secure distributed key distribution schemes for general access structures. Theoretical Computer Science, 320(2– 3):269–291, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Domaratzki:2004:DAT

[2001] Michael Domaratzki. Deletion along trajectories. Theoretical Computer Science, 320(2–3):293–313, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2004:SCR

[2002] Arto Salomaa, Derick Wood, and Sheng Yu. On the state complexity of reversals of regular languages. Theoretical Computer Science, 320(2–3):

315–329, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Healy:2004:AML

[2003] Patrick Healy and Ago Kuusik. Algorithms for multi-level graph planarity testing and layout. Theoretical Computer Science, 320(2–3):331– 344, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cocco:2004:HAC

[2004] Simona Cocco and R. Rémi Monasson. Heuristic average-case analysis of the backtrack resolution of random 3-satisfiability instances. Theoretical Computer Science, 320(2–3):345– 372, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dahllof:2004:AFV

[2005] Vilhelm Dahllöf, Peter Jonsson, and Richard Beigel. Algorithms for four variants of the exact satisfiability problem. Theoretical Computer Science, 320(2–3):373–394, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beal:2004:AVG

[2006] M.-P. Marie-Pierre Béal, Anne Bergeron, Sylvie Corteel, and Mathieu Raffinot. An algorithmic view of gene teams. Theoretical Computer Science, 320(2–3):395–418, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2004:NNL

[2007] Alexander Okhotin. On the number of nonterminals in linear conjunctive grammars. Theoretical Computer Science, 320(2–3):419–448, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ben-Hur:2004:PAA

[2008] Asa Ben-Hur, Alexander Roitershtein, and H. T. Hava T. Siegelmann. On probabilistic analog automata. Theoretical Computer Science, 320(2–3): 449–464, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hunziker:2004:CPF

[2009] Markus Hunziker, António Machiavelo, and Jihun Park. Chebyshev polynomials over finite fields and reversibility of σ-automata on square grids. Theoretical Computer Science, 320(2–3): 465–483, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kang:2004:PDI

[2010] Liying Kang, M. Y. Moo Young Sohn, and T. C. E. Cheng. Paired-domination in inflated graphs. Theoretical Computer Science, 320(2–3):485–494, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hitchcock:2004:SS

[2011] J. M. John M. Hitchcock. The size of SPP. Theoretical Computer Science, 320(2–3):495–503, June 14, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIk

[2012] Anonymous. Author index. Theoretical Computer Science, 320(2–3):505– 506, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:MIb

[2013] Anonymous. Master index. Theoretical Computer Science, 320(2–3):507– 518, June 14, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garay:2004:P

[2014] J. A. Juan A. Garay and Sergio Rajsbaum. Preface. Theoretical Computer Science, 321(1):1–3, June 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bender:2004:LAP

[2015] Michael A. Bender and M. Martín Farach-Colton. The Level Ancestor Problem simplified. Theoretical Computer Science, 321(1):5–12, June 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bornstein:2004:FM

[2016] C. F. Claudson F. Bornstein and Santosh Vempala. Flow metrics. Theoretical Computer Science, 321(1):13– 24, June 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bronnimann:2004:SEP

[2017] Hervé Brönnimann, John Iacono, Jyrki Katajainen, Pat Morin, Jason Morrison, and Godfried Toussaint. Spaceefficient planar convex hull algorithms.

Theoretical Computer Science, 321(1): 25–40, June 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carmo:2004:SRP

[2018] R. Carmo, J. Donadelli, Y. Kohayakawa, and E. Laber. Searching in random partially ordered sets. Theoretical Computer Science, 321(1):41– 57, June 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garefalakis:2004:GWP

[2019] Theodoulos Garefalakis. The generalized Weil pairing and the discrete logarithm problem on elliptic curves. Theoretical Computer Science, 321(1): 59–72, June 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hevia:2004:EJV

[2020] Alejandro Hevia and Marcos Kiwi. Electronic jury voting protocols. Theoretical Computer Science, 321(1):73– 94, June 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muthukrishnan:2004:EAS

[2021] S. Muthukrishnan and S. Cenk Sahinalp. An efficient algorithm for sequence comparison with block reversals. Theoretical Computer Science, 321(1):95–101, June 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shachnai:2004:TBO

[2022] Hadas Shachnai and Tami Tamir. Tight bounds for online class-constrained packing. Theoretical Computer Science, 321(1):103–123, June 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stevens:2004:PA

[2023] Brett Stevens and Eric Mendelsohn.

Packing arrays. Theoretical Computer Science, 321(1):125–148, June 16, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szegedy:2004:CBF

[2024] Mario Szegedy and Xiaomin Chen.

Computing Boolean functions from multiple faulty copies of input bits. Theoretical Computer Science, 321(1): 149–170, June 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBl

[2025] Anonymous. Editorial board. Theoretical Computer Science, 321(1):v– ix, June 16, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sales:2004:DFP

[2026] Cláudia Linhares Sales and Frédéric Maffray. On dart-free perfectly contractile graphs. Theoretical Computer Science, 321(2–3):171–194, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2004:ARL

[2027] Keqin Li. Analysis of randomized load distribution for reproduction trees in linear arrays and rings. Theoretical Computer Science, 321(2–3):195– 214, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holdsworth:2004:GTG

[2028] Jason J. Holdsworth. Graph traversal and graph transformation. Theoretical Computer Science, 321(2–3): 215–231, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cavaliere:2004:EON

[2029] Matteo Cavaliere and Peter Leupold. Evolution and observation—a nonstandard way to generate formal languages. Theoretical Computer Science, 321(2–3):233–248, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Downey:2004:KR

[2030] Rodney G. Downey, Evan J. Griffiths, and Stephanie Reid. On Kurtz randomness. Theoretical Computer Science, 321(2–3):249–270, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klazar:2004:LEG

[2031] Martin Klazar. On the least exponential growth admitting uncountably many closed permutation classes. Theoretical Computer Science, 321(2–3): 271–281, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Navarro:2004:ACE

[2032] Gonzalo Navarro and Kimmo Fredriksson. Average complexity of exact and approximate multiple string matching. Theoretical Computer Science, 321(2– 3):283–290, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouyer:2004:UTA

[2033] Patricia Bouyer, Catherine Dufourd, Emmanuel Fleury, and Antoine Petit. Updatable timed automata. Theoretical Computer Science, 321(2–3): 291–345, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esteban:2004:CRB

[2034] Juan Luis Esteban, Nicola Galesi, and Jochen Messner. On the complexity of resolution with bounded conjunctions. Theoretical Computer Science, 321(2– 3):347–370, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nock:2004:DPI

[2035] Richard Nock and Frank Nielsen. On domain-partitioning induction criteria: worst-case bounds for the worst-case based. Theoretical Computer Science, 321(2–3):371–382, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ravikumar:2004:PSS

[2036] B. Ravikumar. Peg-solitaire, string rewriting systems and finite automata. Theoretical Computer Science, 321(2– 3):383–394, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chi:2004:SWS

[2037] Dong Pyo Chi and DoYong Kwon.

Sturmian words, β-shifts, and transcendence. Theoretical Computer Science, 321(2–3):395–404, August 2004.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Albert:2004:LIS

[2038] Michael H. Albert, Alexander Golynski, Angèle M. Hamel, Alejandro López-Ortiz, S. Srinivasa Rao, and Mohammad Ali Safari. Longest increasing subsequences in sliding windows. Theoretical Computer Science, 321(2– 3):405–414, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Woeginger:2004:SLO

[2039] Gerhard J. Woeginger. Seventeen lines and one-hundred-and-one points. Theoretical Computer Science, 321(2–3): 415–421, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIl

[2040] Anonymous. Author index. Theoretical Computer Science, 321(2–3): 423–424, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cegielski:2004:F

[2041] Patrick Cégielski and Malika More. Foreword. Theoretical Computer Science, 322(1):1–3, August 2004. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamowicz:2004:WBP

[2042] Zofia Adamowicz and Leszek Aleksander Kołodziejczyk. Well-behaved principles alternative to bounded induction. Theoretical Computer Science, 322(1):5–16, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beltiukov:2004:SIS

[2043] Anatoly Petrivich Beltiukov. A strong induction scheme that leads to polynomially computable realizations. Theoretical Computer Science, 322(1):17– 39, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chateau:2004:CDU

[2044] A. Chateau. On the complexity of decision using destinies in H-bounded structures. Theoretical Computer Science, 322(1):41–67, August 2004. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Finkel:2004:CPL

[2045] Olivier Finkel. Closure properties of locally finite ω-languages. Theoretical Computer Science, 322(1):69–84, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becher:2004:RTP

[2046] Verónica Becher and Serge Grigorieff.

Recursion and topology on 2^{≤ω} for possibly infinite computations. Theoretical Computer Science, 322(1):85– 136, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hachaichi:2004:ADC

[2047] Yassine Hachaïchi. Arithmetical definability and computational complexity. Theoretical Computer Science, 322(1): 137–146, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Formisano:2004:TVS

[2048] Andrea Formisano, Eugenio G. Omodeo, and Alberto Policriti. Three-variable statements of set-pairing. Theoretical Computer Science, 322(1):147–173, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ressayre:2004:WSI

[2049] Jean-Pierre Ressayre. Winning strategies for infinite games: from large cardinals to computer science extended abstract. Theoretical Computer Science, 322(1):175–179, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shahbazyan:2004:ELA

[2050] K. Shahbazyan and Yu. Shoukourian. EMSO-logic and automata related to homogeneous flow event structures. Theoretical Computer Science, 322(1): 181–201, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Woods:2004:SSC

[2051] Alan R. Woods. Subset sum “cubes” and the complexity of primality testing. Theoretical Computer Science, 322(1): 203–219, August 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zaslavsky:2004:SGP

[2052] I. D. Zaslavsky. On some generalizations of the primitive recursive arithmetic. Theoretical Computer Science, 322(1):221–230, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBm

[2053] Anonymous. Editorial board. Theoretical Computer Science, 322(1):v–ix, August 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demongeot:2004:DAP

[2054] J. Demongeot and J. Mazoyer. Discrete applied problems, cellular automata and lattices: florilegium for E. Goles. Theoretical Computer Science, 322(2):231–232, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nivat:2004:AEG

[2055] M. Nivat. For the 50th anniversary of Eric Goles: A few words by Maurice Nivat. Theoretical Computer Science, 322(2):233–235, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aracena:2004:LCM

[2056] Julio Aracena, Jacques Demongeot, and Eric Goles. On limit cycles of monotone functions with symmetric connection graph. Theoretical Computer Science, 322(2):237–244, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baiou:2004:SAF

[2057] Mourad Baïou and Michel Balinski.

Student admissions and faculty recruitment. Theoretical Computer Science, 322(2):245–265, August 30, 2004.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gajardo:2004:DCA

[2058] A. Gajardo and E. Goles. Dynamics of a class of ants on a one-dimensional lattice. Theoretical Computer Science, 322(2):267–283, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goles:2004:FT

[2059] Eric Goles. Folding and tiling. Theoretical Computer Science, 322(2):285– 296, August 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moreira:2004:GAI

[2060] Andrés Moreira. Genetic algorithms for the imitation of genomic styles in protein backtranslation. Theoretical Computer Science, 322(2):297– 312, August 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Weil:2004:CGC

[2061] Georges Weil, Kamel Heus, Thomas Faraut, and Jacques Demongeot. The cyclic genetic code as a constraint satisfaction problem. Theoretical Computer Science, 322(2):313–334, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delorme:2004:RTR

[2062] M. Delorme and J. Mazoyer. Real-time recognition of languages on an twodimensional Archimedean thread. Theoretical Computer Science, 322(2):335– 354, August 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durr:2004:CAC

[2063] Christoph Dürr, Ivan Rapaport, and Guillaume Theyssier. Cellular automata and communication complexity. Theoretical Computer Science, 322 (2):355–368, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matamala:2004:DCA

[2064] Martín Matamala and Eduardo Moreno.

Dynamic of cyclic automata over Z^{2}. Theoretical Computer Science, 322(2): 369–381, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goles:2004:SML

[2065] Éric Goles, Matthieu Latapy, Clémence Magnien, Michel Morvan, and Ha Duong Phan. Sandpile models and lattices: a comprehensive survey. Theoretical Computer Science, 322(2):383– 407, August 30, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Remila:2004:LSS

[2066] Eric Rémila. The lattice structure of the set of domino tilings of a polygon. Theoretical Computer Science, 322(2): 409–422, August 30, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sassone:2004:P

[2067] Vladimiro Sassone. Preface. Theoretical Computer Science, 322(3):423– 426, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abadi:2004:PA

[2068] Martín Abadi and Cédric Fournet. Private authentication. Theoretical Computer Science, 322(3):427–476, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busi:2004:EPM

[2069] Nadia Busi and Gianluigi Zavattaro. On the expressive power of movement and restriction in pure mobile ambients. Theoretical Computer Science, 322(3):477–515, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caires:2004:SLC

[2070] Luís Caires and Luca Cardelli. A spatial logic for concurrency—II. Theoretical Computer Science, 322(3):517– 565, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chothia:2004:AFT

[2071] Tom Chothia and Dominic Duggan. Abstractions for fault-tolerant global computing. Theoretical Computer Science, 322(3):567–613, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hennessy:2004:TBT

[2072] Matthew Hennessy, Massimo Merro, and Julian Rathke. Towards a behavioural theory of access and mobility control in distributed systems. Theoretical Computer Science, 322(3):615– 669, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIm

[2073] Anonymous. Author index. Theoretical Computer Science, 322(3):671– 672, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jeffrey:2004:TBF

[2074] Alan Jeffrey and Julian Rathke. A theory of bisimulation for a fragment of concurrent ML with local names. Theoretical Computer Science, 323(1–3): 1–48, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Raffalli:2004:GRP

[2075] C. Raffalli. Getting results from programs extracted from classical proofs. Theoretical Computer Science, 323(1– 3):49–70, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cerrito:2004:PMC

[2076] Serenella Cerrito and Delia Kesner.

Pattern matching as cut elimination. Theoretical Computer Science, 323(1– 3):71–127, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baldan:2004:DES

[2077] P. Baldan, N. Busi, A. Corradini, and G. M. Pinna. Domain and event structure semantics for Petri nets with read and inhibitor arcs. Theoretical Computer Science, 323(1–3):129– 189, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hierons:2004:TCD

[2078] R. M. Hierons and M. Harman. Testing conformance of a deterministic implementation against a nondeterministic stream X-machine. Theoretical Computer Science, 323(1–3): 191–233, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Porter:2004:ISK

[2079] Timothy Porter. Interpreted systems and Kripke models for multiagent systems from a categorical perspective. Theoretical Computer Science, 323(1– 3):235–266, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peltier:2004:FOT

[2080] Nicolas Peltier. The first order theory of primal grammars is decidable. Theoretical Computer Science, 323(1– 3):267–320, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2004:RPD

[2081] Guo-Qiang Zhang and William C. Rounds. Reasoning with power defaults. Theoretical Computer Science, 323(1–3):321–350, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dehlinger:2004:FGM

[2082] Christophe Dehlinger and Jean-François Dufourd. Formalizing generalized maps in Coq. Theoretical Computer Science, 323(1–3):351–397, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dehlinger:2004:FTT

[2083] Christophe Dehlinger and Jean-François Dufourd. Formalizing the trading theorem in Coq. Theoretical Computer Science, 323(1–3):399– 442, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Urso:2004:SGM

[2084] Pascal Urso and Emmanuel Kounalis.

Sound generalizations in mathematical induction. Theoretical Computer Science, 323(1–3):443–471, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Urban:2004:NU

[2085] Christian Urban, Andrew M. Pitts, and Murdoch J. Gabbay. Nominal unification. Theoretical Computer Science, 323(1–3):473–497, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIn

[2086] Anonymous. Author index. Theoretical Computer Science, 323(1–3): 499, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBn

[2087] Anonymous. Editorial board. Theoretical Computer Science, 323(1–3): v–ix, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ito:2004:F

[2088] Masami Ito. Foreword. Theoretical Computer Science, 324(1):1, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:2004:IS

[2089] Zoltán Ésik and Werner Kuich. Inductive ^{∗}-semirings. Theoretical Computer Science, 324(1):3–33, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2004:MAD

[2090] Tero Harju and Juhani Karhumäki. Many aspects of defect theorems. Theoretical Computer Science, 324(1):35– 54, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ogawa:2004:WQO

[2091] Mizuhito Ogawa. Well-quasi-orders and regularω-languages. Theoretical Computer Science, 324(1):55–60, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paun:2004:PMD

[2092] Gheorghe Păun, Yasuhiro Suzuki, Hiroshi Tanaka, and Takashi Yokomori. On the power of membrane division in P systems. Theoretical Computer Science, 324(1):61–85, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Petkovic:2004:UAS

[2093] Tatjana Petković, Miroslav Ćirić, and Stojan Bogdanović. Unary algebras, semigroups and congruences on free semigroups. Theoretical Computer Science, 324(1):87–105, September 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schott:2004:TOP

[2094] René Schott and Jean-Claude Spehner.

Two optimal parallel algorithms on the commutation class of a word. Theoretical Computer Science, 324(1):107– 131, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBo

[2095] Anonymous. Editorial board. Theoretical Computer Science, 324(1):v– ix, September 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiat:2004:F

[2096] Amos Fiat and Sandy Irani. Foreword. Theoretical Computer Science, 324(2–3):133–135, September 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blum:2004:OLO

[2097] Avrim Blum, Vijay Kumar, Atri Rudra, and Felix Wu. Online learning in online auctions. Theoretical Computer Science, 324(2–3):137–146, September 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Augustine:2004:LTA

[2098] John E. Augustine and Steven Seiden.

Linear time approximation schemes for vehicle scheduling problems. Theoretical Computer Science, 324(2–3):147– 160, September 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kesselman:2004:HBM

[2099] Alexander Kesselman and Yishay Mansour. Harmonic buffer management policy for shared memory switches. Theoretical Computer Science, 324(2– 3):161–182, September 20, 2004. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mendel:2004:OCC

[2100] M. Mendel and Steven S. Seiden. Online companion caching. Theoretical Computer Science, 324(2–3):183–200, September 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feder:2004:CRS

[2101] Tomás Feder, Rajeev Motwani, Rina Panigrahy, Steve Seiden, Rob van Stee, and An Zhu. Combining request scheduling with web caching. Theoretical Computer Science, 324(2–3):201– 218, September 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fleischer:2004:NRO

[2102] Rudolf Fleischer, Włodzimierz Głazek, and Steve Seiden. New results for online page replication. Theoretical Computer Science, 324(2–3):219–251, September 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lempel:2004:CCQ

[2103] Ronny Lempel and Shlomo Moran. Competitive caching of query results in search engines. Theoretical Computer Science, 324(2–3):253–271, September 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bose:2004:COR

[2104] Prosenjit Bose and Pat Morin. Competitive online routing in geometric graphs. Theoretical Computer Science, 324(2–3):273–288, September 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:2004:WSP

[2105] Marek Chrobak and Jiří Sgall. The weighted 2-server problem. Theoretical Computer Science, 324(2–3):289– 312, September 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Awerbuch:2004:LGS

[2106] Baruch Awerbuch, Yossi Azar, and Yair Bartal. On-line generalized Steiner problem. Theoretical Computer Science, 324(2–3):313–324, September 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becchetti:2004:SCS

[2107] Luca Becchetti, Stefano Leonardi, Alberto Marchetti-Spaccamela, and Kirk Pruhs. Semi-clairvoyant scheduling. Theoretical Computer Science, 324(2– 3):325–335, September 20, 2004. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartal:2004:CRW

[2108] Yair Bartal and Elias Koutsoupias. On the competitive ratio of the work function algorithm for the k-server problem. Theoretical Computer Science, 324(2–3):337–345, September 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koutsoupias:2004:CPO

[2109] Elias Koutsoupias and David Scot Taylor. The CNN problem and other kserver variants. Theoretical Computer Science, 324(2–3):347–359, September 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2004:BEV

[2110] Peter Chen and Guoli Ding. The best expert versus the smartest algorithm. Theoretical Computer Science, 324(2–3):361–380, September 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIo

[2111] Anonymous. Author index. Theoretical Computer Science, 324(2–3):381– 382, September 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Priami:2004:P

[2112] Corrado Priami. Preface. Theoretical Computer Science, 325(1):1–2, September 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eveillard:2004:MSC

[2113] Damien Eveillard, Delphine Ropers, Hidde de Jong, Christiane Branlant, and Alexander Bockmayr. A multiscale constraint programming model of alternative splicing regulation. Theoretical Computer Science, 325(1):3–24, September 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chabrier-Rivier:2004:MQB

[2114] Nathalie Chabrier-Rivier, Marc Chiaverini, Vincent Danos, François Fages, and Vincent Schächter. Modeling and querying biomolecular interaction networks. Theoretical Computer Science, 325(1):25–44, September 28, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Antoniotti:2004:TCB

[2115] M. Antoniotti, C. Piazza, A. Policriti, M. Simeoni, and B. Mishra. Taming the complexity of biochemical models through bisimulation and collapsing: theory and practice. Theoretical Computer Science, 325(1):45–67, September 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danos:2004:FMB

[2116] Vincent Danos and Cosimo Laneve.

Formal molecular biology. Theoretical Computer Science, 325(1):69–110, September 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Curti:2004:MBP

[2117] M. Curti, P. Degano, C. Priami, and C. T. Baldari. Modelling biochemical pathways through enhanced π-calculus. Theoretical Computer Science, 325(1): 111–140, September 28, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Regev:2004:BAB

[2118] Aviv Regev, Ekaterina M. Panina, William Silverman, Luca Cardelli, and Ehud Shapiro. Bioambients: an abstraction for biological compartments. Theoretical Computer Science, 325(1): 141–167, September 28, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBp

[2119] Anonymous. Editorial board. Theoretical Computer Science, 325(1):v– ix, September 28, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2004:P

[2120] Jarkko Kari. Preface. Theoretical Computer Science, 325(2):169, October 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barbe:2004:PMH

[2121] A. Barbé and F. von Haeseler. The Pascal matroid as a home for generating sets of cellular automata configurations defined by quasigroups. Theoretical Computer Science, 325(2):171– 214, October 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boykett:2004:EEL

[2122] Tim Boykett. Efficient exhaustive listings of reversible one dimensional cellular automata. Theoretical Computer Science, 325(2):215–247, October 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cattaneo:2004:SSC

[2123] Gianpiero Cattaneo, Alberto Dennunzio, and Luciano Margara. Solution of some conjectures about topological properties of linear cellular automata. Theoretical Computer Science, 325(2): 249–271, October 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2004:SIN

[2124] Eugen Czeizler. On the size of the inverse neighborhoods for onedimensional reversible cellular automata. Theoretical Computer Science, 325(2):273–284, October 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moreira:2004:CMO

[2125] Andrés Moreira, Nino Boccara, and Eric Goles. On conservative and monotone one-dimensional cellular automata and their particle representation. Theoretical Computer Science, 325(2):285– 316, October 1, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sutner:2004:CRC

[2126] K. Sutner. The complexity of reversible cellular automata. Theoretical Computer Science, 325(2):317–328, October 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Toffoli:2004:HTS

[2127] Tommaso Toffoli, Silvio Capobianco, and Patrizia Mentrasti. How to turn a second-order cellular automaton into a lattice gas: a new inversion scheme. Theoretical Computer Science, 325(2):

329–344, October 1, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Warnow:2004:P

[2128] Tandy Warnow and Binhai Zhu. Preface. Theoretical Computer Science, 325 (3):345–346, October 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marron:2004:GDU

[2129] Mark Marron, Krister M. Swenson, and Bernard M. E. Moret. Genomic distances under deletions and insertions. Theoretical Computer Science, 325(3): 347–360, October 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2004:SEA

[2130] Zhi-Zhong Chen, Yong Gao, Guohui Lin, Robert Niewiadomski, Yang Wang, and Junfeng Wu. A spaceefficient algorithm for sequence alignment with inversions and reversals. Theoretical Computer Science, 325(3): 361–372, October 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutz:2004:CBM

[2131] Martin Kutz. The complexity of Boolean matrix root computation. Theoretical Computer Science, 325(3): 373–390, October 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schwabe:2004:EDM

[2132] Eric J. Schwabe and Ian M. Sutherland. Efficient data mappings for parity-declustered data layouts. Theoretical Computer Science, 325(3):391– 407, October 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chin:2004:ADR

[2133] Francis Y. L. Chin, Xiaotie Deng, Qizhi Fang, and Shanfeng Zhu. Approximate and dynamic rank aggregation. Theoretical Computer Science, 325(3):409– 424, October 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asano:2004:SNG

[2134] Tetsuo Asano, Naoki Katoh, Hisao Tamaki, and Takeshi Tokuyama. The structure and number of global roundings of a graph. Theoretical Computer Science, 325(3):425–437, October 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Halldorsson:2004:RAS

[2135] Magnús M. Halldórsson, Kazuo Iwama, Shuichi Miyazaki, and Hiroki Yanagisawa. Randomized approximation of the stable marriage problem. Theoretical Computer Science, 325(3):439– 465, October 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chin:2004:ICA

[2136] Francis Y. L. Chin and Stanley P. Y.

Fung. Improved competitive algorithms for online scheduling with partial job values. Theoretical Computer Science, 325(3):467–478, October 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2004:SBD

[2137] Jae-Hoon Kim and Kyung-Yong Chwa. Scheduling broadcasts with deadlines. Theoretical Computer Science, 325(3): 479–488, October 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIp

[2138] Anonymous. Author index. Theoretical Computer Science, 325(3):489– 490, October 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2004:LBH

[2139] Lane A. Hemaspaandra and Mayur Thakur. Lower bounds and the hardness of counting properties. Theoretical Computer Science, 326(1–3):1– 28, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chauve:2004:MIO

[2140] Cedric Chauve and Guillaume Fertin. On maximal instances for the original syntenic distance. Theoretical Computer Science, 326(1–3):29–43, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michel:2004:STM

[2141] Pascal Michel. Small Turing machines and generalized busy beaver competition. Theoretical Computer Science, 326(1–3):45–56, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eisenbrand:2004:CFP

[2142] Friedrich Eisenbrand and Fabrizio Grandoni. On the complexity of fixed parameter clique and dominating set. Theoretical Computer Science, 326(1– 3):57–67, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fill:2004:LDA

[2143] James Allen Fill and Nevin Kapur. Limiting distributions for additive functionals on Catalan trees. Theoretical Computer Science, 326(1–3):69– 102, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ganjali:2004:CNS

[2144] Yashar Ganjali and MohammadTaghi Hajiaghayi. Characterization of networks supporting multi-dimensional linear interval routing schemes. Theoretical Computer Science, 326(1–3): 103–116, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matsui:2004:RGC

[2145] Tomomi Matsui, Yasuko Matsui, and Yoko Ono. Random generation of 2 × 2 ⋯ × 2 × J contingency tables. Theoretical Computer Science, 326(1–3): 117–135, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bockenhauer:2004:HCM

[2146] Hans-Joachim Böckenhauer, Dirk Bongartz, Juraj Hromkovič, Ralf Klasing, Guido Proietti, Sebastian Seibert, and Walter Unger. On the hardness of constructing minimal 2connected spanning subgraphs in complete graphs with sharpened triangle inequality. Theoretical Computer Science, 326(1–3):137–153, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gonzalez:2004:CPS

[2147] Teofilo F. Gonzalez and David Serena. Complexity of pairwise shortest path routing in the grid. Theoretical Computer Science, 326(1–3):155– 185, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ziegler:2004:CLA

[2148] Martin Ziegler and Vasco Brattka. Computability in linear algebra. Theoretical Computer Science, 326(1–3): 187–211, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janson:2004:ASC

[2149] Svante Janson, Stefano Lonardi, and Wojciech Szpankowski. On average sequence complexity. Theoretical Computer Science, 326(1–3):213– 227, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duval:2004:LTC

[2150] Jean-Pierre Duval, Roman Kolpakov, Gregory Kucherov, Thierry Lecroq, and Arnaud Lefebvre. Linear-time computation of local periods. Theoretical Computer Science, 326(1–3):229– 240, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edson:2004:RPI

[2151] Marcia Edson and Luca Q. Zamboni.

On representations of positive integers in the Fibonacci base. Theoretical Computer Science, 326(1–3):241– 260, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiala:2004:DCL

[2152] Jiří Fiala, Aleksei V. Fishkin, and Fedor Fomin. On distance constrained labeling of disk graphs. Theoretical Computer Science, 326(1–3):261– 292, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jelenkovic:2004:LRU

[2153] Predrag R. Jelenković and Ana Radovanović. Least-recently-used caching with dependent requests. Theoretical Computer Science, 326(1–3): 293–327, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kiwi:2004:CHP

[2154] Marcos Kiwi and Alexander Russell.

The Chilean Highway Problem. Theoretical Computer Science, 326(1–3): 329–342, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dessmark:2004:OGE

[2155] Anders Dessmark and Andrzej Pelc.

Optimal graph exploration without good maps. Theoretical Computer Science, 326(1–3):343–362, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yi:2004:AKA

[2156] Xun Yi. Authenticated key agreement in dynamic peer groups. Theoretical Computer Science, 326(1–3):363– 382, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bernholt:2004:CCM

[2157] Thorsten Bernholt and Paul Fischer. The complexity of computing the MCD-estimator. Theoretical Computer Science, 326(1–3):383–398, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2004:PEE

[2158] Minming Li, Shawn L. Huang, Xiaoming Sun, and Xiao Huang. Performance evaluation for energy efficient topologic control in ad hoc wireless networks. Theoretical Computer Science, 326(1–3):399–408, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gibbons:2004:RSE

[2159] Alan Gibbons and Paul Sant. Rotation sequences and edge-colouring of binary tree pairs. Theoretical Computer Science, 326(1–3):409–418, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2004:UCF

[2160] Qi Cheng. On the ultimate complexity of factorials. Theoretical Computer Science, 326(1–3):419–429, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leonardi:2004:CMC

[2161] Stefano Leonardi and Guido Schäfer.

Cross-monotonic cost sharing methods for connected facility location games. Theoretical Computer Science, 326(1– 3):431–442, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIq

[2162] Anonymous. Author index. Theoretical Computer Science, 326(1–3):443– 444, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBq

[2163] Anonymous. Editorial board. Theoretical Computer Science, 326(1–3):v– ix, October 20, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gumm:2004:P

[2164] H. Peter Gumm. Preface. Theoretical Computer Science, 327(1–2):1– 2, October 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartels:2004:HPS

[2165] Falk Bartels, Ana Sokolova, and Erik de Vink. A hierarchy of probabilistic system types. Theoretical Computer Science, 327(1–2):3–22, October 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2004:CAK

[2166] Hubie Chen and Riccardo Pucella. A coalgebraic approach to Kleene algebra with tests. Theoretical Computer Science, 327(1–2):23–44, October 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cirstea:2004:CAD

[2167] Corina Cîrstea. A compositional approach to defining logics for coalgebras. Theoretical Computer Science, 327(1– 2):45–69, October 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hughes:2004:SC

[2168] Jesse Hughes and Bart Jacobs. Simulations in coalgebra. Theoretical Computer Science, 327(1–2):71–108, October 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kupke:2004:SC

[2169] Clemens Kupke, Alexander Kurz, and Yde Venema. Stone coalgebras. Theoretical Computer Science, 327(1–2): 109–134, October 25, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lenisa:2004:CTO

[2170] Marina Lenisa, John Power, and Hiroshi Watanabe. Category theory for operational semantics. Theoretical Computer Science, 327(1–2):135– 154, October 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matthes:2004:SNW

[2171] Ralph Matthes and Tarmo Uustalu. Substitution in non-wellfounded syntax with variable binding. Theoretical Computer Science, 327(1–2):155– 174, October 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Palmigiano:2004:CVP

[2172] Alessandra Palmigiano. A coalgebraic view on positive modal logic. Theoretical Computer Science, 327(1–2):175– 195, October 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rosu:2004:BAH

[2173] Grigore Roşu. Behavioral abstraction is hiding information. Theoretical Computer Science, 327(1–2):197– 221, October 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBr

[2174] Anonymous. Editorial board. Theoretical Computer Science, 327(1–2):v– ix, October 25, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:2004:F

[2175] Zoltán Ésik and Zoltán Fülöp. Foreword. Theoretical Computer Science, 327(3):223–224, November 2, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ananichev:2004:SMA

[2176] D. S. Ananichev and M. V. Volkov.

Synchronizing monotonic automata. Theoretical Computer Science, 327(3): 225–239, November 2, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Champarnaud:2004:NRA

[2177] J.-M. Champarnaud and F. Coulon. NFA reduction algorithms by means of regular inequalities. Theoretical Computer Science, 327(3):241–253, November 2, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [2226].

DAlessandro:2004:WQO

[2178] Flavio D’Alessandro and Stefano Varricchio. Well quasi-orders and contextfree grammars ^{,}. Theoretical Computer Science, 327(3):255–268, November 2, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deFalco:2004:FSO

[2179] Diego de Falco, Massimiliano Goldwurm, and Violetta Lonati. Frequency of symbol occurrences in bicomponent stochastic models. Theoretical Computer Science, 327(3):269– 300, November 2, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hofbauer:2004:DSR

[2180] Dieter Hofbauer and Johannes Waldmann. Deleting string rewriting systems preserve regularity. Theoretical Computer Science, 327(3):301–317, November 2, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holzer:2004:DFA

[2181] Markus Holzer and Barbara König. On deterministic finite automata and syntactic monoid size. Theoretical Computer Science, 327(3):319–347, November 2, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klimann:2004:DUS

[2182] Ines Klimann, Sylvain Lombardy, Jean Mairesse, and Christophe Prieur. Deciding unambiguity and sequentiality from a finitely ambiguous max-plus automaton. Theoretical Computer Science, 327(3):349–373, November 2, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malcher:2004:MFA

[2183] Andreas Malcher. Minimizing finite automata is computationally hard. Theoretical Computer Science, 327(3):375– 390, November 2, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIr

[2184] Anonymous. Author index. Theoretical Computer Science, 327(3):391– 392, November 2, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2004:E

[2185] Oscar H. Ibarra. Editorial. Theoretical Computer Science, 328(1–2):1, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allauzen:2004:OPD

[2186] Cyril Allauzen and Mehryar Mohri.

An optimal pre-determinization algorithm for weighted transducers. Theoretical Computer Science, 328(1–2):3– 18, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2004:CPF

[2187] Xiang Fu, Tevfik Bultan, and Jianwen Su. Conversation protocols: a formalism for specification and verification of reactive electronic services. Theoretical Computer Science, 328(1–2):19– 37, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guingne:2004:ANM

[2188] Franck Guingne, Florent Nicart, and André Kempe. Acyclic networks maximizing the printing complexity. Theoretical Computer Science, 328(1–2):39– 51, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuske:2004:BAC

[2189] Dietrich Kuske and Ingmar Meinecke. Branching automata with costs — a way of reflecting parallelism in costs. Theoretical Computer Science, 328(1– 2):53–75, November 29, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lombardy:2004:IVA

[2190] Sylvain Lombardy, Yann Régis-Gianas, and Jacques Sakarovitch. Introducing VAUCANSON. Theoretical Computer Science, 328(1–2):77–96, November 29, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Miyamoto:2004:TDA

[2191] Satoru Miyamoto, Shunsuke Inenaga, Masayuki Takeda, and Ayumi Shinohara. Ternary directed acyclic word graphs. Theoretical Computer Science, 328(1–2):97–111, November 29, 2004.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ravikumar:2004:WMD

[2192] B. Ravikumar and G. Eisman. Weak minimization of DFA — an algorithm and applications. Theoretical Computer Science, 328(1–2):113–133, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tamm:2004:BAM

[2193] Hellis Tamm and Esko Ukkonen. Bideterministic automata and minimal representations of regular languages. Theoretical Computer Science, 328(1–2): 135–149, November 29, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Trahtman:2004:RTC

[2194] A. N. Trahtman. Reducing the time complexity of testing for local threshold testability. Theoretical Computer Science, 328(1–2):151–160, November 29, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanZijl:2004:BNS

[2195] Lynette van Zijl. On binary ⊕-NFAs and succinct descriptions of regular languages. Theoretical Computer Science, 328(1–2):161–170, November 29, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vilares:2004:FFR

[2196] M. Vilares, V. M. Darriba, J. Vilares, and F. J. Ribadas. A formal frame for robust parsing. Theoretical Computer Science, 328(1–2):171– 186, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2004:RSC

[2197] Farn Wang and Hsu-Chun Yen. Reachability solution characterization of parametric real-time systems. Theoretical Computer Science, 328(1–2):187– 201, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xie:2004:LRP

[2198] Gaoyan Xie, Cheng Li, and Zhe Dang. Linear reachability problems and minimal solutions to linear Diophantine equation systems. Theoretical Computer Science, 328(1–2):203– 219, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBVa

[2199] Anonymous. Editorial board. Theoretical Computer Science, 328(1–2):v– ix, November 29, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alvarez-Manilla:2004:PPS

[2200] Mauricio Alvarez-Manilla, Achim Jung, and Klaus Keimel. The probabilistic powerdomain for stably compact spaces. Theoretical Computer Science, 328(3):221–244, December 3, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2004:SBP

[2201] Hejiao Huang, To yat Cheung, and Wai Ming Mak. Structure and behavior preservation by Petri-net-based refinements in system design. Theoretical Computer Science, 328(3):245– 269, December 3, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shen:2004:EGS

[2202] Yi-Dong Shen, Jia-Huai You, and Li-Yan Yuan. Enhancing global SLS-resolution with loop cutting and tabling mechanisms. Theoretical Computer Science, 328(3):271–287, December 3, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baillot:2004:TIL

[2203] Patrick Baillot. Type inference for light affine logic via constraints on words. Theoretical Computer Science, 328(3): 289–323, December 3, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hahnle:2004:LRN

[2204] Reiner Hähnle, Neil V. Murray, and Erik Rosenthal. Linearity and regularity with negation normal form. Theoretical Computer Science, 328(3):325– 354, December 3, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Foy:2004:DSW

[2205] John Foy. A dynamical system which must be stable whose stability cannot be proved. Theoretical Computer Science, 328(3):355–361, December 3, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIs

[2206] Anonymous. Author index. Theoretical Computer Science, 328(3):363– 364, December 3, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Coja-Oghlan:2004:TCA

[2207] Amin Coja-Oghlan, Andreas Goerdt, André Lanka, and Frank Schädlich. Techniques from combinatorial approximation algorithms yield efficient algorithms for random 2k-SAT. Theoretical Computer Science, 329(1–3):1– 45, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drmota:2004:RCB

[2208] Michael Drmota. On Robson’s convergence and boundedness conjectures concerning the height of binary search trees. Theoretical Computer Science, 329(1–3):47–70, December 13, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chockler:2004:RLT

[2209] Hana Chockler and Orna Kupferman. ω-Regular languages are testable with a constant number of queries. Theoretical Computer Science, 329(1–3):71– 92, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jonsson:2004:RFV

[2210] Peter Jonsson and Andrei Krokhin. Recognizing frozen variables in constraint satisfaction problems. Theoretical Computer Science, 329(1–3):93– 113, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hornung:2004:STT

[2211] Tamás Hornung and Sándor Vágvölgyi.

Storage-to-tree transducers with lookahead. Theoretical Computer Science, 329(1–3):115–158, December 13, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Strahm:2004:PTC

[2212] Thomas Strahm. A proof-theoretic characterization of the basic feasible functionals. Theoretical Computer Science, 329(1–3):159–176, December 13, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet-Sadri:2004:COP

[2213] F. Blanchet-Sadri. Codes, orderings, and partial words. Theoretical Computer Science, 329(1–3):177–202, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tajima:2004:PTL

[2214] Yasuhiro Tajima, Etsuji Tomita, Mitsuo Wakatsuki, and Matsuaki Terada. Polynomial time learning of simple deterministic languages via queries and a representative sample. Theoretical Computer Science, 329(1–3):203–221, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bostan:2004:CME

[2215] Alin Bostan and Éric Schost. On the complexities of multipoint evaluation and interpolation. Theoretical Computer Science, 329(1–3):223–235, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morohashi:2004:SMM

[2216] Gembu Morohashi and Shigeki Iwata. Some minimum merging networks. Theoretical Computer Science, 329(1– 3):237–250, December 13, 2004. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peyriere:2004:FCM

[2217] Jacques Peyrière, Bo Tan, Zhi-Xiong Wen, and Jun Wu. The factor composition matrix of sequences. Theoretical Computer Science, 329(1–3):251– 269, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kontorovich:2004:UDG

[2218] Leonid Kontorovich. Uniquely decodable n-gram embeddings. Theoretical Computer Science, 329(1–3):271– 284, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malewicz:2004:TAN

[2219] Grzegorz Malewicz. A tight analysis and near-optimal instances of the algorithm of Anderson and Woll. Theoretical Computer Science, 329(1–3):285– 301, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brueggemann:2004:IDL

[2220] Tobias Brueggemann and Walter Kern. An improved deterministic local search algorithm for 3-SAT. Theoretical Computer Science, 329(1–3):303–313, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dalmau:2004:CCH

[2221] Víctor Dalmau and Peter Jonsson.

The complexity of counting homomorphisms seen from the other side. Theoretical Computer Science, 329(1–3): 315–323, December 13, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ruan:2004:GAM

[2222] Lu Ruan, Hongwei Du, Xiaohua Jia, Weili Wu, Yingshu Li, and Ker-I Ko. A greedy approximation for minimum connected dominating sets. Theoretical Computer Science, 329(1–3):325– 330, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zheng:2004:GXL

[2223] Yuliang Zheng and Xian-Mo Zhang.

The Generalized XOR Lemma. Theoretical Computer Science, 329(1–3): 331–337, December 13, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:AIt

[2224] Anonymous. Author index. Theoretical Computer Science, 329(1–3):339– 340, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2004:EBVb

[2225] Anonymous. Editorial board. Theoretical Computer Science, 329(1–3):v– ix, December 13, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Champarnaud:2005:ENR

[2226] J.-M. Champarnaud and F. Coulon. Erratum to “NFA reduction algorithms by means of regular inequalities” [Theoret. Comput. Sci. 327 (2004) 241– 253]. Theoretical Computer Science, 347(1–2):437–440, November 30, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [2177].

Krumke:2006:ENO

[2227] Sven O. Krumke, Willem E. de Paepe, Diana Poensgen, and Leen Stougie. Erratum to “News from the online traveling repairman” [Theoret. Comput. Sci. 295 (1–3) (2003) 279–294]. Theoretical Computer Science, 352(1–3):347– 348, March 7, 2006. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1339].

Yokomori:2007:EPT

[2228] Takashi Yokomori. Erratum to “Polynomial-time identification of very simple grammars from positive data” [Theoret. Comput. Sci. 298 (2003) 179–206]. Theoretical Computer Science, 377(1–3):282–283, May 31, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1413].

Blanqui:2020:CID

[2229] Frédéric Blanqui, Jean-Pierre Jouannaud, and Mitsuhiro Okada. Corrigendum to “Inductive-data-type systems” [Theoret. Comput. Sci. 272 (1–2) (2002) 41–68]. Theoretical Computer Science, 817(??):81– 82, May 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300306. See